### v2

# Semiclassical measure of the propagation between two topological insulators

Abstract
We study the propagation of a family of initial conditions in the presence of two topological insulators without magnetic field where the interface is a smooth connected curve not compact without boundaries. The solution is governed by an adiabatic modulation of a Dirac operator with a variable mass. We determine the evolution of the semiclassical measure of the solution with a two-scale Wigner measure method by reducing the Dirac operator to a normal form.
Keywords : propagation, Dirac operator, semiclassical analysis, Wigner measure, semiclassical measure valued in a separable Hilbert space, time dependent two-scale semiclassical measure.
$2024$ Mathematics Subject Classification : 35B40, 35F05, 35Q40.
Acknowledgments. The author is very grateful to his advisors Clotilde Fermanian-Kammerer and Nicolas Raymond for many insightful discussions and valuable suggestions. The author acknowledges the support of the Région Pays de la Loire via the Connect Talent Project HiFrAn 2022 07750, and from the France 2030 program, Centre Henri Lebesgue ANR-11-LABX-0020-01.
1 Introduction
1.1 Topological insulators and Dirac type equations
Topological insulators are materials that are insulating in their bulk but support robust currents along their boundaries, meaning that electrons can only move along the surface of the material. The Dirac equation has been derived in the context of topological insulators for honeycomb structures [9, 8, 5] such as graphene [22, 18, 7]. From a global point of view, Dirac-type equations are often a simpler continuum macroscopic description of transport in a narrow energy band near the band crossing [2, 12, 24].
In this paper we consider the following system of the Dirac evolution equation on $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$,
$\left\{\begin{array}[]{ll}\Big{(}\varepsilon\mathrm{D}_{t}+\mathrm{m}(x)\sigma_{3}+\varepsilon\mathrm{D}_{1}\sigma_{1}+\varepsilon\mathrm{D}_{2}\sigma_{2}\Big{)}\psi^{\varepsilon}(t,x)=0,&\forall(t,x)\in\mathbb{R}\times\mathbb{R}^{2},\\ \psi^{\varepsilon}(0,x)=\psi^{\varepsilon}_{0}(x),&\forall x\in\mathbb{R}^{2},\end{array}\right.$ (1)
where ${\varepsilon}$ is a small positive parameter, called semiclassical parameter. The function $\psi^{\varepsilon}(t,\cdot)$ also denoted $\psi^{\varepsilon}_{t}$ is interpreted as the wave function of an electron at time $t$ in a material consisting in two topological insulators. The family ${\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}}$ is a uniformly bounded (in $\varepsilon$) family of ${\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}$ of initial wave functions of the electron. The function $x\mapsto\mathrm{m}(x)$ is a smooth real-valued function (i.e. $\mathrm{m}\in\mathscr{C}^{\infty}\left(\mathbb{R}^{2},\mathbb{R}\right)$) and all its derivatives are bounded. The matrices ${\left(\sigma_{j}\right)_{1\leqslant j\leqslant 3}}$ are the Pauli matrices
$\sigma_{1}\coloneqq\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\;,\;\sigma_{2}\coloneqq\begin{pmatrix}0&-\mathrm{i}\\ \mathrm{i}&0\end{pmatrix}\;,\;\sigma_{3}\coloneqq\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}.$
We shall use the notation
$\mathrm{H}_{\varepsilon}\coloneqq\begin{pmatrix}\mathrm{m}(x)&\varepsilon\mathrm{D}_{1}-\mathrm{i}\varepsilon\mathrm{D}_{2}\\ \varepsilon\mathrm{D}_{1}+\mathrm{i}\varepsilon\mathrm{D}_{2}&-\mathrm{m}(x)\end{pmatrix}.$
Then the equation (1) rewrites $\varepsilon\mathrm{D}_{t}\psi_{t}^{\varepsilon}+\mathrm{H}_{\varepsilon}\psi_{t}^{\varepsilon}=0$.
In this model, the function $\mathrm{m}$ splits the space into two parts, which are the two topological insulators. From a modeling point of view, the function $\mathrm{m}$ is positive on one of the topological insulators and negative on the other so the interface between the two topological insulators is defined by
$\mathrm{E}\coloneqq\left\{x\in\mathbb{R}^{2}\ |\ \mathrm{m}(x)=0\right\}.$
We assume the set $\mathrm{E}$ is a connected curve not compact without boundaries. From a physical point of view, the sign of $\mathrm{m}$ is the insulating phase of a material. This generates a unidirectional propagation along the interface $\mathrm{E}$. This propagation is one of the most important issue of physical applications of topological insulators and superconductors [2, 24], photonics [19, 22, 23], acoustics [21, 20] and fluid mechanics [13].
Our aim is to describe the solution $\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}$ to equation (1) for initial conditions in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$. At least, we can understand limits as $\varepsilon$ goes to 00 by constructing semiclassical defect measures, also called Wigner measures. Then, we can deduce various properties of this kind of measures and thereby indirectly recover information about asymptotics as $\varepsilon$ goes to 00.
1.2 Edge states generated by coherent states
The system (1) has already been studied in [1] where are performed precise computation. The edge mode computed is generated by a very specific shape of semiclassical wave packet.
In [6], it was proved that some terms of the solution to equation (1) are negligible in $\mathrm{L}^{\infty}$-norm as $\varepsilon$ goes to 00. It has been proven, for initial conditions, that semiclassical wave packets concentrate at a point on the $\mathrm{E}$ curve with no constraint on orientation. However, these dispersive terms are not necessarily negligible in the $\mathrm{L}^{2}$ norm. Let us recall these results.
Let $x_{0}\in\mathrm{E}$ and $t\mapsto x_{t}$ the solution to the ODE
$\left\{\begin{array}[]{l}\dot{x}_{t}=\frac{\nabla\mathrm{m}(x_{t})^{\perp}}{\left|\nabla\mathrm{m}(x_{t})\right|},\\ \left.x_{t}\right|_{t=0}=x_{0}.\end{array}\right.$
Let us define, for all $t\in\mathbb{R}$, $r_{t}\in\mathbb{R}_{>0}$ and $\theta_{t}\in\mathbb{R}$, such that $r_{t}\begin{pmatrix}-\sin\theta_{t}\\ \cos\theta_{t}\end{pmatrix}=\nabla\mathrm{m}(x_{t})$ and
$\vec{V}_{\theta_{t}}\coloneqq\begin{pmatrix}e^{\frac{-\mathrm{i}\theta_{t}}{2}}\\ -e^{\frac{\mathrm{i}\theta_{t}}{2}}\end{pmatrix}.$
Following [6], we define wave packet concentrated at $(x_{0},\xi_{0})\in\mathbb{R}^{4}$ as a $\varepsilon$-dependent function satisfying for all positive $\varepsilon$ and for all $x\in\mathbb{R}^{2}$,
$\mathrm{WP}^{\varepsilon}_{x_{0},\xi_{0}}[\vec{f}](x)\coloneqq\frac{e^{\frac{\mathrm{i}}{\varepsilon}\xi_{0}\cdot(x-x_{0})}}{\sqrt{\varepsilon}}\vec{f}\left(\frac{x-x_{0}}{\sqrt{\varepsilon}}\right),$
with $\vec{f}\in\mathscr{S}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ independent of $\varepsilon$. We say that $\mathrm{WP}^{\varepsilon}_{x_{0},\xi_{0}}[\vec{f}]$ is oriented along $\vec{V}\in\mathbb{C}^{2}$ if there exists $f\in\mathscr{S}\left(\mathbb{R}^{2},\mathbb{C}\right)$ such that for all $x\in\mathbb{R}^{2},\vec{f}(x)=f(x)\vec{V}$. The parameter $\xi_{0}$ is interpreted as the impulsion (or the speed) of an electron of wave function $\mathrm{WP}^{\varepsilon}_{x_{0},\xi_{0}}[\vec{f}]$.
Theorem 1.1 ([1] Theorem 2 and [6] Theorem A.1).
Let $\vec{f}\in\mathscr{S}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ independent of $\varepsilon$. If ${\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}}$ solves (1) with for all $\varepsilon\in(0,1]$ and for all $x\in\mathbb{R}^{2}$,
$\psi_{0}^{\varepsilon}(x)=\frac{1}{\sqrt{\varepsilon}}\vec{f}\left(\frac{x-x_{0}}{\sqrt{\varepsilon}}\right),$
then there exists $T>0$, such that for all $t\in(0,T)$, uniformly for $\varepsilon\in(0,1]$,
$\psi_{t}^{\varepsilon}=\mathrm{WP}^{\varepsilon}_{x_{t},0}+\mathcal{O}_{\mathrm{L}^{\infty}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\left(\varepsilon^{-1/4}\right)+\mathcal{O}_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\left(\varepsilon^{1/2}\right).$ (2)
where $\mathrm{WP}^{\varepsilon}_{x_{t},0}$ is a wave packet concentrated at $(x_{t},(0,0))$, oriented along $\vec{V}_{\theta_{t}}$. Moreover, we have the following.
  1. (i)
    If $\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}$ is oriented along $\vec{V}_{\theta_{0}}$, then the equality (2) holds for all $t>0$ without $\mathcal{O}_{\mathrm{L}^{\infty}}\left(\varepsilon^{-1/4}\right)$.
  2. (ii)
    If there exists $f\in\mathscr{S}\left(\mathbb{R},\mathbb{C}\right)$ such that for all $\varepsilon\in(0,1]$ and for all $x\in\mathbb{R}^{2}$,
    $\psi_{0}^{\varepsilon}(x)\coloneqq\frac{r_{0}^{1/4}}{\sqrt{\varepsilon}}f\left(\frac{\left(\mathrm{R}_{\theta_{0}}\left(x-x_{0}\right)\right)_{1}}{\sqrt{\varepsilon}}\right)e^{\frac{-r_{0}\left(\mathrm{R}_{\theta_{0}}\left(x-x_{0}\right)\right)_{2}^{2}}{2\varepsilon}}\vec{V}_{\theta_{0}},$
    where $\mathrm{R}_{\theta}$ is the rotation of angle $\theta$ in $\mathbb{R}^{2}$, then uniformly for $\varepsilon\in(0,1]$, for all $(t,x)\in\mathbb{R}\times\mathbb{R}^{2}$,
    $\psi_{t}^{\varepsilon}(x)=\frac{r_{t}^{1/4}}{\sqrt{\varepsilon}}f\left(\frac{\left(\mathrm{R}_{\theta_{t}}\left(x-x_{t}\right)\right)_{1}}{\sqrt{\varepsilon}}\right)e^{\frac{-r_{t}\left(\mathrm{R}_{\theta_{t}}\left(x-x_{t}\right)\right)_{2}^{2}}{2\varepsilon}}\vec{V}_{\theta_{t}}\ +\ \mathcal{O}\left(\varepsilon^{1/2}\langle t\rangle\right).$
  3. (iii)
    If $\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}$ is oriented along $\vec{V}_{\theta_{0}}^{\perp}$, then $\mathrm{WP}^{\varepsilon}_{x_{t},0}=0$.
The initial wave packet is concentrated on the curve $\mathcal{C}\coloneqq\mathrm{E}\times\{(0,0)\}\subset\mathbb{R}^{4}$ and so is the wave packet at time $t$. Physically, the wave packets under consideration have no impulse.
Let us informally describe these wave packets. In the point (ii), the variables $\left(\mathrm{R}_{\theta_{t}}\left(x-x_{t}\right)\right)_{1}$ and $\left(\mathrm{R}_{\theta_{t}}\left(x-x_{t}\right)\right)_{2}$ represent the tangent and the normal variable to the curve $\mathrm{E}$, respectively. They do not play the same role in the considered wave packet. In particular, for all $t\in[0,T)$, the wave packet $\mathrm{WP}^{\varepsilon}_{x_{t},0}$ have a Gaussian behavior in the normal variable and is localized near the point $(x_{t},(0,0))$ in the phase space.
If the orientation is not along $\vec{V}_{\theta_{0}}$ or the decay with respect to the normal variable is different, then the $\mathrm{L}^{\infty}$-remaining part is not necessarily negligible in $\mathrm{L}^{2}$-norm. Therefore Theorem 1.1 does not give a full description of $\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}$ in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$.
To prove this theorem, the Hamiltonian $\mathrm{H}_{\varepsilon}$ is reduced to a normal form through Fourier Integral Operator in [6]. In the same way, we will reduce it through a change of variable to have a model problem which can be understand as if the curve $\mathrm{E}$ was a line. Our objective is two-fold: we aim at describing the evolution of these dispersive terms and we want to consider initial data in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$.
1.3 Main result
Instead of computing the asymptotics of the wave function itself, we focus on the probability densities such as the position one defined by $\left|\psi^{\varepsilon}_{t}(x)\right|_{\mathbb{C}^{2}}^{2}\mathrm{d}x$. Its physical interpretation is the probability of finding the particle at time $t$ in the position $x$. This approach generalizes the previous strategy and catches the information on the quantities that are not negligible in $\mathrm{L}^{2}$-norm. More precisely, we want to describe the evolution as $\varepsilon$ goes to 00 of the quadratic quantities
$a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathbb{C}^{2,2}\right)\longmapsto\left\langle a\psi^{\varepsilon}_{t},\psi^{\varepsilon}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\coloneqq\int_{\mathbb{R}^{2}}\left\langle a(x)\psi^{\varepsilon}_{t}(x),\psi^{\varepsilon}_{t}(x)\right\rangle_{\mathbb{C}^{2}}\mathrm{d}x,$ (3)
where $\mathbb{C}^{2,2}$ is the set of $2$ by $2$ matrices and the scalar product of $(v,w)\in\left(\mathbb{C}^{2}\right)^{2}$ is $\left\langle v,w\right\rangle_{\mathbb{C}^{2}}\coloneqq v\cdot\overline{w}$.
For example, if $a$ is equal to a projector on a direction $\mathbb{R}\vec{V}$ for $\vec{V}\in\mathbb{R}^{2}\setminus\{0\}$ on a subspace $\Omega$, then $\left\langle a\psi^{\varepsilon}_{t},\psi^{\varepsilon}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}$ describes the probability that $\psi^{\varepsilon}_{t}$ is collinear to $\vec{V}$ at time $t$ on the subspace $\Omega$.
We now introduce our assumptions on the function $\mathrm{m}$. Physically, the assumptions describe the interface between the two topological insulators.
Assumption 1.2.
The function $\mathrm{m}$ satisfies
$\inf_{x\in\mathrm{E}}{\Big{|}\nabla\mathrm{m}(x)\Big{|}}>0.$
To simplify the presentation and technical details, we will work with a global tubular neighborhood of the curve $\mathrm{E}$. To ensure the existence of such a neighborhood, we assume the following assumption. Actually, if we consider a neighborhood of a point of the curve $\mathrm{E}$, the local version of the following assumption will always be satisfied.
Let $x_{0}\in\mathrm{E}$. Let us consider the parametrization $\gamma$ of the curve $\mathrm{E}$ and the vector-valued function $\nu$ such that $\gamma(0)=x_{0}$ and for all $s\in\mathbb{R}$,
$\gamma^{\prime}(s)\coloneqq\frac{-{\nabla\mathrm{m}(\gamma(s))}^{\perp}}{\left|\nabla\mathrm{m}(\gamma(s))\right|}\;\;,\;\;\nu(s)\coloneqq\frac{\nabla\mathrm{m}(\gamma(s))}{\left|\nabla\mathrm{m}(\gamma(s))\right|},$
where denotes the $\pi/2$-counterclockwise rotation. Then the curvature of the curve $\mathrm{E}$ at a point $\gamma(s)$ is the real number $\kappa(s)$ given by Frenet’s formula
$\kappa(s)\coloneqq\gamma^{\prime\prime}(s)\cdot\nu(s).$
Let us denote ${\mathrm{I}\coloneqq\left(-\frac{1}{2||\kappa||_{\infty}},\frac{1}{2||\kappa||_{\infty}}\right)}$.
Assumption 1.3.
The map
$\begin{matrix}\Phi:&\mathbb{R}\times\mathrm{I}&\longrightarrow&\mathbb{R}^{2}\\ &(s,y)&\longmapsto&\gamma(s)+y\nu(s)\end{matrix}$ (4)
is a global diffeomorphism from ×$\mathbb{R}\times\mathrm{I}$ to
$\Omega\coloneqq\left\{\gamma(s)+y\nu(s)|\ (s,y)\in\mathbb{R}\times\mathrm{I}\right\}.$
[…]
$\frac{\nabla\mathrm{m}(\gamma(s))}{\left|\nabla\mathrm{m}(\gamma(s))\right|}=\begin{pmatrix}\cos\theta(s)\\ \sin\theta(s)\end{pmatrix}.$
We will denote the dual variable of $(s,y)$ by $(\sigma,\eta)$. Let us consider for all $(s,y,\sigma)\in\mathbb{R}^{3}$ and for all $n\in\mathbb{Z}_{\neq 0}$,
$\lambda_{0}(\sigma)\coloneqq-\sigma\;\;,\;\;\lambda_{n}(s,\sigma)\coloneqq\mathrm{sgn}(n)\sqrt{\sigma^{2}+2|n||\nabla\mathrm{m}(\gamma(s))|}.$ (5)
Although the function $\lambda_{0}$ does not depend on $s$, we also will denote it by $\lambda_{0}(s,\sigma)$.
A consequence of our main result (in Section 3.3) is the following asymptotic description of the position density of the solution of equation (1) for initial data which are wave packets concentrating on the curve $\mathcal{C}$.
Theorem 1.4.
Assume Assumption 1.2 and 1.3. Let $\vec{f}\in\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$. If ${\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}}$ solves (1) with for all $\varepsilon\in(0,1]$ and for all $x\in\mathbb{R}^{2}$,
$\psi_{0}^{\varepsilon}(x)=\frac{1}{\sqrt{\varepsilon}}\vec{f}\left(\frac{x-x_{0}}{\sqrt{\varepsilon}}\right),$
then, there exist a vanishing sequence ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ of positive numbers and a family of measurable maps ${t\mapsto\left(\mathrm{d}\gamma^{t}_{n}\right)_{n\in\mathbb{Z}}}$ valued in the set of scalar non negative Radon measures on $\mathbb{R}^{2}$, such that for all $(a,\Xi)\in\mathscr{C}_{c}^{\infty}\left(\Omega,\mathbb{C}^{2,2}\right)\times\mathscr{C}_{c}^{\infty}(\mathbb{R},\mathbb{C})$,
$\int_{\mathbb{R}}\Xi(t)\left\langle a\psi^{\varepsilon_{k}}_{t},\psi^{\varepsilon_{k}}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\mathrm{d}t\underset{k\to+\infty}{\longrightarrow}\int_{\mathbb{R}}\int_{\mathbb{R}}\Xi(t)\mathrm{tr}\Big{(}a\big{(}\gamma(s)\big{)}\mathrm{d}\rho_{t}(s)\Big{)}\mathrm{d}t$
with for all $s\in\mathbb{R}$,
$\mathrm{d}\rho_{t}(s)=\frac{1}{2}\sum_{n\in\mathbb{Z}}\int_{\mathbb{R}}\begin{pmatrix}1&\frac{\mathrm{i}\sigma}{\lambda_{n}(s,\sigma)}e^{-\mathrm{i}\theta(s)}\\ \frac{-\mathrm{i}\sigma}{\lambda_{n}(s,\sigma)}e^{\mathrm{i}\theta(s)}&1\end{pmatrix}\mathrm{d}\gamma^{t}_{n}(s,\mathrm{d}\sigma),$
where for all $n\in\mathbb{Z}$, $\mathrm{d}\gamma_{n}^{t}$ satisfies the following equation
$\left\{\begin{array}[]{rll}\partial_{t}\mathrm{d}\gamma_{n}^{t}&=&\partial_{\sigma}\mathrm{d}\gamma_{n}^{t}\ \partial_{s}\lambda_{n}-\partial_{s}\mathrm{d}\gamma_{n}^{t}\ \partial_{\sigma}\lambda_{n}\\ \mathrm{d}\gamma_{n}^{t}|_{t=0}&=&\mathrm{d}\gamma_{n,0}\end{array}\right.,$ (6)
and $\mathrm{d}\gamma_{n,0}$ is determined by the initial condition.
We will see in Section 3.3 a more general result (Theorem 3.10), with initial data in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$.
Remark (Link between Theorem 1.1 and Theorem 1.4).
With the notation of Theorem 1.4, the measure $\mathrm{d}\gamma_{0}^{t}$ is the measure associated to the part $\mathrm{WP}^{\varepsilon}_{x_{t},0}$ of the solution of Theorem 1.1. Likewise, the measures $\left(\mathrm{d}\gamma_{n}^{t}\right)_{n\in\mathbb{Z}_{\neq 0}}$ are associated to the $\mathrm{L}^{\infty}$-remainder part of the solution of Theorem 1.1. Notably, if for all $n\in\mathbb{Z}_{\neq 0},\mathrm{d}\gamma_{n,0}=0$ and $\mathrm{d}\gamma_{0,0}=1$, then for all $(t,s,\sigma)\in\mathbb{R}^{3}$ and for all $n\in\mathbb{Z}_{\neq 0}$,
$\mathrm{d}\gamma_{0}^{t}(s,\sigma)=\delta(s-t)\delta(\sigma)\quad,\quad\mathrm{d}\gamma_{n}^{t}(s,\sigma)=0,$
where $\delta$ is the Dirac mass over 00.
Theorem 1.4 gives a description of the evolution of the position density when initial data are wave packets. But Theorem 3.10 allow us to consider any initial data in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ concentrating near the curve $\mathcal{C}$ to describe their asymptotic position density behavior. A consequence of this theorem is the following one.
Theorem 1.5.
Assume Assumption 1.2 and 1.3. Let ${\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}}$ a uniformly bounded family of $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$, $\nu$ a non negative scalar Radon measure on $\mathbb{R}_{t}\times\mathbb{R}_{x}^{2}$ and ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ a vanishing sequence such that for all $(a,\Xi)\in\mathscr{C}_{c}^{\infty}\left(\Omega,\mathbb{C}\right)\times\mathscr{C}_{c}^{\infty}(\mathbb{R},\mathbb{C})$,
$\int_{\mathbb{R}}\int_{\Omega}\Xi(t)a(x)\left|\psi_{t}^{\varepsilon_{k}}(x)\right|^{2}\mathrm{d}x\mathrm{d}t\underset{k\to+\infty}{\longrightarrow}\int_{\mathbb{R}}\int_{\Omega}\Xi(t)a(x)\mathrm{d}\nu(t,x),$
where ${\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}}$ solves (1) with ${\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}}$ as initial condition, then,
$\mathrm{d}\nu\mathds{1}_{\mathrm{E}}\geqslant\sum_{n\in\mathbb{Z}}\Phi^{*}\left(\mathrm{d}\gamma^{t}_{n}\right)\mathds{1}_{\mathrm{E}}\mathrm{d}t$
where $\left(\mathrm{d}\gamma_{n}^{t}\right)_{n\in\mathbb{Z}}$ is defined by (6).
This theorem gives a lower bound estimation of the position density at time $t$ of the evolution of any initial data concentrating over the curve $\mathrm{E}$.
1.4 Main ideas of the proofs and organisation of the article
There are two main ideas in the proof of Theorem 3.10.
The first step is to transform the operator $\mathrm{H}_{\varepsilon}$ into a model operator as if the curve $\mathrm{E}$ were exactly a line. Let us discuss the case of the line.
If the interface is a line defined by a mass $\mathrm{m}$ such that for all $x\coloneqq(x_{1},x_{2})\in\mathbb{R}^{2},\mathrm{m}(x)=\widetilde{m}(x_{1})x_{2}$, then the operator $\mathrm{H}_{\varepsilon}$ is unitary equivalent to the following one,
$\sqrt{\varepsilon}\begin{pmatrix}\sqrt{\varepsilon}\mathrm{D}_{1}&x_{2}\widetilde{m}(x_{1})+i\mathrm{D}_{2}\\ x_{2}\widetilde{m}(x_{1})-i\mathrm{D}_{2}&-\sqrt{\varepsilon}\mathrm{D}_{1}\end{pmatrix}.$
This operator is $\sqrt{\varepsilon}$-semiclassical with respect to $x_{1}$ only. Thus, it will be considered as the quantization of an operator-valued symbol. It leads us to the second idea of the proof.
The second step is to build a two-scale quantization that fits the scale of our new operator. In fact, our analysis relies on a microlocal approach, replacing densities $\left(\left|\psi_{t}^{\varepsilon}(x)\right|^{2}\mathrm{d}x\right)_{\varepsilon>0}$ by the Wigner transform [26, 15]. The Wigner transform provides us with a description of the dynamics for a very general class of equations and initial data, allowing us to deal with quantum observables. Therefore, we will study the quadratic quantity (3), with an observable $a$ depending of position and momentum, i.e.
$a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2}_{x}\times\mathbb{R}^{2}_{\xi},\mathbb{C}^{2,2}\right)\longmapsto\left\langle a\left(\cdot{},\varepsilon\mathrm{D}_{x}\right)\psi^{\varepsilon}_{t},\psi^{\varepsilon}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}.$
We recall basic facts about these notions in Section 2. However, to precisely analyze what happens near the curve $\mathcal{C}$, we have to take into account how $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ concentrates on the curve $\mathcal{C}$ with respect to the scale $\sqrt{\varepsilon}$. We introduce in Section 3 the two-scale Wigner measure that be of crucial use for proving our main results Theorems 3.8 and 3.10. Finally, we prove Theorems 1.41.53.8 and 3.10 in Section 4.
1.5 Notation
We denote by $\mathbb{S}^{2}$ the unit sphere of $\mathbb{R}^{3}$.
We denote by $v\otimes w$ the $2$ by $2$ matrix $v\,^{t}w$ for $(v,w)\in\left(\mathbb{C}^{2}\right)^{2}$.
We denote by ${\rm tr}$ the trace in $\mathcal{L}\left(\mathbb{C}^{2}\right)$ and by $\mathrm{Tr}$ the trace in $\mathcal{L}\left(\mathrm{L}^{2}\left(\mathbb{R},\mathbb{C}^{2}\right)\right)$.
Let $\mathcal{H}$ a separable Hilbert space and $n$ an integer, for $f,g\in\mathrm{L}^{2}\left(\mathbb{R}^{n},\mathcal{H}\right)$, we denote the usual inner product by
$\left\langle f,g\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{n},\mathcal{H}\right)}\coloneqq\int_{\mathbb{R}^{n}}\left\langle f(x),g(x)\right\rangle_{\mathcal{H}}\,\mathrm{d}x.$
For $A$ an operator acting on $\mathrm{L}^{2}\left(\mathbb{R}^{n},\mathcal{H}\right)$, we denote by $A^{\dagger}$ the formal adjoint operator of $A$ associated to the usual scalar product of $\mathrm{L}^{2}\left(\mathbb{R}^{n},\mathcal{H}\right)$.
We define the semiclassical Fourier transform for all $\varepsilon>0$ and for $f\in\mathrm{L}^{1}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ by
$\mathcal{F}_{\varepsilon}f(\xi)\coloneqq\frac{1}{2\pi\varepsilon}\int_{\mathbb{R}^{2}}e^{-\mathrm{i}\frac{x\cdot\xi}{\varepsilon}}f(x)\mathrm{d}x,\quad\xi\in\mathbb{R}^{2},$
and we denote by $\widehat{f}$ the Fourier transform of $f$ defined by $\widehat{f}\coloneqq\mathcal{F}_{1}f$.
2 Semiclassical analysis
Our problem requires an analysis in the phase space. We consider the Wigner transform of the solution $\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}$ of equation (1) and its weak limits in the space of distributions that are called Wigner measures. We recall the definition and the main properties of the Wigner transform in Section 2.1 below. Then, we apply these results to the solution of our problem in Section 2.2.
2.1 Wigner transform and Wigner measure
Let ${(f^{\varepsilon})_{\varepsilon>0}}$ uniformly bounded in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$. The Wigner transform of ${(f^{\varepsilon})_{\varepsilon>0}}$ is defined as follow, for all $(x,\xi)\in\mathbb{R}^{2}\times\mathbb{R}^{2}$,
$\mathrm{W}^{\varepsilon}\left[f^{\varepsilon}\right](x,\xi)\coloneqq\int_{\mathbb{R}^{2}}e^{i\xi\cdot y}f^{\varepsilon}\left(x+\frac{\varepsilon y}{2}\right)\otimes\overline{f^{\varepsilon}\left(x-\frac{\varepsilon y}{2}\right)}\frac{\mathrm{d}y}{(2\pi)^{2}}.$
It was introduced by Wigner in [25] and in [26]. One can check that $\mathrm{W}^{\varepsilon}\left[f^{\varepsilon}\right]$ is valued in the set of Hermitian matrices and $\mathrm{W}^{\varepsilon}\left[f^{\varepsilon}\right]\in\mathrm{L}^{2}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)\cap\mathscr{C}^{0}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)$. Moreover, one recovers formally the density in position through the relation
$\int_{\mathbb{R}^{2}}{\rm tr}\mathrm{W}^{\varepsilon}\left[f^{\varepsilon}\right](x,\xi)\mathrm{d}\xi=\left|f^{\varepsilon}(x)\right|_{\mathbb{C}^{2}}^{2}.$
The function $\mathrm{W}^{\varepsilon}\left[f^{\varepsilon}\right]$ is usually considered as a distribution on the phase space $\mathbb{R}^{4}=\mathbb{R}^{2}_{x}\times\mathbb{R}^{2}_{\xi}$. Moreover, we have for all $a\in\mathscr{C}^{\infty}_{c}(\mathbb{R}^{4},\mathbb{C}^{2,2})$,
$\mathrm{I}^{\varepsilon}_{f}(a)\coloneqq\int_{\mathbb{R}^{4}}{\rm tr}\Big{(}a(x,\xi)\mathrm{W}^{\varepsilon}\left[f^{\varepsilon}\right](x,\xi)\Big{)}\mathrm{d}x\mathrm{d}\xi=\left\langle\mathrm{Op}_{\varepsilon}(a)f^{\varepsilon},f^{\varepsilon}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)},$
where $\mathrm{Op}_{\varepsilon}(a)$ is the semiclassical pseudo-differential operator of symbol $a$ obtained by the Weyl quantization
$\mathrm{Op}_{\varepsilon}(a)f(x)\coloneqq\int_{\mathbb{R}^{4}}a\left(\frac{x+y}{2},\xi\right)f(y)e^{\mathrm{i}\frac{\xi\cdot(x-y)}{\varepsilon}}\frac{\mathrm{d}y\mathrm{d}\xi}{(2\pi\varepsilon)^{2}},\quad f\in\mathscr{S}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right),\ x\in\mathbb{R}^{2}.$
The Calderón-Vaillancourt Theorem [4] asserts that the family of operators $\left(\mathrm{Op}_{\varepsilon}(a)\right)_{\varepsilon>0}$ is uniformly bounded with respect to $\varepsilon$ in $\mathcal{L}\left(L^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)$: there exists $(C,N)\in\mathbb{R}_{>0}\times\mathbb{N}_{\geqslant 0}$ such that for all $a=(a_{i,j})_{1\leqslant i,j\leqslant 2}\in\mathscr{S}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)$, for all $f\in\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right),$
$\left\lVert\mathrm{Op}_{\varepsilon}(a)f\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\leqslant C\,\max_{1\leqslant i,j\leqslant 2}\,\max_{\left|\alpha\right|+\left|\beta\right|\leqslant N}\left\lVert\partial^{\alpha}_{x}\partial^{\beta}_{\xi}a_{i,j}\right\rVert_{\mathrm{L}^{\infty}\left(\mathbb{R}^{4}\right)}\left\lVert f\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}.$ (7)
The Wigner transform is not a positive distribution. However, its weak limits are positive measures. This is due to the weak Gårding inequality. Adapting [11] Proposition 2.13 to the matrix case, for all $a\in\mathscr{C}^{\infty}_{c}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)$ non negative, for all $\delta>0$, there exists $C_{\delta}>0$, for all $f\in\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$,
$\left\langle\mathrm{Op}_{\varepsilon}(a)f,f\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\geqslant-\left(\delta+C_{\delta}\varepsilon\right)\left\lVert f\right\rVert^{2}_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}.$
There exists a subsequence $\left(f_{\varepsilon_{n}}\right)_{n\in\mathbb{N}}$ and a positive matrix of Radon measures $\mathrm{d}\mu=(\mathrm{d}\mu_{i,j})_{1\leqslant i,j\leqslant 2}$ such that for all $a\in\mathscr{S}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)$,
$\mathrm{I}^{\varepsilon_{n}}_{f}(a)\underset{n\to+\infty}{\longrightarrow}\int_{\mathbb{R}^{4}}{\rm tr}\Big{(}a(x,\xi)\mathrm{d}\mu(x,\xi)\Big{)}.$ (8)
The measure $\mathrm{d}\mu$ is called the Wigner measure or the semiclassical measure of $\left(f^{\varepsilon_{n}}\right)_{n\in\mathbb{N}}$. A proof of existence is recalled in Section B.1 for families valued in a separable Hilbert space.
Example 2.1.
Let us consider $\alpha\in(0,1]$, $f\in\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ and for all $x\in\mathbb{R}^{2}$,
$f^{\varepsilon}(x)\coloneqq\frac{1}{\varepsilon^{\alpha/2}}f\left(\frac{x}{\varepsilon^{\alpha}}\right),$
so ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$ has only one Wigner measure that depends of $\alpha$.
If $\alpha=1/2$, then for all $(x,\xi)\in\mathbb{R}^{4}$, ${\mathrm{d}\mu(x,\xi)=\left(\int_{\mathbb{R}^{2}}f(x)\otimes\overline{f(x)}\mathrm{d}x\right)\delta_{0}(x)\otimes\delta_{0}(\xi)}$.
If $\alpha=1$, then for all $(x,\xi)\in\mathbb{R}^{4}$, ${\mathrm{d}\mu(x,\xi)=\widehat{f}(\xi)\otimes\overline{\widehat{f}(\xi)}\delta_{0}(x)\frac{\mathrm{d}\xi}{2\pi}}$.
We say a bounded family $\left(f^{\varepsilon}\right)_{\varepsilon>0}$ of $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ is $\varepsilon$-oscillating if
$\limsup_{\varepsilon\to 0}\int_{\left|\xi\right|>R/\varepsilon}\left|\widehat{f^{\varepsilon}}(\xi)\right|^{2}_{\mathbb{C}^{2}}\mathrm{d}\xi\underset{R\to+\infty}{\longrightarrow}0.$
We say a bounded family $\left(f^{\varepsilon}\right)_{\varepsilon>0}$ of $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ is compact at infinity if
$\limsup_{\varepsilon\to 0}\int_{\left|x\right|>R}\left|f^{\varepsilon}(x)\right|^{2}_{\mathbb{C}^{2}}\mathrm{d}x\underset{R\to+\infty}{\longrightarrow}0.$
If $\left(f^{\varepsilon}\right)_{\varepsilon>0}$ is $\varepsilon$-oscillating and $\left(\left|f^{\varepsilon}\right|^{2}_{\mathbb{C}^{2}}\mathrm{d}x\right)_{\varepsilon>0}$ converges weakly to $\nu$, then for all semiclassical measure $\mathrm{d}\mu$ of $\left(f^{\varepsilon}\right)_{\varepsilon>0}$,
$\nu(x)=\mathrm{tr}\int_{\mathbb{R}^{2}}\mathrm{d}\mu(x,\mathrm{d}\xi).$
If $\left(f^{\varepsilon}\right)_{\varepsilon>0}$ is compact at infinity and $\left(\left|\widehat{f^{\varepsilon}}\right|^{2}_{\mathbb{C}^{2}}\mathrm{d}\xi\right)_{\varepsilon>0}$ converges weakly to $\widetilde{\nu}$, then for all semiclassical measure $\mathrm{d}\mu$,
$\widetilde{\nu}(\xi)=\mathrm{tr}\int_{\mathbb{R}^{2}}\mathrm{d}\mu(\mathrm{d}x,\xi).$
Now, we apply these first properties on solutions to equation (1).
2.2 Application to the Dirac equation
In this section, we consider $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ the solution to equation (1) with the initial normalized condition $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$. We revisit in this paragraph the notions introduced in the preceding one for this special family.
2.2.1 $\varepsilon$-oscillation of solution to Dirac equation
Lemma 2.2 (Conservation of $\varepsilon$-oscillating).
Let $T>0$. If $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ is $\varepsilon$-oscillating then $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ is $\varepsilon$-oscillating uniformly for $t\in[-T,T]$. In other words, uniformly for $t\in[-T,T]$,
$\limsup_{\varepsilon\to 0}\int_{\left|\xi\right|>R/\varepsilon}\left|\widehat{\psi^{\varepsilon}_{t}}(\xi)\right|^{2}\mathrm{d}\xi\underset{R\to+\infty}{\longrightarrow}0.$
Proof of Lemma 2.2.
Let $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ be $\varepsilon$-oscillating. The proof relies on the analysis of
$\psi^{\varepsilon,R}_{t}\coloneqq\chi^{R}(\varepsilon\mathrm{D}_{x})\psi_{t}^{\varepsilon},$
where ${\chi^{R}(\xi)=\chi\left(\frac{\xi}{R}\right)}$ with $\chi\in\mathscr{C}^{\infty}\left(\mathbb{R}^{2},[0,1]\right),\chi\equiv 0$ for $\left|\xi\right|<1$ and $\chi\equiv 1$ for $\left|\xi\right|>2$. We observe that $\psi^{\varepsilon,R}_{t}$ solves
$\mathrm{i}\varepsilon\partial_{t}\psi^{\varepsilon,R}_{t}=\mathrm{H}_{\varepsilon}\psi^{\varepsilon,R}_{t}+\left[\chi_{R}(\varepsilon\mathrm{D}_{x}),\begin{pmatrix}\mathrm{m}(x)&0\\ 0&-\mathrm{m}(x)\end{pmatrix}\right]\psi^{\varepsilon,R}_{t}.$
Nevertheless, the derivatives of $\mathrm{m}$ are bounded therefore there exists $C>0$ such that
$\left\lVert\frac{1}{\varepsilon}\left[\chi^{R}(\varepsilon\mathrm{D}_{x}),\begin{pmatrix}\mathrm{m}(x)&0\\ 0&-\mathrm{m}(x)\end{pmatrix}\right]\right\rVert_{\mathcal{L}\left(\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)}\leqslant\frac{C}{R}.$
Then, by an energy argument, for all $t\in[-T,T]$,
$\frac{\mathrm{d}}{\mathrm{d}\mathrm{t}}\left\lVert\psi^{\varepsilon,R}_{t}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\leqslant\frac{C}{R},$
so,
$\limsup_{\varepsilon\rightarrow 0}\left\lVert\psi^{\varepsilon,R}_{t}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}$ $\leqslant\limsup_{\varepsilon\rightarrow 0}\left\lVert\psi^{\varepsilon,R}_{0}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}+\frac{C}{R}\left|T\right|,$
$\underset{R\to+\infty}{\longrightarrow}0.$
because $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ is uniformly bounded in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ and $\varepsilon$-oscillating. ∎
The $\varepsilon$-oscillation conservation property justifies the semiclassical approach of solutions to equation (1) for analyzing the limits of the quantities (3). Note that similar arguments give that there is also no loss of mass at infinity in configuration space in the sense that if $\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}$ is compact at infinity, then $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ is compact at infinity uniformly in time $t\in[-T,T]$, for some $T>0$. The proof is in Appendix A.
As a consequence, if $\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}$ is normalized, $\varepsilon$-oscillating and compact at infinity, then for all $\chi\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},[0,1]\right)$, with $\chi\equiv 1$ for $\left|x\right|<1$, $\chi\equiv 0$ for $\left|x\right|>2$,
$\lim_{\varepsilon\to 0}\left\langle\mathrm{Op}_{\varepsilon}\left(\chi\left(\frac{x}{R}\right)\chi\left(\frac{\varepsilon\mathrm{D}_{x}}{R}\right)\right)\psi^{\varepsilon}_{0},\psi^{\varepsilon}_{0}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\;\underset{R\to+\infty}{\longrightarrow}\;\lim_{\varepsilon\to 0}\left\lVert\psi_{0}^{\varepsilon}\right\rVert^{2}_{\mathrm{L}^{2}\left(R^{2},\mathbb{C}^{2}\right)}=1.$
So for all semiclassical measure $\mathrm{d}\mu$ of $\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}$ normalized, $\varepsilon$-oscillating and compact at infinity,
$\mathrm{tr}\int_{\mathbb{R}^{4}}\mathrm{d}\mu(x,\xi)=1.$
2.2.2 Time-averaged Wigner measures
For taking into account the time dependence of the family $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$, we will consider time-averaged Wigner measure. We test the Wigner transform $\mathrm{W}^{\varepsilon}[\psi^{\varepsilon}_{t}]$ against time-dependent observables, so for all $\Xi\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R},\mathbb{C}\right)$,
$\int_{\mathbb{R}}\int_{\mathbb{R}^{4}}\Xi(t){\rm tr}_{\mathbb{C}^{2}}\Big{(}a(x,\xi)\mathrm{W}^{\varepsilon}[\psi^{\varepsilon}_{t}](x,\xi)\Big{)}\mathrm{d}x\mathrm{d}\xi\mathrm{d}t\;=\;\int_{\mathbb{R}}\Xi(t)\left\langle\mathrm{Op}_{\varepsilon}(a)\psi^{\varepsilon}_{t},\psi^{\varepsilon}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\mathrm{d}t.$
For the same reasons as before, we obtain the existence of a vanishing sequence $\left(\varepsilon_{k}\right)_{k\in\mathbb{N}}$ of positive numbers and of a measure $\mathrm{d}\mu$ on $\mathbb{R}^{5}=\mathbb{R}_{t}\times\mathbb{R}^{2}_{x}\times\mathbb{R}^{2}_{\xi}$ such that
$\int_{\mathbb{R}}\Xi(t)\left\langle\mathrm{Op}_{\varepsilon_{k}}(a)\psi^{\varepsilon_{k}}_{t},\psi^{\varepsilon_{k}}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\mathrm{d}t\;\underset{k\to+\infty}{\longrightarrow}\;\int_{\mathbb{R}^{5}}\Xi(t)a(x,\xi)\mathrm{d}\mu(t,x,\xi).$
Moreover, thanks to the control
$\left|\int_{\mathbb{R}^{5}}\Xi(t)a(x,\xi)\mathrm{d}\mu(t,x,\xi)\right|\leqslant C\ \left\lVert\Xi\right\rVert_{\mathrm{L}^{1}(\mathbb{R})}\left\lVert\mathrm{Op}_{\varepsilon}(a)\right\rVert_{\mathcal{L}\left(\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)},$
we can write $\mathrm{d}\mu(t,x,\xi)=\mathrm{d}\mu_{t}(x,\xi)\mathrm{d}t$. In other words, the measure $\mathrm{d}\mu$ is absolutely continuous in time with respect to Lebesgue measure.
2.2.3 First properties of the Wigner measure of the solution of the Dirac equation
It is known from [15, 11] that any time averaged Wigner measure, outside the curve $\mathcal{C}$, of $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ is completely characterized by those of the initial data. Let us first introduce some notations. Let us consider the symbol of the operator $\mathrm{H}_{\varepsilon}$ according to the Weyl quantization; for all $(x,\xi)\in\mathbb{R}^{4}$,
$\mathrm{H}(x,\xi)\coloneqq\begin{pmatrix}\mathrm{m}(x)&\xi_{1}-\mathrm{i}\xi_{2}\\ \xi_{1}+\mathrm{i}\xi_{2}&-\mathrm{m}(x)\end{pmatrix}.$
We remark that ${\det\mathrm{H}(x,\xi)=-\mathrm{m}(x)^{2}-|\xi|^{2}}$. Let $(x,\xi)\in\mathbb{R}^{4}$ such that $\det\mathrm{H}(x,\xi)\neq 0$, then $\mathrm{H}(x,\xi)$ has two different eigenvalues $\lambda_{+}(x,\xi)$ and $\lambda_{-}(x,\xi)$ defined by
[…]
Let us denote the projectors on the corresponding eigenspaces by
$\Pi_{+}(x,\xi)\coloneqq\mathrm{Id}-\frac{1}{\lambda_{+}(x,\xi)}\mathrm{H}(x,\xi)\quad,\quad\Pi_{-}(x,\xi)\coloneqq\mathrm{Id}-\frac{1}{\lambda_{-}(x,\xi)}\mathrm{H}(x,\xi).$
According to the definition of $\mathcal{C}$, we remark that
$(x,\xi)\in\mathcal{C}\iff\det\mathrm{H}(x,\xi)\neq 0.$
We recall the definition of the Poisson bracket
$\left\{f,g\right\}\coloneqq\nabla_{\xi}f\nabla_{x}g-\nabla_{x}f\nabla_{\xi}g\;,\;(f,g)\in\mathscr{C}^{1}\left(\mathbb{R}^{2}_{x}\times\mathbb{R}^{2}_{\xi}\right)^{2}.$ (9)
Lemma 2.3 (Evolution of semiclassical measure outside $\mathcal{C}$).
Let $\mathrm{d}\mu_{t}$ be a semiclassical measure of $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ associated to $(\varepsilon_{n})_{n\in\mathbb{N}}$, then the measure $\mathrm{d}\mu_{t}$ can be decomposed outside $\mathcal{C}$ as follows : for all $(x,\xi)\in\mathbb{R}^{4}\setminus\mathcal{C}$
$\mathrm{d}\mu_{t}(x,\xi)=\Pi_{+}(x,\xi)\mathrm{d}\mu_{t}^{+}(x,\xi)+\Pi_{-}(x,\xi)\mathrm{d}\mu_{t}^{-}(x,\xi),$
where the scalar measures $\mathrm{d}\mu_{t}^{\pm}$ satisfy
$\partial_{t}\ \mathrm{d}\mu_{t}^{+}=\left\{\mathrm{d}\mu_{t}^{+},\lambda_{+}\right\}$ $\;\;,\;\;\mathrm{d}\mu_{0}^{+}=\mathrm{tr}\Big{(}\Pi_{+}\mathrm{d}\mu_{0}\Big{)},$
$\partial_{t}\ \mathrm{d}\mu_{t}^{-}=\left\{\mathrm{d}\mu_{t}^{-},\lambda_{-}\right\}$ $\;\;,\;\;\mathrm{d}\mu_{0}^{-}=\mathrm{tr}\Big{(}\Pi_{-}\mathrm{d}\mu_{0}\Big{)},$
and $\mathrm{d}\mu_{0}$ is the semiclassical measure of $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ associated to $(\varepsilon_{n})_{n\in\mathbb{N}}$.
This theorem is proved in [15, section 6]. The proof uses two major arguments. First, we have the decomposition of the measure $\mathrm{d}\mu_{t}$ as a sum $\Pi_{+}\mathrm{d}\mu_{t}^{+}+\Pi_{-}\mathrm{d}\mu_{t}^{-}$. Secondly, the equation for $\mathrm{d}\mu_{t}^{\pm}$ comes from the analysis for all $a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{4},\mathbb{C}\right)$ of
$\frac{1}{\mathrm{i}\varepsilon}\Big{[}\mathrm{Op}_{\varepsilon}\left(a\Pi_{\pm}\right),\mathrm{Op}_{\varepsilon}\left(\mathrm{H}_{\varepsilon}\right)\Big{]}.$
This theorem implies that for all initial data $\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}$ localized around $(x_{0},\xi_{0})\notin\mathcal{C}$, the semiclassical measure evolved along the Hamiltonian trajectories associated to the functions $\lambda_{+}$ and $\lambda_{-}$. These trajectories are included in the hypersurface $\det\mathrm{H}(x,\xi)=\det\mathrm{H}(x_{0},\xi_{0})$. It means that if we localize the initial data far from the curve $\mathcal{C}$, then the semiclassical measure never approaches the curve $\mathcal{C}$ in the phase space. Our aim in the next sections is to describe the evolution of the measure $\mathrm{d}\mu_{t}$ above the curve $\mathcal{C}$.
3 Two-scale analysis of concentrations
To study the evolution of the measure $\mathrm{d}\mu_{t}$ near the curve $\mathcal{C}$, we use a two-scale analysis near $\mathcal{C}$. We use the geometric assumptions on $\mathcal{C}$ stated in the introduction to define our two-scale quantization. In geometric terms, the measure $\mathrm{d}\mu_{t}$ is a measure on the cotangent space to $\mathbb{R}^{4}$ and, as such, have the associated geometric invariance. In particular, we can operate a change of coordinates. We work locally in a tubular neighborhood of $\mathcal{C}$ where we can use normal geodesic coordinates that we introduce in Section 3.1. In these coordinates $(s,y)$, the curve $\mathrm{E}$ writes $\{y=0\}$. We use these coordinates to transform the operator and obtain a normal form of the operator $\mathrm{H}_{\varepsilon}$. In Section 3.2 we introduce two-scale Wigner measures. Finally, we state our main results, describing these two-scale semiclassical measures for families of solutions to equation (1) in Section 3.3.
3.1 Normal geodesic coordinates
We assume assumptions 1.2 and 1.3. We will work in the normal geodesic coordinates denoted $(s,y)$ defined by (4). The curve $\mathcal{C}$ in these new coordinates is reduced to $\left\{y=\sigma=\eta=0\right\}$ where $(\sigma,\eta)$ play the role of $\xi$.
Let us define ${\mathrm{I}_{\delta_{0}}\coloneqq\left(-\frac{\delta_{0}}{2||\kappa||_{\infty}},\frac{\delta_{0}}{2||\kappa||_{\infty}}\right)}$ for $\delta_{0}\in\left(0,1\right]$ and let us consider the map
$\begin{matrix}\Phi_{\delta_{0}}:&\mathbb{R}\times\mathrm{I}_{\delta_{0}}&\longrightarrow&\mathbb{R}^{2}\\ &(s,y)&\longmapsto&\gamma(s)+y\nu(s)\end{matrix}$
A consequence of Assumption 1.3 is that for all $\delta_{0}\in\left(0,1\right]$, the function $\Phi_{\delta_{0}}$ is a diffeomorphism from ×$\mathbb{R}\times\mathrm{I}_{\delta_{0}}$ to
$\Omega_{\delta_{0}}$$\coloneqq\left\{\gamma(s)+y\nu(s)|\ (s,y)\in\mathbb{R}\times\mathrm{I}_{\delta_{0}}\right\},$
=$=\left\{x\in\mathbb{R}^{2}\Big{|}\ \mathrm{d}\left(x,\mathrm{E}\right)<\frac{\delta_{0}}{2||\kappa||_{\infty}}\right\}.$
We now use $\Phi$ to obtain a normal form of the operator $\mathrm{H}_{\varepsilon}$ as in [3].
Proposition 3.1 (Hamiltonian Normal form near the interface).
There exists a unitary operator $\mathrm{U}$ from $\mathrm{L}^{2}(\Omega,\mathbb{C}^{2})$ to $\mathrm{L}^{2}(\mathbb{R}\times\mathrm{I},\mathbb{C}^{2})$ such that $\mathrm{U}\mathrm{H}_{\varepsilon}\mathrm{U}^{-1}=\mathrm{H}_{\varepsilon}^{\mathrm{E}}$ where the operator $\mathrm{H}_{\varepsilon}^{\mathrm{E}}$ is acting on $\mathrm{L}^{2}(\mathbb{R}\times\mathrm{I},\mathbb{C}^{2})$ and defined as
$\mathrm{H}_{\varepsilon}^{\mathrm{E}}\coloneqq\mathrm{m}\Big{(}\Phi(s,y)\Big{)}\sigma_{1}-\varepsilon\mathrm{D}_{y}\sigma_{2}+\frac{\varepsilon}{1-y\kappa(s)}\mathrm{D}_{s}\sigma_{3}-\frac{\mathrm{i}\varepsilon y\kappa^{\prime}(s)}{2(1-y\kappa(s))^{2}}\sigma_{3}.$
Remark.
We have the following properties.
  1. 1.
    For all $(s,y)\in\mathbb{R}\times\mathrm{I},\left|1-y\kappa(s)\right|>1/2$ therefore the previous quantities are well-defined.
  2. 2.
    The operator $\mathrm{U}$ is a matrix-valued Fourier Integral Operator. It contains the pull-back by the change of variables $\Phi$ and the multiplication by an adequate $s$-dependent matrix (a rotation and a multiplication by a $s$-phase function).
  3. 3.
    In the following, we will more generally apply the operator $\mathrm{U}$ to function of $\mathrm{L}^{2}(\mathbb{R}^{2},\mathbb{C}^{2})$ but compactly supported in $\Omega$.
Proof of Proposition 3.1.
Let us define for all $(s,y)\in\mathbb{R}\times\mathrm{I}$ and for all $u\in\mathrm{L}^{2}(\Omega,\mathbb{C}^{2})$,
$\mathrm{U}u(s,y)\coloneqq\sqrt{\frac{1-y\kappa(s)}{2}}\begin{pmatrix}e^{\mathrm{i}\frac{\pi}{4}+\mathrm{i}\frac{\theta(s)}{2}}&-e^{-\mathrm{i}\frac{\pi}{4}-\mathrm{i}\frac{\theta(s)}{2}}\\ e^{\mathrm{i}\frac{\pi}{4}+\mathrm{i}\frac{\theta(s)}{2}}&e^{-\mathrm{i}\frac{\pi}{4}-\mathrm{i}\frac{\theta(s)}{2}}\end{pmatrix}u\Big{(}\Phi(s,y)\Big{)}.$
In order to facilitate the computation, we split the operator $\mathrm{U}$ into the composition of two operators. Let us define for all $(s,y)\in\mathbb{R}\times\mathrm{I}$, for all $\theta\in\mathbb{R}$, for all $u\in\mathrm{L}^{2}(\Omega,\mathbb{C}^{2})$ and for all $v\in\mathrm{L}^{2}\left(\mathbb{R}\times\mathrm{I},\mathbb{C}^{2}\right)$,
$\mathrm{V}_{\Phi}u(s,y)$ $\coloneqq\sqrt{1-y\kappa(s)}u\Big{(}\Phi(s,y)\Big{)},$
$\mathrm{U}_{\theta}v(s,y)$ $\coloneqq\frac{1}{\sqrt{2}}\begin{pmatrix}e^{\mathrm{i}\frac{\pi}{4}+\mathrm{i}\frac{\theta}{2}}&-e^{-\mathrm{i}\frac{\pi}{4}-\mathrm{i}\frac{\theta}{2}}\\ e^{\mathrm{i}\frac{\pi}{4}+\mathrm{i}\frac{\theta}{2}}&e^{-\mathrm{i}\frac{\pi}{4}-\mathrm{i}\frac{\theta}{2}}\end{pmatrix}v(s,y).$ (10)
The operator $\mathrm{V}_{\Phi}$ is a unitary operator from $\mathrm{L}^{2}\left(\Omega,\mathbb{C}^{2}\right)$ to $\mathrm{L}^{2}\left(\mathbb{R}\times\mathrm{I},\mathbb{C}^{2}\right)$ and the operator $\mathrm{U}_{\theta}$ is a unitary operator from $\mathrm{L}^{2}\left(\mathbb{R}\times\mathrm{I},\mathbb{C}^{2}\right)$ to $\mathrm{L}^{2}\left(\mathbb{R}\times\mathrm{I},\mathbb{C}^{2}\right)$. Notably, $\mathrm{U}=\mathrm{U}_{\theta(s)}\mathrm{V}_{\Phi}$. Moreover,
$\mathrm{V}_{\Phi}\mathrm{H}_{\varepsilon}\mathrm{V}_{\Phi}^{-1}=$ $\ \mathrm{m}\Big{(}\Phi(s,y)\Big{)}\sigma_{3}+\frac{\varepsilon}{1-y\kappa(s)}(\gamma^{\prime}\odot\sigma)\mathrm{D}_{s}+\varepsilon(\nu\odot\sigma)\mathrm{D}_{y}$
$\ -\ \frac{\mathrm{i}\varepsilon y\kappa^{\prime}(s)}{2(1-y\kappa(s))^{2}}(\gamma^{\prime}\odot\sigma)-\frac{\mathrm{i}\varepsilon\kappa(s)}{2(1-y\kappa(s))}(\nu\odot\sigma),$
with the notation $x\odot\sigma\coloneqq x_{1}\sigma_{1}+x_{2}\sigma_{2}$ for $x\in\mathbb{C}^{2}$ and $\sigma_{1},\sigma_{2}$ Pauli matrices. We conclude with the identities
$\begin{split}\mathrm{U}_{\theta(s)}\sigma_{3}\mathrm{U}_{\theta(s)}^{-1}=\sigma_{1},\quad&\mathrm{U}_{\theta(s)}(\gamma^{\prime}\odot\sigma)\mathrm{U}_{\theta(s)}^{-1}=\sigma_{3},\\ \mathrm{U}_{\theta(s)}(\nu\odot\sigma)\mathrm{U}_{\theta(s)}^{-1}=-\sigma_{2},\quad&\mathrm{U}_{\theta(s)}(\gamma^{\prime}\odot\sigma)\left(\mathrm{D}_{s}\mathrm{U}_{\theta(s)}^{-1}\right)=-\frac{\kappa(s)}{2}\mathrm{i}\sigma_{2}.\end{split}$
Let $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ solves (1) with the initial normalized condition $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$. Let $\chi\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},[0,1]\right)$, with $\chi\equiv 1$ for $\mathrm{d}(x,\mathrm{E})<\frac{1}{4||\kappa||_{\infty}}$ and $\chi\equiv 0$ for $\mathrm{d}(x,\mathrm{E})>\frac{1}{2||\kappa||_{\infty}}$, then we define for all $\varepsilon>0$,
$\varphi^{\varepsilon}_{t}\coloneqq\mathrm{U}\chi\psi^{\varepsilon}_{t}.$ (11)
Lemma 3.2 (Dirac equation in normal geodesic coordinates).
The family $\left(\varphi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ satisfies on $\mathbb{R}_{t}\times\mathbb{R}_{s}\times\mathrm{I}_{1/2}$ the following equation
$\varepsilon\mathrm{D}_{t}\varphi^{\varepsilon}_{t}+\mathrm{H}^{\mathrm{E}}_{\varepsilon}\varphi^{\varepsilon}_{t}=0.$ (12)
Proof of Lemma 3.2.
We apply $\mathrm{U}\chi$ to equation (1),
$\varepsilon\mathrm{D}_{t}\varphi_{t}^{\varepsilon}+\mathrm{H}^{\mathrm{E}}_{\varepsilon}\varphi_{t}^{\varepsilon}=\mathrm{U}\left[\mathrm{H}_{\varepsilon},\chi\right]\psi_{t}^{\varepsilon}.$
We compute
$\left[\mathrm{H}_{\varepsilon},\chi\right]=-\mathrm{i}\varepsilon\begin{pmatrix}0&\partial_{1}\chi-\mathrm{i}\partial_{2}\chi\\ \partial_{1}\chi+\mathrm{i}\partial_{2}\chi&0\end{pmatrix},$
where for all $j\in\left\{1,2\right\},\partial_{j}\chi\in\mathscr{C}_{c}^{\infty}(\mathbb{R}^{2},\mathbb{C})$ is supported in $\left\{x\in\mathbb{R}^{2}\ \Big{|}\ \frac{1}{4||\kappa||_{\infty}}\leqslant\mathrm{d}(x,\mathrm{E})\leqslant\frac{1}{2||\kappa||_{\infty}}\right\}$, then for all $(t,s,y)\in\mathbb{R}_{t}\times\mathbb{R}_{s}\times\mathrm{I}_{1/2}$,
$\Big{(}\mathrm{U}\left[\mathrm{H}_{\varepsilon},\chi\right]\psi_{t}^{\varepsilon}\Big{)}(s,y)=0.$
3.2 Two-scale Wigner measures
In this section, we introduce observables that allow to concentrate on the set $\left\{y=0,\sigma=0,\eta=0\right\}$ at a precise scale, that we choose as $\sqrt{\varepsilon}$. As we will see later and as expected in view of [1], this scale is natural.
3.2.1 Two-scale symbols and their quantization
Following the course [11], we introduce
Definition (Two-scale observable).
We define the set $\mathfrak{A}$ as follows
$\mathfrak{A}\coloneqq\left\{a\in\mathscr{C}^{\infty}\Big{(}\mathbb{R}^{4}\times\mathbb{R}^{3},\mathbb{C}^{2,2}\Big{)}\ \Bigg{|}\ \begin{array}[]{c}\exists K\subset\mathbb{R}^{4}\ \text{compact},\forall z\in\mathbb{R}^{3},a(\cdot,\cdot,z)\ \text{is supported in K}\\ \exists R_{0}>0,\forall|z|>R_{0},a(\cdot,\cdot,z)=a_{\infty}\left(\cdot,\cdot,\frac{z}{|z|}\right)\end{array}\right\}.$
We can identify $\mathscr{C}^{\infty}_{c}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)$ as a subset of $\mathfrak{A}$. We define a second scale quantization near the interface $\mathcal{C}$ for these observable by
$\begin{array}[]{ccccc}\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}&:&\mathfrak{A}&\to&\mathcal{L}\big{(}\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\big{)}\\ &&a&\mapsto&\mathrm{Op}_{1}\left(a\left(s,y,\varepsilon\sigma,\varepsilon\eta,\frac{y}{\sqrt{\varepsilon}},\sqrt{\varepsilon}\sigma,\sqrt{\varepsilon}\eta\right)\right)\end{array}$ (13)
In particular, we have, for all $a\in\mathfrak{A}$,
$\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}(a)$ $=\mathrm{Op}_{\varepsilon}\left(a\left(s,y,\sigma,\eta,\frac{(y,\sigma,\eta)}{\sqrt{\varepsilon}}\right)\right)$
$=\mathfrak{T}_{\sqrt{\varepsilon}}^{\dagger}\mathrm{Op}_{1}\Big{(}a\left(s,\sqrt{\varepsilon}y,\varepsilon\sigma,\sqrt{\varepsilon}\eta,y,\sqrt{\varepsilon}\sigma,\eta\right)\Big{)}\mathfrak{T}_{\sqrt{\varepsilon}},\ \left(\text{semiclassical in }\sqrt{\varepsilon}\right)$ (14)
with for all $f\in\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ and for all $(s,y)\in\mathbb{R}^{2},$
$\mathfrak{T}_{\sqrt{\varepsilon}}^{\dagger}[f](s,y)=\frac{1}{\varepsilon^{1/4}}f\left(s,\frac{y}{\sqrt{\varepsilon}}\right).$ (15)
Notably, $\mathfrak{T}_{\sqrt{\varepsilon}}$ is a unitary operator from $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ to $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$. Moreover, the quantization map $\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}$ associates to $a$ a bounded operator on $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ according to the Calderón-Vaillancourt Theorem (7) applied to expression (14).
These observables allow us to differentiate three regions of the phase space $\mathbb{R}_{s}\times\mathrm{I}_{1/2}\times\mathbb{R}^{2}_{\sigma,\eta}$. Let us explain this fact. Let $a\in\mathfrak{A}$ and $\chi\in\mathscr{C}^{\infty}_{c}(\mathbb{R},[0,1])$ supported in $[-1,1]$ with $\chi\equiv 1$ on $[-1/2,1/2]$. Let $\delta>0$, $R>R_{0}$ with $R_{0}$ associated to $a$ and let us consider $\varepsilon$ small enough such that
$2R\sqrt{\varepsilon}<\delta.$ (16)
Let us denote $x\coloneqq(s,y),\xi\coloneqq(\sigma,\eta)$ and $z\coloneqq\left(y,\sigma,\eta\right)$ for simplicity. We now split $a$ into three observables,
$a\left(x,\xi,z\right)=$ $\ a\left(x,\xi,z\right)\left(1-\chi\left(\frac{\sqrt{\varepsilon}\left|z\right|}{\delta}\right)\right)$ $\left(\eqqcolon a^{\delta}_{\varepsilon}(x,\xi,z)\right)$ (17)
$+\ a\left(x,\xi,z\right)\chi\left(\frac{\sqrt{\varepsilon}\left|z\right|}{\delta}\right)\left(1-\chi\left(\frac{\left|z\right|}{R}\right)\right)$ $\left(\eqqcolon a_{\varepsilon,\delta}^{R}(x,\xi,z)\right)$ (18)
$+\ a\left(x,\xi,z\right)\chi\left(\frac{\left|z\right|}{R}\right)$ $\left(\eqqcolon a_{\varepsilon,R}(x,\xi,z)\right).$ (19)
Let us define $\chi_{R}\coloneqq\chi\left(\frac{\left|\cdot\right|}{R}\right)$ and $\chi^{R}\coloneqq 1-\chi_{R}$. The symbol $a_{\varepsilon}^{\delta}$ through the second scale quantization is supported outside $\mathcal{C}$, at a distance $\delta$ of $\mathcal{C}$. Because of (16), $\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a_{\varepsilon}^{\delta}\right)$ is a semiclassical operator which satisfies
$\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a_{\varepsilon}^{\delta}\right)=\mathrm{Op}_{1}\left(a_{\infty}\left(s,y,\varepsilon\sigma,\varepsilon\eta,\frac{(y,\varepsilon\sigma,\varepsilon\eta)}{|(y,\varepsilon\sigma,\varepsilon\eta)|}\right)\left(1-\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{\delta}\right)\right)\right).$
So, we can apply the semiclassical theory recalled in Section 2 to the observable $a_{\varepsilon}^{\delta}$.
In the second scale quantization, the symbol $a_{\varepsilon,R}$ is supported at distance of order $\sqrt{\varepsilon}$ of $\mathcal{C}$ and $a_{\varepsilon,\delta}^{R}$ is supported at distance larger than $\sqrt{\varepsilon}$ of $\mathcal{C}$ but smaller than $\delta$. We are particularly interested by the two regions covered by $a_{\varepsilon,\delta}^{R}$ and $a_{\varepsilon,R}$. They will describe the measure near the curve $\mathcal{C}$.
For the observable $a_{\varepsilon,\delta}^{R}$, we have
$\begin{split}\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a_{\varepsilon,\delta}^{R}\right)=&\ \mathrm{Op}_{1}\left(a\left(s,0,0,0,\frac{y}{\sqrt{\varepsilon}},\sqrt{\varepsilon}\sigma,\sqrt{\varepsilon}\eta\right)\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{\delta}\right)\left(1-\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{R\sqrt{\varepsilon}}\right)\right)\right)\\ &+\mathcal{O}\left(\delta\right).\end{split}$ (20)
But we choose $R>R_{0}$ then
$\begin{split}\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a_{\varepsilon,\delta}^{R}\right)=&\ \mathrm{Op}_{1}\left(a_{\infty}\left(s,0,0,0,\frac{(y,\varepsilon\sigma,\varepsilon\eta)}{\left|(y,\varepsilon\sigma,\varepsilon\eta)\right|}\right)\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{\delta}\right)\left(1-\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{R\sqrt{\varepsilon}}\right)\right)\right)\\ &+\mathcal{O}\left(\delta\right).\end{split}$
By using (14) on $a_{\varepsilon,R}$, we have
$\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a_{\varepsilon,R}\right)$ $=\mathrm{Op}_{1}\left(a(s,0,0,0,\frac{y}{\sqrt{\varepsilon}},\sqrt{\varepsilon}\sigma,\sqrt{\varepsilon}\eta)\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{R\sqrt{\varepsilon}}\right)\right)\ +\ \mathcal{O}\left(\sqrt{\varepsilon}\right),$
$=\mathfrak{T}_{\sqrt{\varepsilon}}^{\dagger}\mathrm{Op}_{1}\left(a(s,0,0,0,y,\sqrt{\varepsilon}\sigma,\eta)\chi\left(\frac{|(y,\sqrt{\varepsilon}\sigma,\eta)|}{R}\right)\right)\mathfrak{T}_{\sqrt{\varepsilon}}\ +\ \mathcal{O}\left(\sqrt{\varepsilon}\right).$ (21)
So we will consider the separable Hilbert space $\mathcal{H}\coloneqq\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)$ and the setting of Appendix B: the functional space $\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathcal{L}\left(\mathcal{H}\right)\right)$, and the two-scale quantization defined for $b\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathcal{L}\left(\mathcal{H}\right)\right)$,
$\mathrm{Op}_{\sqrt{\varepsilon},1}(b)\coloneqq\mathrm{Op}_{1}\Big{(}b\left(s,\sqrt{\varepsilon}\sigma\right)\Big{)}.$ (22)
Let us denote, for all $(s,\sigma)\in\mathbb{R}^{2}$, the bounded operator on $\mathcal{H}$ obtained by the Weyl quantization of $(y,\eta)\mapsto a(s,0,0,0,y,\sigma,\eta)$ by $a^{W}(s,0,0,0,y,\sigma,\mathrm{D}_{y})$. Then, for the operator $b(s,\sigma)$ defined by for all $(s,\sigma)\in\mathbb{R}^{2}$, $b(s,\sigma)\coloneqq a^{W}(s,0,0,0,y,\sigma,\mathrm{D}_{y})\chi\left(\frac{|(y,\sigma,\mathrm{D}_{y})|}{R}\right)$, we have
$\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a_{\varepsilon,R}\right)=\mathfrak{T}_{\sqrt{\varepsilon}}^{\dagger}\mathrm{Op}_{\sqrt{\varepsilon},1}(b)\mathfrak{T}_{\sqrt{\varepsilon}}\ +\ \mathcal{O}\left(\sqrt{\varepsilon}\right).$ (23)
In the following sections, we will take advantage of these properties of $a_{\varepsilon,R}$ and $a_{\varepsilon,\delta}^{R}$.
3.2.2 Two-scale semiclassical measures
Let us first consider the stationary case. We use the observable of $\mathfrak{A}$ to analyze the oscillations at the scale $\sqrt{\varepsilon}$ and $\delta$ of uniformly bounded families $\left(f^{\varepsilon}\right)_{\varepsilon>0}$ of $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ through $a_{\varepsilon,R}$ and $a_{\varepsilon,\delta}^{R}$, respectively. Let us define for $a\in\mathfrak{A}$ the quantity
$\mathcal{I}^{\varepsilon}_{f}(a)\coloneqq\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}(a)f^{\varepsilon},f^{\varepsilon}\right\rangle_{\mathrm{L}^{2}(\mathbb{R}^{2},\mathbb{C}^{2})}.$
Theorem 3.3 (Existence of the two-scale semiclassical measure).
Let ${(f^{\varepsilon})_{\varepsilon>0}}$ uniformly bounded in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$, then, there exist ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ a vanishing sequence of positive numbers and
  • $\bullet$
    Two measurable map $\mathrm{d}\nu_{\infty}$ and $\mathrm{d}\gamma$, valued in the set of scalar non negative measures on $\mathcal{C}\times\mathbb{S}^{2}$ and $\mathbb{R}^{2}$, respectively,
  • $\bullet$
    A $\mathrm{d}\nu_{\infty}$-measurable map $(s,\omega)\mapsto M_{\infty}(s,\omega)\in\mathbb{C}^{2,2}$, valued in the set of Hermitian trace $1$ positive matrices,
  • $\bullet$
    A $\mathrm{d}\gamma(s,\sigma)$ measurable map $(s,\sigma)\mapsto\Gamma(s,\sigma)$, valued in the set of trace $1$ positive operator on $\mathrm{L}^{2}\left(\mathbb{R},\mathbb{C}^{2}\right)$,
such that for all $a\in\mathfrak{A}$ we have
$\mathcal{I}^{\varepsilon_{k}}_{f}(a)\underset{k\to+\infty}{\longrightarrow}$ $\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\mathrm{tr}\left(a_{\infty}\left(s,y,\sigma,\eta,\frac{(y,\sigma,\eta)}{|(y,\sigma,\eta)|}\right)\mathrm{d}\mu(s,y,\sigma,\eta)\right)$ (24)
+×$+\ \int_{\mathbb{R}\times\mathbb{S}^{2}}\mathrm{tr}\Big{(}a_{\infty}(s,0,0,0,\omega)M_{\infty}(s,\omega)\Big{)}\mathrm{d}\nu_{\infty}(s,\omega)$ (25)
$+\ \int_{\mathbb{R}^{2}}\mathrm{Tr}\Big{(}a^{W}(s,0,0,0,y,\sigma,\mathrm{D}_{y})\Gamma(s,\sigma)\Big{)}\mathrm{d}\gamma(s,\sigma),$ (26)
where $\mathrm{d}\mu$ the semiclassical measure of ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$ associated to ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ and for all $(s,\sigma)\in\mathbb{R}^{2}$, $a^{W}(s,0,0,0,y,\sigma,\mathrm{D}_{y})$ denotes the bounded operator on $\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)$ obtained by the Weyl quantization of $(y,\eta)\mapsto a(s,0,0,0,y,\sigma,\eta)$.
The proof of this theorem is in Appendix B.
We denote by $\Gamma\mathrm{d}\gamma$ the pair $\left(\Gamma,\mathrm{d}\gamma\right)$ with $\Gamma$ a trace one positive operator $\mathrm{d}\gamma$-measurable and $\mathrm{d}\gamma$ a non negative measure. We denote by $M_{\infty}\mathrm{d}\nu_{\infty}$ the pair $\left(M_{\infty},\mathrm{d}\nu_{\infty}\right)$ with $M_{\infty}$ a Hermitian trace one positive matrix $\mathrm{d}\nu_{\infty}$-measurable and $\mathrm{d}\nu_{\infty}$ a non negative measure.
Definition (Two-scale semiclassical measure).
We call two-scale semiclassical measures, associated with the concentration of ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$ near the curve $\mathcal{C}$ of order between $\sqrt{\varepsilon}$ and $\delta$ and of order $\sqrt{\varepsilon}$, respectively, the pair ${\left(M_{\infty}\mathrm{d}\nu_{\infty},\Gamma\mathrm{d}\gamma\right)}$. We will say that $\Gamma\mathrm{d}\gamma$ is a two-scale semiclassical measure at finite distance and $M_{\infty}\mathrm{d}\nu_{\infty}$ is a two-scale semiclassical measure at infinity.
Remark 3.4 (Link with the limit of the previous observable).
Let ${(f^{\varepsilon})_{\varepsilon>0}}$ uniformly bounded in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ and ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ a vanishing sequence given by Theorem 3.3, then, with the same notations, for all $a\in\mathfrak{A}$,
$\limsup_{\delta\to 0}\limsup_{R\to+\infty}\lim_{k\to+\infty}\mathcal{I}^{\varepsilon_{k}}_{f}\left(a_{\varepsilon_{k}}^{\delta}\right)=$ $\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\mathrm{tr}\left(a_{\infty}\left(s,y,\sigma,\eta,\frac{(y,\sigma,\eta)}{|(y,\sigma,\eta)|}\right)\mathrm{d}\mu(s,y,\sigma,\eta)\right),$
$\limsup_{\delta\to 0}\limsup_{R\to+\infty}\lim_{k\to+\infty}\mathcal{I}^{\varepsilon_{k}}_{f}\left(a_{\varepsilon_{k},\delta}^{R}\right)=$$\int_{\mathbb{R}\times\mathbb{S}^{2}}\mathrm{tr}\Big{(}a_{\infty}(s,0,0,0,\omega)M_{\infty}(s,\omega)\Big{)}\mathrm{d}\nu_{\infty}(s,\omega),$
$\limsup_{\delta\to 0}\limsup_{R\to+\infty}\lim_{k\to+\infty}\mathcal{I}^{\varepsilon_{k}}_{f}\left(a_{\varepsilon_{k},R}\right)=$ $\int_{\mathbb{R}^{2}}\mathrm{Tr}\Big{(}a^{W}(s,0,0,0,y,\sigma,\mathrm{D}_{y})\Gamma(s,\sigma)\Big{)}\mathrm{d}\gamma(s,\sigma).$
Example.
Let us consider, for all $(s,y)\in\mathbb{R}^{2}$,
$\phi_{\varepsilon}(s,y)=\frac{1}{\varepsilon^{1/4}}u_{\varepsilon}\left(s\right)f\left(\frac{y}{\sqrt{\varepsilon}}\right)\vec{V}$
with $\vec{V}\in\mathbb{C}^{2}$, $(f,u_{\varepsilon})\in\mathscr{S}(\mathbb{R},\mathbb{C})^{2}$ and the knowledge of $\mathrm{d}\rho$ the semiclassical measure of $(u_{\varepsilon})_{\varepsilon>0}$ at the scale $\sqrt{\varepsilon}$, then for all $a\in\mathscr{C}^{\infty}_{c}(\mathbb{R}^{7},\mathbb{C}^{2,2})$,
$\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}(a)\phi_{\varepsilon},\phi_{\varepsilon}\right\rangle\underset{\varepsilon\to 0}{\longrightarrow}\int_{\mathbb{R}^{2}}\mathrm{Tr}\Big{(}a^{W}(s,0,0,0,y,\sigma,\mathrm{D}_{y})\Gamma(s,\sigma)\Big{)}\mathrm{d}\gamma(s,\sigma),$
where
$\Gamma(s,\sigma)\mathrm{d}\gamma(s,\sigma)=\frac{1}{2\pi}\Pi_{f\vec{V}}\mathrm{d}\rho(s,\sigma),$
with $\Pi_{f\vec{V}}$ the projector on $\mathrm{L}^{2}\left(\mathbb{R},\mathbb{C}^{2}\right)$ over the function $f\vec{V}$. Notably, if $u_{\varepsilon}$ is a wave packet, which means for all $s\in\mathbb{R}$,
$u_{\varepsilon}\left(s\right)\coloneqq\frac{1}{\varepsilon^{1/4}}u\left(\frac{s}{\sqrt{\varepsilon}}\right)$
with $u\in\mathscr{S}(\mathbb{R},\mathbb{C})$, then the semiclassical measure of $(u_{\varepsilon})_{\varepsilon>0}$ at the scale $\sqrt{\varepsilon}$ is $\mathrm{d}\rho(\sigma)=\left|\mathcal{F}u(\sigma)\right|^{2}\mathrm{d}\sigma$ (according to the case $\alpha=1$ of Example 2.1 but with a $\sqrt{\varepsilon}$-Weyl quantization) so
$\Gamma(s,\sigma)\mathrm{d}\gamma(s,\sigma)=\left|\mathcal{F}u(\sigma)\right|^{2}\Pi_{f\vec{V}}\frac{\mathrm{d}\sigma}{2\pi}.$
Remark 3.5.
Let ${(f^{\varepsilon})_{\varepsilon>0}}$ uniformly bounded in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ and ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ a vanishing sequence given by Theorem 3.3. According to the notations of Theorem 3.3, the two-scale semiclassical measures ${\left(M_{\infty}\mathrm{d}\nu_{\infty},\Gamma\mathrm{d}\gamma\right)}$ satisfy for all $a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)$,
$\mathcal{I}^{\varepsilon_{k}}_{f}(a)\underset{k\to+\infty}{\longrightarrow}$ $\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\mathrm{tr}\Big{(}a\left(s,y,\sigma,\eta\right)\mathrm{d}\mu(s,y,\sigma,\eta)\Big{)}$
+×$+\ \int_{\mathbb{R}\times\mathbb{S}^{2}}\mathrm{tr}\Big{(}a(s,0,0,0)M_{\infty}(s,\omega)\Big{)}\mathrm{d}\nu_{\infty}(s,\omega)$
$+\ \int_{\mathbb{R}^{2}}\mathrm{Tr}\Big{(}a(s,0,0,0)\Gamma(s,\sigma)\Big{)}\mathrm{d}\gamma(s,\sigma).$
Corollary 3.6 (A condition to have $M_{\infty}\mathrm{d}\nu_{\infty}=0$).
Let ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$ uniformly bounded in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ such that there exists $\chi$ a cut off function with $\chi\in\mathscr{C}^{\infty}_{c}(\mathbb{R},\mathbb{R})$ supported in $[-1,1]$ and $\chi\equiv 1$ on $[-1/2,1/2]$ such that
$\limsup_{\varepsilon\to 0}\left\lVert\mathrm{Op}_{1}\left(1-\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{R\sqrt{\varepsilon}}\right)\right)f^{\varepsilon}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\underset{R\to+\infty}{\longrightarrow}0.$
Then, for all two-scale semiclassical measure $M_{\infty}\mathrm{d}\nu_{\infty}$ at infinity of ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$,
$M_{\infty}\mathrm{d}\nu_{\infty}=0.$
Proof of Corollary 3.6.
Let ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$ uniformly bounded in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ and $\chi$ a cut off function, $\chi\in\mathscr{C}^{\infty}_{c}(\mathbb{R},[0,1])$ supported in $[-1,1]$ and $\chi\equiv 1$ on $[-1/2,1/2]$ such that
$\limsup_{\varepsilon\to 0}\left\lVert\mathrm{Op}_{1}\left(1-\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{R\sqrt{\varepsilon}}\right)\right)f^{\varepsilon}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\underset{R\to+\infty}{\longrightarrow}0.$
For all $a\in\mathfrak{A}$, we consider,
$\alpha(s,y,\sigma,\eta)$ $\coloneqq a\left(s,y,\varepsilon\sigma,\varepsilon\eta,\frac{y}{\sqrt{\varepsilon}},\sqrt{\varepsilon}\sigma,\sqrt{\varepsilon}\eta\right)\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{\delta}\right),$
$\beta(y,\sigma,\eta)$ $\coloneqq 1-\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{R\sqrt{\varepsilon}}\right).$
According to pseudo-differential calculus and of the inequality (16), we fix $\delta$ and let $\varepsilon$ goes to 00 then $R$ goes to +$\infty$+ ∞, so
$\mathrm{Op}_{1}\Big{(}\alpha(s,y,\sigma,\eta)\beta(y,\sigma,\eta)\Big{)}=$ $\ \mathrm{Op}_{1}\Big{(}\alpha(s,y,\sigma,\eta)\Big{)}\mathrm{Op}_{1}\Big{(}\beta(y,\sigma,\eta)\Big{)}$
$+\ \mathcal{O}_{\mathcal{L}\left(\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)}\left(\sqrt{\varepsilon}\right)+\mathcal{O}_{\mathcal{L}\left(\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)}\left(\frac{1}{R}\right).$
Moreover,
$\mathrm{Op}_{1}\Big{(}\alpha(s,y,\sigma,\eta)\Big{)}=\mathfrak{T}_{\sqrt{\varepsilon}}^{\dagger}\mathrm{Op}_{1}\left(a\left(s,\sqrt{\varepsilon}y,\varepsilon\sigma,\sqrt{\varepsilon}\eta,y,\sqrt{\varepsilon}\sigma,\eta\right)\chi\left(\frac{\sqrt{\varepsilon}|(y,\sqrt{\varepsilon}\sigma,\eta)|}{\delta}\right)\right)\mathfrak{T}_{\sqrt{\varepsilon}}.$
By application of the Calderón-Vaillancourt Theorem (7),
$\left\lVert\mathrm{Op}_{1}\Big{(}\alpha(s,y,\sigma,\eta)\Big{)}\right\rVert_{\mathcal{L}\left(\mathrm{L}^{2}\right)}$ $=\left\lVert\mathrm{Op}_{1}\left(a\left(s,\sqrt{\varepsilon}y,\varepsilon\sigma,\sqrt{\varepsilon}\eta,y,\sqrt{\varepsilon}\sigma,\eta\right)\chi\left(\frac{\sqrt{\varepsilon}|(y,\sqrt{\varepsilon}\sigma,\eta)|}{\delta}\right)\right)\right\rVert_{\mathcal{L}\left(\mathrm{L}^{2}\right)}$
$\leqslant C(a)$
where $C$ does not depend of $\varepsilon$. Then
$\mathcal{I}^{\varepsilon}_{f}\left(a_{\varepsilon,\delta}^{R}\right)=$ $\ \Big{\langle}\mathrm{Op}_{1}\Big{(}\alpha(s,y,\sigma,\eta)\Big{)}f^{\varepsilon},\mathrm{Op}_{1}\Big{(}\beta(y,\sigma,\eta)\Big{)}f^{\varepsilon}\Big{\rangle}_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}$
$+\ \mathcal{O}\left(\sqrt{\varepsilon}\right)+\mathcal{O}\left(\frac{1}{R}\right),$
$\left|\mathcal{I}^{\varepsilon}_{f}\left(a_{\varepsilon,\delta}^{R}\right)\right|\leqslant$ $\ \left\lVert\mathrm{Op}_{1}\Big{(}\alpha(s,y,\sigma,\eta)\Big{)}f^{\varepsilon}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\left\lVert\mathrm{Op}_{1}\Big{(}\beta(y,\sigma,\eta)\Big{)}f^{\varepsilon}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}$
$+\ \mathcal{O}\left(\sqrt{\varepsilon}\right)+\mathcal{O}\left(\frac{1}{R}\right),$
$\leqslant$ $\ C(a)\left\lVert f^{\varepsilon}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\left\lVert\mathrm{Op}_{1}\left(1-\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{R\sqrt{\varepsilon}}\right)\right)f^{\varepsilon}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}$
$+\ \mathcal{O}\left(\sqrt{\varepsilon}\right)+\mathcal{O}\left(\frac{1}{R}\right).$
Then,
$\limsup_{\delta\to 0}\limsup_{R\to+\infty}\limsup_{\varepsilon\to 0}\mathcal{I}^{\varepsilon}_{f}\left(a_{\varepsilon,\delta}^{R}\right)=0.$
We conclude by using Remark 3.4. ∎
3.2.3 Time dependent two-scale semiclassical measures
As in Section 2.2.2, one can extend the definition of two-scale semiclassical measure to time dependent families. More precisely, we consider families ${\left(f^{\varepsilon}_{t}\right)_{\varepsilon>0}}$ that are uniformly bounded in $\mathrm{L}^{\infty}\left(\mathbb{R},\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)$. Let $a\in\mathfrak{A},\Xi\in\mathscr{C}_{c}^{\infty}(\mathbb{R},\mathbb{C})$, we define the quantity
$\mathcal{I}^{\varepsilon}_{f}(\Xi,a)\coloneqq\int_{\mathbb{R}}\Xi(t)\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}(a)f_{t}^{\varepsilon},f_{t}^{\varepsilon}\right\rangle_{\mathrm{L}^{2}(\mathbb{R}^{2},\mathbb{C}^{2})}\mathrm{d}t.$
As before, we are interested by the limit of $\mathcal{I}^{\varepsilon}_{f}(\Xi,a)$ when $\varepsilon$ goes to 00. Moreover, if $a$ does not depend of $z$, then the limit will be the semiclassical measure as defined in (8). Adapting the context of  [10] to time dependent families, one defines two-scale semiclassical measures over the curve $\mathcal{C}$. Indeed, by the same argument of Section 2.2.2, we can extend Theorem 3.3 to time families. Therefore, such as before, the limit measures will be absolutely continuous in time variable with respect to Lebesgue measure.
Theorem 3.7 (Existence of the time-averaged two-scale semiclassical measure).
Let ${(f^{\varepsilon}_{t})_{\varepsilon>0}}$ uniformly bounded in $\mathrm{L}^{\infty}\left(\mathbb{R},\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)$, then, there exist ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ a vanishing sequence and
  • $\bullet$
    Two measurables map $t\mapsto\mathrm{d}\nu^{t}_{\infty}$ and $t\mapsto\mathrm{d}\gamma^{t}$, valued in the set of scalar non negative measures on $\mathcal{C}\times\mathbb{S}^{2}$ and $\mathbb{R}^{2}$, respectively,
  • $\bullet$
    A $\mathrm{d}\nu^{t}_{\infty}\otimes\mathrm{d}t$-measurable map $(t,s,\omega)\mapsto M^{t}_{\infty}(s,\omega)\in\mathbb{C}^{2,2}$, valued in the set of Hermitian trace $1$ positive matrices,
  • $\bullet$
    A $\mathrm{d}\gamma^{t}(s,\sigma)\otimes\mathrm{d}t$ measurable map $(t,s,\sigma)\mapsto\Gamma^{t}(s,\sigma)$, valued in the set of trace $1$ positive operator on $\mathrm{L}^{2}\left(\mathbb{R},\mathbb{C}^{2}\right)$,
such that for all $(a,\Xi)\in\mathfrak{A}\times\mathscr{C}_{c}^{\infty}(\mathbb{R},\mathbb{C})$,
$\mathcal{I}^{\varepsilon_{k}}_{f}(\Xi,a)\underset{k\to+\infty}{\longrightarrow}$ $\int_{\mathbb{R}}\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\Xi(t)\mathrm{tr}\left(a_{\infty}\left(s,y,\sigma,\eta,\frac{(y,\sigma,\eta)}{|(y,\sigma,\eta)|}\right)\mathrm{d}\mu_{t}(s,y,\sigma,\eta)\right)\mathrm{d}t$
+×$+\ \int_{\mathbb{R}}\int_{\mathbb{R}\times\mathbb{S}^{2}}\Xi(t)\mathrm{tr}\Big{(}a_{\infty}(s,0,0,0,\omega)M_{\infty}^{t}(s,\omega)\Big{)}\mathrm{d}\nu_{\infty}^{t}(s,\omega)\mathrm{d}t$
$+\ \int_{\mathbb{R}}\int_{\mathbb{R}^{2}}\Xi(t)\mathrm{Tr}\Big{(}a^{W}(s,0,0,0,y,\sigma,\mathrm{D}_{y})\Gamma^{t}(s,\sigma)\Big{)}\mathrm{d}\gamma^{t}(s,\sigma)\mathrm{d}t,$
where $\mathrm{d}\mu_{t}$ is the semiclassical measure of ${(f^{\varepsilon}_{t})_{\varepsilon>0}}$ associated to ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ and for all $(s,\sigma)\in\mathbb{R}^{2}$, $a^{W}(s,0,0,0,y,\sigma,\mathrm{D}_{y})$ denotes the bounded operator on $\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)$ obtained by the Weyl quantization of $(y,\eta)\mapsto a(s,0,0,0,y,\sigma,\eta)$.
Remark (Geometrical point of view).
In the parametrization $(s,y)$ of ×$\mathbb{R}\times\mathrm{I}$, the variables $(s,y,\sigma,\eta)$ parametrized $\mathrm{T}^{*}\mathbb{R}^{2}$. Moreover the variable $s$ parametrizes the curve $\mathrm{E}\subset\mathbb{R}^{2}$. The variable $y$ parametrizes the normal fiber to the curve $\mathrm{E}$ at the point $\gamma(s)$. In the computation of $\mathrm{Tr}$, the space $\mathrm{L}^{2}(\mathbb{R}_{y},\mathbb{C}^{2})$ can be understand as the space $\mathrm{L}^{2}(N_{\gamma(s)}\mathrm{E},\mathbb{C}^{2})$. The variable $\sigma$ denotes the variable of the fiber of the cotangent bundle $\mathrm{T}^{*}\mathrm{E}$ over $\mathrm{E}$. For $\gamma(s)\in\mathrm{E}$ and $\mathrm{T}_{s}\mathrm{E}$, the tangent to $\mathrm{E}$ above $s$, the fiber of the cotangent bundle $\mathrm{T}^{*}\mathrm{E}$ over $\gamma(s)$ is $(\mathrm{T}_{s}\mathrm{E})^{*}\sim\mathbb{R}$. The measure $\mathrm{d}\gamma^{t}$ defines a measure on $\mathrm{T}^{*}\mathrm{E}$.
Since $\mathcal{C}=\{y=\eta=\sigma=0\}\subset\mathrm{T}^{*}\mathbb{R}^{2}$, $\mathcal{C}\times\mathbb{S}^{2}$ coincides to the normal bundle in sphere to $\mathcal{C}$: $\mathrm{S}\mathcal{C}$. The fibers of $\mathrm{S}\mathcal{C}$ over $\rho\in\mathcal{C}$ is obtained as follows : let $\mathrm{T}_{\rho}\mathcal{C}$ be the tangent space to $\mathcal{C}$ over $\rho$, define $\mathrm{N}_{\rho}\mathcal{C}$ the subspace of $\left(\mathrm{T}_{\rho}\left(\mathrm{T}^{*}\mathbb{R}^{2}\right)\right)^{*}$ consisting of the linear forms on $\mathrm{T}_{\rho}\left(\mathrm{T}^{*}\mathbb{R}^{2}\right)$ which vanish on $\mathrm{T}_{\rho}\mathcal{C}$, $S_{\rho}\mathcal{C}$ is obtained by taking the quotient of $\mathrm{N}_{\rho}\mathcal{C}$ by the action of $\mathbb{R}_{\geqslant 0}$ by homotheties.
3.3 Main result
Now we will describe the semiclassical measure of Theorem 3.7 for the solution to equation (1). Let us define some quantities involved in the main theorem. Let us define the Hermite functions $\left(\mathfrak{h}_{m}\right)_{m\in\mathbb{N}_{\geqslant 0}}$ as the Hilbertian basis of $\mathrm{L}^{2}\left(\mathbb{R},\mathbb{C}\right)$ satisfying for all $y\in\mathbb{R}$ and for all $m\in\mathbb{N}_{\geqslant 0}$,
$-\frac{\mathrm{d}^{2}\mathfrak{h}_{m}}{\mathrm{d}y^{2}}(y)+y^{2}\mathfrak{h}_{m}(y)=(2m+1)\mathfrak{h}_{m}(y).$ (27)
Let us consider for all $(s,y,\sigma)\in\mathbb{R}^{3}$,
$g_{0}^{s}(y)\coloneqq\frac{1}{\sqrt{r(s)}}\begin{pmatrix}0\\ \mathfrak{h}_{0}\big{(}r(s)y\big{)}\end{pmatrix}\;\;$ $,\;\;g_{n}^{s,\sigma}(y)\coloneqq\frac{1}{\sqrt{2r(s)}}\sqrt{1+\frac{\sigma}{\lambda_{n}(s,\sigma)}}\begin{pmatrix}\mathfrak{h}_{|n|-1}\big{(}r(s)y\big{)}\\ \frac{\sqrt{2|n|}r(s)}{\lambda_{n}(s,\sigma)+\sigma}\mathfrak{h}_{|n|}\big{(}r(s)y\big{)}\end{pmatrix},$ (28)
$\Pi_{0}(s)\coloneqq g_{0}^{s}\otimes g_{0}^{s}\;\;$ $,\;\;\Pi_{n}(s,\sigma)\coloneqq g_{n}^{s,\sigma}\otimes g_{n}^{s,\sigma},$ (29)
where $\left(\lambda_{n}\right)_{n\in\mathbb{Z}}$ were defined in (5), $r(s)\coloneqq|\nabla\mathrm{m}(\gamma(s))|^{1/2}$. Although the functions $g_{0}^{s}$ and $\Pi_{0}$ do not depend on $\sigma$, we also will denote them by $g_{0}^{s,\sigma}$ and $\Pi_{0}(s,\sigma)$.
Let us consider for all $s\in\mathbb{R}$ and for all $z\in\mathbb{R}^{3}$,
$\Pi_{\pm}^{\infty}(s,z)\coloneqq\frac{1}{2}\mathrm{Id}\pm\frac{1}{2\sqrt{z_{\sigma}^{2}+z_{y}^{2}r(s)^{4}+z_{\eta}^{2}}}\begin{pmatrix}z_{\sigma}&z_{y}r(s)^{2}+iz_{\eta}\\ z_{y}r(s)^{2}-iz_{\eta}&-z_{\sigma}\end{pmatrix}$ (30)
We are going to use the two-scale semiclassical measures of the family $\left(\varphi_{t}^{\varepsilon}\right)_{\varepsilon>0}$ to describe the semiclassical measure of $\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}$ near the curve $\mathcal{C}$. We already know its semiclassical measure and its evolution outside $\mathcal{C}$. Therefore, we will focus on terms (25) and (26).
Theorem 3.8 (Evolution of the two-scale semiclassical measure in normal geodesic coordinates).
Assume Assumption 1.2 and 1.3. Let $\left(\varphi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ uniformly bounded in $\mathrm{L}^{\infty}\left(\mathbb{R},\mathrm{L}^{2}\left(\mathbb{R}\times\mathrm{I},\mathbb{C}^{2}\right)\right)$ defined in (11). Let ${\left(M^{t}_{\infty}\mathrm{d}\nu^{t}_{\infty},\Gamma^{t}\mathrm{d}\gamma^{t}\right)}$ be a pair of two-scale semiclassical measure of $\left(\varphi^{\varepsilon}_{t}\right)_{\varepsilon>0}$, then we have the following properties.
  1. 1.
    There exists a family of non negative scalar Radon measure $\left(\mathrm{d}\gamma_{n}^{t}\right)_{n\in\mathbb{Z}}$ such that
    $\Gamma^{t}\mathrm{d}\gamma^{t}=\sum_{n\in\mathbb{Z}}\Pi_{n}\mathrm{d}\gamma_{n}^{t}.$ (31)
    Moreover, for all $n\in\mathbb{Z}$, $\mathrm{d}\gamma_{n}^{t}$ satisfies the following equation
    $\left\{\begin{array}[]{rll}\partial_{t}\ \mathrm{d}\gamma_{n}^{t}&=&\left\{\mathrm{d}\gamma_{n}^{t},\lambda_{n}\right\}\\ \mathrm{d}\gamma_{n}^{t}|_{t=0}&=&\mathrm{d}\gamma_{n,0}\end{array}\right.,$
    with $\mathrm{d}\gamma_{n,0}=\mathrm{tr}\Big{(}\Pi_{n}\Gamma^{0}\Big{)}\mathrm{d}\gamma^{0}$ and $\Gamma^{0}\mathrm{d}\gamma^{0}$ is a two-scale semiclassical measure associated to the initial data.
  2. 2.
    There exists a family of non negative scalar Radon measure $\mathrm{d}\nu^{t}_{\pm}$ such that
    $M^{t}_{\infty}\mathrm{d}\nu^{t}_{\infty}=\Pi_{+}^{\infty}\mathrm{d}\nu^{t}_{+}+\Pi_{-}^{\infty}\mathrm{d}\nu^{t}_{-}.$ (32)
    Moreover, $\mathrm{d}\nu_{\pm}^{t}$ satisfies the invariance $\mathrm{div}_{\omega}\left(\vec{V}^{\infty}(s,\omega)\mathrm{d}\nu_{\pm}^{t}(s,\omega)\right)=0,$ with for all $s\in\mathbb{R}$ and for all $\omega\coloneqq\left(\omega_{y},\omega_{\sigma},\omega_{\eta}\right)\in\mathbb{S}^{2}$,
    $\vec{V}^{\infty}(s,\omega)\coloneqq\begin{pmatrix}\left(\omega_{y}\omega_{\eta}\left(1-r(s)^{4}\right)-1\right)\omega_{\eta}\\ \left(1-r(s)^{4}\right)\omega_{y}\omega_{\eta}\omega_{\sigma}\\ \left(\omega_{y}\omega_{\eta}\left(1-r(s)^{4}\right)+r(s)^{4}\right)\omega_{y}\end{pmatrix}.$
The proof of this theorem is in Section 4.3.
Remark (Continuity of the two-scale semiclassical measure at finite distance).
From the evolution (6), we can deduce that the measure $t\mapsto\Gamma^{t}\mathrm{d}\gamma^{t}$ is continuous and entirely determined by the evolution equation and the initial data.
Remark (Value at zero of the two-scale semiclassical measure at finite distance).
The value at zero of $\Gamma^{t}\mathrm{d}\gamma^{t}$ is not necessarily the two-scale semiclassical measure at finite distance of the initial condition. As we will see in the proof, $\Gamma^{t}(s,\sigma)$ commutes with the operator $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$ because of the PDE satisfied by $\left(\varphi^{\varepsilon}_{t}\right)_{\varepsilon>0}$. Nevertheless, there exists initial condition such that its two-scale semiclassical measure at finite distance does not commute with the operator $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$. We will see an example in Remark 4.4.
Corollary 3.9 (Evolution consequence).
With the notation of Theorem 3.8, if
$\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\mathrm{tr}\Big{(}\mathrm{d}\mu_{0}(x,\xi)\Big{)}+\int_{\mathbb{R}^{2}}\mathrm{Tr}\Big{(}\Gamma^{0}(s,\sigma)\Big{)}\mathrm{d}\gamma^{0}(s,\sigma)=1,$
with $\mathrm{d}\mu_{0}$ a semiclassical measure of $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ associated to the same vanishing sequence $\left(\varepsilon_{n}\right)_{n\in\mathbb{N}}$ of the two-scale semiclassical measure $\left(M^{0}_{\infty}\mathrm{d}\nu^{0}_{\infty},\Gamma^{0}\mathrm{d}\gamma^{0}\right)$, then, for almost all $t$,
$M^{t}_{\infty}(s,\omega)\mathrm{d}\nu^{t}_{\infty}(s,\omega)=0.$
Proof of Corollary 3.9.
With the notation of Theorem 3.8, the evolution of $(\mathrm{d}\mu_{t},\Gamma^{t}\mathrm{d}\gamma^{t})$ implies that for all $t\in\mathbb{R}$,
$\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\mathrm{tr}\Big{(}\mathrm{d}\mu_{t}(x,\xi)\Big{)}$ $=\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\mathrm{tr}\Big{(}\mathrm{d}\mu_{0}(x,\xi)\Big{)},$
$\int_{\mathbb{R}^{2}}\mathrm{Tr}\Big{(}\Gamma^{t}(s,\sigma)\Big{)}\mathrm{d}\gamma^{t}(s,\sigma)$ $=\int_{\mathbb{R}^{2}}\mathrm{Tr}\Big{(}\Gamma^{0}(s,\sigma)\Big{)}\mathrm{d}\gamma^{0}(s,\sigma).$
By taking $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}\in\mathrm{L}^{2}(\mathbb{R}^{2},\mathbb{C}^{2})$ normalized such that
$\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\mathrm{tr}\Big{(}\mathrm{d}\mu_{0}(x,\xi)\Big{)}+\int_{\mathbb{R}^{2}}\mathrm{Tr}\Big{(}\Gamma^{0}(s,\sigma)\Big{)}\mathrm{d}\gamma^{0}(s,\sigma)=1=\left\lVert\psi^{\varepsilon}_{0}\right\rVert_{2}^{2},$
we have the following equality because of the preservation of the $\mathrm{L}^{2}$-norm
$\left\lVert\psi^{\varepsilon}_{t}\right\rVert_{2}^{2}=$ $\ \int_{\mathbb{R}^{4}\backslash\mathcal{C}}\mathrm{tr}\Big{(}\mathrm{d}\mu_{t}(x,\xi)\Big{)}+\int_{\mathbb{R}\times\mathbb{S}^{2}}\mathrm{tr}\Big{(}M^{t}_{\infty}(s,\omega)\Big{)}\mathrm{d}\nu^{t}_{\infty}(s,\omega)$
$+\int_{\mathbb{R}^{2}}\mathrm{Tr}\Big{(}\Gamma^{t}(s,\sigma)\Big{)}\mathrm{d}\gamma^{t}(s,\sigma),$
$=$ $\ 1$
Therefore, for almost all $t$,
$M^{t}_{\infty}(s,\omega)\mathrm{d}\nu^{t}_{\infty}(s,\omega)=0.$
The evolution of the two-scale semiclassical measure in normal geodesic coordinates allows us to describe the general evolution of any initial normalized condition for equation (1).
Theorem 3.10 (Evolution of the semiclassical measure of solution to equation (1)).
Assume Assumption 1.2 and 1.3. Let $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ uniformly bounded in $\mathrm{L}^{\infty}\left(\mathbb{R},\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)$ solves (1) with initial normalized condition $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$. Then, there exist ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ a vanishing sequence of positive numbers and
  • $\bullet$
    a semiclassical measure $\mathrm{d}\mu_{t}$, evolving as described in Lemma 2.3,
  • a family of non negative scalar Radon measure $\left(\mathrm{d}\gamma_{n}^{t}\right)_{n\in\mathbb{Z}}$ satisfying (6),
  • $\bullet$
    a two-scale semiclassical measure ${M^{t}_{\infty}\mathrm{d}\nu^{t}_{\infty}}$ at infinity described in Theorem 3.8,
such that for all $(a,\Xi)\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)\times\mathscr{C}_{c}^{\infty}(\mathbb{R},\mathbb{C})$,
$\mathcal{I}^{\varepsilon_{k}}_{\psi}(\Xi,a)\underset{k\to+\infty}{\longrightarrow}\ \int_{\mathbb{R}}\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\Xi(t)\mathrm{tr}\Big{(}a(x,\xi)\mathrm{d}\mu_{t}(x,\xi)\Big{)}\mathrm{d}t+\int_{\mathbb{R}}\int_{\mathbb{R}}\Xi(t)\mathrm{tr}\Big{(}a\big{(}\gamma(s),0\big{)}\mathrm{d}\rho_{t}(s)\Big{)}\mathrm{d}t,$
with for all $s\in\mathbb{R}$,
$\mathrm{d}\rho_{t}(s)=\int_{\mathbb{S}^{2}}\mathrm{U}_{\theta(s)}^{\dagger}M^{t}_{\infty}(s,\omega)\mathrm{U}_{\theta(s)}\mathrm{d}\nu^{t}_{\infty}(s,\omega)+\frac{1}{2}\sum_{n\in\mathbb{Z}}\int_{\mathbb{R}}\begin{pmatrix}1&\frac{\mathrm{i}\sigma}{\lambda_{n}(s,\sigma)}e^{-\mathrm{i}\theta(s)}\\ \frac{-\mathrm{i}\sigma}{\lambda_{n}(s,\sigma)}e^{\mathrm{i}\theta(s)}&1\end{pmatrix}\mathrm{d}\gamma^{t}_{n}(s,\mathrm{d}\sigma),$
where $\mathrm{U}_{\theta(s)}$ is defined in (10).
The proof of this theorem is in Section 4.3.
4 Proofs of the main result
4.1 Change of variable on semiclassical measure
First of all, we will see the consequences of the change of variable from Proposition 3.1 on the semiclassical measure of $\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}$ solution to equation (1) and the link with the two-scale semiclassical measure of the family $\left(\varphi_{t}^{\varepsilon}\right)_{\varepsilon>0}$ defined in (11).
Lemma 4.1 (Semiclassical measure and normal geodesic coordinates).
All semiclassical measure $\mathrm{d}\mu_{t}\mathrm{d}t$ of $(\psi^{\varepsilon}_{t})_{\varepsilon>0}$ solution to equation (1) associated to $(\varepsilon_{k})_{k\in\mathbb{N}}$ satisfies
$\mathrm{d}\mu_{t}\mathrm{d}t\mathds{1}_{\mathcal{C}}=\Phi^{*}\Big{(}\mathrm{U}_{\theta(s)}^{\dagger}\mathrm{d}\rho_{t}\mathrm{U}_{\theta(s)}\Big{)}\mathrm{d}t\mathds{1}_{\mathcal{C}}$
where $\mathrm{U}_{\theta(s)}$ is defined in (10) and $\mathrm{d}\rho_{t}$ is a semiclassical measure on $\mathbb{R}_{s}\times\mathrm{I}_{1/2}\times\mathbb{R}_{\sigma,\eta}^{2}\times\mathbb{R}_{t}$ associated to $\left(\varphi_{t}^{\varepsilon}\right)_{\varepsilon>0}$ and $\left(\varepsilon_{k}\right)_{k\in\mathbb{N}}$.
Proof of Lemma 4.1.
Let $a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)$ such that $\mathrm{supp}(a)\subset\Omega_{1/2}\times\mathbb{R}^{2}$ and $\chi$ the cut-off function taken in (11), then for all $N\in\mathbb{N}$,
$\left\langle\mathrm{Op}_{\varepsilon}\left(a\right)\psi^{\varepsilon}_{t},\psi^{\varepsilon}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}$ $=\left\langle\mathrm{Op}_{\varepsilon}\left(a\right)\chi\psi^{\varepsilon}_{t},\chi\psi^{\varepsilon}_{t}\right\rangle_{\mathrm{L}^{2}\left(\Omega_{1/2},\mathbb{C}^{2}\right)}+\mathcal{O}\left(\varepsilon^{N}\right)$
$=\left\langle\mathrm{U}\mathrm{Op}_{\varepsilon}\left(a\right)\mathrm{U}^{\dagger}\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}\times\mathrm{I}_{1/2},\mathbb{C}^{2}\right)}+\mathcal{O}\left(\varepsilon^{N}\right).$
Since $\rm U=U_{\theta(s)}V_{\Phi}$, $\rm U_{\theta(s)}$ being the multiplication by a function depending only of the variable $s$ and $\rm V_{\Phi}$ a change of variables, standard results on semiclassical pseudodifferential calculus ([27, Theorem 9.3]) imply
$\left\langle\mathrm{Op}_{\varepsilon}\left(a\right)\psi^{\varepsilon}_{t},\psi^{\varepsilon}_{t}\right\rangle=$ $\ \left\langle\mathrm{U}_{\theta(s)}\mathrm{Op}_{\varepsilon}\Big{(}\mathrm{V}_{\Phi}a\mathrm{V}_{\Phi}^{\dagger}\Big{)}\mathrm{U}_{\theta(s)}^{\dagger}\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}\times\mathrm{I}_{1/2},\mathbb{C}^{2}\right)}+\mathcal{O}\left(\varepsilon\right)$
$=$ $\ \left\langle\mathrm{Op}_{\varepsilon}\left(\mathrm{U}_{\theta(s)}a\mathrm{U}_{\theta(s)}^{\dagger}\left(\Phi(s,y),{}^{t}\nabla\left(\Phi^{-1}\right)\Big{(}\Phi(s,y)\Big{)}\left[\sigma,\eta\right]\right)\right)\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}\times\mathrm{I}_{1/2},\mathbb{C}^{2}\right)}$
$+\mathcal{O}\left(\varepsilon\right).$
For simplicity, we assume $\left(\varepsilon\right)_{\varepsilon>0}$ realizes the semiclassical measure $\mathrm{d}\mu_{t}\mathrm{d}t$ of $(\psi^{\varepsilon}_{t})_{\varepsilon>0}$. Passing to the limit, we obtain
$\left\langle\mathrm{Op}_{\varepsilon}\left(a\right)\psi^{\varepsilon}_{t},\psi^{\varepsilon}_{t}\right\rangle$ $\underset{\varepsilon\to 0}{\longrightarrow}\int_{\mathbb{R}\times\mathrm{I}_{1/2}\times\mathbb{R}^{2}}\mathrm{Tr}\left(\mathrm{U}_{\theta(s)}a\mathrm{U}_{\theta(s)}^{\dagger}\left(\Phi(s,y),{}^{t}\nabla\left(\Phi^{-1}\right)\Big{(}\Phi(s,y)\Big{)}\left[\sigma,\eta\right]\right)\mathrm{d}\rho_{t}(s,y,\sigma,\eta)\right)$
=×$=\int_{\mathbb{R}\times\mathrm{I}_{1/2}\times\mathbb{R}^{2}}\mathrm{Tr}\left(a\left(\Phi(s,y),{}^{t}\nabla\left(\Phi^{-1}\right)\Big{(}\Phi(s,y)\Big{)}\left[\sigma,\eta\right]\right)\Big{(}\mathrm{U}_{\theta(s)}^{\dagger}\mathrm{d}\rho_{t}(s,y,\sigma,\eta)\mathrm{U}_{\theta(s)}\Big{)}\right)$
$=\int_{\Omega_{1/2}\times\mathbb{R}^{2}}\mathrm{Tr}\left(a(x,\xi)\Phi^{*}\Big{(}\mathrm{U}_{\theta(s)}^{\dagger}\mathrm{d}\rho_{t}\mathrm{U}_{\theta(s)}\Big{)}(x,\xi)\right).$
We conclude by taking the restriction of the semiclassical measure $\mathrm{d}\rho_{t}$ to the curve $\mathcal{C}$. ∎
According to the definition (15) of the rescaling operator $\mathfrak{T}_{\sqrt{\varepsilon}}$, we define the family $\left(u^{\varepsilon}_{t}\right)_{\varepsilon>0}$ by
$\mathfrak{T}_{\sqrt{\varepsilon}}^{\dagger}\left[u^{\varepsilon}_{t}\right](s,y)=\varphi_{t}^{\varepsilon}(s,y)\quad,\quad\forall(t,s,y)\in\mathbb{R}\times\mathbb{R}\times\mathrm{I}_{1/2}.$
The family $\left(u^{\varepsilon}_{t}\right)_{\varepsilon>0}$ is defined on ×$\mathbb{R}\times\mathrm{I}_{\frac{1}{2\sqrt{\varepsilon}}}$ but we will study it on ×$\mathbb{R}\times\mathrm{I}_{\sqrt{\varepsilon}}^{\delta}$ with
$\mathrm{I}_{\sqrt{\varepsilon}}^{\delta}\coloneqq\left(-\frac{\delta}{\sqrt{\varepsilon}},\frac{\delta}{\sqrt{\varepsilon}}\right)$
where we consider $\delta\leqslant\frac{1}{4\left|\left|\kappa\right|\right|_{\infty}}$.
Lemma 4.2 (Double scaling operator).
The family $\left(u^{\varepsilon}_{t}\right)_{\varepsilon>0}$ satisfies on ×$\mathbb{R}\times\mathrm{I}_{\sqrt{\varepsilon}}^{\delta}$,
$\begin{split}\left(\sqrt{\varepsilon}\mathrm{D}_{t}+y\left|\nabla\mathrm{m}(\gamma(s))\right|\sigma_{1}-\mathrm{D}_{y}\sigma_{2}+\frac{\sqrt{\varepsilon}}{1-\sqrt{\varepsilon}y\kappa(s)}\mathrm{D}_{s}\sigma_{3}-\varepsilon y\frac{\mathrm{i}\kappa^{\prime}(s)}{2(1-\sqrt{\varepsilon}y\kappa(s))^{2}}\sigma_{3}\right)u_{t}^{\varepsilon}=-\sqrt{\varepsilon}\mathfrak{R}^{\varepsilon}_{2}u_{t}^{\varepsilon},\end{split}$ (33)
where for all $(s,y)\in\mathbb{R}\times\mathrm{I}_{\sqrt{\varepsilon}}^{\delta},$
$\begin{split}\big{(}\mathfrak{R}_{2}^{\varepsilon}u_{t}^{\varepsilon}\big{)}(s,y)\coloneqq&y^{2}\int_{0}^{1}\nabla^{2}\mathrm{m}\Big{(}\Phi(s,\tau\sqrt{\varepsilon}y)\Big{)}\cdot\Big{(}\nu(s),\nu(s)\Big{)}\mathrm{d}\tau\sigma_{1}u_{t}^{\varepsilon}(s,y).\end{split}$
Proof of Lemma 4.2.
On the one hand, we rewrite the equation (12) with the definition of $\left(u^{\varepsilon}_{t}\right)_{\varepsilon>0}$ on ×$\mathbb{R}\times\mathrm{I}_{\sqrt{\varepsilon}}^{\delta}$,
$\left[\varepsilon\mathrm{D}_{t}+\mathrm{m}\left(\Phi(s,\sqrt{\varepsilon}y)\right)\sigma_{1}-\sqrt{\varepsilon}\mathrm{D}_{y}\sigma_{2}+\frac{\varepsilon}{1-\sqrt{\varepsilon}y\kappa(s)}\mathrm{D}_{s}\sigma_{3}-\frac{\mathrm{i}\varepsilon^{3/2}y\kappa^{\prime}(s)}{2(1-\sqrt{\varepsilon}y\kappa(s))^{2}}\sigma_{3}\right]u^{\varepsilon}_{t}=0.$
On the other hand, for all $(s,y)\in\mathbb{R}\times\mathrm{I}_{\sqrt{\varepsilon}}^{\delta}$,
$\mathrm{m}\left(\Phi(s,\sqrt{\varepsilon}y)\right)=$ $\ \underbrace{\mathrm{m}(\gamma(s))}_{=0}+\sqrt{\varepsilon}y\underbrace{\nabla\mathrm{m}(\gamma(s))\cdot\nu(s)}_{=|\nabla\mathrm{m}(\gamma(s))|}$
$+\ \varepsilon y^{2}\int_{0}^{1}(1-\tau)\nabla^{2}\mathrm{m}\Big{(}\Phi(s,\tau\sqrt{\varepsilon}y)\Big{)}\cdot\Big{(}\nu(s),\nu(s)\Big{)}\mathrm{d}\tau.$
The operator $\sqrt{\varepsilon}\mathfrak{R}^{\varepsilon}_{2}$ will play a role of rest in the determination of $\Gamma\mathrm{d}\gamma$, the two-scale semiclassical measure at finite distance.
For now, we do some remarks about the two-scale operator. Indeed, we remark that $\sqrt{\varepsilon}$ is in front of $\mathrm{D}_{s}$ but there is no $\sqrt{\varepsilon}$ in front of $\mathrm{D}_{y}$. It fits with the two-scale operators that we have introduced in Section 3.2. It is a semiclassical operator for the scale $\sqrt{\varepsilon}$ in the variable $s$, without scale in $y$. Because of the second scale of the problem, we will consider the symbol of the principal term of equation (33) according only to the first variable $s$, keeping the operator character over the variable $y$. Let us define, for all $(s,\sigma)\in\mathbb{R}^{2}$,
$\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)\coloneqq\begin{pmatrix}\sigma&y|\nabla\mathrm{m}(\gamma(s))|+\mathrm{i}\mathrm{D}_{y}\\ y|\nabla\mathrm{m}(\gamma(s))|-\mathrm{i}\mathrm{D}_{y}&-\sigma\end{pmatrix},$ (34)
the symbol of an operator acting on $\mathrm{L}^{2}\left(\mathbb{R}_{s},\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)\right)$ as in the end of the Section 3.2.1 in (22). From a spectral analysis (Appendix C), the eigenvalues of the operator $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$ were defined in (5) with the associated normalized eigenvectors (28) and the associated eigenprojectors (29).
We also define the symbol of $\mathfrak{R}^{\varepsilon}_{2}$ for the two-scale quantization (22) as the operator acting on $\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)$ such that for all $(s,\sigma)\in\mathbb{R}^{2}$,
$\mathfrak{r}^{\varepsilon}_{2}(s,\sigma)\coloneqq y^{2}\int_{0}^{1}\nabla^{2}\mathrm{m}\Big{(}\Phi(s,\tau\sqrt{\varepsilon}y)\Big{)}\cdot\Big{(}\nu(s),\nu(s)\Big{)}\mathrm{d}\tau\sigma_{1}.$
Moreover, we define the operator acting on $\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)$ such that for all $(s,\sigma)\in\mathbb{R}^{2}$,
$\mathfrak{r}^{\varepsilon}_{3}(s,\sigma)\coloneqq\frac{y\kappa(s)\sigma}{1-\sqrt{\varepsilon}y\kappa(s)}\sigma_{3}.$
We define $\mathfrak{r}^{\varepsilon}\coloneqq\mathfrak{r}^{\varepsilon}_{2}+\mathfrak{r}^{\varepsilon}_{3}$. With the notation (22), the equation (33) writes
$\sqrt{\varepsilon}\mathrm{D}_{t}u^{\varepsilon}_{t}+\mathrm{Op}_{\sqrt{\varepsilon},1}\Big{(}\mathrm{T}_{\mathrm{E}}^{\rm ds}+\sqrt{\varepsilon}\mathfrak{r}^{\varepsilon}\Big{)}u^{\varepsilon}_{t}=0.$
4.2 The two-scale semiclassical measures in normal geodesic coordinates
In this section we will describe the evolution of the two-scale semiclassical measures defined in Theorem 3.7 for $\left(\varphi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ solution to equation (12). In fact, in Section 4.2.1, we describe the evolution of the measure at finite distance $\Gamma^{t}\mathrm{d}\gamma^{t}$ and in Section 4.2.2, we describe the evolution of the measure at infinity $M_{\infty}^{t}\mathrm{d}\nu_{\infty}^{t}$. In Section 4.3, we prove Theorems 3.8 and 3.10. In Section 4.4, we prove Theorem 1.4.
4.2.1 Two-scale Wigner measure at finite distance
We prove here the part one of Theorem 3.8. We recall the following lemma of [16].
Lemma 4.3 (Poisson bracket and projector properties).
Let $\left(\Pi_{n}\right)_{n\in\mathbb{N}}$ be a family of orthogonal projectors such that for all $(n,k)\in\mathbb{N}^{2},\Pi_{n}\Pi_{k}=\delta_{k,n}$. Then for all $(j,k,\ell)\in\mathbb{N}^{3}$
$\Pi_{k}\Big{(}\left\{\Pi_{j},\Pi_{\ell}\right\}-\left\{\Pi_{\ell},\Pi_{j}\right\}\Big{)}\Pi_{k}=0$
Proof of part one of Theorem 3.8.
Assume Assumption 1.2 and 1.3. For convenience, we assume that the sequence $\left(\varepsilon\right)_{\varepsilon>0}$ realizes the measure $\Gamma^{t}\mathrm{d}\gamma^{t}$. Let $a\in\mathfrak{A}$ and $R,\delta$ as defined in Section 3.2. We consider the part (19) of $a$ denoted $a_{\varepsilon,R}$. Let us consider $b(s,\sigma)\coloneqq a(s,0,0,0,y,\sigma,\mathrm{D}_{y})\chi\left(\frac{|(y,\sigma,\mathrm{D}_{y})|}{R}\right)$, the operator acting on $\mathcal{H}\coloneqq\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)$ and obtained by Weyl quantization. The link between the second scale quantization of $a_{\varepsilon,R}$ (13) and the two-scale quantization of $b$ (22) was established in (23). By considering $\left(u^{\varepsilon}_{t}\right)_{\varepsilon>0}$ solution to equation (33) as a family of $\mathrm{L}^{2}\left(\mathbb{R}_{s},\mathcal{H}\right)$, we have
$\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b)u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle=$ $\ \frac{1}{\mathrm{i}\sqrt{\varepsilon}}\left\langle\Big{[}\mathrm{Op}_{\sqrt{\varepsilon},1}(b),\mathrm{Op}_{\sqrt{\varepsilon},1}\Big{(}\mathrm{T}_{\mathrm{E}}^{\rm ds}\Big{)}\Big{]}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle$
$+\frac{1}{\mathrm{i}}\left\langle\Big{[}\mathrm{Op}_{\sqrt{\varepsilon},1}(b),\mathrm{Op}_{\sqrt{\varepsilon},1}\big{(}\mathfrak{r}^{\varepsilon}\big{)}\Big{]}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle$
Because of (9), we have
$\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b)u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle=$ $\ \frac{1}{\mathrm{i}\sqrt{\varepsilon}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big{(}\left[b,\mathrm{T}_{\mathrm{E}}^{\rm ds}\right]\Big{)}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle$
$+\frac{1}{\mathrm{i}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\left(\frac{1}{2\mathrm{i}}\left\{b,\mathrm{T}_{\mathrm{E}}^{\rm ds}\right\}-\frac{1}{2\mathrm{i}}\left\{\mathrm{T}_{\mathrm{E}}^{\rm ds},b\right\}+\left[b,\mathfrak{r}^{\varepsilon}\right]\right)u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle$
$+\mathcal{O}_{R}\left(\sqrt{\varepsilon}\right).$
  • $\bullet$
    Step 1 : Structure of $\Gamma^{t}$
Let $\Xi\in\mathscr{C}^{\infty}_{c}(\mathbb{R})$, then
$\int_{\mathbb{R}}\Xi(t)\mathrm{i}\sqrt{\varepsilon}\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b)u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\mathrm{d}t=$ $\int_{\mathbb{R}}\Xi(t)\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big{(}\left[b,\mathrm{T}_{\mathrm{E}}^{\rm ds}\right]\Big{)}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\mathrm{d}t$
$+\frac{\sqrt{\varepsilon}}{2\mathrm{i}}\int_{\mathbb{R}}\Xi(t)\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big{(}\left\{b,\mathrm{T}_{\mathrm{E}}^{\rm ds}\right\}-\left\{\mathrm{T}_{\mathrm{E}}^{\rm ds},b\right\}\Big{)}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\mathrm{d}t$
$+\sqrt{\varepsilon}\int_{\mathbb{R}}\Xi(t)\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big{(}\left[b,\mathfrak{r}^{\varepsilon}\right]\Big{)}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\mathrm{d}t+\mathcal{O}_{R}\left(\varepsilon\right)$
$=$ $-\mathrm{i}\sqrt{\varepsilon}\int_{\mathbb{R}}\Xi^{\prime}(t)\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b)u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\mathrm{d}t+\mathcal{O}_{R}\left(\varepsilon\right).$
But $b$ is compactly supported in $\left\{\left|(y,\sigma)\right|\leqslant R\right\}$ and for all $(s,\sigma,y)\in\mathbb{R}\times\left(-R,R\right)^{2}$,
$y\underbrace{\frac{\kappa(s)}{1-\sqrt{\varepsilon}y\kappa(s)}}_{\in\mathrm{L}^{\infty}\big{(}\mathbb{R}\times\left(-R,R\right)\big{)}}\sigma=\mathcal{O}\left(R^{2}\right),\quad\quad y^{2}\underbrace{\int_{0}^{1}\nabla^{2}\mathrm{m}\Big{(}\Phi(s,\tau\sqrt{\varepsilon}y)\Big{)}\cdot\Big{(}\nu(s),\nu(s)\Big{)}\mathrm{d}\tau}_{\in\mathrm{L}^{\infty}\big{(}\mathbb{R}\times\left(-R,R\right)\big{)}}=\mathcal{O}\left(R^{2}\right).$
Then
$\sqrt{\varepsilon}\int_{\mathbb{R}}\Xi(t)\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big{(}[b,\mathfrak{r}^{\varepsilon}]\Big{)}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\mathrm{d}t=\mathcal{O}_{R}\left(\sqrt{\varepsilon}\right).$
By passing to the limit, we have, for all $\Xi\in\mathscr{C}^{\infty}_{c}(\mathbb{R})$ and for all $a\in\mathfrak{A}$,
00 0 $=\int_{\mathbb{R}^{3}}\Xi(t)\mathrm{Tr}_{\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)}\Bigg{(}\Big{[}a(s,0,0,y,\sigma,\mathrm{D}_{y}),\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)\Big{]}\Gamma^{t}(s,\sigma)\Bigg{)}\mathrm{d}\gamma^{t}(s,\sigma)\mathrm{d}t$
$=\int_{\mathbb{R}^{3}}\Xi(t)\mathrm{Tr}\Bigg{(}a(s,0,0,y,\sigma,\mathrm{D}_{y})\Big{[}\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma),\Gamma^{t}(s,\sigma)\Big{]}\Bigg{)}\mathrm{d}\gamma^{t}(s,\sigma)\mathrm{d}t.$
So, $t$ a.e. $\mathrm{d}\gamma^{t}$ a.e. $\Big{[}\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma),\Gamma^{t}(s,\sigma)\Big{]}=0$, which implies (31).
  • $\bullet$
    Step 2 : Equation of $\left(\mathrm{d}\gamma_{n}^{t}\right)_{n\in\mathbb{Z}}$
By Step $1$, for all $(k,\ell)\in\mathbb{Z}^{2}$, with $k\neq\ell$ and for all $\beta$ function from $\mathbb{R}_{s}\times\mathbb{R}_{\sigma}$ to $\mathcal{L}\left(\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)\right)$ such that $\Pi_{k}\beta\Pi_{\ell}\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathcal{L}\left(\mathcal{H}\right)\right)$, we have
$\limsup_{\delta\to 0}\limsup_{R\to+\infty}\lim_{\varepsilon\to 0}\int_{\mathbb{R}}\Xi(t)\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big{(}\Pi_{k}\beta\Pi_{\ell}\Big{)}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\mathrm{d}t=0.$
Then we can consider symbol of the form $\Pi_{\ell}\beta\Pi_{\ell}$. Since the projector $\Pi_{\ell}$ has rank one, we have
$\Pi_{\ell}\beta\Pi_{\ell}=\mathrm{Tr}\Big{(}\Pi_{\ell}\beta\Pi_{\ell}\Big{)}\Pi_{\ell},$
where $\mathrm{Tr}_{\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)}\big{(}\Pi_{\ell}\beta\Pi_{\ell}\big{)}\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathbb{C}\right)$. Therefore we will consider symbols of the form $\beta\Pi_{\ell}$ with $\beta\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathbb{C}\right)$. Let $n\in\mathbb{Z}$ and $\beta\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathbb{C}\right)$. Let us define for all $(s,\sigma)\in\mathbb{R}^{2}$,
$b_{n}(s,\sigma)\coloneqq\beta(s,\sigma)\Pi_{n}(s,\sigma).$
The form of $b_{n}$ implies $[b_{n},\mathrm{T}_{\mathrm{E}}^{\rm ds}]=0$. Then
$\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b_{n})u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle=$ $\ \frac{1}{2}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big{(}\{\mathrm{T}_{\mathrm{E}}^{\rm ds},b_{n}\}-\{b_{n},\mathrm{T}_{\mathrm{E}}^{\rm ds}\}\Big{)}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle$
$+\frac{1}{\mathrm{i}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big{(}[b_{n},\mathfrak{r}^{\varepsilon}]\Big{)}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle+\mathcal{O}_{R}\left(\sqrt{\varepsilon}\right).$
We first show that the term with $\mathfrak{r}^{\varepsilon}$ will vanish as $\varepsilon$ goes to 00. By definition of $\Gamma^{t}\mathrm{d}\gamma^{t}$, we have
$\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\big{(}[b_{n},\mathfrak{r}^{\varepsilon}]\big{)}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle$ $\underset{\varepsilon\to 0}{\longrightarrow}\int_{\mathbb{R}^{2}}\mathrm{Tr}\Bigg{(}\Big{[}a(s,0,0,y,\sigma,D_{y}),\mathfrak{r}^{0}(s,\sigma)\Big{]}\Gamma^{t}(s,\sigma)\Bigg{)}\mathrm{d}\gamma^{t}(s,\sigma)$
$=\sum_{m\in\mathbb{Z}}\int_{\mathbb{R}^{2}}\mathrm{Tr}\Bigg{(}\Big{[}\beta(s,\sigma)\Pi_{n}(s,\sigma),\mathfrak{r}^{0}(s,\sigma)\Big{]}\Pi_{m}(s,\sigma)\Bigg{)}\mathrm{d}\gamma^{t}_{m}(s,\sigma),$
$=\sum_{m\in\mathbb{Z}}\int_{\mathbb{R}^{2}}\beta(s,\sigma)\mathrm{Tr}\Bigg{(}\Big{[}\Pi_{n}(s,\sigma),\mathfrak{r}^{0}(s,\sigma)\Big{]}\Pi_{m}(s,\sigma)\Bigg{)}\mathrm{d}\gamma^{t}_{m}(s,\sigma),$
where the second equality holds because of Step $1$ and the chosen shape of $a$ and the last equality holds because $\beta$ is scalar and $\mathfrak{r}^{0}$ is the value at zero of $\mathfrak{r}^{\varepsilon}$ given by for all $(s,\sigma)\in\mathbb{R}^{2}$,
$\mathfrak{r}^{0}(s,\sigma)=\mathfrak{r}^{0}_{2}(s,\sigma)+\mathfrak{r}^{0}_{3}(s,\sigma)=y^{2}\nabla^{2}\mathrm{m}\Big{(}\Phi(s,0)\Big{)}\cdot\Big{(}\nu(s),\nu(s)\Big{)}\sigma_{1}+y\kappa(s)\sigma\sigma_{3}.$
Then, we recall, for all $m\in\mathbb{Z}$,
$\mathrm{Tr}\Big{(}\big{[}\Pi_{n},\mathfrak{r}^{0}\big{]}\Pi_{m}\Big{)}=\mathrm{Tr}\Big{(}\Pi_{n}\mathfrak{r}^{0}\Pi_{m}\Big{)}-\mathrm{Tr}\Big{(}\mathfrak{r}^{0}\Pi_{n}\Pi_{m}\Big{)}=\mathrm{Tr}\Big{(}\mathfrak{r}^{0}\Pi_{m}\Pi_{n}\Big{)}-\mathrm{Tr}\Big{(}\mathfrak{r}^{0}\Pi_{n}\Pi_{m}\Big{)}\\ =\delta_{n=m}\left(\mathrm{Tr}\Big{(}\mathfrak{r}^{0}\Pi_{n}\Big{)}-\mathrm{Tr}\Big{(}\mathfrak{r}^{0}\Pi_{n}\Big{)}\right)=0.\$
Let $k\in\mathbb{Z}$, let us compute $\Pi_{k}\{\mathrm{T}_{\mathrm{E}}^{\rm ds},b_{n}\}\Pi_{k}$,
$\Pi_{k}\{\mathrm{T}_{\mathrm{E}}^{\rm ds},b_{n}\}\Pi_{k}=$ $\ \Pi_{k}\left\{\sum_{m\in\mathbb{Z}}\lambda_{m}\Pi_{m}\,,\,\beta\Pi_{n}\right\}\Pi_{k}$ (35)
$=$ $\ \Pi_{k}\sum_{m\in\mathbb{Z}}\Bigg{(}\Pi_{m}\left\{\lambda_{m},\beta\right\}\Pi_{n}+\lambda_{m}\left\{\Pi_{m},\beta\right\}\Pi_{n}+\lambda_{m}\beta\left\{\Pi_{m},\Pi_{n}\right\}\Bigg{)}\Pi_{k}$
$\ +\Pi_{k}\sum_{m\in\mathbb{Z}}\Big{(}\beta\Pi_{m}\left\{\lambda_{m},\Pi_{n}\right\}\Big{)}\Pi_{k}.$
Let $m\in\mathbb{Z}$,
$\Pi_{k}\Pi_{m}\left\{\lambda_{m},\beta\right\}\Pi_{n}\Pi_{k}$ $=\mathds{1}_{k=m=n}\left\{\lambda_{n},\beta\right\}\Pi_{n},$
[…]
$\Pi_{k}\lambda_{m}\beta\left\{\Pi_{m},\Pi_{n}\right\}\Pi_{k}$ $=\lambda_{m}\beta\Pi_{k}\left\{\Pi_{m},\Pi_{n}\right\}\Pi_{k},$
[…]
where the two vanishing equations hold because $\lambda_{n}$ and $\beta$ are scalar and $\Pi_{k}\partial\Pi_{m}\Pi_{k}$ is always equal to zero. Moreover, according to Lemma 4.3,
$\lambda_{m}\beta\Pi_{k}\Big{(}\left\{\Pi_{m},\Pi_{n}\right\}-\left\{\Pi_{n},\Pi_{m}\right\}\Big{)}\Pi_{k}=0\,,\,\text{for all}\ m\in\mathbb{Z}.$
Therefore,
$\{\mathrm{T}_{\mathrm{E}}^{\rm ds},b_{n}\}-\{b_{n},\mathrm{T}_{\mathrm{E}}^{\rm ds}\}=\Big{(}\left\{\lambda_{n},\beta\right\}-\left\{\beta,\lambda_{n}\right\}\Big{)}\Pi_{n}.$
Then,
$\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b_{n})u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle=\frac{1}{2}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big{(}\Big{(}\left\{\lambda_{n},\beta\right\}-\left\{\beta,\lambda_{n}\right\}\Big{)}\Pi_{n}\Big{)}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\ +\ \mathcal{O}_{R}\left(\sqrt{\varepsilon}\right).$
By passing to the limit,
$\int_{\mathbb{R}}\int_{\mathbb{R}^{2}}\Xi^{\prime}(t)\mathrm{Tr}\Big{(}a^{W}(s,0,0,0,y,\sigma,\mathrm{D}_{y})\Gamma^{t}(s,\sigma)\Big{)}\mathrm{d}\gamma^{t}(s,\sigma)\mathrm{d}t\\ =\int_{\mathbb{R}}\int_{\mathbb{R}^{2}}\Xi(t)\mathrm{Tr}\Big{(}\left\{\lambda_{n},\beta\right\}\Pi_{n}(s,0,0,0,y,\sigma,\mathrm{D}_{y})\Gamma^{t}(s,\sigma)\Big{)}\mathrm{d}\gamma^{t}(s,\sigma)\mathrm{d}t\\ =\int_{\mathbb{R}}\int_{\mathbb{R}^{2}}\Xi(t)\mathrm{Tr}\Big{(}\left\{\lambda_{n},\beta\right\}\Pi_{n}(s,0,0,0,y,\sigma,\mathrm{D}_{y})\Big{)}\mathrm{d}\gamma^{t}_{n}(s,\sigma)\mathrm{d}t\\ =\int_{\mathbb{R}}\int_{\mathbb{R}^{2}}\Xi(t)\left\{\lambda_{n},\beta\right\}(s,\sigma)\mathrm{d}\gamma^{t}_{n}(s,\sigma)\mathrm{d}t.\\ =-\int_{\mathbb{R}}\int_{\mathbb{R}^{2}}\Xi(t)\beta(s,\sigma)\left\{\lambda_{n},\mathrm{d}\gamma^{t}_{n}\right\}(s,\sigma)\mathrm{d}t.\$
Then $\mathrm{d}\gamma_{n}^{t}$ satisfies the equation (6) and is notably continuous on $\mathbb{R}$. It remains to prove that the value at zero of $\mathrm{d}\gamma_{n}^{t}$ coincides with the measure of the initial condition $\mathrm{d}\gamma_{n,0}$. Let $T\in\mathbb{R}$, on the one hand,
$\int_{0}^{T}\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b_{n})u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\mathrm{d}t=\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b_{n})u^{\varepsilon}_{T},u^{\varepsilon}_{T}\right\rangle-\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b_{n})u^{\varepsilon}_{0},u^{\varepsilon}_{0}\right\rangle.$
On the other hand,
$\left|\int_{0}^{T}\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b_{n})u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\mathrm{d}t\right|$ $\leqslant\frac{1}{2}\int_{0}^{T}\left|\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big{(}\left\{\lambda_{n},\beta\right\}\Pi_{n}\Big{)}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\right|\mathrm{d}t\ +\ \mathcal{O}_{T,R}\left(\sqrt{\varepsilon}\right),$
$\leqslant\frac{C\left(\left\{\lambda_{n},\beta\right\}\Pi_{n}\right)T}{2}\left\lVert u^{\varepsilon}_{0}\right\rVert_{\mathrm{L}^{2}}^{2}\ +\ \mathcal{O}_{T,R}\left(\sqrt{\varepsilon}\right).$
where the last inequality holds because of the Calderón-Vaillancourt Theorem (7) and the preservation of the $\mathrm{L}^{2}$-norm of the solution. Then, by passing to the limit, we have
$\begin{split}\limsup_{T\to 0}\limsup_{\delta\to 0}\limsup_{R\to+\infty}\lim_{\varepsilon\to 0}\mathcal{I}^{\varepsilon}_{\varphi_{t}^{\varepsilon}}\left(a_{\varepsilon,R}\right)&=\int_{\mathbb{R}^{2}}\mathrm{Tr}\Big{(}a^{W}(s,0,0,0,y,\sigma,\mathrm{D}_{y})\Pi_{n}(s,\sigma)\Big{)}\mathrm{d}\gamma_{n}^{0}(s,\sigma),\\ &=\limsup_{T\to 0}\limsup_{\delta\to 0}\limsup_{R\to+\infty}\lim_{\varepsilon\to 0}\mathcal{I}^{\varepsilon}_{\varphi_{0}^{\varepsilon}}\left(a_{\varepsilon,R}\right).\end{split}$
Therefore, for all $n\in\mathbb{Z},\mathrm{d}\gamma_{n}^{t}$ is continuous and the value at zero coincides with the initial condition. Moreover, the equality $\Gamma^{t}\mathrm{d}\gamma^{t}=\sum_{n\in\mathbb{Z}}\Pi_{n}\mathrm{d}\gamma_{n}^{t}$ holds as trace class operator so, thanks to Lebesgue dominated convergence Theorem, $\Gamma^{t}\mathrm{d}\gamma^{t}$ is continuous on $\mathbb{R}$. ∎
Remark 4.4 (Zero value of the two-scale semiclassical measure at finite distance).
We can construct the initial data with inspiration of the proof of the Lemma B.1 where we construct semiclassical measures for families valued in a separable Hilbert space.
Let $\left(f^{\varepsilon}\right)_{\varepsilon>0}$ be a bounded family of $\mathrm{L}^{2}\left(\mathbb{R}_{s},\mathbb{C}\right)$ with the knowledge of $\mathrm{d}\rho$, the semiclassical measure of $\left(f^{\varepsilon}\right)_{\varepsilon>0}$ at the scale $\sqrt{\varepsilon}$. The function $(s,\sigma)\mapsto g_{1}^{s,\sigma}$ is smooth, bounded and with all its derivatives bounded uniformly in $y$, therefore we can consider its Weyl $\sqrt{\varepsilon}$-quantization according to the variable $s$. Let us consider initial data $\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}$ such that for all $(s,y)\in\mathbb{R}\times\mathrm{I}_{1/2}$,
$u_{0}^{\varepsilon}(s,y)=g_{0}^{s}(y)f(s)+\Big{(}\mathrm{Op}_{\sqrt{\varepsilon}}^{(s)}\big{(}g_{1}^{s,\sigma}(y)\big{)}f\Big{)}(s),$
where the Weyl $\sqrt{\varepsilon}$-quantization of $g_{1}^{s,\sigma}$ is taken only in the $s$ variable. Then the two-scale semiclassical measure at finite distance of this initial data is for all $(s,\sigma)\in\mathbb{R}^{2}$,
$\frac{1}{2\pi}\Big{(}\Pi_{0}(s)+\Pi_{1}(s,\sigma)+\Pi_{1,0}(s)+\Pi_{0,1}(s,\sigma)\Big{)}\mathrm{d}\rho(s,\sigma),$
where for all $(k,\ell)\in\left\{0,1\right\}^{2}$ and for all $(s,\sigma)\in\mathbb{R}^{2},\Pi_{k,\ell}(s,\sigma)=g^{s,\sigma}_{k}\otimes g^{s,\sigma}_{\ell}$. Notably, it does not commute with $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$.
4.2.2 Two-scale Wigner measure at infinity
We prove here the part two of Theorem 3.8. We consider the symbol of $\mathrm{H}^{\mathrm{E}}_{\varepsilon}$ according to the second scale quantization (13), given by for all $(s,y)\in\mathbb{R}^{2}$ and $z\coloneqq(z_{y},z_{\sigma},z_{\eta})\in\mathbb{R}^{3}$,
$\mathrm{T}^{\mathrm{E}}(s,y,z)\coloneqq z_{y}\int_{0}^{1}\nabla\mathrm{m}\Big{(}\Phi(s,\tau y)\Big{)}\cdot\nu(s)\mathrm{d}\tau\sigma_{1}-z_{\eta}\sigma_{2}+\frac{z_{\sigma}}{1-y\kappa(s)}\sigma_{3}$
which has two eigenvalues :
$\lambda_{\pm}^{\mathrm{E}}(s,y,z)\coloneqq\pm\sqrt{\frac{1}{(1-y\kappa(s))^{2}}z_{\sigma}^{2}+z_{y}^{2}\left|\int_{0}^{1}\nabla\mathrm{m}\Big{(}\Phi(s,\tau y)\Big{)}\cdot\nu(s)\mathrm{d}\tau\right|^{2}+z_{\eta}^{2}}$
which are homogeneous functions of order 1 in $z$. The eigenprojector $\Pi_{\pm}^{\mathrm{E}}(s,y,z)$ defined by
$\Pi_{\pm}^{\mathrm{E}}(s,y,z)\coloneqq\frac{1}{2}\mathrm{Id}+\frac{1}{2\lambda_{\pm}^{\mathrm{E}}(s,y,z)}\mathrm{T}^{\mathrm{E}}(s,y,z)$
are homogeneous functions of order 0 in $z$.
Moreover, we denote the principal symbol of the operator $\mathrm{H}_{\varepsilon}^{\mathrm{E}}$ associated to the second scale quantization (13) by for all $(s,y)\in\mathbb{R}^{2}$ and for all $z\coloneqq(z_{y},z_{\sigma},z_{\eta})\in\mathbb{R}^{3}$,
$\mathcal{T}^{\infty}_{\mathrm{E}}(s,z)\coloneqq\mathrm{T}^{\mathrm{E}}(s,0,z)=\begin{pmatrix}z_{\sigma}&z_{y}r(s)^{2}+iz_{\eta}\\ z_{y}r(s)^{2}-iz_{\eta}&-z_{\sigma}\end{pmatrix}.$
Its eigenvalues are
$\lambda_{\pm}^{\infty}(s,z)\coloneqq\pm\sqrt{z_{\sigma}^{2}+z_{y}^{2}r(s)^{4}+z_{\eta}^{2}},$
and the associated eigenprojectors $\Pi_{\pm}^{\infty}(s,z)$ were defined in (30).
Then, the equation (12) writes on $\mathbb{R}_{t}\times\mathbb{R}_{s}\times\mathrm{I}_{1/2}$
$\sqrt{\varepsilon}\mathrm{D}_{t}\varphi^{\varepsilon}_{t}+\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\Big{(}\mathrm{T}^{\mathrm{E}}\Big{)}\varphi^{\varepsilon}_{t}=0.$
Proof of part two of Theorem 3.8.
Assume Assumption 1.2 and 1.3. For convenience, we assume that the sequence $\left(\varepsilon\right)_{\varepsilon>0}$ realizes the measure $M^{t}_{\infty}\mathrm{d}\nu^{t}_{\infty}$. Let $a\in\mathfrak{A}$ and $R,\delta$ as defined in Section 3.2. We consider the part (18) of $a$ divided by $\lambda^{\infty}_{+}$. The symbol $\mathrm{T}^{\mathrm{E}}$ is homogeneous in $z$ of order $1$ and the symbol $a_{\varepsilon,\delta}^{R}/\lambda^{\infty}_{+}$ is homogeneous in $z$ of order $-1$. Because of (18), we consider symbols on a subset of $\left\{z\in\mathbb{R}^{3}|\frac{R}{2}\leqslant\left|z\right|\leqslant\frac{2\delta}{\sqrt{\varepsilon}}\right\}$. Moreover, we have the following properties.
  1. 1.
    $\left\{\mathrm{T}^{\mathrm{E}},\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right\}_{z_{y},z_{\eta}}$ is homogeneous in $z$ of order $-2$.
  2. 2.
    $\left\{\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathrm{T}^{\mathrm{E}}\right\}_{s,z_{\sigma}}$ is homogeneous in $z$ of order $-1$.
  3. 3.
    $\left\{\mathrm{T}^{\mathrm{E}},\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right\}_{(s,y),(\sigma,\eta)}$ is homogeneous in $z$ of order 00.
Then, according to standard semiclassical calculus,
$\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\mathrm{T}^{\mathrm{E}},\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right\}_{z_{y},z_{\eta}}-\left\{\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathrm{T}^{\mathrm{E}}\right\}_{z_{y},z_{\eta}}\right)$ $=\mathcal{O}\left(\frac{1}{R^{2}}\right),$
$\sqrt{\varepsilon}\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\mathrm{T}^{\mathrm{E}},\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right\}_{s,z_{\sigma}}-\left\{\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathrm{T}^{\mathrm{E}}\right\}_{s,z_{\sigma}}\right)$ $=\mathcal{O}_{R}\left(\sqrt{\varepsilon}\right),$
$\varepsilon\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\mathrm{T}^{\mathrm{E}},\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right\}_{(s,y),(\sigma,\eta)}-\left\{\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathrm{T}^{\mathrm{E}}\right\}_{(s,y),(\sigma,\eta)}\right)$ $=\mathcal{O}_{R}\left(\varepsilon\right).$
So
$\left[\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right),\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\mathrm{T}^{\mathrm{E}}\right)\right]=$ $\ \mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left[\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathrm{T}^{\mathrm{E}}\right]\right)$
$+\frac{1}{2\mathrm{i}}\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathrm{T}^{\mathrm{E}}\right\}_{z_{y},z_{\eta}}-\left\{\mathrm{T}^{\mathrm{E}},\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right\}_{z_{y},z_{\eta}}\right)+\mathcal{O}\left(\frac{1}{R^{4}}\right)$
$+\frac{\sqrt{\varepsilon}}{2\mathrm{i}}\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathrm{T}^{\mathrm{E}}\right\}_{s,z_{\sigma}}-\left\{\mathrm{T}^{\mathrm{E}},\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right\}_{s,z_{\sigma}}\right)+\mathcal{O}_{R}\left(\varepsilon\right).$
Considering $\left(\varphi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ solution to (12), we have
$\mathrm{i}\sqrt{\varepsilon}\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right)\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle=\left\langle\left[\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right),\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\mathrm{T}^{\mathrm{E}}\right)\right]\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle,\\ =\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left[\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathrm{T}^{\mathrm{E}}\right]\right)\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle+\frac{\sqrt{\varepsilon}}{2\mathrm{i}}\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathrm{T}^{\mathrm{E}}\right\}_{s,z_{\sigma}}-\left\{\mathrm{T}^{\mathrm{E}},\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right\}_{s,z_{\sigma}}\right)\\ +\mathcal{O}_{R}\left(\varepsilon\right)+\mathcal{O}\left(\frac{1}{R^{2}}\right).\$
  • $\bullet$
    Step 1 : Structure of $M_{\infty}^{t}$
Let $\Xi\in\mathscr{C}^{\infty}_{c}(\mathbb{R})$, then
$\int_{\mathbb{R}}\Xi(t)\mathrm{i}\sqrt{\varepsilon}\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right)\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle\mathrm{d}t=\int_{\mathbb{R}}\Xi(t)\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left[\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathrm{T}^{\mathrm{E}}\right]\right)\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle\mathrm{d}t\\ +\frac{\sqrt{\varepsilon}}{2\mathrm{i}}\int_{\mathbb{R}}\Xi(t)\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathrm{T}^{\mathrm{E}}\right\}_{s,z_{\sigma}}-\left\{\mathrm{T}^{\mathrm{E}},\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right\}_{s,z_{\sigma}}\right)\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle\mathrm{d}t+\ \mathcal{O}_{R}\left(\varepsilon\right)+\mathcal{O}\left(\frac{1}{R^{2}}\right)\\ =-\mathrm{i}\sqrt{\varepsilon}\int_{\mathbb{R}}\Xi^{\prime}(t)\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right)\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle\mathrm{d}t.\$
But $\left[\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathrm{T}^{\mathrm{E}}\right]=\left[a,\frac{\mathrm{T}^{\mathrm{E}}}{\lambda^{\infty}_{+}}\right]_{\varepsilon,\delta}^{R}$ and $\left[a,\frac{\mathrm{T}^{\mathrm{E}}}{\lambda^{\infty}_{+}}\right]\in\mathfrak{A}$, then by passing to the limit, we deduce that for all $\Xi\in\mathscr{C}^{\infty}_{c}(\mathbb{R})$ and for all $a\in\mathfrak{A}$,
00 0 $=\int_{\mathbb{R}^{2}\times\mathbb{S}^{2}}\Xi(t)\mathrm{tr}\Bigg{(}\left[a,\frac{\mathrm{T}^{\mathrm{E}}}{\lambda^{\infty}_{+}}\right]_{\infty}(s,0,0,0,\omega)M_{\infty}^{t}(s,\omega)\Bigg{)}\mathrm{d}\nu_{\infty}^{t}(s,\omega)\mathrm{d}t$
$=\int_{\mathbb{R}^{2}\times\mathbb{S}^{2}}\Xi(t)\mathrm{tr}\Bigg{(}\left[a_{\infty}(s,0,0,0,\omega),\frac{1}{\lambda^{\infty}_{+}(s,\omega)}\mathrm{T}^{\mathrm{E}}(s,0,\omega)\right]M_{\infty}^{t}(s,\omega)\Bigg{)}\mathrm{d}\nu_{\infty}^{t}(s,\omega)\mathrm{d}t$
$=\int_{\mathbb{R}^{2}\times\mathbb{S}^{2}}\Xi(t)\frac{1}{\lambda^{\infty}_{+}(s,\omega)}\mathrm{tr}\Bigg{(}a_{\infty}(s,0,0,0,\omega)\Big{[}\mathcal{T}^{\infty}_{\mathrm{E}}(s,\omega),M_{\infty}^{t}(s,\omega)\Big{]}\Bigg{)}\mathrm{d}\nu_{\infty}^{t}(s,\omega)\mathrm{d}t.$
So, for almost all $t\in\mathbb{R}$, we have $\mathrm{d}\nu_{\infty}^{t}\ \mathrm{a.e.}$ $\Big{[}\mathcal{T}^{\infty}_{\mathrm{E}}(s,\omega),M_{\infty}^{t}(s,\omega)\Big{]}=0$ which implies (32).
  • $\bullet$
    Step 2 : Equation of $\mathrm{d}\nu_{\infty}^{t}$.
Let $\beta\in\mathfrak{A}$ and $a$ defined by for all $(s,y,\sigma,\eta,z)\in\mathbb{R}^{7}$,
$a(s,y,\sigma,\eta,z)\coloneqq\lambda^{\infty}_{\pm}(s,z)\beta(s,y,\sigma,\eta,z)\Pi_{\pm}^{\mathrm{E}}(s,y,z).$
The form of $a$ implies $\left[a,\mathrm{T}^{\mathrm{E}}\right]=0$. Then
$\sqrt{\varepsilon}\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a_{\varepsilon,\delta}^{R}\right)\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle=$ $\ \frac{1}{2}\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\mathrm{T}^{\mathrm{E}},a_{\varepsilon,\delta}^{R}\right\}_{z_{y},z_{\eta}}-\left\{a_{\varepsilon,\delta}^{R},\mathrm{T}^{\mathrm{E}}\right\}_{z_{y},z_{\eta}}\right)\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle+\mathcal{O}\left(\frac{1}{R^{2}}\right)$
$+\frac{\sqrt{\varepsilon}}{2}\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\mathrm{T}^{\mathrm{E}},a_{\varepsilon,\delta}^{R}\right\}_{s,z_{\sigma}}-\left\{a_{\varepsilon,\delta}^{R},\mathrm{T}^{\mathrm{E}}\right\}_{s,z_{\sigma}}\right)\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle+\mathcal{O}_{R}\left(\varepsilon\right).$
Because of the homogeneity in $z$, we have $\left\{\mathrm{T}^{\mathrm{E}},a\right\}_{z_{y},z_{\eta}}\in\mathfrak{A}$. Moreover,
$\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\mathrm{T}^{\mathrm{E}},a_{\varepsilon,\delta}^{R}\right\}_{z_{y},z_{\eta}}\right)=$ $\ \mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left(\left\{\mathrm{T}^{\mathrm{E}},a\right\}_{z_{y},z_{\eta}}\right)_{\varepsilon,\delta}^{R}\right)+\frac{1}{R}\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a\chi_{\frac{\delta}{\sqrt{\varepsilon}}}\left\{\mathrm{T}^{\mathrm{E}},\chi^{R}\right\}_{z_{y},z_{\eta}}\right)$
$+\frac{\sqrt{\varepsilon}}{\delta}\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a\chi^{R}\left\{\mathrm{T}^{\mathrm{E}},\chi_{\frac{\delta}{\sqrt{\varepsilon}}}\right\}_{z_{y},z_{\eta}}\right).$
The function $\partial\chi^{R}$ is compactly supported in the ring $\left\{z\in\mathbb{R}^{3}|\frac{R}{2}\leqslant\left|z\right|\leqslant R\right\}$ and the function $a\chi_{\frac{\delta}{\sqrt{\varepsilon}}}\left\{\mathrm{T}^{\mathrm{E}},\chi^{R}\right\}_{z_{y},z_{\eta}}$ belongs to $\mathfrak{A}$ so
$\left(a\chi_{\frac{\delta}{\sqrt{\varepsilon}}}\left\{\mathrm{T}^{\mathrm{E}},\chi^{R}\right\}_{z_{y},z_{\eta}}\right)_{\infty}=0.$
The operator $\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a\chi^{R}\left\{\mathrm{T}^{\mathrm{E}},\chi_{\frac{\delta}{\sqrt{\varepsilon}}}\right\}_{z_{y},z_{\eta}}\right)$ is uniformly bounded in $\varepsilon$ then
$\lim_{\varepsilon\to 0}\frac{\sqrt{\varepsilon}}{\delta}\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a\chi^{R}\left\{\mathrm{T}^{\mathrm{E}},\chi_{\frac{\delta}{\sqrt{\varepsilon}}}\right\}_{z_{y},z_{\eta}}\right)\varphi_{t}^{\varepsilon},\varphi_{t}^{\varepsilon}\right\rangle=0.$
As in the computation of (35), we consider the quantity $\Pi_{+}^{\mathrm{E}}\left\{a,\mathrm{T}^{\mathrm{E}}\right\}_{z_{y},z_{\eta}}\Pi_{+}^{\mathrm{E}}$ then
$\Pi_{+}^{\mathrm{E}}\left\{a,\mathrm{T}^{\mathrm{E}}\right\}_{z_{y},z_{\eta}}\Pi_{+}^{\mathrm{E}}=\lambda_{+}^{\infty}\left\{\beta,\lambda_{+}^{\infty}\right\}_{z_{y},z_{\eta}}\Pi_{+}^{\mathrm{E}}.$
Then, by passing to the limit, we have
00 0 $=\int_{\mathbb{R}^{2}\times\mathbb{S}^{2}}\Xi(t)\mathrm{tr}\Bigg{(}\left(\lambda_{+}^{\infty}\left\{\beta,\lambda_{+}^{\infty}\right\}_{z_{y},z_{\eta}}\Pi_{+}^{\mathrm{E}}\right)_{\infty}(s,0,0,0,\omega)M_{\infty}^{t}(s,\omega)\Bigg{)}\mathrm{d}\nu_{\infty}^{t}(s,\omega)\mathrm{d}t,$
$=\int_{\mathbb{R}^{2}\times\mathbb{S}^{2}}\Xi(t)\mathrm{tr}\Bigg{(}\lambda_{+}^{\infty}\left\{\beta_{\infty},\lambda_{+}^{\infty}\right\}_{\omega_{y},\omega_{\eta}}(s,0,0,0,\omega)\Pi_{+}^{\infty}(s,\omega)\Bigg{)}\mathrm{d}\nu_{+}^{t}(s,\omega)\mathrm{d}t,$
$=\int_{\mathbb{R}^{2}\times\mathbb{S}^{2}}\Xi(t)\lambda_{+}^{\infty}\left\{\beta_{\infty},\lambda_{+}^{\infty}\right\}_{\omega_{y},\omega_{\eta}}(s,0,0,0,\omega)\mathrm{d}\nu_{+}^{t}(s,\omega)\mathrm{d}t,$
but for all $(s,\omega)\in\mathbb{R}\times\mathbb{S}^{2}$,
$\lambda_{+}^{\infty}(s,\omega)\frac{\partial\lambda_{+}^{\infty}}{\partial z_{y}}(s,\omega)=\omega_{y}r(s)^{4},\quad\quad\quad\lambda_{+}^{\infty}(s,\omega)\frac{\partial\lambda_{+}^{\infty}}{\partial z_{\eta}}(s,\omega)=\omega_{\eta}.$
Then
$\int_{\mathbb{R}^{2}\times\mathbb{S}^{2}}\Xi(t)\Big{(}\vec{V}_{+}(s,\omega)\cdot\nabla_{\omega}\beta_{\infty}(s,\omega)\Big{)}\mathrm{d}\nu_{+}^{t}(s,\omega)\mathrm{d}t=0,$
with
$\vec{V}_{+}(s,\omega)=\begin{pmatrix}-\omega_{\eta}\\ 0\\ \omega_{y}r(s)^{4}\end{pmatrix}$
and $\beta_{\infty}$ a function on the sphere with $\nabla_{\omega}\beta_{\infty}$ is a vector field over the sphere. Therefore we can rewrite $\vec{V}_{\pm}\cdot\nabla_{\omega}\beta=\vec{V}_{\pm}^{\infty}\cdot\nabla_{\omega}\beta$ where $V_{\pm}^{\infty}$ is the component of $\vec{V}_{\pm}$ which is tangential to the sphere. The vector $\vec{V}_{\pm}^{\infty}$ defines a vector field over the sphere and is given by for all $(s,\omega)\in\mathbb{R}\times\mathbb{S}^{2}$,
$\vec{V}^{\infty}(s,\omega)\coloneqq\vec{V}_{+}(s,\omega)-\Big{(}\vec{V}_{+}(s,\omega)\cdot\omega\Big{)}\omega=\begin{pmatrix}\left(\left(1-r(s)^{4}\right)\omega_{y}^{2}-1\right)\omega_{\eta}\\ \left(1-r(s)^{4}\right)\omega_{y}\omega_{\eta}\omega_{\sigma}\\ \left(\left(1-r(s)^{4}\right)\omega_{\eta}^{2}+r(s)^{4}\right)\omega_{y}\end{pmatrix}.$
Because of the equality $\lambda^{\infty}_{+}=-\lambda^{\infty}_{-}$, we have the same invariance for $\mathrm{d}\nu_{-}^{t}$. So, $t$ almost everywhere,
×$\int_{\mathbb{R}\times\mathbb{S}^{2}}\Big{(}\vec{V}^{\infty}(s,\omega)\cdot\nabla_{\omega}\beta_{\infty}(s,\omega)\Big{)}\mathrm{d}\nu_{\infty}^{t}(s,\omega)=0.$
So, for almost all $(t,s,\omega)\in\mathbb{R}^{2}\times\mathbb{S}^{2}$, we have $\mathrm{div}_{\omega}\Big{(}\vec{V}^{\infty}(s,\omega)\mathrm{d}\nu^{t}(s,\omega)\Big{)}=0$. ∎
4.3 The Wigner measure
Proof of Theorem 3.8.
Assume Assumption 1.2 and 1.3. Let $\left(\varphi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ a uniformly bounded family in $\mathrm{L}^{\infty}\left(\mathbb{R},\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)$ be a solution to equation (12) with initial normalized condition $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$. According to Theorem 3.7, we can consider ${\left(M^{t}_{\infty}\mathrm{d}\nu^{t}_{\infty},\Gamma^{t}\mathrm{d}\gamma^{t}\right)}$ a pair of two-scale semiclassical measure of $\left(\varphi^{\varepsilon}_{t}\right)_{\varepsilon>0}$. Then, according to proof of Section 4.2.1 and Section 4.2.2, we conclude. ∎
Proof of Theorem 3.10.
Assume Assumption 1.2 and 1.3. Let $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ a uniformly bounded family in $\mathrm{L}^{\infty}\left(\mathbb{R},\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)$ a solution to equation (1) with initial normalized condition $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$.
On the one hand, according to Lemma 2.3, we know the evolution of any semiclassical measure of $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ outside the curve $\mathcal{C}$.
On the other hand, according to Lemma 3.2, we consider the family $\left(\varphi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ uniformly bounded in $\mathrm{L}^{\infty}\left(\mathbb{R},\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)$ solution to equation (12) with initial normalized condition $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ a family of $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$. Then, by applying the Lemma 4.1, we have the link between any semiclassical measure of $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ over the curve $\mathcal{C}$ and the associated measure of $\left(\varphi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ on $\mathbb{R}_{s}$. Then, by applying Theorem 3.8, we have the evolution of the semiclassical measure over the curve $\mathcal{C}$.
Finally, by applying Remark 3.5 and by calculating the shape of the two-scale semiclassical measure at finite distance according to expressions (5) and (28), we conclude. ∎
4.4 Proof of Theorem 1.4 and Theorem 1.5
Proof of Theorem 1.4.
Assume Assumption 1.2 and 1.3. Let $\vec{f}\in\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$. Let ${\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}}$ be the solution to equation (1) with for all $\varepsilon\in(0,1]$ and for all $x\in\mathbb{R}^{2}$,
$\psi_{0}^{\varepsilon}(x)=\frac{1}{\sqrt{\varepsilon}}\vec{f}\left(\frac{x-x_{0}}{\sqrt{\varepsilon}}\right),$
then, applying Theorem 3.10, according to Remark 3.5 and Corollary 3.6, there exist ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ a vanishing sequence and ${\Gamma^{t}\mathrm{d}\gamma^{t}}$ a two-scale semiclassical measure at finite distance, such that for all $(a,\Xi)\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)\times\mathscr{C}_{c}^{\infty}(\mathbb{R},\mathbb{C})$,
$\mathcal{I}^{\varepsilon_{k}}_{\psi}(\Xi,a)\underset{k\to+\infty}{\longrightarrow}\frac{1}{2}\sum_{n\in\mathbb{Z}}\int_{\mathbb{R}^{3}}\Xi(t)\mathrm{tr}\left(a\big{(}\gamma(s)\big{)}\begin{pmatrix}1&\frac{\mathrm{i}\sigma}{\lambda_{n}(s,\sigma)}e^{-\mathrm{i}\theta(s)}\\ \frac{-\mathrm{i}\sigma}{\lambda_{n}(s,\sigma)}e^{\mathrm{i}\theta(s)}&1\end{pmatrix}\mathrm{d}\gamma^{t}_{n}(s,\sigma)\right)\mathrm{d}t,$
with for all $n\in\mathbb{Z}$, $\mathrm{d}\gamma_{n}^{t}$ satisfies the following equation
$\left\{\begin{array}[]{rll}\partial_{t}\ \mathrm{d}\gamma_{n}^{t}&=&\left\{\mathrm{d}\gamma_{n}^{t},\lambda_{n}\right\}\\ \mathrm{d}\gamma_{n}^{t}|_{t=0}&=&\mathrm{d}\gamma_{n,0}\end{array}\right.,$
with $\mathrm{d}\gamma_{n,0}=\mathrm{tr}\Big{(}\Pi_{n}\Gamma^{0}\Big{)}\mathrm{d}\gamma^{0}$ and $\Gamma^{0}\mathrm{d}\gamma^{0}$ is a two-scale semiclassical measure associated to the initial data.
We conclude thanks to the conservation of $\varepsilon$-oscillation (Lemma 2.2). ∎
Proof of Theorem 1.5.
Assume Assumption 1.2 and 1.3. Let ${\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}}$ a uniformly bounded family of $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$, $\nu$ a non negative scalar Radon measure on $\mathbb{R}_{t}\times\mathbb{R}_{x}^{2}$ and ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ a vanishing sequence such that for all $(a,\Xi)\in\mathscr{C}_{c}^{\infty}\left(\Omega,\mathbb{C}\right)\times\mathscr{C}_{c}^{\infty}(\mathbb{R},\mathbb{C})$,
$\int_{\mathbb{R}}\int_{\Omega}\Xi(t)a(x)\left|\psi_{t}^{\varepsilon_{k}}(x)\right|^{2}\mathrm{d}x\mathrm{d}t\underset{k\to+\infty}{\longrightarrow}\int_{\mathbb{R}}\int_{\Omega}\Xi(t)a(x)\mathrm{d}\nu(t,x),$
where ${\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}}$ solves (1) with ${\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}}$ as initial condition. A consequence of Theorem 3.8 and Remark 3.5 is
$\mathrm{d}\nu\mathds{1}_{\mathcal{C}}\geqslant\sum_{n\in\mathbb{Z}}\Phi^{*}\left(\mathrm{Tr}\Big{(}\Pi_{n}(s,\sigma)\Big{)}\mathrm{d}\gamma^{t}_{n}\right)\mathds{1}_{\mathcal{C}}\mathrm{d}t$
where $\left(\mathrm{d}\gamma_{n}^{t}\right)_{n\in\mathbb{Z}}$ is defined by (6). We conclude with $\mathrm{Tr}\left(\Pi_{n}(s,\sigma)\right)=1$. ∎
Appendix A Proof of compactness at infinity
Lemma A.1 (Conservation of compactness at infinity).
Let $T>0$. If $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ is compact at infinity then $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ solution to (1) is compact at infinity uniformly for $t\in[-T,T]$.
Proof of Lemma A.1.
Let $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ be compact at infinity. The proof relies on the analysis of
$\psi_{R}^{\varepsilon}(t)\coloneqq\chi_{R}(x)\psi_{t}^{\varepsilon},$
where ${\chi_{R}(x)=\chi\left(\frac{x}{R}\right)}$ with $\chi\in\mathscr{C}^{\infty}\left(\mathbb{R}^{2},[0,1]\right),\chi\equiv 0$ for $\left|x\right|<1$ and $\chi\equiv 1$ for $\left|x\right|>2$. We observe that $\psi_{R}^{\varepsilon}$ solves
$\mathrm{i}\varepsilon\partial_{t}\psi_{R}^{\varepsilon}=\mathrm{H}_{\varepsilon}\psi_{R}^{\varepsilon}+\varepsilon\left[\chi_{R}(x),\begin{pmatrix}0&\mathrm{D}_{1}-\mathrm{i}\mathrm{D}_{2}\\ \mathrm{D}_{1}+\mathrm{i}\mathrm{D}_{2}&0\end{pmatrix}\right]\psi_{t}^{\varepsilon}.$
Nevertheless, the first derivatives of $\chi$ are bounded therefore there exists $C>0$ such that
$\left\lVert\left[\chi_{R}(x),\begin{pmatrix}0&\mathrm{D}_{1}-\mathrm{i}\mathrm{D}_{2}\\ \mathrm{D}_{1}+\mathrm{i}\mathrm{D}_{2}&0\end{pmatrix}\right]\right\rVert_{\mathcal{L}(\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right))}\leqslant\frac{C}{R}.$
Then, by an energy argument,
$\frac{\mathrm{d}}{\mathrm{d}\mathrm{t}}\left\lVert\psi_{R}^{\varepsilon}(t)\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\leqslant\frac{C}{R}\left\lVert\psi^{\varepsilon}_{0}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)},$
so
$\left\lVert\psi_{R}^{\varepsilon}(t)\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}$ $\leqslant\left\lVert\psi_{R}^{\varepsilon}(0)\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}+\frac{C}{R}\left|t\right|\left\lVert\psi^{\varepsilon}_{0}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)},$
$\underset{R\to+\infty}{\longrightarrow}0.$
Appendix B Construction of semiclassical measures
B.1 Semiclassical measures for families valued in a separable Hilbert space
We will use the more general framework of [14] and consider families valued in a separable Hilbert space $\mathcal{H}$. More precisely, in this section, $\mathcal{H}=\mathrm{L}^{2}\left(\mathbb{R},\mathbb{C}^{2}\right)$. Note that in section 2.1, $\mathcal{H}=\mathbb{C}^{2}$.
Lemma B.1 (Operator valued measure).
Let ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$ a bounded family in $\mathrm{L}^{2}\left(\mathbb{R},\mathcal{H}\right)$, then there exists ${(\varepsilon_{k})_{k>0}}$ a vanishing sequence and a positive operator valued measure $\Gamma\mathrm{d}\gamma$ such that for all $a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathcal{K}\left(\mathcal{H}\right)\right)$,
$\left\langle\mathrm{Op}_{\varepsilon_{k}}(a)f^{\varepsilon_{k}},f^{\varepsilon_{k}}\right\rangle\underset{k\to+\infty}{\longrightarrow}\int_{\mathbb{R}^{2}}\mathrm{Tr}_{\mathcal{L}(\mathcal{H})}\Big{(}a(x,\xi)\Gamma(x,\xi)\Big{)}\mathrm{d}\gamma(x,\xi),$
with $\mathcal{K}\left(\mathcal{H}\right)$ the set of compact operators on $\mathcal{H}$.
Proof of Lemma B.1.
Let $\mathcal{H}$ be a separable Hilbert space, ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$ uniformly bounded in $\mathrm{L}^{2}\left(\mathbb{R},\mathcal{H}\right)$, $(h_{n})_{n\in\mathbb{N}}$ a Hermitian basis of $\mathcal{H}$ and $\left(\left(f^{\varepsilon}_{n}\right)_{\varepsilon>0}\right)_{n\in\mathbb{N}}$ families of $\mathrm{L}^{2}\left(\mathbb{R}\right)$ such that $x$ a.e.
$f^{\varepsilon}(x)=\sum_{n\in\mathbb{N}}\underbrace{f^{\varepsilon}_{n}(x)}_{\in\mathbb{C}}h_{n}.$
Then the families $\left(f^{\varepsilon}_{n}\right)_{\varepsilon>0}$ satisfy, for some constant $C>0$,
$\sum_{n\in\mathbb{N}}\left\lVert f^{\varepsilon}_{n}\right\rVert_{\mathrm{L}^{2}(\mathbb{R})}^{2}=\left\lVert f^{\varepsilon}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R},\mathcal{H}\right)}^{2}\leqslant C.$ (36)
For all $n\in\mathbb{N}$, $\left(f^{\varepsilon}_{n}\right)_{\varepsilon>0}$ are bounded in $\mathrm{L}^{2}\left(\mathbb{R}\right)$. By a diagonal extraction process, we can find a sequence ${(\varepsilon_{k})_{k>0}}$ and a family of Radon measures $\left(\mathrm{d}\mu_{n,m}\right)_{(n,m)\in\mathbb{N}^{2}}$ such that for all $b\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2}\right)$ and for all $(n,m)\in\mathbb{N}^{2}$,
$\left\langle\mathrm{Op}_{\varepsilon_{k}}(b)f^{\varepsilon_{k}}_{n},f^{\varepsilon_{k}}_{m}\right\rangle\underset{k\to+\infty}{\longrightarrow}\int_{\mathbb{R}^{2}}\mathrm{tr}\Big{(}b(x,\xi)\mathrm{d}\mu_{n,m}(x,\xi)\Big{)},$
with ${\mathrm{d}\mu_{n,n}\left(\mathbb{R}^{2}\right)\leqslant\limsup_{k\to+\infty}\left\lVert f^{\varepsilon_{k}}_{n}\right\rVert_{\mathrm{L}^{2}(\mathbb{R})}^{2}}$ and $\mathrm{d}\mu_{n,m}$ absolutely continuous with respect to $\mathrm{d}\mu_{n,n}$ and $\mathrm{d}\mu_{m,m}$. Moreover, for all $(n,m)\in\mathbb{N}^{2}$, $\mathrm{d}\mu_{n,n}$ is a positive Radon measure and $\mathrm{d}\mu_{m,n}=\overline{\mathrm{d}\mu_{n,m}}$.
First of all, we construct the operator valued measure $\Gamma\mathrm{d}\gamma$. By (36),
$\limsup_{k\to+\infty}\sum_{n\in\mathbb{N}}\left\lVert f^{\varepsilon_{k}}_{n}\right\rVert_{\mathrm{L}^{2}}^{2}<+\infty,$
then ${\mathrm{d}\gamma\coloneqq\sum_{n\in\mathbb{N}}\mathrm{d}\mu_{n,n}}$ defines a positive Radon measure of finite mass. By the Radon-Nykodym Theorem, there exists a sequence of $\mathrm{d}\gamma$-measurable functions $\left(\Gamma_{n,m}\right)_{(n,m)\in\mathbb{N}^{2}}$ such that $\forall(n,m)\in\mathbb{N}^{2}$,
$\mathrm{d}\mu_{n,m}=\Gamma_{n,m}\mathrm{d}\gamma$
with $0\leqslant\Gamma_{n,n}\leqslant 1$ and $\Gamma_{n,m}=\overline{\Gamma_{m,n}}$ $\mathrm{d}\gamma$-almost everywhere. Let us define the operator valued function $\Gamma:(x,\xi)\mapsto\Gamma(x,\xi)$ such that, for all $(n,m)\in\mathbb{N}^{2}$, for all $(x,\xi)\in\mathbb{R}^{2}$,
$\left\langle\Gamma(x,\xi)h_{n},h_{m}\right\rangle_{\mathcal{H}}=\Gamma_{n,m}(x,\xi).$
Because of the positivity of $\left(\mathrm{d}\mu_{n,n}\right)_{n\in\mathbb{N}}$ and the adjoint relation of $\left(\mathrm{d}\mu_{n,m}\right)_{(n,m)\in\mathbb{N}^{2}}$, $\Gamma(x,\xi)$ is a positive self-adjoint operator and its trace satisfies
$\int_{\mathbb{R}^{2}}\mathrm{Tr}_{\mathcal{L}(\mathcal{H})}\Big{(}\Gamma(x,\xi)\Big{)}\mathrm{d}\gamma(x,\xi)=\sum_{n\in\mathbb{N}}\int_{\mathbb{R}^{2}}\Gamma_{n,n}(x,\xi)\mathrm{d}\gamma(x,\xi)=\sum_{n\in\mathbb{N}}\mathrm{d}\gamma_{n,n}\left(\mathbb{R}^{2}\right)<+\infty.$
Therefore, the pair $\left(\Gamma,\mathrm{d}\gamma\right)$ generates an operator measure $\Gamma\mathrm{d}\gamma$. We now have to prove that $\Gamma\mathrm{d}\gamma$ plays the expected role.
Let $a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathcal{K}\left(\mathcal{H}\right)\right)$, $\Pi_{\leqslant n}$ the projector on the finite dimensional subspace of $\mathcal{H}$ generated by $(h_{m})_{m\leqslant n}$ and consider the finite rank operator-valued function $a_{n}\coloneqq\Pi_{\leqslant n}a\Pi_{\leqslant n}$ which belongs to $\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathcal{K}\left(\mathcal{H}\right)\right)$, then
$\sup_{(x,\xi)\in\mathrm{supp}(a)}\left\lVert a(x,\xi)-a_{n}(x,\xi)\right\rVert_{\mathcal{L}\left(\mathcal{H}\right)}\underset{n\to+\infty}{\longrightarrow}0.$ (37)
By passing to the limit, thanks to the construction of $\left(\varepsilon_{k}\right)_{k\in\mathbb{N}}$,
$\left\langle\mathrm{Op}_{\varepsilon_{k}}\left(a_{n}\right)f^{\varepsilon_{k}},f^{\varepsilon_{k}}\right\rangle$ $=\sum_{m,\ell\leqslant n}\left\langle\mathrm{Op}_{\varepsilon_{k}}\Big{(}\left\langle a_{n}h_{m},h_{\ell}\right\rangle_{\mathcal{H}}\Big{)}f^{\varepsilon_{k}}_{m},f^{\varepsilon_{k}}_{\ell}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R},\mathbb{C}^{2}\right)}$
$\underset{k\to+\infty}{\longrightarrow}\sum_{m,\ell\leqslant n}\int_{\mathbb{R}^{2}}\left\langle a_{n}(x,\xi)h_{m},h_{\ell}\right\rangle_{\mathcal{H}}\mathrm{d}\gamma_{m,\ell}(x,\xi)$
$=\sum_{m,\ell\leqslant n}\int_{\mathbb{R}^{2}}\left\langle a_{n}h_{m},h_{\ell}\right\rangle_{\mathcal{H}}\Gamma_{m,\ell}\mathrm{d}\gamma(x,\xi)$
$=\int_{\mathbb{R}^{2}}\mathrm{Tr}_{\mathcal{L}(\mathcal{H})}\Big{(}a_{n}(x,\xi)\Gamma(x,\xi)\Big{)}\mathrm{d}\gamma(x,\xi).$
Then, with (37) and Lebesgue dominated convergence Theorem, we have
$\int_{\mathbb{R}^{2}}\mathrm{Tr}_{\mathcal{L}(\mathcal{H})}\Big{(}\big{(}a(x,\xi)-a_{n}(x,\xi)\big{)}\Gamma(x,\xi)\Big{)}\mathrm{d}\gamma(x,\xi)\underset{n\to+\infty}{\longrightarrow}0.$
B.2 Two-scale semiclassical measures
In this section, we prove Theorem 3.3.
Proof of Theorem 3.3.
Let ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$ uniformly bounded in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$. For simplicity, we assume that the sequence $\left(\varepsilon\right)_{\varepsilon>0}$ realizes the semiclassical measure $\mathrm{d}\mu$ of ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$. We use the decomposition (17). First we analysis of the concentration of $a^{\delta}$.
  • $\bullet$
    Construction of $\mathrm{d}\mu$ outside $\mathcal{C}$
For $a\in\mathfrak{A}$ such that $a=0$ in a neighborhood of $\mathcal{C}$, then for $\varepsilon$ small enough, we have
$\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a\right)$ $=\mathrm{Op}_{\varepsilon}\left(a\left(s,y;\sigma,\eta;\frac{y}{\sqrt{\varepsilon}},\frac{\sigma}{\sqrt{\varepsilon}},\frac{\eta}{\sqrt{\varepsilon}}\right)\right)$
$=\mathrm{Op}_{1}\left(a_{\infty}\left(s,y,\varepsilon\sigma,\varepsilon\eta,\frac{(y,\varepsilon\sigma,\varepsilon\eta)}{|(y,\varepsilon\sigma,\varepsilon\eta)|}\right)\right).$
After extraction of a subsequence that we still denote by $(\varepsilon)_{\varepsilon>0}$, we have
$\limsup_{\delta\to 0}\limsup_{R\to+\infty}\limsup_{\varepsilon\to 0}\mathcal{I}^{\varepsilon}_{f}\left(a^{\delta}\right)=\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\mathrm{tr}\left(a_{\infty}\left(s,y,\sigma,\eta,\frac{(y,\sigma,\eta)}{|(y,\sigma,\eta)|}\right)\mathrm{d}\mu(s,y,\sigma,\eta)\right).$
Then we construct the pair $\Gamma\mathrm{d}\gamma$.
  • $\bullet$
    Construction of $\Gamma\mathrm{d}\gamma$
Let $a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{7},\mathbb{C}^{2,2}\right)$, since $a$ is compactly supported, we recall (21)
$\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a\right)=\mathfrak{T}_{\sqrt{\varepsilon}}^{\dagger}\mathrm{Op}_{\sqrt{\varepsilon},1}(b)\mathfrak{T}_{\sqrt{\varepsilon}}+\mathcal{O}\left(\sqrt{\varepsilon}\right),$
The function $b$ is valued in the set of compact operators on $\mathcal{H}$. We can see ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$ as a bounded family of $\mathrm{L}^{2}\left(\mathbb{R}_{s},\mathcal{H}\right)$. We are thus doing classical calculus on $\mathrm{L}^{2}\left(\mathbb{R}_{s},\mathcal{H}\right)$. We apply Lemma B.1, then there exists $\left(\varepsilon_{k}\right)_{k\in\mathbb{N}}$ a vanishing sequence such that for all $a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{7},\mathbb{C}^{2,2}\right)$,
$\lim_{k\to+\infty}\left\langle\mathfrak{Op}_{\varepsilon_{k}}\left(a\right)f^{\varepsilon_{k}},f^{\varepsilon_{k}}\right\rangle_{\mathrm{L}^{2}(\mathbb{R}^{2},\mathbb{C}^{2})}=\int_{\mathbb{R}^{2}}\mathrm{Tr}\Big{(}a^{W}(s,0;0,0;y,\sigma,\mathrm{D}_{y})\Gamma(s,\sigma)\Big{)}\mathrm{d}\gamma(s,\sigma).$
Then, we apply this property to $a_{\varepsilon,R}$. Finally we construct the pair $M_{\infty}\mathrm{d}\nu_{\infty}$.
  • $\bullet$
    Construction of $M_{\infty}\mathrm{d}\nu_{\infty}$
Let $a\in\mathfrak{A}$, then we recall (20)
$\begin{split}\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a_{\varepsilon,\delta}^{R}\right)=&\ \mathfrak{T}_{\sqrt{\varepsilon}}^{\dagger}\mathrm{Op}_{1}\Bigg{(}a(s,\sqrt{\varepsilon}y,\varepsilon\sigma,\sqrt{\varepsilon}\eta,y,\sqrt{\varepsilon}\sigma,\eta)\chi\left(\frac{|(\sqrt{\varepsilon}y,\varepsilon\sigma,\sqrt{\varepsilon}\eta)|}{\delta}\right)\\ &\times\left(1-\chi\left(\frac{|(y,\sqrt{\varepsilon}\sigma,\eta)|}{R}\right)\right)\Bigg{)}\mathfrak{T}_{\sqrt{\varepsilon}}\\ =&\ \mathfrak{T}_{\sqrt{\varepsilon}}^{\dagger}\mathrm{Op}_{1}\underbrace{\left(a(s,0,0,0,y,\sqrt{\varepsilon}\sigma,\eta)\chi\left(\frac{|(\sqrt{\varepsilon}y,\varepsilon\sigma,\sqrt{\varepsilon}\eta)|}{\delta}\right)\left(1-\chi\left(\frac{|(y,\sqrt{\varepsilon}\sigma,\eta)|}{R}\right)\right)\right)}_{\eqqcolon\widetilde{a}^{R}_{\delta}(s,0,0,0,y,\sqrt{\varepsilon}\sigma,\eta)}\mathfrak{T}_{\sqrt{\varepsilon}}\\ &+\mathcal{O}\left(\delta\right).\end{split}$
therefore we consider the family $\widetilde{f}^{\varepsilon}\coloneqq\mathfrak{T}_{\sqrt{\varepsilon}}f^{\varepsilon}$ which is bounded in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$. Then the quantity
$\mathscr{I}^{\varepsilon_{k}}_{\widetilde{f}}\left(\widetilde{a}^{R}_{\delta}\right)\coloneqq\left\langle\mathrm{Op}_{1}\left(\widetilde{a}^{R}_{\delta}(s,0,0,0,y,\sqrt{\varepsilon_{k}}\sigma,\eta)\right)\widetilde{f}^{\varepsilon_{k}},\widetilde{f}^{\varepsilon_{k}}\right\rangle_{\mathrm{L}^{2}(\mathbb{R}^{2},\mathbb{C}^{2})}.$
is uniformly bounded in $\varepsilon>0$ and $R>1$. $a^{R}_{\delta}$ is supported at distance at least $R\sqrt{\varepsilon}$ of $\mathcal{C}$, so
$\mathscr{I}^{\varepsilon_{k}}_{\widetilde{f}}\left(\widetilde{a}^{R}_{\delta}\right)=\left\langle\mathrm{Op}_{1}\left(a_{\infty}(s,0,0,0,y,\sqrt{\varepsilon_{k}}\sigma,\eta)\right)\widetilde{f}^{\varepsilon_{k}},\widetilde{f}^{\varepsilon_{k}}\right\rangle_{\mathrm{L}^{2}(\mathbb{R}^{2},\mathbb{C}^{2})},$
as soon as $R$ is large enough, depending of $\varepsilon$. We then deduce by a diagonal extraction argument that there exists a subsequence $\left(\varepsilon_{n}\right)_{n\in\mathbb{N}}$ of $\left(\varepsilon_{k}\right)_{k\in\mathbb{N}}$, a sequence $\left(R_{n}\right)_{n\in\mathbb{N}}$ and a sequence $\left(\delta_{n}\right)_{n\in\mathbb{N}}$ such that for all $a\in\mathfrak{A}$,
$\mathscr{I}^{\varepsilon_{n}}_{\widetilde{f}}\left(\widetilde{a}^{R_{n}}_{\delta_{n}}\right)\underset{n\to+\infty}{\longrightarrow}\mathscr{I}\left(a_{\infty}\right).$
It remains to prove that $a_{\infty}\mapsto\mathscr{I}\left(a_{\infty}\right)$ is a measure, which will allow us to define $M_{\infty}\mathrm{d}\nu_{\infty}$. First, we prove that $a_{\infty}\mapsto\mathscr{I}\left(a_{\infty}\right)$ is a non negative matrix-valued distribution. We observe that there exists $C>0$ such that for all $a\in\mathfrak{A}$, for all $n\in\mathbb{N}$,
$\mathscr{I}^{\varepsilon_{n}}_{\widetilde{f}}\left(\widetilde{a}^{R_{n}}_{\delta_{n}}\right)\leqslant C\mathrm{N}_{d}\left(\widetilde{a}^{R_{n}}_{\delta_{n}}\right),$
with $\mathrm{N}_{d}$ a Schwarz semi-norm. We have $\mathrm{N}_{d}\left(\widetilde{a}^{R_{n}}_{\delta_{n}}\right)\to\mathrm{N}_{d}\left(a_{\infty}\right)$ then $\mathscr{I}\left(a_{\infty}\right)\leqslant C\mathrm{N}_{d}\left(a_{\infty}\right)$. Therefore $\mathscr{I}$ is a matrix-valued distribution. Secondly, we observe that the operators $a\mapsto\mathrm{Op}_{1}\left(a^{R}_{\delta}(s,0,0,0,y,\sqrt{\varepsilon}\sigma,\eta)\right)$ satisfy a semiclassical calculus in the parameters $\varepsilon$, $1/R$ and $\delta$. We have for all $a,a_{1},a_{2}\in\mathfrak{A}$,
  • $\bullet$
    Adjoint : $\mathrm{Op}_{1}\left(\widetilde{a}^{R}_{\delta}(s,0,0,0,y,\sqrt{\varepsilon}\sigma,\eta)\right)^{\dagger}=\mathrm{Op}_{1}\left(\overline{\widetilde{a}^{R}_{\delta}}(s,0,0,0,y,\sqrt{\varepsilon}\sigma,\eta)\right)$
  • $\bullet$
    Symbolic Calculus : as operator of $\mathcal{L}\left(\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)$,
    $\mathrm{Op}_{1}\Big{(}\widetilde{a}^{R}_{1,\delta}\Big{)}\mathrm{Op}_{1}\Big{(}\widetilde{a}^{R}_{2,\delta}\Big{)}=\mathrm{Op}_{1}\Big{(}\left(\widetilde{a}_{1}\widetilde{a}_{2}\right)^{R}_{1,\delta}\Big{)}+\mathcal{O}\left(\sqrt{\varepsilon}\right)+\mathcal{O}\left(\frac{1}{R}\right),$
  • $\bullet$
    Weak Gårding inequality : if $a\geqslant 0$, then for all $\widetilde{\delta}$, there exists $C_{\widetilde{\delta}}>0$, such that for all $f\in\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$,
    $\left\langle\mathrm{Op}_{1}\Big{(}\widetilde{a}^{R}_{\delta}\Big{)}\widetilde{f},\widetilde{f}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\geqslant-\left(\widetilde{\delta}+C_{\widetilde{\delta}}\left(\varepsilon+\frac{1}{R}\right)\right)\left\lVert f\right\rVert^{2}_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)},$
Therefore $a_{\infty}\mapsto\mathscr{I}\left(a_{\infty}\right)$ is a non negative matrix-valued measure defined on ×$\mathbb{R}\times\mathbb{S}^{2}$ that we denote $M\mathrm{d}\nu$. ∎
Appendix C Spectral Analysis of the operator $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$
In this appendix, we study the spectrum of the operator $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$ introduced in (34).
Theorem C.1 (Spectrum of $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$).
For all $(s,\sigma)\in\mathbb{R}^{2}$, the spectrum of $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$ consists in multiplicity one eigenvalues
$\mathrm{Sp}\Big{(}\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)\Big{)}=\left\{\lambda_{0}(\sigma)\right\}\bigcup\left\{\lambda_{n}(s,\sigma)\ |\ n\in\mathbb{Z}_{\neq 0}\right\},$
with for all $n\in\mathbb{Z}_{\neq 0}$,
$\lambda_{0}(\sigma)\coloneqq-\sigma\;,\;\lambda_{n}(s,\sigma)\coloneqq\mathrm{sgn}(n)\sqrt{\sigma^{2}+2|n|r(s)^{2}}.$
Moreover, an Hilbertian basis of $\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)$, composed of eigenfunctions of $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$, is given by ${\left(g_{n}^{s,\sigma}\right)_{n\in\mathbb{Z}}}$
$g_{0}^{s}(y)$ $\coloneqq\frac{1}{\sqrt{r(s)}}\begin{pmatrix}0\\ \mathfrak{h}_{0}\big{(}r(s)y\big{)}\end{pmatrix},$
$g_{n}^{s,\sigma}(y)$ $\coloneqq\alpha_{n}(s,\sigma)\begin{pmatrix}\mathfrak{h}_{|n|-1}\big{(}r(s)y\big{)}\\ \frac{\sqrt{2|n|}r(s)}{\lambda_{n}(\sigma)+\sigma}\mathfrak{h}_{|n|}\big{(}r(s)y\big{)}\end{pmatrix},\quad\quad n\in\mathbb{Z}_{\neq 0}$
$\mbox{with}\quad\quad\alpha_{n}(s,\sigma)$ $\coloneqq\frac{1}{\sqrt{2r(s)}}\sqrt{1+\frac{\sigma}{\lambda_{n}(s,\sigma)}},$
and $\left(\mathfrak{h}_{n}\right)_{n\in\mathbb{N}}$ are Hilbert functions defined in (27). Moreover, the projectors over the eigenspaces of $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$ defined by
$\Pi_{0}(s,\sigma)\coloneqq g_{0}^{s}\otimes g_{0}^{s},\quad\quad\Pi_{n}(s,\sigma)\coloneqq g_{n}^{s,\sigma}\otimes g_{n}^{s,\sigma},$
are in $\mathscr{C}^{\infty}\left(\mathbb{R}^{2},\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)\right)$.
Notably,
$\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)=\sum_{n\in\mathbb{Z}}\lambda_{n}(s,\sigma)\Pi_{n}(s,\sigma).$
Proof of Theorem C.1.
First, we will rescale the function in order to have the form $f\left(s,r(s)y\right)$. We can do it because $r$ is smooth ($\mathrm{m}$ is in $\mathscr{C}^{\infty}\left(\mathbb{R}^{2},\mathbb{R}\right)$) and strictly non negative (Assumption 1.2). Then we dilate with $r(s)^{-1}$ and we denote by $\zeta$ the ratio $\sigma/r(s)$. For all $f\in\mathrm{L}^{2}\left(\mathbb{R},\mathbb{C}^{2}\right)$,
$\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)\Bigg{[}f\Big{(}r(s)y\Big{)}\Bigg{]}$ $=\left[\begin{pmatrix}\sigma&r(s)(y+\mathrm{i}\mathrm{D}_{y})\\ r(s)(y-\mathrm{i}\mathrm{D}_{y})&-\sigma\end{pmatrix}f\right]\Big{(}r(s)y\Big{)}$
$=r(s)\left[\begin{pmatrix}\zeta&y+\mathrm{i}\mathrm{D}_{y}\\ y-\mathrm{i}\mathrm{D}_{y}&-\zeta\end{pmatrix}f\right]\Big{(}r(s)y\Big{)}.$
Let us define
$\mathscr{T}\coloneqq\begin{pmatrix}\zeta&\mathfrak{a}\\ \mathfrak{a}^{\dagger}&-\zeta\end{pmatrix},$
where $\mathfrak{a}\coloneqq y+\mathrm{i}\mathrm{D}_{y}$ denotes the usual annihilation operator on the variable $y$. Then the link between the spectrum is
$\mathrm{Sp}\Big{(}\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)\Big{)}=r(s)\mathrm{Sp}(\mathscr{T})$
So we can study $\mathscr{T}$ instead of $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$. Let $\lambda_{\zeta}\in\mathrm{Sp}(\mathscr{T}),\begin{pmatrix}f\\ g\end{pmatrix}\in E_{\lambda_{\zeta}}(\mathscr{T})$ if and only if
$\left\{\begin{matrix}2\zeta f&+&\mathfrak{a}g&=&\left(\lambda_{\zeta}+\zeta\right)f\\ &&\mathfrak{a}^{\dagger}f&=&\left(\lambda_{\zeta}+\zeta\right)g\end{matrix}\right.$
If $\lambda_{\zeta}=-\zeta$, the injectivity of $\mathfrak{a}^{\dagger}$ on $\mathrm{L}^{2}\left(\mathbb{R}\right)$ implies $f=0$ whence $\mathfrak{a}g=0$ so $g=A\mathfrak{h}_{0}$ for some $A\in\mathbb{C}$. So $-\zeta\in\mathrm{Sp}(\mathscr{T})$ and is a simple eigenvalue with ${\begin{pmatrix}0\\ \mathfrak{h}_{0}\end{pmatrix}}$ a normalized eigenfunction. If $\lambda_{\zeta}\neq-\zeta$, then
$\left\{\begin{matrix}\mathfrak{a}\mathfrak{a}^{\dagger}f&=&\left(\lambda_{\zeta}^{2}-\zeta^{2}\right)f\\ g&=&\frac{1}{\lambda_{\zeta}+\zeta}\mathfrak{a}^{\dagger}f\end{matrix}\right.$
and eigenvalues of $\mathfrak{a}\mathfrak{a}^{\dagger}$ are known so we deduce that there exists $n\in\mathbb{N}_{>0},\lambda_{\zeta}^{2}-\zeta^{2}=2n$ so
$\lambda_{\zeta}\in\left\{\mathrm{sgn}(n)\sqrt{\zeta^{2}+2|n|}\ |\ n\in\mathbb{Z}_{\neq 0}\right\}=\mathrm{Sp}\left(\mathfrak{a}\mathfrak{a}^{\dagger}\right),$
and for some $B\in\mathbb{C}$,
$\forall(\zeta,y)\in\mathbb{R}^{2},\ \begin{pmatrix}f_{n}\\ g_{n}\end{pmatrix}(\zeta,y)\coloneqq B\begin{pmatrix}\mathfrak{h}_{|n|-1}(y)\\ \frac{\sqrt{2|n|}}{\mathrm{sgn}(n)\sqrt{\zeta^{2}+2|n|}+\zeta}\mathfrak{h}_{|n|}(y)\end{pmatrix}.$
Then
$\mathrm{Sp}\Big{(}\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)\Big{)}=\left\{\lambda_{0}(\sigma)\right\}\bigcup\left\{\lambda_{n}(s,\sigma)\ |\ n\in\mathbb{Z}_{\neq 0}\right\}.$
The associated projectors $\left(\Pi_{n}\right)_{n\in\mathbb{Z}}$ are smooth because $r$ is smooth ($\mathrm{m}$ is in $\mathscr{C}^{\infty}\left(\mathbb{R}^{2},\mathbb{R}\right)$) and strictly non negative (Assumption 1.2); for all $n\in\mathbb{Z}_{\neq 0}$, $\lambda_{n}$ and $\lambda_{n}+\sigma$ is smooth and never vanishes. ∎

### v3

@@
-Semiclassical measure of the propagation between two topological insulators
+Propagation of Semiclassical Measures Between Two Topological Insulators
 Abstract
-We study the propagation of a family of initial conditions in the presence of two topological insulators without magnetic field where the interface is a smooth connected curve not compact without boundaries.
-The solution is governed by an adiabatic modulation of a Dirac operator with a variable mass.
-We determine the evolution of the semiclassical measure of the solution with a two-scale Wigner measure method by reducing the Dirac operator to a normal form.
-Keywords : propagation, Dirac operator, semiclassical analysis, Wigner measure, semiclassical measure valued in a separable Hilbert space, time dependent two-scale semiclassical measure.
+We study propagation in a system consisting of two topological insulators without a magnetic field, whose interface is a non-compact, smooth, and connected curve without boundary.
+The dynamics are governed by an adiabatic modulation of a Dirac operator with a smooth, effective variable mass.
+We determine the evolution of the semiclassical measure of the solution using a two-scale Wigner measure method, after reducing the Hamiltonian to a normal form.
+Keywords: propagation, Dirac operator, semiclassical analysis, Wigner measures, semiclassical measures valued in a separable Hilbert space, time-dependent two-scale semiclassical measures.
 $2024$ Mathematics Subject Classification : 35B40, 35F05, 35Q40.
-Acknowledgments.
-The author is very grateful to his advisors Clotilde Fermanian-Kammerer and Nicolas Raymond for many insightful discussions and valuable suggestions.
-The author acknowledges the support of the Région Pays de la Loire via the Connect Talent Project HiFrAn 2022 07750, and from the France 2030 program, Centre Henri Lebesgue ANR-11-LABX-0020-01.
 1 Introduction
-1.1 Topological insulators and Dirac type equations
-Topological insulators are materials that are insulating in their bulk but support robust currents along their boundaries, meaning that electrons can only move along the surface of the material.
-The Dirac equation has been derived in the context of topological insulators for honeycomb structures [9, 8, 5] such as graphene [22, 18, 7].
-From a global point of view, Dirac-type equations are often a simpler continuum macroscopic description of transport in a narrow energy band near the band crossing [2, 12, 24].
-In this paper we consider the following system of the Dirac evolution equation on $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$,
-$\left\{\begin{array}[]{ll}\Big{(}\varepsilon\mathrm{D}_{t}+\mathrm{m}(x)\sigma_{3}+\varepsilon\mathrm{D}_{1}\sigma_{1}+\varepsilon\mathrm{D}_{2}\sigma_{2}\Big{)}\psi^{\varepsilon}(t,x)=0,&\forall(t,x)\in\mathbb{R}\times\mathbb{R}^{2},\\ \psi^{\varepsilon}(0,x)=\psi^{\varepsilon}_{0}(x),&\forall x\in\mathbb{R}^{2},\end{array}\right.$ (1)
-where ${\varepsilon}$ is a small positive parameter, called semiclassical parameter.
-The function $\psi^{\varepsilon}(t,\cdot)$ also denoted $\psi^{\varepsilon}_{t}$ is interpreted as the wave function of an electron at time $t$ in a material consisting in two topological insulators.
-The family ${\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}}$ is a uniformly bounded (in $\varepsilon$) family of ${\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}$ of initial wave functions of the electron.
-The function $x\mapsto\mathrm{m}(x)$ is a smooth real-valued function (i.e.
-$\mathrm{m}\in\mathscr{C}^{\infty}\left(\mathbb{R}^{2},\mathbb{R}\right)$) and all its derivatives are bounded.
-The matrices ${\left(\sigma_{j}\right)_{1\leqslant j\leqslant 3}}$ are the Pauli matrices
-$\sigma_{1}\coloneqq\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\;,\;\sigma_{2}\coloneqq\begin{pmatrix}0&-\mathrm{i}\\ \mathrm{i}&0\end{pmatrix}\;,\;\sigma_{3}\coloneqq\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}.$
-We shall use the notation
-$\mathrm{H}_{\varepsilon}\coloneqq\begin{pmatrix}\mathrm{m}(x)&\varepsilon\mathrm{D}_{1}-\mathrm{i}\varepsilon\mathrm{D}_{2}\\ \varepsilon\mathrm{D}_{1}+\mathrm{i}\varepsilon\mathrm{D}_{2}&-\mathrm{m}(x)\end{pmatrix}.$
-Then the equation (1) rewrites $\varepsilon\mathrm{D}_{t}\psi_{t}^{\varepsilon}+\mathrm{H}_{\varepsilon}\psi_{t}^{\varepsilon}=0$.
-In this model, the function $\mathrm{m}$ splits the space into two parts, which are the two topological insulators.
-From a modeling point of view, the function $\mathrm{m}$ is positive on one of the topological insulators and negative on the other so the interface between the two topological insulators is defined by
-$\mathrm{E}\coloneqq\left\{x\in\mathbb{R}^{2}\ |\ \mathrm{m}(x)=0\right\}.$
-We assume the set $\mathrm{E}$ is a connected curve not compact without boundaries.
-From a physical point of view, the sign of $\mathrm{m}$ is the insulating phase of a material.
-This generates a unidirectional propagation along the interface $\mathrm{E}$.
-This propagation is one of the most important issue of physical applications of topological insulators and superconductors [2, 24], photonics [19, 22, 23], acoustics [21, 20] and fluid mechanics [13].
-Our aim is to describe the solution $\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}$ to equation (1) for initial conditions in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$.
-At least, we can understand limits as $\varepsilon$ goes to 00 by constructing semiclassical defect measures, also called Wigner measures.
-Then, we can deduce various properties of this kind of measures and thereby indirectly recover information about asymptotics as $\varepsilon$ goes to 00.
+Topological insulators and Dirac-type equations.
+Topological insulators are electronic materials that possess a bulk band gap, like ordinary insulators, but exhibit protected conducting states at their edges.
+The Dirac equation arises naturally in the context of topological insulators for honeycomb structures Fefferman et al.
+(2016b, a); Drouot (2019), such as graphene Raghu and Haldane (2008); Lee-Thorp et al.
+(2019); Drouot and Weinstein (2020).
+In this paper, we study the following system of Dirac evolution equation on $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$,
+$\left\{\begin{array}[]{ll}\left(\varepsilon\mathrm{D}_{t}+\begin{pmatrix}\mathrm{m}(x)&\varepsilon\mathrm{D}_{1}-\mathrm{i}\varepsilon\mathrm{D}_{2}\\ \varepsilon\mathrm{D}_{1}+\mathrm{i}\varepsilon\mathrm{D}_{2}&-\mathrm{m}(x)\end{pmatrix}\right)\psi^{\varepsilon}(t,x)=0,&\forall(t,x)\in\mathbb{R}\times\mathbb{R}^{2},\\ \psi^{\varepsilon}(0,x)=\psi^{\varepsilon}_{0}(x),&\forall x\in\mathbb{R}^{2},\end{array}\right$.
+(1.1)
+where ${\mathrm{D}_{\#}\coloneqq-\mathrm{i}\partial_{\#}}$ for $\in\{t,1,2\}$ with $\partial_{j}\coloneqq\partial_{x_{j}}$ for $j\in\{1,2\}$, and where ${\varepsilon}$ is a small positive parameter.
+The family ${\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}}$ is uniformly bounded (with respect to $\varepsilon$) in ${\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}$.
+The function $x\mapsto\mathrm{m}(x)$ is a smooth real-valued function, that is, $\mathrm{m}\in\mathscr{C}^{\infty}\left(\mathbb{R}^{2},\mathbb{R}\right)$, whose derivatives are all bounded.
+The interface between the two topological insulators is defined by
+$\mathrm{E}\coloneqq\left\{x\in\mathbb{R}^{2}\mid\mathrm{m}(x)=0\right\}.$
+To ensure that $\mathrm{m}$ separates $\mathbb{R}^{2}$ into two distinct regions, we assume that $\mathrm{E}$ is a smooth, connected curve without boundary.
+Under these conditions, there exists a unique family ${\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}}$ of solutions in ${\mathscr{C}\left(\mathbb{R},\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)}$ to (1.1).
+Physical interpretation.
+The conducting states at the edge of topological insulators are called edge states.
+Their propagation is one of the most important problems in the physical applications of topological insulators and superconductors Bernevig (2013); Volovik (1989), photonics Lu et al.
+(2014); Raghu and Haldane (2008); Rechtsman et al.
+(2013), acoustics Peano et al.
+(2015); Ma et al.
+(2019), and fluid mechanics Graf et al.
+(2021).
+These states are physically realized through the combined effects of spin–-orbit interactions and time-reversal symmetry.
+More precisely, two-dimensional topological insulators are quantum spin Hall insulators, which are closely related to the integer quantum Hall state Hasan and Kane (2010).
+In such materials, there exists an energy gap between the valence and conduction bands; however, unlike in a trivial insulator, these bands are ”twisted” or ”crossing” Zhu et al.
+(2012).
+From a global point of view, Dirac-type equations often provide a simpler continuum macroscopic description of transport in a narrow energy band near a band crossing Bernevig (2013); Fruchart and Carpentier (2013); Volovik (1989).
+Such models typically describe the evolution of the electronic wave function in topological insulators, such as graphene monolayers subjected to an external field or to non-homogeneous deformations.
+Here, the topological insulators under consideration are characterized by the sign of $\mathrm{m}$: the region where $\mathrm{m}$ is positive corresponds to a topological insulator of index $1$, while the region where $\mathrm{m}$ is negative corresponds to a topological insulator of index 1-1.
+The function $\psi^{\varepsilon}(t,\cdot)$, also denoted by $\psi^{\varepsilon}_{t}$, is interpreted as the wave function of an electron at time $t$ in a material consisting of two topological insulators.
+Figure 1: Illustration of asymmetric transport for the mass $\mathrm{m}(x)=\arctan(x_{1})-x_{2}$.
+Since equation (1.1) originates from quantum mechanics through the quantum spin Hall effect, we are naturally led to study the problem in phase space $\mathrm{T}^{*}\mathbb{R}^{2}=\mathbb{R}^{2}_{x}\times\mathbb{R}^{2}_{\xi}$.
+In this setting, $\xi\in\mathbb{R}^{2}$ is interpreted as the momentum of the particle.
 1.1 Edge states generated by coherent states
-The system (1) has already been studied in [1] where are performed precise computation. The edge mode computed is generated by a very specific shape of semiclassical wave packet.
-In [6], it was proved that some terms of the solution to equation (1) are negligible in $\mathrm{L}^{\infty}$-norm as $\varepsilon$ goes to 00.
-It has been proven, for initial conditions, that semiclassical wave packets concentrate at a point on the $\mathrm{E}$ curve with no constraint on orientation.
-However, these dispersive terms are not necessarily negligible in the $\mathrm{L}^{2}$ norm.
-Let us recall these results.
-Let $x_{0}\in\mathrm{E}$ and $t\mapsto x_{t}$ the solution to the ODE
-$\left\{\begin{array}[]{l}\dot{x}_{t}=\frac{\nabla\mathrm{m}(x_{t})^{\perp}}{\left|\nabla\mathrm{m}(x_{t})\right|},\\ \left.x_{t}\right|_{t=0}=x_{0}.\end{array}\right.$
-Let us define, for all $t\in\mathbb{R}$, $r_{t}\in\mathbb{R}_{>0}$ and $\theta_{t}\in\mathbb{R}$, such that $r_{t}\begin{pmatrix}-\sin\theta_{t}\\ \cos\theta_{t}\end{pmatrix}=\nabla\mathrm{m}(x_{t})$ and
-$\vec{V}_{\theta_{t}}\coloneqq\begin{pmatrix}e^{\frac{-\mathrm{i}\theta_{t}}{2}}\\ -e^{\frac{\mathrm{i}\theta_{t}}{2}}\end{pmatrix}.$
-Following [6], we define wave packet concentrated at $(x_{0},\xi_{0})\in\mathbb{R}^{4}$ as a $\varepsilon$-dependent function satisfying for all positive $\varepsilon$ and for all $x\in\mathbb{R}^{2}$,
-$\mathrm{WP}^{\varepsilon}_{x_{0},\xi_{0}}[\vec{f}](x)\coloneqq\frac{e^{\frac{\mathrm{i}}{\varepsilon}\xi_{0}\cdot(x-x_{0})}}{\sqrt{\varepsilon}}\vec{f}\left(\frac{x-x_{0}}{\sqrt{\varepsilon}}\right),$
+Coherent states.
+The system (1.1) has already been studied in Bal et al.
+(2021), where the computed edge mode is generated by a very specific semiclassical wave packet.
+Following Drouot (2022), we define a wave packet concentrated at $(x_{0},\xi_{0})\in\mathbb{R}^{4}$, a point in phase space, as an $\varepsilon$-dependent function satisfying, for all $\varepsilon>0$,
+$\mathrm{WP}^{\varepsilon}_{x_{0},\xi_{0}}[\vec{f}](x)\coloneqq\frac{e^{\frac{\mathrm{i}}{\varepsilon}\xi_{0}\cdot(x-x_{0})}}{\sqrt{\varepsilon}}\vec{f}\left(\frac{x-x_{0}}{\sqrt{\varepsilon}}\right),\quad x\in\mathbb{R}^{2},$
 with $\vec{f}\in\mathscr{S}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ independent of $\varepsilon$.
-We say that $\mathrm{WP}^{\varepsilon}_{x_{0},\xi_{0}}[\vec{f}]$ is oriented along $\vec{V}\in\mathbb{C}^{2}$ if there exists $f\in\mathscr{S}\left(\mathbb{R}^{2},\mathbb{C}\right)$ such that for all $x\in\mathbb{R}^{2},\vec{f}(x)=f(x)\vec{V}$.
-The parameter $\xi_{0}$ is interpreted as the impulsion (or the speed) of an electron of wave function $\mathrm{WP}^{\varepsilon}_{x_{0},\xi_{0}}[\vec{f}]$.
-Theorem 1.1 ([1] Theorem 2 and [6] Theorem A.1).
-Let $\vec{f}\in\mathscr{S}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ independent of $\varepsilon$.
-If ${\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}}$ solves (1) with for all $\varepsilon\in(0,1]$ and for all $x\in\mathbb{R}^{2}$,
-$\psi_{0}^{\varepsilon}(x)=\frac{1}{\sqrt{\varepsilon}}\vec{f}\left(\frac{x-x_{0}}{\sqrt{\varepsilon}}\right),$
-then there exists $T>0$, such that for all $t\in(0,T)$, uniformly for $\varepsilon\in(0,1]$,
-$\psi_{t}^{\varepsilon}=\mathrm{WP}^{\varepsilon}_{x_{t},0}+\mathcal{O}_{\mathrm{L}^{\infty}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\left(\varepsilon^{-1/4}\right)+\mathcal{O}_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\left(\varepsilon^{1/2}\right).$ (2)
-where $\mathrm{WP}^{\varepsilon}_{x_{t},0}$ is a wave packet concentrated at $(x_{t},(0,0))$, oriented along $\vec{V}_{\theta_{t}}$. Moreover, we have the following.
-  1. (i)
-    If $\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}$ is oriented along $\vec{V}_{\theta_{0}}$, then the equality (2) holds for all $t>0$ without $\mathcal{O}_{\mathrm{L}^{\infty}}\left(\varepsilon^{-1/4}\right)$.
-  2. (ii)
-    If there exists $f\in\mathscr{S}\left(\mathbb{R},\mathbb{C}\right)$ such that for all $\varepsilon\in(0,1]$ and for all $x\in\mathbb{R}^{2}$,
-    $\psi_{0}^{\varepsilon}(x)\coloneqq\frac{r_{0}^{1/4}}{\sqrt{\varepsilon}}f\left(\frac{\left(\mathrm{R}_{\theta_{0}}\left(x-x_{0}\right)\right)_{1}}{\sqrt{\varepsilon}}\right)e^{\frac{-r_{0}\left(\mathrm{R}_{\theta_{0}}\left(x-x_{0}\right)\right)_{2}^{2}}{2\varepsilon}}\vec{V}_{\theta_{0}},$
-    where $\mathrm{R}_{\theta}$ is the rotation of angle $\theta$ in $\mathbb{R}^{2}$, then uniformly for $\varepsilon\in(0,1]$, for all $(t,x)\in\mathbb{R}\times\mathbb{R}^{2}$,
-    $\psi_{t}^{\varepsilon}(x)=\frac{r_{t}^{1/4}}{\sqrt{\varepsilon}}f\left(\frac{\left(\mathrm{R}_{\theta_{t}}\left(x-x_{t}\right)\right)_{1}}{\sqrt{\varepsilon}}\right)e^{\frac{-r_{t}\left(\mathrm{R}_{\theta_{t}}\left(x-x_{t}\right)\right)_{2}^{2}}{2\varepsilon}}\vec{V}_{\theta_{t}}\ +\ \mathcal{O}\left(\varepsilon^{1/2}\langle t\rangle\right).$
-  3. (iii)
-    If $\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}$ is oriented along $\vec{V}_{\theta_{0}}^{\perp}$, then $\mathrm{WP}^{\varepsilon}_{x_{t},0}=0$.
-The initial wave packet is concentrated on the curve $\mathcal{C}\coloneqq\mathrm{E}\times\{(0,0)\}\subset\mathbb{R}^{4}$ and so is the wave packet at time $t$.
-Physically, the wave packets under consideration have no impulse.
-Let us informally describe these wave packets.
-In the point (ii), the variables $\left(\mathrm{R}_{\theta_{t}}\left(x-x_{t}\right)\right)_{1}$ and $\left(\mathrm{R}_{\theta_{t}}\left(x-x_{t}\right)\right)_{2}$ represent the tangent and the normal variable to the curve $\mathrm{E}$, respectively.
-They do not play the same role in the considered wave packet.
-In particular, for all $t\in[0,T)$, the wave packet $\mathrm{WP}^{\varepsilon}_{x_{t},0}$ have a Gaussian behavior in the normal variable and is localized near the point $(x_{t},(0,0))$ in the phase space.
-If the orientation is not along $\vec{V}_{\theta_{0}}$ or the decay with respect to the normal variable is different, then the $\mathrm{L}^{\infty}$-remaining part is not necessarily negligible in $\mathrm{L}^{2}$-norm.
-Therefore Theorem 1.1 does not give a full description of $\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}$ in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$.
-To prove this theorem, the Hamiltonian $\mathrm{H}_{\varepsilon}$ is reduced to a normal form through Fourier Integral Operator in [6].
-In the same way, we will reduce it through a change of variable to have a model problem which can be understand as if the curve $\mathrm{E}$ was a line.
-Our objective is two-fold: we aim at describing the evolution of these dispersive terms and we want to consider initial data in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$.
-1.3 Main result
-Instead of computing the asymptotics of the wave function itself, we focus on the probability densities such as the position one defined by $\left|\psi^{\varepsilon}_{t}(x)\right|_{\mathbb{C}^{2}}^{2}\mathrm{d}x$.
-Its physical interpretation is the probability of finding the particle at time $t$ in the position $x$.
-This approach generalizes the previous strategy and catches the information on the quantities that are not negligible in $\mathrm{L}^{2}$-norm.
-More precisely, we want to describe the evolution as $\varepsilon$ goes to 00 of the quadratic quantities
-$a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathbb{C}^{2,2}\right)\longmapsto\left\langle a\psi^{\varepsilon}_{t},\psi^{\varepsilon}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\coloneqq\int_{\mathbb{R}^{2}}\left\langle a(x)\psi^{\varepsilon}_{t}(x),\psi^{\varepsilon}_{t}(x)\right\rangle_{\mathbb{C}^{2}}\mathrm{d}x,$ (3)
-where $\mathbb{C}^{2,2}$ is the set of $2$ by $2$ matrices and the scalar product of $(v,w)\in\left(\mathbb{C}^{2}\right)^{2}$ is $\left\langle v,w\right\rangle_{\mathbb{C}^{2}}\coloneqq v\cdot\overline{w}$.
-For example, if $a$ is equal to a projector on a direction $\mathbb{R}\vec{V}$ for $\vec{V}\in\mathbb{R}^{2}\setminus\{0\}$ on a subspace $\Omega$, then $\left\langle a\psi^{\varepsilon}_{t},\psi^{\varepsilon}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}$ describes the probability that $\psi^{\varepsilon}_{t}$ is collinear to $\vec{V}$ at time $t$ on the subspace $\Omega$.
-We now introduce our assumptions on the function $\mathrm{m}$. Physically, the assumptions describe the interface between the two topological insulators.
-Assumption 1.2.
+We say that $\mathrm{WP}^{\varepsilon}_{x_{0},\xi_{0}}[\vec{f}]$ is oriented along $\vec{V}\in\mathbb{C}^{2}$ if there exists $f\in\mathscr{S}\left(\mathbb{R}^{2},\mathbb{C}\right)$ such that, for all $x\in\mathbb{R}^{2},\vec{f}(x)=f(x)\vec{V}$.
+Geometrical setup.
+We first introduce the geometrical setup.
+Assumption 1.1 (Transversality condition).
 The function $\mathrm{m}$ satisfies
 $\inf_{x\in\mathrm{E}}{\Big|\nabla\mathrm{m}(x)\Big|}>0.$
-To simplify the presentation and technical details, we will work with a global tubular neighborhood of the curve $\mathrm{E}$.
-To ensure the existence of such a neighborhood, we assume the following assumption.
-Actually, if we consider a neighborhood of a point of the curve $\mathrm{E}$, the local version of the following assumption will always be satisfied.
-Let $x_{0}\in\mathrm{E}$. Let us consider the parametrization $\gamma$ of the curve $\mathrm{E}$ and the vector-valued function $\nu$ such that $\gamma(0)=x_{0}$ and for all $s\in\mathbb{R}$,
-$\gamma^{\prime}(s)\coloneqq\frac{-{\nabla\mathrm{m}(\gamma(s))}^{\perp}}{\left|\nabla\mathrm{m}(\gamma(s))\right|}\;\;,\;\;\nu(s)\coloneqq\frac{\nabla\mathrm{m}(\gamma(s))}{\left|\nabla\mathrm{m}(\gamma(s))\right|},$
-where denotes the $\pi/2$-counterclockwise rotation. Then the curvature of the curve $\mathrm{E}$ at a point $\gamma(s)$ is the real number $\kappa(s)$ given by Frenet’s formula
-$\kappa(s)\coloneqq\gamma^{\prime\prime}(s)\cdot\nu(s).$
-Let us denote ${\mathrm{I}\coloneqq\left(-\frac{1}{2||\kappa||_{\infty}},\frac{1}{2||\kappa||_{\infty}}\right)}$.
+Let $x_{0}\in\mathrm{E}$. In view of Assumption 1.1, we consider an arc-length parametrization $\mathbf{t}$ of the curve $\mathrm{E}$, such that $\mathbf{t}(0)=x_{0}$, so
+$\mathbf{t}^{\prime}(s)\coloneqq\frac{{\nabla\mathrm{m}(\mathbf{t}(s))}^{\perp}}{\left|\nabla\mathrm{m}(\mathbf{t}(s))\right|},\quad s\in\mathbb{R},$
+where denotes the counterclockwise rotation by $\pi/2$. We also introduce the vector-valued function $\mathbf{n}$ defined by
+$\mathbf{n}(s)\coloneqq\frac{\nabla\mathrm{m}(\mathbf{t}(s))}{\left|\nabla\mathrm{m}(\mathbf{t}(s))\right|},\quad s\in\mathbb{R}.$
+The curvature of the curve $\mathrm{E}$ at $\mathbf{t}(s)$ is the real number $\kappa(s)$, defined by Frenet’s formula
+$\kappa(s)\coloneqq-\mathbf{t}^{\prime\prime}(s)\cdot\mathbf{n}(s).$
+For all $s\in\mathbb{R}$, we set $r(s)\coloneqq\left|\nabla\mathrm{m}(\mathbf{t}(s))\right|^{1/2}$ and ${\theta(s)\coloneqq\theta_{0}+\int_{0}^{s}\kappa(\eta)d\eta}$, where $\theta_{0}\in[0,2\pi)$ is chosen such that ${\mathbf{n}(0)=\begin{pmatrix}-\sin\theta_{0}\\ \cos\theta_{0}\end{pmatrix}}$.
+Then
+$\frac{\nabla\mathrm{m}(\mathbf{t}(s))}{\left|\nabla\mathrm{m}(\mathbf{t}(s))\right|}=\begin{pmatrix}-\sin\theta(s)\\ \cos\theta(s)\end{pmatrix}$.
+Generated edge states (Theorem 2 in Bal et al. (2021) and Theorem A.1 in Drouot (2022)).
+Let $\vec{f}\in\mathscr{S}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ be independent of $\varepsilon$. We consider the following initial condition, where for all $\varepsilon\in(0,1]$,
+$\psi_{0}^{\varepsilon}(x)\coloneqq\frac{1}{\sqrt{\varepsilon}}\vec{f}\left(\frac{x-x_{0}}{\sqrt{\varepsilon}}\right)=\mathrm{WP}^{\varepsilon}_{\mathbf{t}(0),0}\left[\vec{f}\right](x),\quad x\in\mathbb{R}^{2}.$ (1.2)
+This initial wave packet is concentrated on the curve $\mathcal{C}\coloneqq\mathrm{E}\times\{(0,0)\}\subset\mathbb{R}^{2}_{x}\times\mathbb{R}^{2}_{\xi}$, which physically means that the wave packets carry no momentum.
+Classically, this implies that they are not expected to move.
+However, if ${\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}}$ solves (1.1) with the initial condition (1.2), then there exists $T>0$, such that, for all $t\in(0,T)$, the following holds uniformly for $\varepsilon\in(0,1]$,
+$\psi_{t}^{\varepsilon}=\mathrm{WP}^{\varepsilon}_{\mathbf{t}(t),0}\left[F(t,\cdot)\vec{V}_{\theta(t)}\right]+\mathcal{O}_{\mathrm{L}^{\infty}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\left(\varepsilon^{-1/4}\right)+\mathcal{O}_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\left(\varepsilon^{1/2}\right).$ (1.3)
+where $F\in\mathscr{C}^{\infty}\left(\mathbb{R}_{t},\mathscr{S}\left(\mathbb{R}^{2},\mathbb{C}\right)\right)$ is determined by the initial data and ${\vec{V}_{\theta}\coloneqq\begin{pmatrix}e^{\frac{-\mathrm{i}\theta}{2}}\\ -e^{\frac{\mathrm{i}\theta}{2}}\end{pmatrix}}$.
+Moreover,
+  1. (i)
+    If $\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}$ is oriented along $\vec{V}_{\theta_{0}}$, then the equality (1.3) holds for all $t>0$ without the $\mathcal{O}_{\mathrm{L}^{\infty}}\left(\varepsilon^{-1/4}\right)$ term.
+  2. (ii)
+    If there exists $f\in\mathscr{S}\left(\mathbb{R},\mathbb{C}\right)$ such that, for all $\varepsilon\in(0,1]$,
+    $\psi_{0}^{\varepsilon}(x)\coloneqq\frac{r(0)^{1/2}}{\sqrt{\varepsilon}}f\left(\frac{\left(\mathrm{R}_{\theta_{0}}\left(x-x_{0}\right)\right)_{1}}{\sqrt{\varepsilon}}\right)\exp\left(\frac{-r(0)^{2}\left(\mathrm{R}_{\theta_{0}}\left(x-x_{0}\right)\right)_{2}^{2}}{2\varepsilon}\right)\vec{V}_{\theta_{0}},\quad x\in\mathbb{R}^{2},$
+    where $\mathrm{R}_{\theta}$ denotes the rotation of angle $\theta$ in $\mathbb{R}^{2}$, then, uniformly for $\varepsilon\in(0,1]$ and for all $(t,x)\in\mathbb{R}\times\mathbb{R}^{2}$,
+    $\psi_{t}^{\varepsilon}(x)=\frac{r_{t}^{1/2}}{\sqrt{\varepsilon}}f\left(\frac{\left(\mathrm{R}_{\theta_{t}}\left(x-\mathbf{t}(t)\right)\right)_{1}}{\sqrt{\varepsilon}}\right)\exp\left(\frac{-r_{t}^{2}\left(\mathrm{R}_{\theta_{t}}\left(x-\mathbf{t}(t)\right)\right)_{2}^{2}}{2\varepsilon}\right)\vec{V}_{\theta_{t}}\ +\ \mathcal{O}\left(\varepsilon^{1/2}\langle t\rangle\right).$
+    Here, the variables $\left(\mathrm{R}_{\theta_{t}}\left(x-\mathbf{t}(t)\right)\right)_{1}$ and $\left(\mathrm{R}_{\theta_{t}}\left(x-\mathbf{t}(t)\right)\right)_{2}$ represent tangent and normal coordinates to the curve $\mathrm{E}$, respectively.
+They play different roles in the wave packet.
+In particular, for all $t\in[0,T)$, the wave packet part exhibits Gaussian behavior in the normal coordinate and remains localized near the point $(\mathbf{t}(t),(0,0))$ in the phase space.
+  3. (iii)
+    If $\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}$ is oriented along $\vec{V}_{\theta_{0}}^{\perp}$, then $F=0$.
+This implies that if the orientation is not aligned with $\vec{V}_{\theta_{0}}$, or if the decay in the normal variable is orthogonal in $\mathrm{L}^{2}$ to a Gaussian profile, then the $\mathrm{L}^{\infty}$ remainder is not necessarily negligible in $\mathrm{L}^{2}$-norm.
+Furthermore, the terms in the $\mathrm{L}^{\infty}$ remainder are not wave packets.
+Figure 2: Example of geometric setup.
+In Drouot (2022), the $\mathrm{L}^{\infty}$ remainder is described using a WKB expansion, for initial data that are wave packets, where the Hamiltonian is reduced to a normal form via a Fourier Integral Operator.
+In Bal (2024), a parametrix is constructed for wave packets used as initial conditions.
+Our objective is twofold: we aim to describe the evolution of dispersive terms, and we also consider initial data in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$, not necessarily coherent states as in (1.2).
+We will focus on computing the Wigner measure of the solution to (1.1) for initial data in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$.
+1.2 Wigner transform and semiclassical measure
+Wigner transform.
+In the present paper, instead of computing the asymptotics of the wave function itself, as in (1.3), we focus on probability densities, in particular the position density defined by $\left|\psi^{\varepsilon}_{t}(x)\right|_{\mathbb{C}^{2}}^{2}\mathrm{d}x$.
+Its physical interpretation is the probability of finding the particle at time $t$ in position $x$.
+This approach generalizes the previous strategy and captures information on all quantities that are not negligible in $\mathrm{L}^{2}$ norm.
+Since our problem requires an analysis in phase space, we go one step further and consider the Wigner transform of $\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}$ and its weak limits in the space of distributions, known as Wigner measures.
+These were introduced by Wigner in Wigner (1932) and Wigner (1971).
+The Wigner transform of ${(\psi_{t}^{\varepsilon})_{\varepsilon>0}}$ is defined as follows,
+$\mathrm{W}^{\varepsilon}\left[\psi_{t}^{\varepsilon}\right](x,\xi)\coloneqq\int_{\mathbb{R}^{2}}e^{\mathrm{i}\xi\cdot y}\psi_{t}^{\varepsilon}\left(x+\frac{\varepsilon y}{2}\right)\otimes\overline{\psi_{t}^{\varepsilon}\left(x-\frac{\varepsilon y}{2}\right)}\frac{\mathrm{d}y}{(2\pi)^{2}},\quad(x,\xi)\in\mathbb{R}^{2}\times\mathbb{R}^{2},$
+where $v\otimes w$ denotes the $2\times 2$ matrix $v\,^{t}w$ for $(v,w)\in\mathbb{C}^{2}\times\mathbb{C}^{2}$.
+The Wigner transform $\mathrm{W}^{\varepsilon}\left[\psi_{t}^{\varepsilon}\right]$ belongs to $\mathrm{L}^{2}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)\cap\mathscr{C}^{0}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)$ and takes values in the set of Hermitian matrices.
+Moreover, the position density can be formally recovered through the relation
+$\int_{\mathbb{R}^{2}}{\rm tr}\mathrm{W}^{\varepsilon}\left[\psi_{t}^{\varepsilon}\right](x,\xi)\mathrm{d}\xi=\left|\psi_{t}^{\varepsilon}(x)\right|_{\mathbb{C}^{2}}^{2}.$
+It is also possible to recover the momentum density by integrating over position rather than momentum.
+Distributional Properties.
+The function $\mathrm{W}^{\varepsilon}\left[\psi_{t}^{\varepsilon}\right]$ is a distribution on the phase space $\mathbb{R}^{2}_{x}\times\mathbb{R}^{2}_{\xi}$. In fact,
+$\mathrm{I}^{\varepsilon}_{\psi_{t}^{\varepsilon}}(a)\coloneqq\int_{\mathbb{R}^{4}}{\rm tr}\Big(a(x,\xi)\mathrm{W}^{\varepsilon}\left[\psi_{t}^{\varepsilon}\right](x,\xi)\Big)\mathrm{d}x\mathrm{d}\xi=\Big\langle\mathrm{Op}_{\varepsilon}(a)\psi_{t}^{\varepsilon},\psi_{t}^{\varepsilon}\Big\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)},\quad a\in\mathscr{C}^{\infty}_{c}(\mathbb{R}^{4},\mathbb{C}^{2,2})$ (1.4)
+where $\mathrm{Op}_{\varepsilon}(a)$ denotes the semiclassical pseudodifferential operator with symbol $a$, obtained via Weyl quantization. That is,
+$\mathrm{Op}_{\varepsilon}(a)\psi_{t}^{\varepsilon}(x)\coloneqq\int_{\mathbb{R}^{4}}a\left(\frac{x+y}{2},\xi\right)\psi_{t}^{\varepsilon}(y)e^{\mathrm{i}\frac{\xi\cdot(x-y)}{\varepsilon}}\frac{\mathrm{d}y\mathrm{d}\xi}{(2\pi\varepsilon)^{2}},\quad x\in\mathbb{R}^{2}.$
+The Calderón-Vaillancourt Theorem Calderón and Vaillancourt (1971) asserts that the family of operators $\left(\mathrm{Op}_{\varepsilon}(a)\right)_{\varepsilon>0}$ is uniformly bounded with respect to $\varepsilon$ in $\mathcal{L}\left(\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)$: there exists constants $C>0$ and $N\in\mathbb{N}$ such that, for all $a=(a_{i,j})_{1\leqslant i,j\leqslant 2}\in\mathscr{C}^{\infty}_{c}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)$ and all $f\in\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$,
+$\left\lVert\mathrm{Op}_{\varepsilon}(a)f\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\leqslant C\,\max_{1\leqslant i,j\leqslant 2}\,\max_{\left|\alpha\right|+\left|\beta\right|\leqslant N}\left\lVert\partial^{\alpha}_{x}\partial^{\beta}_{\xi}a_{i,j}\right\rVert_{\mathrm{L}^{\infty}\left(\mathbb{R}^{4}\right)}\left\lVert f\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}.$ (1.5)
+However, the Wigner transform is not a positive distribution.
+Nevertheless, its weak limits are positive measures.
+This follows from the weak Gårding inequality.
+Adapting (Kammerer, 2023, Proposition 2.13) to the matrix-valued case, we have: for all non-negative $a\in\mathscr{C}^{\infty}_{c}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)$ and for all $\delta>0$, there exists a constant $C_{\delta}>0$ such that,
+$\left\langle\mathrm{Op}_{\varepsilon}(a)f,f\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\geqslant-\left(\delta+C_{\delta}\varepsilon\right)\left\lVert f\right\rVert^{2}_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)},\quad f\in\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right).$
+Time-averaged Wigner measure.
+To account for the time dependence of the family $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$, we consider time-averaged Wigner measures.
+More precisely, we test the Wigner transform $\mathrm{W}^{\varepsilon}[\psi^{\varepsilon}_{t}]$ against separated time-dependent observables: for all $\Xi\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R},\mathbb{C}\right)$, for all $a\in\mathscr{C}^{\infty}_{c}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)$,
+$\mathcal{I}^{\varepsilon}_{\psi_{t}^{\varepsilon}}(\Xi,a)\coloneqq\int_{\mathbb{R}}\int_{\mathbb{R}^{4}}\Xi(t){\rm tr}_{\mathbb{C}^{2}}\Big(a(x,\xi)\mathrm{W}^{\varepsilon}[\psi^{\varepsilon}_{t}](x,\xi)\Big)\mathrm{d}x\mathrm{d}\xi\mathrm{d}t\;=\;\int_{\mathbb{R}}\Xi(t)\left\langle\mathrm{Op}_{\varepsilon}(a)\psi^{\varepsilon}_{t},\psi^{\varepsilon}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\mathrm{d}t.$
+Semiclassical measure.
+Our goal is to describe the solution $\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}$ to equation (1.1) for initial data in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$.
+At the very least, we can characterize the limit as $\varepsilon$ goes to 0 by constructing semiclassical measures.
+More precisely, there exists a vanishing sequence of positive numbers $\left(\varepsilon_{k}\right)_{k\in\mathbb{N}}$ and a positive matrix of Radon measures $\mu$ on $\mathbb{R}_{t}\times\mathbb{R}^{2}_{x}\times\mathbb{R}^{2}_{\xi}$, such that,
+$\mathcal{I}^{\varepsilon_{k}}_{\psi_{t}^{\varepsilon_{k}}}(\Xi,a)\;\underset{k\to+\infty}{\longrightarrow}\;\int_{\mathbb{R}^{5}}\Xi(t){\rm tr}\Big(a(x,\xi)\mathrm{d}\mu(t,x,\xi)\Big),\quad(a,\Xi)\in\mathscr{C}^{\infty}_{c}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)\times\mathscr{C}_{c}^{\infty}(\mathbb{R},\mathbb{C}).$ (1.6)
+The measure $\mu$ is called the time-averaged Wigner measure of the family $\left(\psi_{t}^{\varepsilon_{k}}\right)_{k\in\mathbb{N}}$.
+A proof of existence is recalled in Appendix A.2 for families taking values in a separable Hilbert space, as this will be a main topic of Section 2.3 (while observables and quantization will be precised in Section 2.2).
+Moreover, thanks to the control
+$\left|\int_{\mathbb{R}^{5}}\Xi(t)a(x,\xi)\mathrm{d}\mu(t,x,\xi)\right|\leqslant C\ \left\lVert\Xi\right\rVert_{\mathrm{L}^{1}(\mathbb{R})}\left\lVert\mathrm{Op}_{\varepsilon}(a)\right\rVert_{\mathcal{L}\left(\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)},$
+we can write $\mathrm{d}\mu(t,x,\xi)=\mathrm{d}\mu_{t}(x,\xi)\mathrm{d}t$.
+In other words, the measure $\mu$ is absolutely continuous in time with respect to the Lebesgue measure.
+Our goal is to describe the propagation of these measures.
+1.3 Bulk propagation of the semiclassical measure
+It is known from Gérard et al.
+(1997); Kammerer (2023) that any time-averaged Wigner measure of the family $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$, outside the curve $\mathcal{C}$, is completely determined by the corresponding measures of the initial data.
+Let us first introduce some notations.
+Let us consider the symbol of the Hamiltonian in the Weyl quantization defined as
+$\mathrm{H}(x,\xi)\coloneqq\begin{pmatrix}\mathrm{m}(x)&\xi_{1}-\mathrm{i}\xi_{2}\\ \xi_{1}+\mathrm{i}\xi_{2}&-\mathrm{m}(x)\end{pmatrix},\quad(x,\xi)\in\mathbb{R}^{4}.$
+Notably, ${\det\mathrm{H}(x,\xi)=-\mathrm{m}(x)^{2}-|\xi|^{2}}$ so that $\mathcal{C}=\{(x,\xi)\in\mathbb{R}^{4}\mid\det\mathrm{H}(x,\xi)=0\}$.
+Let $(x,\xi)\in\mathbb{R}^{4}\setminus\mathcal{C}$.
+Then, the matrix $\mathrm{H}(x,\xi)$ has two distinct eigenvalues, denoted by $\lambda_{+}(x,\xi)$ and $\lambda_{-}(x,\xi)$ defined by
+$\lambda_{+}(x,\xi)\ =\ -\lambda_{-}(x,\xi)\ \coloneqq\ \sqrt{\mathrm{m}(x)^{2}+|\xi|^{2}}\ >\ 0,$
+and the projectors onto the corresponding eigenspaces are given by
+$\Pi_{+}(x,\xi)\coloneqq\mathrm{Id}-\frac{1}{\lambda_{+}(x,\xi)}\mathrm{H}(x,\xi),\quad\Pi_{-}(x,\xi)\coloneqq\mathrm{Id}-\frac{1}{\lambda_{-}(x,\xi)}\mathrm{H}(x,\xi).$
+We recall the definition of the Poisson bracket
+$\left\{f,g\right\}\coloneqq\nabla_{\xi}f\nabla_{x}g-\nabla_{x}f\nabla_{\xi}g\;,\;(f,g)\in\mathscr{C}^{1}\left(\mathbb{R}^{2}_{x}\times\mathbb{R}^{2}_{\xi}\right)^{2}.$ (1.7)
+Following (Gérard et al., 1997, section 6), any semiclassical measure $\mu$ can be decomposed outside $\mathcal{C}$.
+Lemma 1.2 (Evolution of semiclassical measure outside $\mathcal{C}$).
+Let $\mu_{t}$ be a semiclassical measure of the family $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ associated with the sequence $(\varepsilon_{n})_{n\in\mathbb{N}}$. Then,
+$\mathrm{d}\mu_{t}(x,\xi)=\Pi_{+}(x,\xi)\mathrm{d}\mu_{t}^{+}(x,\xi)\mathrm{d}t+\Pi_{-}(x,\xi)\mathrm{d}\mu_{t}^{-}(x,\xi)\mathrm{d}t,\quad(x,\xi)\in\mathbb{R}^{4}\setminus\mathcal{C},$
+where the scalar measures $\mu_{t}^{\pm}$ satisfy
+$\partial_{t}\mu_{t}^{\pm}=\left\{\mu_{t}^{\pm},\lambda_{\pm}\right\},\quad\mu_{0}^{\pm}=\mathrm{tr}\Big(\Pi_{\pm}\mu_{0}\Big),\\$
+and $\mu_{0}$ is the semiclassical measure of the family $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ associated with the sequence $(\varepsilon_{n})_{n\in\mathbb{N}}$.
+This theorem implies that, for all initial data $\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}$ microlocalized around $(x_{0},\xi_{0})\notin\mathcal{C}$, the semiclassical measures evolve along the Hamiltonian trajectories associated with the functions $\lambda_{+}$ and $\lambda_{-}$.
+These trajectories lie on the hypersurface $\det\mathrm{H}(x,\xi)=\det\mathrm{H}(x_{0},\xi_{0})$.
+In other words, if the initial data are localized far from the curve $\mathcal{C}$, then the semiclassical measure never approaches the curve $\mathcal{C}$ in phase space.
+Our aim in this paper is to describe the evolution of the measure $\mu$ above the curve $\mathcal{C}$.
+As previously noted in Bal et al.
+(2021) and Drouot (2022), the concentration of $\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}$ is a two-scale problem involving the scale $\sqrt{\varepsilon}$.
+Therefore, we will adopt a two-scale Wigner approach at scale $\sqrt{\varepsilon}$, as initiated in Kammerer and Gérard (2002) (see also Kammerer (2023); Nier (1996); Miller (1996)).
+To achieve this objective, we work in a neighborhood of $\mathrm{E}$.
+1.4 Two-scale analysis above the interface
+Geometrical setup (2).
+To simplify the presentation, we will impose an assumption on $\mathrm{m}$ to provide a global tubular neighborhood of $\mathrm{E}$. Let us denote
+[…]
+We assume $\mathrm{E}$ is non-compact. The following assumption ensures the existence of a global tubular neighborhood of size $|\mathrm{I}|$.
 Assumption 1.3.
 The map
-$\begin{matrix}\Phi:&\mathbb{R}\times\mathrm{I}&\longrightarrow&\mathbb{R}^{2}\\ &(s,y)&\longmapsto&\gamma(s)+y\nu(s)\end{matrix}$ (4)
-is a global diffeomorphism from ×$\mathbb{R}\times\mathrm{I}$ to
-$\Omega\coloneqq\left\{\gamma(s)+y\nu(s)|\ (s,y)\in\mathbb{R}\times\mathrm{I}\right\}.$
+$\begin{matrix}\Phi:&\mathbb{R}\times\mathrm{I}&\longrightarrow&\mathbb{R}^{2}\\ &(s,y)&\longmapsto&\mathbf{t}(s)+y\mathbf{n}(s)\end{matrix}$ (1.8)
+is a global diffeomorphism from $\mathbb{R}\times\mathrm{I}$ to
+$\Omega\coloneqq\left\{\mathbf{t}(s)+y\mathbf{n}(s)|\ (s,y)\in\mathbb{R}\times\mathrm{I}\right\}.$
+The set $\Omega$ is a global tubular neighborhood of the curve $\mathrm{E}$ Hirsch (1976). Actually, this assumption is always satisfied locally.
+Let us assume Assumption 1.1 and Assumption 1.3.
+We will work in the normal geodesic coordinates, denoted $(s,y)$, defined by (1.8), and denote $(\sigma,\eta)$ as the dual variables of $(s,y)$, which play the role of $\xi$.
+In these new coordinates, the curve $\mathcal{C}$ is reduced to $\left\{y=\sigma=\eta=0\right\}$.
+Therefore, the measure $\mu$ over $\mathcal{C}$ can be rewritten as
+$\mathrm{d}\mu\mathrm{d}t\mathds{1}_{\mathcal{C}}=\Phi^{*}\Big(\mathrm{U}_{\theta(s)}^{\dagger}\mathrm{d}\rho_{t}\mathrm{U}_{\theta(s)}\Big)\mathrm{d}t$ (1.9)
+where $\rho_{t}$ is a semiclassical measure on $\mathbb{R}_{s}$ and the unitary operator $\mathrm{U}_{\theta}$ is the multiplication by a matrix defined by
+$\mathrm{U}_{\theta}\coloneqq\frac{1}{\sqrt{2}}\begin{pmatrix}e^{\mathrm{i}\frac{\theta}{2}}&-e^{-\mathrm{i}\frac{\theta}{2}}\\ e^{\mathrm{i}\frac{\theta}{2}}&e^{-\mathrm{i}\frac{\theta}{2}}\end{pmatrix},\quad\theta\in\mathbb{R}.$ (1.10)
+Two-scale quantization.
+Based on the geometric assumptions on $\mathcal{C}$, we introduce a two-scale quantization procedure.
+We follow a two-scale Wigner approach at scale $\sqrt{\varepsilon}$, as initiated in Kammerer and Gérard (2002) (see also Kammerer (2023), Nier (1996) and Miller (1996)).
+The approach using tubular coordinates is inspired by Chabu et al.
+(2022).
+We say that $a\in\mathscr{C}^{\infty}\Big(\mathbb{R}^{4}\times\mathbb{R}^{3},\mathbb{C}^{2,2}\Big)$ is a two-scale observable if, and only if,
+  1. (i)
+    there exists a compact set $K\subset\mathbb{R}^{4}$ such that for all $z\in\mathbb{R}^{3}$, $a(\cdot,\cdot,z)$ is supported in $K$,
+  2. (ii)
+    there exists a constant $R_{0}>0$ and $a_{\infty}\in\mathscr{C}^{\infty}\Big(\mathbb{R}^{4}\times\mathbb{S}^{2},\mathbb{C}^{2,2}\Big)$, such that for all $|z|>R_{0}$, $a(\cdot,\cdot,z)=a_{\infty}\left(\cdot,\cdot,\frac{z}{|z|}\right)$.
+We denote by $\mathcal{A}$ the set of two-scale observables. We define a second-scale quantization near the interface $\mathcal{C}$ for these observables by
+$\begin{array}[]{ccccc}\mathbf{op}^{\mathcal{C}}_{\varepsilon}&:&\mathcal{A}&\longrightarrow&\mathcal{L}\big(\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\big)\\ &&a&\longmapsto&\mathrm{Op}_{1}\left(a\left(s,y,\varepsilon\sigma,\varepsilon\eta,\frac{y}{\sqrt{\varepsilon}},\sqrt{\varepsilon}\sigma,\sqrt{\varepsilon}\eta\right)\right)\end{array}$ (1.11)
+The quantization map $\mathbf{op}^{\mathcal{C}}_{\varepsilon}$ associates to $a$ a bounded operator on $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ according to the Calderón-Vaillancourt Theorem Calderón and Vaillancourt (1971) as we will see in Section 2.2.
+We can identify our observables in $\mathscr{C}^{\infty}_{c}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)$ as a subset of $\mathcal{A}$.
+With an abuse of notation, for all $\in\mathscr{C}^{\infty}_{c}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right), aa$ belongs to $\mathcal{A}$ with $a_{\infty}=a$ and moreover
+$\mathbf{op}^{\mathcal{C}}_{\varepsilon}(a)=\mathrm{Op}_{\varepsilon}(a).$
+By analyzing the limits (1.6) for observables supported on $\Omega$, we can replace $\mathrm{Op}_{\varepsilon}(a)$ by $\mathbf{op}^{\mathcal{C}}_{\varepsilon}(a)$, which will gives us (according to Section 2) the following decomposition where $\mathbb{S}^{2}$ denotes the unit sphere of $\mathbb{R}^{3}$, for all $(s,\sigma)\in\mathbb{R}^{2}$, $a^{W}(s;0;y,\sigma,\mathrm{D}_{y})$ denotes the bounded operator on $\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)$ obtained by the Weyl quantization of $(y,\eta)\mapsto a(s;0;y,\sigma,\eta)$ and $\mathrm{Tr}$ denotes the trace in $\mathcal{L}\left(\mathrm{L}^{2}\left(\mathbb{R},\mathbb{C}^{2}\right)\right)$.
+Theorem 1.4 (Time-averaged two-scale semiclassical measure above the interface).
+Let ${(\psi^{\varepsilon}_{t})_{\varepsilon>0}}$ be a solution to (1.1) with normalized initial condition in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$.
+Then, there exist a vanishing sequence of positive numbers ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ and
+  1. (i)
+    Two measurable maps $t\mapsto\mathrm{d}\nu^{t}_{\infty}$ and $t\mapsto\mathrm{d}\nu^{t}$, valued in the set of non-negative scalar measures on $\mathcal{C}\times\mathbb{S}^{2}$ and $\mathbb{R}^{2}$, respectively,
+  2. (ii)
+    A $\mathrm{d}\nu^{t}_{\infty}\otimes\mathrm{d}t$-measurable map $(t,s,\omega)\mapsto M^{t}_{\infty}(s,\omega)\in\mathbb{C}^{2,2}$, valued in the set of Hermitian trace $1$ positive matrices,
+  3. (iii)
+    A $\mathrm{d}\nu^{t}(s,\sigma)\otimes\mathrm{d}t$ measurable map $(t,s,\sigma)\mapsto M^{t}(s,\sigma)$, valued in trace-one positive operators on $\mathrm{L}^{2}\left(\mathbb{R},\mathbb{C}^{2}\right)$,
+such that the semiclassical measure $\rho_{t}$ defined in (1.9) associated with ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ satisfy, for all $s\in\mathbb{R}$,
+$\rho_{t}(s)=\int_{\mathbb{S}^{2}}\mathrm{tr}\Big(a_{\infty}(s;0;\omega)M_{\infty}^{t}(s,\omega)\Big)\mathrm{d}\nu_{\infty}^{t}(s,\omega)+\int_{\mathbb{R}}\mathrm{Tr}\Big(a^{W}(s;0;y,\sigma,\mathrm{D}_{y})M^{t}(s,\sigma)\Big)\mathrm{d}\nu^{t}(s,\sigma).$
+This Theorem is a direct consequence of two results.
+The first one is the existence of two-scale semiclassical measures, which is Theorem 2.4 and the second one is the link between semiclassical measures above the interface, which is Lemma 2.3.
+Remark (Geometrical point of view).
+In the parametrization $(s,y)$ of $\mathbb{R}\times\mathrm{I}$, the variables $(s,y,\sigma,\eta)$ parametrize $\mathrm{T}^{*}\mathbb{R}^{2}$.
+Moreover, the variable $s$ parametrizes the curve $\mathrm{E}\subset\mathbb{R}^{2}$.
+The variable $y$ parametrizes the normal fiber to the curve $\mathrm{E}$ at the point $\mathbf{t}(s)$.
+In the computation of the trace $\mathrm{Tr}$, the space $\mathrm{L}^{2}(\mathbb{R}_{y},\mathbb{C}^{2})$ can be understood as the space $\mathrm{L}^{2}(N_{\mathbf{t}(s)}\mathrm{E},\mathbb{C}^{2})$, where $N_{\mathbf{t}(s)}\mathrm{E}$ denotes the normal bundle of the curve $\mathrm{E}$ at the point $\mathbf{t}(s)$.
+The variable $\sigma$ denotes the variable of the fiber of the cotangent bundle $\mathrm{T}^{*}\mathrm{E}$ over $\mathrm{E}$.
+Specifically, for $\mathbf{t}(s)\in\mathrm{E}$ and $\mathrm{T}_{s}\mathrm{E}$, the tangent to $\mathrm{E}$ at $s$, the fiber of the cotangent bundle $\mathrm{T}^{*}\mathrm{E}$ over $\mathbf{t}(s)$ is $(\mathrm{T}_{s}\mathrm{E})^{*}\sim\mathbb{R}$, which is one-dimensional space.
+The measure $\nu^{t}$ defines a measure on $\mathrm{T}^{*}\mathrm{E}$, the cotangent bundle of the curve $\mathcal{E}$.
+Since $\mathcal{C}=\{y=\eta=\sigma=0\}\subset\mathrm{T}^{*}\mathbb{R}^{2}$, the set $\mathcal{C}\times\mathbb{S}^{2}$ coincides with the spherical normal bundle of $\mathcal{C}$ denoted by $\mathrm{S}\mathcal{C}$.
+The fiber of $\mathrm{S}\mathcal{C}$ over a point $\rho\in\mathcal{C}$ is defined as follows.
+Let $\mathrm{T}_{\rho}\mathcal{C}$ be the tangent space to $\mathcal{C}$ at $\rho$, and define $\mathrm{N}_{\rho}\mathcal{C}$ at the subspace of $\left(\mathrm{T}_{\rho}\left(\mathrm{T}^{*}\mathbb{R}^{2}\right)\right)^{*}$ consisting of linear forms on $\mathrm{T}_{\rho}\left(\mathrm{T}^{*}\mathbb{R}^{2}\right)$ that vanish on $\mathrm{T}_{\rho}\mathcal{C}$.
+Then $\mathrm{S}_{\rho}\mathcal{C}$ is obtained by taking the quotient of $\mathrm{N}_{\rho}\mathcal{C}$ by the action of $\mathbb{R}_{\geqslant 0}$ by homotheties.
+Definition (Two-scale semiclassical measure).
+We denote by $M\mathrm{d}\nu$ the pair $\left(M,\nu\right)$, where $M$ is a positive trace-one operator that is $\mathrm{d}\nu$-measurable, and $\nu$ is a non-negative measure.
+We denote by $M_{\infty}\mathrm{d}\nu_{\infty}$ the pair $\left(M_{\infty},\nu_{\infty}\right)$ with $M_{\infty}$ a Hermitian trace one positive matrix $\mathrm{d}\nu_{\infty}$-measurable and $\nu_{\infty}$ a non-negative measure.
+We call two-scale semiclassical measures, associated with the concentration of the family ${\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}}$ near the curve $\mathcal{C}$ at scales between $\sqrt{\varepsilon}$ and $\delta$, and at scale $\sqrt{\varepsilon}$, respectively, the pair ${\left(M_{\infty}\mathrm{d}\nu_{\infty},M\mathrm{d}\nu\right)}$.
+We say that $M\mathrm{d}\nu$ is a two-scale semiclassical measure at finite distance, while $M_{\infty}\mathrm{d}\nu_{\infty}$ is a two-scale semiclassical measure at infinity.
+Figure 3: Illustration of two-scale separation of space before passing to limit.
+1.5 Main result : edge propagation of the semiclassical measure
+Let us specify the effective Hamiltonians associated with the different concentration regimes.
+As in Lemma 1.2, the eigenvalues and eigenprojectors of these Hamiltonians will play a crucial role in the propagation.
+Concentration at finite distance.
+Let us consider the symbol of the principal term of the Hamiltonian, depending only on the tangential variable $s$, while retaining its operator character with respect to the normal variable $y$; namely, the operator acting on $\mathrm{L}^{2}\left(\mathbb{R}_{s},\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)\right)$ of symbol
+$\mathrm{T}_{\mathrm{E}}(s,\sigma)\coloneqq\begin{pmatrix}\sigma&y|\nabla\mathrm{m}(\mathbf{t}(s))|-\mathrm{i}\mathrm{D}_{y}\\ y|\nabla\mathrm{m}(\mathbf{t}(s))|+\mathrm{i}\mathrm{D}_{y}&-\sigma\end{pmatrix},\quad(s,\sigma)\in\mathbb{R}^{2}.$ (1.12)
+A complete spectral analysis of this operator is provided in Appendix B.
+Its eigenvectors depend on the Hermite functions $\left(\mathfrak{h}_{m}\right)_{m\in\mathbb{N}_{\geqslant 0}}$, which form an Hilbertian basis of $\mathrm{L}^{2}\left(\mathbb{R},\mathbb{C}\right)$ and satisfy, for all $y\in\mathbb{R}$ and all $m\in\mathbb{N}_{\geqslant 0}$,
+$\frac{\mathrm{d}^{2}\mathfrak{h}_{m}}{\mathrm{d}y^{2}}(y)+y^{2}\mathfrak{h}_{m}(y)=(2m+1)\mathfrak{h}_{m}(y).$ (1.13)
+Let us consider, for all $(s,y,\sigma)\in\mathbb{R}^{3}$,
+$g_{0}^{s}(y)\coloneqq\frac{1}{\sqrt{r(s)}}\begin{pmatrix}\mathfrak{h}_{0}\big(r(s)y\big)\\ 0\end{pmatrix}$,\;\; $\;\;g_{n}^{s,\sigma}(y)\coloneqq\frac{1}{\sqrt{2r(s)}}\sqrt{1-\frac{\sigma}{\lambda_{n}(s,\sigma)}}\begin{pmatrix}\frac{\sqrt{2|n|}r(s)}{\lambda_{n}(s,\sigma)-\sigma}\mathfrak{h}_{|n|}\big(r(s)y\big)\\ \mathfrak{h}_{|n|-1}\big(r(s)y\big)\end{pmatrix}$, (1.14)
+$\Pi_{0}(s)\coloneqq g_{0}^{s}\otimes g_{0}^{s},\;\;$ $\;\;\Pi_{n}(s,\sigma)\coloneqq g_{n}^{s,\sigma}\otimes g_{n}^{s,\sigma},$ (1.15)
+where
+$\lambda_{0}(\sigma)\coloneqq\sigma,\;\;\;\;\lambda_{n}(s,\sigma)\coloneqq\mathrm{sgn}(n)\sqrt{\sigma^{2}+2|n||\nabla\mathrm{m}(\mathbf{t}(s))|}.$ (1.16)
+Although the functions $g_{0}^{s},\lambda_{0}$ and $\Pi_{0}$ do not depend on $\sigma$, we will also denote them by $g_{0}^{s,\sigma},\lambda_{0}(s,\sigma)$ and $\Pi_{0}(s,\sigma)$.
+Since we retain the operator character in the variable $y$, we are naturally led to consider semiclassical measure on the separable Hilbert space $\mathcal{H}=\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)$.
+Concentration at infinity.
+We consider the symbol of the Hamiltonian corresponding to the second-scale quantization (1.11) at infinity, defined by
+$\mathrm{T}^{\infty}_{\mathrm{E}}(s,z)\coloneqq\begin{pmatrix}z_{\sigma}&z_{y}r(s)^{2}-\mathrm{i}z_{\eta}\\ z_{y}r(s)^{2}+\mathrm{i}z_{\eta}&-z_{\sigma}\end{pmatrix},\quad(s,y)\in\mathbb{R}^{2},\ z\coloneqq(z_{y},z_{\sigma},z_{\eta})\in\mathbb{R}^{3},$ (1.17)
+which has two eigenprojectors
+$\Pi_{\pm}^{\infty}(s,z)\coloneqq\frac{1}{2}\mathrm{Id}\pm\frac{1}{2\sqrt{z_{\sigma}^{2}+z_{y}^{2}r(s)^{4}+z_{\eta}^{2}}}\begin{pmatrix}z_{\sigma}&z_{y}r(s)^{2}-\mathrm{i}z_{\eta}\\ z_{y}r(s)^{2}+\mathrm{i}z_{\eta}&-z_{\sigma}\end{pmatrix}$.
+(1.18)
+Main result.
+We now describe the semiclassical measures of Theorem 1.4 for solutions of equation (1.1).
+Their semiclassical measures and their evolution outside $\mathcal{C}$ are already well understood.
+Therefore, we focus on the two-scale semiclassical measures.
+By applying a method similar to that of Gérard et al.
+(1997) to the operators $\mathrm{T}_{\mathrm{E}}$ and $\mathrm{T}^{\infty}_{\mathrm{E}}$ in normal geodesic coordinates, together with a two-scale approach, we obtain the following main result.
+Theorem 1.5 (Propagation of the two-scale semiclassical measures).
+Assume Assumption 1.1 and Assumption 1.3.
+Let ${\left(M^{t}_{\infty}\mathrm{d}\nu^{t}_{\infty},M^{t}\mathrm{d}\nu^{t}\right)}$ be a pair of two-scale semiclassical measures associated with the family $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$, solution to (1.1) with normalized initial condition.
+Then the following properties hold.
+  1. 1.
+    There exists a family of non-negative scalar Radon measures $\left(\nu_{n}^{t}\right)_{n\in\mathbb{Z}}$ such that
+    $M^{t}\mathrm{d}\nu^{t}=\sum_{n\in\mathbb{Z}}\Pi_{n}\mathrm{d}\nu_{n}^{t}.$ (1.19)
+    Moreover, for all $n\in\mathbb{Z}$, $\nu_{n}^{t}$ satisfies the following equation
+    $\left\{\begin{array}[]{rll}\partial_{t}\nu_{n}^{t}&=&\left\{\nu_{n}^{t},\lambda_{n}\right\}\\ \nu_{n}^{t}|_{t=0}&=&\nu_{n,0}\end{array}\right$., (1.20)
+    where $\nu_{n,0}=\mathrm{tr}\Big(\Pi_{n}M^{0}\Big)\nu^{0}$, and $M^{0}\mathrm{d}\nu^{0}$ is a two-scale semiclassical measure associated with the initial data.
+  2. 2.
+    There exists a family of non-negative scalar Radon measures $\nu^{t}_{\pm}$ such that
+    $M^{t}_{\infty}\mathrm{d}\nu^{t}_{\infty}=\Pi_{+}^{\infty}\mathrm{d}\nu^{t}_{+}+\Pi_{-}^{\infty}\mathrm{d}\nu^{t}_{-}.$ (1.21)
+    Moreover, $\nu_{\pm}^{t}$ satisfies the invariance equation $\mathrm{div}_{\omega}\left(\vec{V}^{\infty}(s,\omega)\nu_{\pm}^{t}(s,\omega)\right)=0$, with
+    $\vec{V}^{\infty}(s,\omega)\coloneqq\begin{pmatrix}\left(\omega_{y}\omega_{\eta}\left(1-r(s)^{4}\right)-1\right)\omega_{\eta}\\ \left(1-r(s)^{4}\right)\omega_{y}\omega_{\eta}\omega_{\sigma}\\ \left(\omega_{y}\omega_{\eta}\left(1-r(s)^{4}\right)+r(s)^{4}\right)\omega_{y}\end{pmatrix},\quad s\in\mathbb{R},\quad\omega\coloneqq\left(\omega_{y},\omega_{\sigma},\omega_{\eta}\right)\in\mathbb{S}^{2}.$ (1.22)
+Figure 4: Illustration of the propagation of semiclassical measures for general initial data.
+Remark (Continuity of the two-scale semiclassical measure at finite distance).
+From the evolution (1.20), it follows that the map $t\mapsto M^{t}\mathrm{d}\nu^{t}$ is continuous and entirely determined by the evolution equation together with the initial data.
+Remark (Value at $t=0$ of the two-scale semiclassical measure at finite distance).
+The value of $M^{t}\mathrm{d}\nu^{t}$ at $t=0$ is not necessarily the two-scale semiclassical measure at finite distance of the initial condition.
+As we will see in the proof, $M^{t}(s,\sigma)$ commutes with the operator $\mathrm{T}_{\mathrm{E}}(s,\sigma)$ due to the PDE satisfied by the solution near the interface.
+Nevertheless, there exist initial conditions whose two-scale semiclassical measure at finite distance does not commute with $\mathrm{T}_{\mathrm{E}}(s,\sigma)$.
+An example of this will be discussed in Remark 3.2.
+Remark.
+Retaining the notation of the previous theorem, the measure $\nu_{0}^{t}$ corresponds to the wave packet associated with the solution studied in Bal et al.
+(2021).
+Similarly, the measures $\left(\nu_{n}^{t}\right)_{n\in\mathbb{Z}_{\neq 0}}$ correspond to the $\mathrm{L}^{\infty}$-remainder part of the solution in Drouot (2022).
+In particular, the mass of the edge mode is given by
+$\int_{\mathbb{R}^{2}}\mathrm{d}\nu_{0,0}(s,\sigma).$
+1.6 Applications
+Our main result describes the evolution of the semiclassical measure for arbitrary initial data, without any restriction on its structure.
+A direct consequence is the following asymptotic description of the position density of the solution to equation (1.1) for initial data given by wave packets concentrating on the curve $\mathcal{C}$.
+Corollary 1.6 (Evolution consequence).
+Retaining the notations of Theorem 1.5, if
+$\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\mathrm{tr}\Big(\mathrm{d}\mu_{0}(x,\xi)\Big)+\int_{\mathbb{R}^{2}}\mathrm{Tr}\Big(M^{0}(s,\sigma)\Big)\mathrm{d}\nu^{0}(s,\sigma)=1,$
+with $\mu_{0}$ a semiclassical measure of $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ associated with the same vanishing sequence $\left(\varepsilon_{n}\right)_{n\in\mathbb{N}}$ as the two-scale semiclassical measure $\left(M^{0}_{\infty}\mathrm{d}\nu^{0}_{\infty},M^{0}\mathrm{d}\nu^{0}\right)$, it follows that, for almost all $t$,
+$M^{t}_{\infty}(s,\omega)\mathrm{d}\nu^{t}_{\infty}(s,\omega)=0.$
+The evolution of the two-scale semiclassical measure in normal geodesic coordinates allows us to describe the general evolution of any normalized initial condition for equation (1.1).
+This yields the following corollary, which is proved in Section 3.3.
+Corollary 1.7 (Propagation of semiclassical measure).
+Assume Assumption 1.1 and Assumption 1.3.
+Let $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ be solutions to (1.1) with normalized initial condition $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$.
+Then, there exist a vanishing sequence of positive numbers ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ and
+  • $\bullet$
+    a semiclassical measure $\mu_{t}$, evolving as described in Lemma 1.2,
+  • a family of non-negative scalar Radon measures $\left(\nu_{n}^{t}\right)_{n\in\mathbb{Z}}$ satisfying (1.20),
+  • $\bullet$
+    a family $\nu^{t}_{\pm}$ of non-negative scalar Radon measures on $\mathcal{C}\times\mathbb{S}^{2}$ such that satisfying the invariance equation $\mathrm{div}_{\omega}\left(\vec{V}^{\infty}(s,\omega)\nu_{\pm}^{t}(s,\omega)\right)=0$, with $\vec{V}^{\infty}$ defined in (1.22),
+such that, for all $(a,\Xi)\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)\times\mathscr{C}_{c}^{\infty}(\mathbb{R},\mathbb{C})$,
+$\mathcal{I}^{\varepsilon_{k}}_{\psi}(\Xi,a)\underset{k\to+\infty}{\longrightarrow}\ \int_{\mathbb{R}}\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\Xi(t)\mathrm{tr}\Big(a(x,\xi)\mathrm{d}\mu_{t}(x,\xi)\Big)\mathrm{d}t+\int_{\mathbb{R}}\int_{\mathbb{R}}\Xi(t)\mathrm{tr}\Big(a\big(\mathbf{t}(s),0\big)\mathrm{d}\rho_{t}(s)\Big)\mathrm{d}t,$
+with, for all $s\in\mathbb{R}$,
+$\rho_{t}(s)=\sum_{\pm}\int_{\mathbb{S}^{2}}\mathrm{U}_{\theta(s)}^{\dagger}\Pi^{\infty}_{\pm}(s,\omega)\mathrm{U}_{\theta(s)}\mathrm{d}\nu^{t}_{\pm}(s,\omega)+\frac{1}{2}\sum_{n\in\mathbb{Z}}\int_{\mathbb{R}}\begin{pmatrix}1&\frac{-\sigma}{\lambda_{n}(s,\sigma)}e^{-\mathrm{i}\theta(s)}\\ \frac{-\sigma}{\lambda_{n}(s,\sigma)}e^{\mathrm{i}\theta(s)}&1\end{pmatrix}\mathrm{d}\nu^{t}_{n}(s,\mathrm{d}\sigma).$
+Link between Drouot (2022), Bal et al. (2021) and Theorem 1.5
+A direct corollary of Theorem 1.5 is the following.
+Corollary 1.8.
+Assume Assumption 1.1 and Assumption 1.3.
+Let $\vec{f}\in\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$.
+If ${\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}}$ solves (1.1) with, for all $\varepsilon\in(0,1]$,
+$\psi_{0}^{\varepsilon}(x)=\frac{1}{\sqrt{\varepsilon}}\vec{f}\left(\frac{x-x_{0}}{\sqrt{\varepsilon}}\right),\quad x\in\mathbb{R}^{2},$
+then there exist a vanishing sequence of positive numbers ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ and a family of measurable maps ${t\mapsto\left(\nu^{t}_{n}\right)_{n\in\mathbb{Z}}}$ valued in the set of non-negative scalar Radon measures on $\mathbb{R}^{2}$, such that for all $(a,\Xi)\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)\times\mathscr{C}_{c}^{\infty}(\mathbb{R},\mathbb{C})$,
+$\int_{\mathbb{R}}\Xi(t)\left\langle\mathrm{Op}_{\varepsilon}(a)\psi^{\varepsilon_{k}}_{t},\psi^{\varepsilon_{k}}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\mathrm{d}t\underset{k\to+\infty}{\longrightarrow}\int_{\mathbb{R}}\int_{\mathbb{R}}\Xi(t)\mathrm{tr}\Big(a\big(\mathbf{t}(s),0\big)\mathrm{d}\rho_{t}(s)\Big)\mathrm{d}t$
+with,
+$\rho_{t}(s)=\frac{1}{2}\sum_{n\in\mathbb{Z}}\int_{\mathbb{R}}\begin{pmatrix}1&\frac{-\sigma}{\lambda_{n}(s,\sigma)}e^{-\mathrm{i}\theta(s)}\\ \frac{-\sigma}{\lambda_{n}(s,\sigma)}e^{\mathrm{i}\theta(s)}&1\end{pmatrix}\mathrm{d}\nu^{t}_{n}(s,\mathrm{d}\sigma),\quad s\in\mathbb{R},$
+where, for all $n\in\mathbb{Z}$, $\nu_{n}^{t}$ satisfies (1.20) and $\nu_{n}^{0}$ is determined by the initial condition.
+With the notation of Corollary 1.8, the measure $\nu_{0}^{t}$ corresponds to the wave packet part of the solution in Bal et al.
+(2021), i.e., the edge mode generated by the initial data (1.2).
+Similarly, the measures $\left(\nu_{n}^{t}\right)_{n\in\mathbb{Z}_{\neq 0}}$ correspond to the $\mathrm{L}^{\infty}$-remainder part of the solution in Drouot (2022).
+Notably, if $\nu_{n}^{0}=0$ for all $n\in\mathbb{Z}_{\neq 0}$ and $\nu_{0}^{0}=1$, then, for all $n\in\mathbb{Z}_{\neq 0}$,
+$\nu_{0}^{t}(s,\sigma)=\delta(s-t)\delta(\sigma),\quad\nu_{n}^{t}(s,\sigma)=0,\quad(t,s,\sigma)\in\mathbb{R}^{3},$
+where $\delta$ denotes the Dirac mass at 0.
+Corollary 1.8 provides a description of the evolution of the position density for initial data given by wave packets.
+However, Theorem 1.5 allows us to consider arbitrary data in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ concentrating near the curve $\mathcal{C}$, and thus to describe their asymptotic position-density behavior.
+Corollary 1.9.
+Assume Assumption 1.1 and Assumption 1.3.
+Let ${\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}}$ be a uniformly bounded family in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$, let $\nu$ be a non-negative scalar Radon measure on $\mathbb{R}_{t}\times\mathbb{R}_{x}^{2}$, and let ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ be a vanishing sequence such that, for all $(a,\Xi)\in\mathscr{C}_{c}^{\infty}\left(\Omega,\mathbb{C}\right)\times\mathscr{C}_{c}^{\infty}(\mathbb{R},\mathbb{C})$,
+$\int_{\mathbb{R}}\int_{\Omega}\Xi(t)a(x)\left|\psi_{t}^{\varepsilon_{k}}(x)\right|^{2}\mathrm{d}x\mathrm{d}t\underset{k\to+\infty}{\longrightarrow}\int_{\mathbb{R}}\int_{\Omega}\Xi(t)a(x)\mathrm{d}\nu(t,x),$
+where ${\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}}$ solves (1.1) with ${\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}}$ as initial condition. Then,
+$\nu\mathds{1}_{\mathrm{E}}\geqslant\sum_{n\in\mathbb{Z}}\Phi^{*}\left(\int_{\mathbb{R}}\nu^{t}_{n}(\cdot,\mathrm{d}\sigma)\right)\mathds{1}_{\mathrm{E}}\mathrm{d}t$
+where $\left(\nu_{n}^{t}\right)_{n\in\mathbb{Z}}$ is defined by (1.20).
+This corollary provides a lower bound for the position density at time $t$ for the evolution of arbitrary initial data concentrating along the curve $\mathrm{E}$.
+1.7 Organization of the paper
+The proof of Theorem 1.5 relies on two main arguments: the decomposition of the two-scale semiclassical measure above the curve $\mathcal{C}$ for arbitrary functions (Theorem 1.4) and the decomposition of each resulting measure (Theorem 1.5) that takes into account equation (1.1).
+First, to study the evolution of the measure $\mu_{t}$ near the curve $\mathcal{C}$, we perform a two-scale analysis in a neighborhood of $\mathcal{C}$.
+Geometrically, $\mu_{t}$ is a measure on the cotangent space $\mathrm{T}^{*}\mathbb{R}^{4}$ and, as such, satisfies the associated geometric invariance properties.
+In particular, this allows us to perform changes of coordinates.
+We work locally in a tubular neighborhood of $\mathcal{C}$, where we can use the normal geodesic coordinates introduced earlier.
+These coordinates $(s,y)$ are then used to transform the Hamiltonian, as detailed in Section 2.1.
+In Section 2.2, we separate and highlight two-scale Wigner observables associated with the different concentration regimes.
+In Section 2.3, we prove the existence of two-scale Wigner measures and describe them for arbitrary families.
+A rescaling is then applied in Section 2.4 to take into account the two-scale Wigner approach of the Hamiltonian.
+We will end up with a principal operator of the following form
+$\sqrt{\varepsilon}\begin{pmatrix}\sqrt{\varepsilon}\mathrm{D}_{1}&x_{2}\widetilde{m}(x_{1})+i\mathrm{D}_{2}\\ x_{2}\widetilde{m}(x_{1})-i\mathrm{D}_{2}&-\sqrt{\varepsilon}\mathrm{D}_{1}\end{pmatrix}$.
+This operator is $\sqrt{\varepsilon}$-semiclassical with respect to $x_{1}$ only which justifies our two-scale approach.
+Secondly, we describe each resulting measure and then prove our main result and its applications in Section 3.
+In fact, in Section 3.1.1, we describe the evolution of the measure at finite distance $M^{t}\mathrm{d}\nu^{t}$ and in Section 3.1.2, we describe the evolution of the measure at infinity $M_{\infty}^{t}\mathrm{d}\nu_{\infty}^{t}$.
+The decomposition of the measure $\mu_{t}$ as a sum of eigenprojectors associated with scalar measures is a consequence of the non-commutativity of matrices since
+$\int_{\mathbb{R}}\Xi(t)\mathrm{i}\sqrt{\varepsilon}\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathrm{Op}_{\varepsilon}(a)\psi^{\varepsilon}_{t},\psi^{\varepsilon}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\mathrm{d}t=\int_{\mathbb{R}}\Xi(t)\left\langle\mathrm{Op}_{\varepsilon}\Big(\left[a,\mathrm{H}\right]\Big)\psi^{\varepsilon}_{t},\psi^{\varepsilon}_{t}\right\rangle\mathrm{d}t+\mathcal{O}(\sqrt{\varepsilon}).$
+The evolution equation satisfied by the scalar measures then follows from the analysis of the next-order terms in the expansion in $\sqrt{\varepsilon}$ which is
+$\int_{\mathbb{R}}\Xi(t)\mathrm{i}\sqrt{\varepsilon}\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathrm{Op}_{\varepsilon}(a)\psi^{\varepsilon}_{t},\psi^{\varepsilon}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\mathrm{d}t=\frac{\sqrt{\varepsilon}}{2\mathrm{i}}\int_{\mathbb{R}}\Xi(t)\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big(\left\{a,\mathrm{H}\right\}-\left\{\mathrm{H},a\right\}\Big)\psi^{\varepsilon}_{t},\psi^{\varepsilon}_{t}\right\rangle\mathrm{d}t.$
+We finally conclude by proving Theorem 1.5 in Section 3.2 and by proving applications in Section 3.3.
+In Appendix A.1, we first apply the main properties of the Wigner transform (introduced in Section 1.2) to the solution of our problem.
+In Appendix A.2, we recall the proof of existence of semiclassical measure for families valued in a separable Hilbert space.
+This is necessarily since we study families valued in $\mathbb{C}$, $\mathbb{C}^{2}$ and $\mathrm{L}^{2}\left(\mathbb{R},\mathbb{C}^{2}\right)$.
+In Appendix B, we give a complete spectral analysis of (1.12).
+1.8 Notations
+We denote by ${\left(\sigma_{j}\right)_{1\leqslant j\leqslant 3}}$ the Pauli matrices, defined as
+$\sigma_{1}\coloneqq\begin{pmatrix}0&1\\ 1&0\end{pmatrix},\;\;\sigma_{2}\coloneqq\begin{pmatrix}0&-\mathrm{i}\\ \mathrm{i}&0\end{pmatrix},\;\;\sigma_{3}\coloneqq\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$.
+We denote $\mathrm{H}_{\varepsilon}$ the operator
+$\mathrm{H}_{\varepsilon}\coloneqq\begin{pmatrix}\mathrm{m}(x)&\varepsilon\mathrm{D}_{1}-\mathrm{i}\varepsilon\mathrm{D}_{2}\\ \varepsilon\mathrm{D}_{1}+\mathrm{i}\varepsilon\mathrm{D}_{2}&-\mathrm{m}(x)\end{pmatrix}=\mathrm{m}(x)\sigma_{3}+\varepsilon\mathrm{D}_{1}\sigma_{1}+\varepsilon\mathrm{D}_{2}\sigma_{2}.$
+Let $\mathcal{H}$ be a separable Hilbert space and $n$ an integer. We denote the usual inner product by
+$\left\langle f,g\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{n},\mathcal{H}\right)}\coloneqq\int_{\mathbb{R}^{n}}\left\langle f(x),g(x)\right\rangle_{\mathcal{H}}\,\mathrm{d}x,\quad(f,g)\in\mathrm{L}^{2}\left(\mathbb{R}^{n},\mathcal{H}\right)^{2}.$
+For an operator $A$ acting on $\mathrm{L}^{2}\left(\mathbb{R}^{n},\mathcal{H}\right)$, we denote by $A^{\dagger}$ its formal adjoint with respect to this inner product.
+For $\varepsilon>0$ and $f\in\mathrm{L}^{1}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$, we define the semiclassical Fourier transform by
+$\mathcal{F}_{\varepsilon}f(\xi)\coloneqq\frac{1}{2\pi\varepsilon}\int_{\mathbb{R}^{2}}e^{-\mathrm{i}\frac{x\cdot\xi}{\varepsilon}}f(x)\mathrm{d}x,\quad\xi\in\mathbb{R}^{2},$
+and we denote by $\widehat{f}$ the usual Fourier transform of $f$, related to $\mathcal{F}_{1}f$ by $\widehat{f}\coloneqq 2\pi\mathcal{F}_{1}f$.
+2 Two-scale analysis approach
+The coordinates $(s,y)$ are used to obtain a normal form of $\mathrm{H}_{\varepsilon}$, as detailed in Section 2.1.
+In Section 2.2, we separate and highlight two-scale Wigner observables corresponding to the different concentration regimes.
+In Section 2.3, we prove the existence of two-scale Wigner measures and describe them for arbitrary families.
+Finally, in Section 2.4, a rescaling is applied to take into account the two-scale Wigner approach for the Hamiltonian.
+2.1 Straightening the edge
+In this section, we consider $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ a family of solutions to (1.1) with normalized initial condition $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$.
+Normal form operator.
+We assume Assumption 1.1 and Assumption 1.3.
+We work in the normal geodesic coordinates denoted $(s,y)$ introduced in (1.8).
+Using the map $\Phi$, we transform $\mathrm{H}_{\varepsilon}$ to a normal form, following the approach of Borrelli et al.
+(2022).
+Proposition 2.1 (Hamiltonian Normal form near the interface).
+There exists a unitary operator $\mathrm{U}:\mathrm{L}^{2}(\Omega,\mathbb{C}^{2})\to\mathrm{L}^{2}(\mathbb{R}\times\mathrm{I},\mathbb{C}^{2})$ such that $\mathrm{U}\mathrm{H}_{\varepsilon}\mathrm{U}^{-1}=\mathrm{H}_{\varepsilon}^{\mathrm{E}}$ where the operator $\mathrm{H}_{\varepsilon}^{\mathrm{E}}$ acts on $\mathrm{L}^{2}(\mathbb{R}\times\mathrm{I},\mathbb{C}^{2})$ and is defined by
+$\mathrm{H}_{\varepsilon}^{\mathrm{E}}\coloneqq\mathrm{m}\Big(\Phi(s,y)\Big)\sigma_{1}+\varepsilon\mathrm{D}_{y}\sigma_{2}-\frac{\varepsilon}{1+y\kappa(s)}\mathrm{D}_{s}\sigma_{3}+\frac{\mathrm{i}\varepsilon y\kappa^{\prime}(s)}{2(1+y\kappa(s))^{2}}\sigma_{3}.$
+Remark.
+The following properties hold.
+  1. 1.
+    For all $(s,y)\in\mathbb{R}\times\mathrm{I}$, we have $\left|1+y\kappa(s)\right|>1/2$; therefore, all the quantities introduced above are well-defined.
+  2. 2.
+    The operator $\mathrm{U}$ is a matrix-valued Fourier Integral Operator.
+It incorporates the pull-back associated with the change of variables $\Phi$ as well as multiplication by an suitable $s$-dependent matrix, which combines a rotation and a multiplication by a $s$-dependent phase factor.
+  3. 3.
+    In what follows, slight abuse of notation, we shall apply the operator $\mathrm{U}$ to functions in $\mathrm{L}^{2}(\mathbb{R}^{2},\mathbb{C}^{2})$ that are compactly supported in $\Omega$.
+Proof of Proposition 2.1.
+Let us define $\mathrm{U}$ as follows
+$\mathrm{U}u(s,y)\coloneqq\sqrt{\frac{1+y\kappa(s)}{2}}\begin{pmatrix}e^{\mathrm{i}\frac{\theta(s)}{2}}&-e^{-\mathrm{i}\frac{\theta(s)}{2}}\\ e^{\mathrm{i}\frac{\theta(s)}{2}}&e^{-\mathrm{i}\frac{\theta(s)}{2}}\end{pmatrix}u\Big(\Phi(s,y)\Big),\quad u\in\mathrm{L}^{2}(\Omega,\mathbb{C}^{2}),\ (s,y)\in\mathbb{R}\times\mathrm{I}.$
+To facilitate computations, we split $\mathrm{U}$ into the composition of two operators. Define
+$\mathrm{V}_{\Phi}u(s,y)\coloneqq\sqrt{1+y\kappa(s)}u\Big(\Phi(s,y)\Big),\quad\quad u\in\mathrm{L}^{2}(\Omega,\mathbb{C}^{2}).$
+The operator $\mathrm{V}_{\Phi}$ is unitary from $\mathrm{L}^{2}\left(\Omega,\mathbb{C}^{2}\right)$ to $\mathrm{L}^{2}\left(\mathbb{R}\times\mathrm{I},\mathbb{C}^{2}\right)$, and the operator $\mathrm{U}_{\theta}$, defined in (1.10), is unitary on $\mathrm{L}^{2}\left(\mathbb{R}\times\mathrm{I},\mathbb{C}^{2}\right)$.
+Notably, $\mathrm{U}=\mathrm{U}_{\theta(s)}\mathrm{V}_{\Phi}$.
+Moreover, we have
+$\mathrm{V}_{\Phi}\mathrm{H}_{\varepsilon}\mathrm{V}_{\Phi}^{-1}=$ $\ \mathrm{m}\Big(\Phi(s,y)\Big)\sigma_{3}-\frac{\varepsilon}{1+y\kappa(s)}(\mathbf{t}^{\prime}\odot\sigma)\mathrm{D}_{s}+\varepsilon(\mathbf{n}\odot\sigma)\mathrm{D}_{y}$
+$\ +\ \frac{\mathrm{i}\varepsilon y\kappa^{\prime}(s)}{2(1+y\kappa(s))^{2}}(\mathbf{t}^{\prime}\odot\sigma)+\frac{\mathrm{i}\varepsilon\kappa(s)}{2(1+y\kappa(s))}(\mathbf{n}\odot\sigma),$
+where $x\odot\sigma\coloneqq x_{1}\sigma_{1}+x_{2}\sigma_{2}$ for $x\in\mathbb{C}^{2}$ and $\sigma_{1},\sigma_{2}$ are the Pauli matrices. Finally, the following identities hold.
+$\begin{split}\mathrm{U}_{\theta(s)}\sigma_{3}\mathrm{U}_{\theta(s)}^{-1}=\sigma_{1},\quad&\mathrm{U}_{\theta(s)}(\mathbf{t}^{\prime}\odot\sigma)\mathrm{U}_{\theta(s)}^{-1}=\sigma_{3},\\ \mathrm{U}_{\theta(s)}(\mathbf{n}\odot\sigma)\mathrm{U}_{\theta(s)}^{-1}=\sigma_{2},\quad&\mathrm{U}_{\theta(s)}(\mathbf{t}^{\prime}\odot\sigma)\left(\mathrm{D}_{s}\mathrm{U}_{\theta(s)}^{-1}\right)=-\frac{\kappa(s)}{2}\mathrm{i}\sigma_{2}.\end{split}$
 […]
-$\frac{\nabla\mathrm{m}(\gamma(s))}{\left|\nabla\mathrm{m}(\gamma(s))\right|}=\begin{pmatrix}\cos\theta(s)\\ \sin\theta(s)\end{pmatrix}.$
-We will denote the dual variable of $(s,y)$ by $(\sigma,\eta)$. Let us consider for all $(s,y,\sigma)\in\mathbb{R}^{3}$ and for all $n\in\mathbb{Z}_{\neq 0}$,
-$\lambda_{0}(\sigma)\coloneqq-\sigma\;\;,\;\;\lambda_{n}(s,\sigma)\coloneqq\mathrm{sgn}(n)\sqrt{\sigma^{2}+2|n||\nabla\mathrm{m}(\gamma(s))|}.$ (5)
-Although the function $\lambda_{0}$ does not depend on $s$, we also will denote it by $\lambda_{0}(s,\sigma)$.
-A consequence of our main result (in Section 3.3) is the following asymptotic description of the position density of the solution of equation (1) for initial data which are wave packets concentrating on the curve $\mathcal{C}$.
-Theorem 1.4.
-Assume Assumption 1.2 and 1.3.
+$\begin{matrix}\Phi_{\delta_{0}}:&\mathbb{R}\times\mathrm{I}_{\delta_{0}}&\longrightarrow&\mathbb{R}^{2}\\ &(s,y)&\longmapsto&\gamma(s)+y\nu(s)\end{matrix}$
+A consequence of Assumption 1.3 is that for all $\delta_{0}\in\left(0,1\right]$, the function $\Phi_{\delta_{0}}$ is a diffeomorphism from $\mathbb{R}\times\mathrm{I}_{\delta_{0}}$ to
+$\Omega_{\delta_{0}}$ $\coloneqq\left\{\gamma(s)+y\nu(s)|\ (s,y)\in\mathbb{R}\times\mathrm{I}_{\delta_{0}}\right\},$
+[…]
+[…]
+$\varphi^{\varepsilon}_{t}\coloneqq\mathrm{U}\chi\psi^{\varepsilon}_{t}.$ (2.1)
+Lemma 2.2 (Dirac equation in normal geodesic coordinates).
+The family $\left(\varphi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ satisfies, on $\mathbb{R}_{t}\times\mathbb{R}_{s}\times\mathrm{I}_{1/2}$, the following equation
+$\varepsilon\mathrm{D}_{t}\varphi^{\varepsilon}_{t}+\mathrm{H}^{\mathrm{E}}_{\varepsilon}\varphi^{\varepsilon}_{t}=0.$ (2.2)
+Proof.
+We apply $\mathrm{U}\chi$ to equation (1.1), obtaining
+$\varepsilon\mathrm{D}_{t}\varphi_{t}^{\varepsilon}+\mathrm{H}^{\mathrm{E}}_{\varepsilon}\varphi_{t}^{\varepsilon}=\mathrm{U}\left[\mathrm{H}_{\varepsilon},\chi\right]\psi_{t}^{\varepsilon}.$
+A direct computation gives
+$\left[\mathrm{H}_{\varepsilon},\chi\right]=-\mathrm{i}\varepsilon\begin{pmatrix}0&\partial_{1}\chi-\mathrm{i}\partial_{2}\chi\\ \partial_{1}\chi+\mathrm{i}\partial_{2}\chi&0\end{pmatrix}$,
+[…]
+$\Big(\mathrm{U}\left[\mathrm{H}_{\varepsilon},\chi\right]\psi_{t}^{\varepsilon}\Big)(s,y)=0,\quad(t,s,y)\in\mathbb{R}_{t}\times\mathbb{R}_{s}\times\mathrm{I}_{1/2}.$
+Pull-back of semiclassical measure.
+We establish the connection between the semiclassical measure of the family $\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}$ and the two-scale semiclassical measure of the family $\left(\varphi_{t}^{\varepsilon}\right)_{\varepsilon>0}$ defined in (2.1) via the change of variables introduced in Proposition 2.1.
+Lemma 2.3 (Semiclassical measure and normal geodesic coordinates).
+All semiclassical measures $\mu_{t}$ of the family $(\psi^{\varepsilon}_{t})_{\varepsilon>0}$ associated with the sequence $(\varepsilon_{k})_{k\in\mathbb{N}}$ satisfy
+$\mathrm{d}\mu_{t}\mathrm{d}t\mathds{1}_{\mathcal{C}}=\Phi^{*}\Big(\mathrm{U}_{\theta(s)}^{\dagger}\mathrm{d}\rho_{t}\mathrm{U}_{\theta(s)}\Big)\mathrm{d}t\mathds{1}_{\mathcal{C}}$
+where $\rho_{t}$ is the semiclassical measure of the family $\left(\varphi_{t}^{\varepsilon}\right)_{\varepsilon>0}$ on $\mathbb{R}_{s}\times\mathrm{I}_{1/2}\times\mathbb{R}_{\sigma,\eta}^{2}\times\mathbb{R}_{t}$ associated with the sequence $\left(\varepsilon_{k}\right)_{k\in\mathbb{N}}$.
+Proof.
+Let $a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)$ with $\mathrm{supp}(a)\subset\Omega_{1/2}\times\mathbb{R}^{2}$. Then, for all $N\in\mathbb{N}$, we have
+$\left\langle\mathrm{Op}_{\varepsilon}\left(a\right)\psi^{\varepsilon}_{t},\psi^{\varepsilon}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}$ $=\left\langle\mathrm{Op}_{\varepsilon}\left(a\right)\chi\psi^{\varepsilon}_{t},\chi\psi^{\varepsilon}_{t}\right\rangle_{\mathrm{L}^{2}\left(\Omega_{1/2},\mathbb{C}^{2}\right)}+\mathcal{O}\left(\varepsilon^{N}\right)$
+$=\left\langle\mathrm{U}\mathrm{Op}_{\varepsilon}\left(a\right)\mathrm{U}^{\dagger}\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}\times\mathrm{I}_{1/2},\mathbb{C}^{2}\right)}+\mathcal{O}\left(\varepsilon^{N}\right).$
+Since $\rm U=U_{\theta(s)}V_{\Phi}$, with $\rm U_{\theta(s)}$ depending only on $s$ and $\rm V_{\Phi}$ a change of variables, standard results in semiclassical pseudodifferential calculus (see (Zworski, 2012, Theorem 9.3)) give
+$\left\langle\mathrm{Op}_{\varepsilon}\left(a\right)\psi^{\varepsilon}_{t},\psi^{\varepsilon}_{t}\right\rangle=$ $\ \left\langle\mathrm{U}_{\theta(s)}\mathrm{Op}_{\varepsilon}\Big(\mathrm{V}_{\Phi}a\mathrm{V}_{\Phi}^{\dagger}\Big)\mathrm{U}_{\theta(s)}^{\dagger}\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}\times\mathrm{I}_{1/2},\mathbb{C}^{2}\right)}+\mathcal{O}\left(\varepsilon\right)$
+$=$ $\ \left\langle\mathrm{Op}_{\varepsilon}\left(\mathrm{U}_{\theta(s)}a\mathrm{U}_{\theta(s)}^{\dagger}\left(\Phi(s,y),{}^{t}\nabla\left(\Phi^{-1}\right)\Big(\Phi(s,y)\Big)\left[\sigma,\eta\right]\right)\right)\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}\times\mathrm{I}_{1/2},\mathbb{C}^{2}\right)}$
+$+\mathcal{O}\left(\varepsilon\right).$
+Assuming the sequence $\left(\varepsilon\right)_{\varepsilon>0}$ realizes the semiclassical measure $\mu_{t}$ of $(\psi^{\varepsilon}_{t})_{\varepsilon>0}$, passing to the limit $\varepsilon\to 0$ gives
+$\left\langle\mathrm{Op}_{\varepsilon}\left(a\right)\psi^{\varepsilon}_{t},\psi^{\varepsilon}_{t}\right\rangle$ $\underset{\varepsilon\to 0}{\longrightarrow}\int_{\mathbb{R}\times\mathrm{I}_{1/2}\times\mathbb{R}^{2}}\mathrm{Tr}\left(\mathrm{U}_{\theta(s)}a\mathrm{U}_{\theta(s)}^{\dagger}\left(\Phi(s,y),{}^{t}\nabla\left(\Phi^{-1}\right)\Big(\Phi(s,y)\Big)\left[\sigma,\eta\right]\right)\mathrm{d}\rho_{t}(s,y,\sigma,\eta)\right)$
+$=\int_{\mathbb{R}\times\mathrm{I}_{1/2}\times\mathbb{R}^{2}}\mathrm{Tr}\left(a\left(\Phi(s,y),{}^{t}\nabla\left(\Phi^{-1}\right)\Big(\Phi(s,y)\Big)\left[\sigma,\eta\right]\right)\Big(\mathrm{U}_{\theta(s)}^{\dagger}\mathrm{d}\rho_{t}(s,y,\sigma,\eta)\mathrm{U}_{\theta(s)}\Big)\right)$
+$=\int_{\Omega_{1/2}\times\mathbb{R}^{2}}\mathrm{Tr}\left(a(x,\xi)\Phi^{*}\Big(\mathrm{U}_{\theta(s)}^{\dagger}\mathrm{d}\rho_{t}\mathrm{U}_{\theta(s)}\Big)(x,\xi)\right).$
+Finally, restricting the measure $\rho_{t}$ to the curve $\mathcal{C}$ gives the semiclassical measure associated with $(\psi^{\varepsilon}_{t})_{\varepsilon>0}$ near $\mathcal{C}$. ∎
+2.2 Two-scale analysis of concentrations
+In this section, we will consider two-scale semiclassical observables and quantization defined in (1.11) at the scale $\sqrt{\varepsilon}$. We first remark that for all $a\in\mathcal{A}$,
+$\mathbf{op}^{\mathcal{C}}_{\varepsilon}(a)$ $=\mathrm{Op}_{\varepsilon}\left(a\left(s,y,\sigma,\eta,\frac{(y,\sigma,\eta)}{\sqrt{\varepsilon}}\right)\right)$
+$=\Lambda_{\sqrt{\varepsilon}}^{\dagger}\mathrm{Op}_{1}\Big(a\left(s,\sqrt{\varepsilon}y,\varepsilon\sigma,\sqrt{\varepsilon}\eta,y,\sqrt{\varepsilon}\sigma,\eta\right)\Big)\Lambda_{\sqrt{\varepsilon}},\ \left(\text{semiclassical in }\sqrt{\varepsilon}\right)$ (2.3)
+with
+$\Lambda_{\sqrt{\varepsilon}}[f](s,y)=\varepsilon^{1/4}f\left(s,\sqrt{\varepsilon}y\right),\quad f\in\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right),\ (s,y)\in\mathbb{R}^{2}.$ (2.4)
+Notably, $\Lambda_{\sqrt{\varepsilon}}$ is a unitary operator on $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$.
+Moreover, the quantization map $\mathbf{op}^{\mathcal{C}}_{\varepsilon}$ associates to $a$ a bounded operator on $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ according to the Calderón-Vaillancourt Theorem Calderón and Vaillancourt (1971) applied to expression (2.3).
+These observables allow us to differentiate three regions of the phase space $\mathbb{R}_{s}\times\mathrm{I}_{1/2}\times\mathbb{R}^{2}_{\sigma,\eta}$.
+Let us explain this fact.
+Let $a\in\mathcal{A}$ and $\chi\in\mathscr{C}^{\infty}_{c}(\mathbb{R},[0,1])$ supported in [1,1][-1,1] with $\chi\equiv 1$ on [1/2,1/2][-1/2,1/2].
+Let $\delta>0$, $R>R_{0}$ with $R_{0}$ associated with $a$ and $\varepsilon$ sufficiently small enough such that
+$2R\sqrt{\varepsilon}<\delta.$ (2.5)
+Let us denote $\mathbf{x}\coloneqq(s,y),\widetilde{\xi}\coloneqq(\sigma,\eta)$ and $z\coloneqq\left(y,\sigma,\eta\right)$ for simplicity. We now split $a$ into three observables,
+$a\left(\mathbf{x},\widetilde{\xi},z\right)=$ $\ a\left(\mathbf{x},\widetilde{\xi},z\right)\left(1-\chi\left(\frac{\sqrt{\varepsilon}\left|z\right|}{\delta}\right)\right)$ $\left(\eqqcolon a^{\delta}_{\varepsilon}\left(\mathbf{x},\widetilde{\xi},z\right)\right)$ (2.6)
+$+\ a\left(\mathbf{x},\widetilde{\xi},z\right)\chi\left(\frac{\sqrt{\varepsilon}\left|z\right|}{\delta}\right)\left(1-\chi\left(\frac{\left|z\right|}{R}\right)\right)$ $\left(\eqqcolon a_{\varepsilon,\delta}^{R}\left(\mathbf{x},\widetilde{\xi},z\right)\right)$ (2.7)
+$+\ a\left(\mathbf{x},\widetilde{\xi},z\right)\chi\left(\frac{\left|z\right|}{R}\right)$ $\left(\eqqcolon a_{\varepsilon,R}\left(\mathbf{x},\widetilde{\xi},z\right)\right).$ (2.8)
+Let us define $\chi_{R}\coloneqq\chi\left(\frac{\left|\cdot\right|}{R}\right)$ and $\chi^{R}\coloneqq 1-\chi_{R}$.
+The symbol $a_{\varepsilon}^{\delta}$ through the second-scale quantization is supported outside $\mathcal{C}$, at a distance $\delta$ of $\mathcal{C}$.
+Because of (2.5), $\mathbf{op}^{\mathcal{C}}_{\varepsilon}\left(a_{\varepsilon}^{\delta}\right)$ is a semiclassical operator which satisfies
+$\mathbf{op}^{\mathcal{C}}_{\varepsilon}\left(a_{\varepsilon}^{\delta}\right)=\mathrm{Op}_{1}\left(a_{\infty}\left(s,y,\varepsilon\sigma,\varepsilon\eta,\frac{(y,\varepsilon\sigma,\varepsilon\eta)}{|(y,\varepsilon\sigma,\varepsilon\eta)|}\right)\left(1-\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{\delta}\right)\right)\right).$
+So, we can apply the semiclassical theory recalled in Appendix A.2 to the observable $a_{\varepsilon}^{\delta}$.
+We are particularly interested in the two regions covered by $a_{\varepsilon,\delta}^{R}$ and $a_{\varepsilon,R}$ as they describe the measure above the curve $\mathcal{C}$.
+In the second-scale quantization, the symbol $a_{\varepsilon,R}$ is supported at a distance of order $\sqrt{\varepsilon}$ of $\mathcal{C}$ and $a_{\varepsilon,\delta}^{R}$ is supported at distance larger than $\sqrt{\varepsilon}$ of $\mathcal{C}$ but smaller than $\delta$.
+For the observable $a_{\varepsilon,\delta}^{R}$, we have
+$\mathbf{op}^{\mathcal{C}}_{\varepsilon}\left(a_{\varepsilon,\delta}^{R}\right)=\mathrm{Op}_{1}\left(a\left(s;0;\frac{y}{\sqrt{\varepsilon}},\sqrt{\varepsilon}\sigma,\sqrt{\varepsilon}\eta\right)\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{\delta}\right)\left(1-\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{R\sqrt{\varepsilon}}\right)\right)\right)+\mathcal{O}\left(\delta\right).$
+But we choose $R>R_{0}$ then
+$\mathbf{op}^{\mathcal{C}}_{\varepsilon}\left(a_{\varepsilon,\delta}^{R}\right)=\mathrm{Op}_{1}\left(a_{\infty}\left(s;0;\frac{(y,\varepsilon\sigma,\varepsilon\eta)}{\left|(y,\varepsilon\sigma,\varepsilon\eta)\right|}\right)\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{\delta}\right)\left(1-\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{R\sqrt{\varepsilon}}\right)\right)\right)+\mathcal{O}\left(\delta\right).$
+By using (2.3) on $a_{\varepsilon,R}$, we have
+$\mathbf{op}^{\mathcal{C}}_{\varepsilon}\left(a_{\varepsilon,R}\right)$ $=\mathrm{Op}_{1}\left(a(s,0,0,0,\frac{y}{\sqrt{\varepsilon}},\sqrt{\varepsilon}\sigma,\sqrt{\varepsilon}\eta)\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{R\sqrt{\varepsilon}}\right)\right)\ +\ \mathcal{O}\left(\sqrt{\varepsilon}\right),$
+$=\Lambda_{\sqrt{\varepsilon}}^{\dagger}\mathrm{Op}_{1}\left(a(s,0,0,0,y,\sqrt{\varepsilon}\sigma,\eta)\chi\left(\frac{|(y,\sqrt{\varepsilon}\sigma,\eta)|}{R}\right)\right)\Lambda_{\sqrt{\varepsilon}}\ +\ \mathcal{O}\left(\sqrt{\varepsilon}\right).$ (2.9)
+So we will consider the separable Hilbert space $\mathcal{H}\coloneqq\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)$ and the setting of Appendix A.2: the functional space $\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathcal{L}\left(\mathcal{H}\right)\right)$, and the two-scale quantization defined for $b\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathcal{L}\left(\mathcal{H}\right)\right)$,
+$\mathrm{Op}_{\sqrt{\varepsilon},1}(b)\coloneqq\mathrm{Op}_{1}\Big(b\left(s,\sqrt{\varepsilon}\sigma\right)\Big).$ (2.10)
+Let us denote, for all $(s,\sigma)\in\mathbb{R}^{2}$, the bounded operator on $\mathcal{H}$ obtained by the Weyl quantization of $(y,\eta)\mapsto a(s,0,0,0,y,\sigma,\eta)$ by $a^{W}(s,0,0,0,y,\sigma,\mathrm{D}_{y})$.
+Then, for the operator $b(s,\sigma)$ defined by for all $(s,\sigma)\in\mathbb{R}^{2}$, $b(s,\sigma)\coloneqq a^{W}(s,0,0,0,y,\sigma,\mathrm{D}_{y})\chi\left(\frac{|(y,\sigma,\mathrm{D}_{y})|}{R}\right)$, we have
+$\mathbf{op}^{\mathcal{C}}_{\varepsilon}\left(a_{\varepsilon,R}\right)=\Lambda_{\sqrt{\varepsilon}}^{\dagger}\mathrm{Op}_{\sqrt{\varepsilon},1}(b)\Lambda_{\sqrt{\varepsilon}}\ +\ \mathcal{O}\left(\sqrt{\varepsilon}\right).$ (2.11)
+In the following sections, we will take advantage of these properties of $a_{\varepsilon,R}$ and $a_{\varepsilon,\delta}^{R}$.
+2.3 Two-scale semiclassical measures
+2.3.1 Existence
+Let us first consider the stationary case.
+We use observables from $\mathcal{A}$ to analyze the oscillations at the scales $\sqrt{\varepsilon}$ and $\delta$ of uniformly bounded families $\left(f^{\varepsilon}\right)_{\varepsilon>0}$ of $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ via $a_{\varepsilon,R}$ and $a_{\varepsilon,\delta}^{R}$, respectively.
+Let us define
+$\mathcal{I}^{\varepsilon}_{f}(a)\coloneqq\left\langle\mathbf{op}^{\mathcal{C}}_{\varepsilon}(a)f^{\varepsilon},f^{\varepsilon}\right\rangle_{\mathrm{L}^{2}(\mathbb{R}^{2},\mathbb{C}^{2})},\quad a\in\mathcal{A}.$
+Theorem 2.4 (Existence of the two-scale semiclassical measure).
+Let ${(f^{\varepsilon})_{\varepsilon>0}}$ uniformly bounded in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$, then, there exist a vanishing sequence of positive numbers ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ and
+  • $\bullet$
+    Two measurable map $\nu_{\infty}$ and $\nu$, valued in the set of non-negative scalar measures on $\mathcal{C}\times\mathbb{S}^{2}$ and $\mathbb{R}^{2}$, respectively,
+  • $\bullet$
+    A $\mathrm{d}\nu_{\infty}$-measurable map $(s,\omega)\mapsto M_{\infty}(s,\omega)\in\mathbb{C}^{2,2}$, valued in the set of Hermitian trace $1$ positive matrices,
+  • $\bullet$
+    A $\mathrm{d}\nu(s,\sigma)$ measurable map $(s,\sigma)\mapsto M(s,\sigma)$, valued in trace-one positive operators on $\mathrm{L}^{2}\left(\mathbb{R},\mathbb{C}^{2}\right)$,
+such that for all $a\in\mathcal{A}$ we have
+$\mathcal{I}^{\varepsilon_{k}}_{f}(a)\underset{k\to+\infty}{\longrightarrow}$ $\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\mathrm{tr}\left(a_{\infty}\left(s,y,\sigma,\eta,\frac{(y,\sigma,\eta)}{|(y,\sigma,\eta)|}\right)\mathrm{d}\mu(s,y,\sigma,\eta)\right)$
+$+\ \int_{\mathbb{R}\times\mathbb{S}^{2}}\mathrm{tr}\Big(a_{\infty}(s,0,0,0,\omega)M_{\infty}(s,\omega)\Big)\mathrm{d}\nu_{\infty}(s,\omega)$
+$+\ \int_{\mathbb{R}^{2}}\mathrm{Tr}\Big(a^{W}(s,0,0,0,y,\sigma,\mathrm{D}_{y})M(s,\sigma)\Big)\mathrm{d}\nu(s,\sigma),$
+where $\mu$ is the semiclassical measure of ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$ associated with the sequence ${(\varepsilon_{k})_{k\in\mathbb{N}}}$.
+Proof.
+Let ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$ uniformly bounded in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$.
+For simplicity, we assume that the sequence $\left(\varepsilon\right)_{\varepsilon>0}$ realizes the semiclassical measure $\mu$ of ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$.
+We use the decomposition (2.6).
+First, we analyze the concentration of $a^{\delta}$.
+  • $\bullet$
+    Construction of $\mu$ outside $\mathcal{C}$
+For $a\in\mathcal{A}$ such that $a=0$ in a neighborhood of $\mathcal{C}$ and for $\varepsilon$ small enough, we have
+$\mathbf{op}^{\mathcal{C}}_{\varepsilon}\left(a\right)=\mathrm{Op}_{\varepsilon}\left(a\left(s,y;\sigma,\eta;\frac{y}{\sqrt{\varepsilon}},\frac{\sigma}{\sqrt{\varepsilon}},\frac{\eta}{\sqrt{\varepsilon}}\right)\right)=\mathrm{Op}_{1}\left(a_{\infty}\left(s,y,\varepsilon\sigma,\varepsilon\eta,\frac{(y,\varepsilon\sigma,\varepsilon\eta)}{|(y,\varepsilon\sigma,\varepsilon\eta)|}\right)\right).$
+After extraction of a subsequence that we still denote by $(\varepsilon)_{\varepsilon>0}$, we have
+$\limsup_{\delta\to 0}\limsup_{R\to+\infty}\limsup_{\varepsilon\to 0}\mathcal{I}^{\varepsilon}_{f}\left(a^{\delta}\right)=\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\mathrm{tr}\left(a_{\infty}\left(s,y,\sigma,\eta,\frac{(y,\sigma,\eta)}{|(y,\sigma,\eta)|}\right)\mathrm{d}\mu(s,y,\sigma,\eta)\right).$
+Then we construct the pair $M\mathrm{d}\nu$.
+  • $\bullet$
+    Construction of $M\mathrm{d}\nu$
+Let $a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{7},\mathbb{C}^{2,2}\right)$.
+Then, the function $b$, given by the relation (2.11), is valued in the set of compact operators on $\mathcal{H}$.
+We can see ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$ as a bounded family of $\mathrm{L}^{2}\left(\mathbb{R}_{s},\mathcal{H}\right)$.
+We are thus doing classical calculus on $\mathrm{L}^{2}\left(\mathbb{R}_{s},\mathcal{H}\right)$.
+We apply Lemma A.4, then there exists a vanishing sequence of positive numbers $\left(\varepsilon_{k}\right)_{k\in\mathbb{N}}$ such that for all $a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{7},\mathbb{C}^{2,2}\right)$,
+$\lim_{k\to+\infty}\left\langle\mathbf{op}^{\mathcal{C}}_{\varepsilon}\left(a\right)f^{\varepsilon_{k}},f^{\varepsilon_{k}}\right\rangle_{\mathrm{L}^{2}(\mathbb{R}^{2},\mathbb{C}^{2})}=\int_{\mathbb{R}^{2}}\mathrm{Tr}\Big(a^{W}(s,0;0,0;y,\sigma,\mathrm{D}_{y})M(s,\sigma)\Big)\mathrm{d}\nu(s,\sigma).$
+By applying this property to $a_{\varepsilon,R}$, we obtain $M\mathrm{d}\nu$. Finally we construct the pair $M_{\infty}\mathrm{d}\nu_{\infty}$.
+  • $\bullet$
+    Construction of $M_{\infty}\mathrm{d}\nu_{\infty}$
+Let $a\in\mathcal{A}$.
+According to the relation (2.9), we are interested in the family $\widetilde{f}^{\varepsilon}\coloneqq\Lambda_{\sqrt{\varepsilon}}f^{\varepsilon}$ which is bounded in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$.
+Then, the quantity
+$\mathscr{I}^{\varepsilon_{k}}_{\widetilde{f}}\left(\widetilde{a}^{R}_{\delta}\right)\coloneqq\left\langle\mathrm{Op}_{1}\left(\widetilde{a}^{R}_{\delta}(s,0,0,0,y,\sqrt{\varepsilon_{k}}\sigma,\eta)\right)\widetilde{f}^{\varepsilon_{k}},\widetilde{f}^{\varepsilon_{k}}\right\rangle_{\mathrm{L}^{2}(\mathbb{R}^{2},\mathbb{C}^{2})}.$
+is uniformly bounded in $\varepsilon>0$ and $R>1$. But the function $a^{R}_{\delta}$ is supported at distance at least $R\sqrt{\varepsilon}$ of $\mathcal{C}$, so
+$\mathscr{I}^{\varepsilon_{k}}_{\widetilde{f}}\left(\widetilde{a}^{R}_{\delta}\right)=\left\langle\mathrm{Op}_{1}\left(a_{\infty}\left(s,0,0,0,\frac{(y,\sqrt{\varepsilon_{k}}\sigma,\eta)}{\left|\left(y,\sqrt{\varepsilon_{k}}\sigma,\eta\right)\right|}\right)\right)\widetilde{f}^{\varepsilon_{k}},\widetilde{f}^{\varepsilon_{k}}\right\rangle_{\mathrm{L}^{2}(\mathbb{R}^{2},\mathbb{C}^{2})},$
+as soon as $R$ is large enough, depending on $\varepsilon$.
+We then deduce by a diagonal extraction argument that there exists a subsequence $\left(\varepsilon_{n}\right)_{n\in\mathbb{N}}$ of $\left(\varepsilon_{k}\right)_{k\in\mathbb{N}}$, a sequence $\left(R_{n}\right)_{n\in\mathbb{N}}$ and a sequence $\left(\delta_{n}\right)_{n\in\mathbb{N}}$ such that for all $a\in\mathcal{A}$,
+$\mathscr{I}^{\varepsilon_{n}}_{\widetilde{f}}\left(\widetilde{a}^{R_{n}}_{\delta_{n}}\right)\underset{n\to+\infty}{\longrightarrow}\mathscr{I}\left(a_{\infty}\right).$
+It remains to prove that $a_{\infty}\mapsto\mathscr{I}\left(a_{\infty}\right)$ is a measure, which will allow us to define $M_{\infty}\mathrm{d}\nu_{\infty}$.
+First, we prove that $a_{\infty}\mapsto\mathscr{I}\left(a_{\infty}\right)$ is a non-negative matrix-valued distribution.
+We observe that there exists $C>0$ such that for all $a\in\mathcal{A}$, for all $n\in\mathbb{N}$,
+$\mathscr{I}^{\varepsilon_{n}}_{\widetilde{f}}\left(\widetilde{a}^{R_{n}}_{\delta_{n}}\right)\leqslant C\mathrm{N}_{d}\left(\widetilde{a}^{R_{n}}_{\delta_{n}}\right),$
+with $\mathrm{N}_{d}$ a Schwarz semi-norm.
+We have $\mathrm{N}_{d}\left(\widetilde{a}^{R_{n}}_{\delta_{n}}\right)\to\mathrm{N}_{d}\left(a_{\infty}\right)$ then $\mathscr{I}\left(a_{\infty}\right)\leqslant C\mathrm{N}_{d}\left(a_{\infty}\right)$.
+Therefore $\mathscr{I}$ is a matrix-valued distribution.
+Secondly, the operators $a\mapsto\mathrm{Op}_{1}\left(a^{R}_{\delta}(s,0,0,0,y,\sqrt{\varepsilon}\sigma,\eta)\right)$ satisfy a semiclassical calculus in the parameters $\varepsilon$, $1/R$ and $\delta$.
+We have for all $a,a_{1},a_{2}\in\mathcal{A}$,
+  • $\bullet$
+    Adjoint : $\mathrm{Op}_{1}\left(\widetilde{a}^{R}_{\delta}(s,0,0,0,y,\sqrt{\varepsilon}\sigma,\eta)\right)^{\dagger}=\mathrm{Op}_{1}\left(\overline{\widetilde{a}^{R}_{\delta}}(s,0,0,0,y,\sqrt{\varepsilon}\sigma,\eta)\right)$
+  • $\bullet$
+    Symbolic Calculus : as operator of $\mathcal{L}\left(\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)$,
+    $\mathrm{Op}_{1}\Big(\widetilde{a}^{R}_{1,\delta}\Big)\mathrm{Op}_{1}\Big(\widetilde{a}^{R}_{2,\delta}\Big)=\mathrm{Op}_{1}\Big(\left(\widetilde{a}_{1}\widetilde{a}_{2}\right)^{R}_{1,\delta}\Big)+\mathcal{O}\left(\sqrt{\varepsilon}\right)+\mathcal{O}\left(\frac{1}{R}\right),$
+  • $\bullet$
+    Weak Gårding inequality : if $a\geqslant 0$, then for all $\widetilde{\delta}$, there exists $C_{\widetilde{\delta}}>0$, such that for all $f\in\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$,
+    $\left\langle\mathrm{Op}_{1}\Big(\widetilde{a}^{R}_{\delta}\Big)\widetilde{f},\widetilde{f}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\geqslant-\left(\widetilde{\delta}+C_{\widetilde{\delta}}\left(\varepsilon+\frac{1}{R}\right)\right)\left\lVert f\right\rVert^{2}_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)},$
+Therefore $a_{\infty}\mapsto\mathscr{I}\left(a_{\infty}\right)$ is a non-negative matrix-valued measure defined on $\mathbb{R}\times\mathbb{S}^{2}$ that we denote $M_{\infty}\mathrm{d}\nu_{\infty}$. ∎
+2.3.2 First properties
+In this section, we first explore properties of two-scale semiclassical measures defined by Theorem 2.4. A direct consequence of the proof is the separated construction of limits as follows.
+Remark 2.5 (Link with the limit of the previous observable).
+Let ${(f^{\varepsilon})_{\varepsilon>0}}$ uniformly bounded in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ and a vanishing sequence of positive numbers ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ given by Theorem 2.4, then, with the same notations, for all $a\in\mathcal{A}$,
+$\limsup_{\delta\to 0}\limsup_{R\to+\infty}\lim_{k\to+\infty}\mathcal{I}^{\varepsilon_{k}}_{f}\left(a_{\varepsilon_{k}}^{\delta}\right)=$ $\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\mathrm{tr}\left(a_{\infty}\left(s,y,\sigma,\eta,\frac{(y,\sigma,\eta)}{|(y,\sigma,\eta)|}\right)\mathrm{d}\mu(s,y,\sigma,\eta)\right),$
+$\limsup_{\delta\to 0}\limsup_{R\to+\infty}\lim_{k\to+\infty}\mathcal{I}^{\varepsilon_{k}}_{f}\left(a_{\varepsilon_{k},\delta}^{R}\right)=$ $\int_{\mathbb{R}\times\mathbb{S}^{2}}\mathrm{tr}\Big(a_{\infty}(s,0,0,0,\omega)M_{\infty}(s,\omega)\Big)\mathrm{d}\nu_{\infty}(s,\omega),$
+$\limsup_{\delta\to 0}\limsup_{R\to+\infty}\lim_{k\to+\infty}\mathcal{I}^{\varepsilon_{k}}_{f}\left(a_{\varepsilon_{k},R}\right)=$ $\int_{\mathbb{R}^{2}}\mathrm{Tr}\Big(a^{W}(s,0,0,0,y,\sigma,\mathrm{D}_{y})M(s,\sigma)\Big)\mathrm{d}\nu(s,\sigma).$
+Example.
+Let us consider,
+$\phi_{\varepsilon}(s,y)=\frac{1}{\varepsilon^{1/4}}u_{\varepsilon}\left(s\right)f\left(\frac{y}{\sqrt{\varepsilon}}\right)\vec{V},\quad(s,y)\in\mathbb{R}^{2},$
+with $\vec{V}\in\mathbb{C}^{2}$, $(f,u_{\varepsilon})\in\mathscr{S}(\mathbb{R},\mathbb{C})^{2}$ and the knowledge of $\rho$ the semiclassical measure of $(u_{\varepsilon})_{\varepsilon>0}$ at the scale $\sqrt{\varepsilon}$, then
+$\left\langle\mathbf{op}^{\mathcal{C}}_{\varepsilon}(a)\phi_{\varepsilon},\phi_{\varepsilon}\right\rangle\underset{\varepsilon\to 0}{\longrightarrow}\int_{\mathbb{R}^{2}}\mathrm{Tr}\Big(a^{W}(s,0,0,0,y,\sigma,\mathrm{D}_{y})M(s,\sigma)\Big)\mathrm{d}\nu(s,\sigma),\quad a\in\mathscr{C}^{\infty}_{c}(\mathbb{R}^{7},\mathbb{C}^{2,2}),$
+where
+$M(s,\sigma)\mathrm{d}\nu(s,\sigma)=\frac{1}{2\pi}\Pi_{f\vec{V}}\mathrm{d}\rho(s,\sigma),$
+with $\Pi_{f\vec{V}}$ the projector on $\mathrm{L}^{2}\left(\mathbb{R},\mathbb{C}^{2}\right)$ over the function $f\vec{V}$. Notably, if $u_{\varepsilon}$ is a wave packet, which means
+$u_{\varepsilon}\left(s\right)\coloneqq\frac{1}{\varepsilon^{1/4}}u\left(\frac{s}{\sqrt{\varepsilon}}\right),\quad u\in\mathscr{S}(\mathbb{R},\mathbb{C}),\quad s\in\mathbb{R},$
+then the semiclassical measure of $(u_{\varepsilon})_{\varepsilon>0}$ at the scale $\sqrt{\varepsilon}$ is $\rho(\sigma)=\left|\mathcal{F}u(\sigma)\right|^{2}$ (according to the case $\alpha=1$ of Example A.1 but with a $\sqrt{\varepsilon}$-Weyl quantization) so
+$M(s,\sigma)\mathrm{d}\nu(s,\sigma)=\left|\mathcal{F}u(\sigma)\right|^{2}\Pi_{f\vec{V}}\frac{\mathrm{d}\sigma}{2\pi}.$
+Another consequence is the application to observables in $\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)$.
+Remark 2.6.
+Let ${(f^{\varepsilon})_{\varepsilon>0}}$ uniformly bounded in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ and a vanishing sequence of positive numbers ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ given by Theorem 2.4.
+According to the notations of Theorem 2.4, the two-scale semiclassical measures ${\left(M_{\infty}\mathrm{d}\nu_{\infty},M\mathrm{d}\nu\right)}$ satisfy for all $a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)$,
+$\mathcal{I}^{\varepsilon_{k}}_{f}(a)\underset{k\to+\infty}{\longrightarrow}$ $\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\mathrm{tr}\Big(a\left(s,y,\sigma,\eta\right)\mathrm{d}\mu(s,y,\sigma,\eta)\Big)+\int_{\mathbb{R}\times\mathbb{S}^{2}}\mathrm{tr}\Big(a(s,0,0,0)M_{\infty}(s,\omega)\Big)\mathrm{d}\nu_{\infty}(s,\omega)$
+$+\ \int_{\mathbb{R}^{2}}\mathrm{Tr}\Big(a(s,0,0,0)M(s,\sigma)\Big)\mathrm{d}\nu(s,\sigma).$
+Even though we cannot describe the evolution of the two-scale semiclassical measures at infinity $M_{\infty}\mathrm{d}\nu_{\infty}$, we have a vanishing condition.
+Corollary 2.7 (A condition to have $M_{\infty}\mathrm{d}\nu_{\infty}=0$).
+Let ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$ uniformly bounded in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ such that there exists $\chi$ a cut off function with $\chi\in\mathscr{C}^{\infty}_{c}(\mathbb{R},\mathbb{R})$ supported in [1,1][-1,1] and $\chi\equiv 1$ on [1/2,1/2][-1/2,1/2] such that
+$\limsup_{\varepsilon\to 0}\left\lVert\mathrm{Op}_{1}\left(1-\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{R\sqrt{\varepsilon}}\right)\right)f^{\varepsilon}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\underset{R\to+\infty}{\longrightarrow}0.$
+Then, for all two-scale semiclassical measure $M_{\infty}\mathrm{d}\nu_{\infty}$ at infinity of ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$,
+$M_{\infty}\mathrm{d}\nu_{\infty}=0.$
+Proof.
+Let ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$ uniformly bounded in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ and $\chi$ a cut off function, $\chi\in\mathscr{C}^{\infty}_{c}(\mathbb{R},[0,1])$ supported in [1,1][-1,1] and $\chi\equiv 1$ on [1/2,1/2][-1/2,1/2] such that
+$\limsup_{\varepsilon\to 0}\left\lVert\mathrm{Op}_{1}\left(1-\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{R\sqrt{\varepsilon}}\right)\right)f^{\varepsilon}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\underset{R\to+\infty}{\longrightarrow}0.$
+For all $a\in\mathcal{A}$, we consider,
+$\alpha(s,y,\sigma,\eta)$ $\coloneqq a\left(s,y,\varepsilon\sigma,\varepsilon\eta,\frac{y}{\sqrt{\varepsilon}},\sqrt{\varepsilon}\sigma,\sqrt{\varepsilon}\eta\right)\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{\delta}\right),$
+$\beta(y,\sigma,\eta)$ $\coloneqq 1-\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{R\sqrt{\varepsilon}}\right).$
+According to pseudodifferential calculus and of the inequality (2.5), we fix $\delta$ and let $\varepsilon$ goes to 0 then $R$ goes to $\infty$, so
+$\mathrm{Op}_{1}\Big(\alpha(s,y,\sigma,\eta)\beta(y,\sigma,\eta)\Big)=\mathrm{Op}_{1}\Big(\alpha(s,y,\sigma,\eta)\Big)\mathrm{Op}_{1}\Big(\beta(y,\sigma,\eta)\Big)+\mathcal{O}\left(\sqrt{\varepsilon}\right)+\mathcal{O}\left(\frac{1}{R}\right),$
+where remainders are taken in $\mathcal{L}\left(\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)$. Moreover,
+$\mathrm{Op}_{1}\Big(\alpha(s,y,\sigma,\eta)\Big)=\Lambda_{\sqrt{\varepsilon}}^{\dagger}\mathrm{Op}_{1}\left(a\left(s,\sqrt{\varepsilon}y,\varepsilon\sigma,\sqrt{\varepsilon}\eta,y,\sqrt{\varepsilon}\sigma,\eta\right)\chi\left(\frac{\sqrt{\varepsilon}|(y,\sqrt{\varepsilon}\sigma,\eta)|}{\delta}\right)\right)\Lambda_{\sqrt{\varepsilon}}.$
+By application of the Calderón-Vaillancourt Theorem Calderón and Vaillancourt (1971),
+$\left\lVert\mathrm{Op}_{1}\Big(\alpha(s,y,\sigma,\eta)\Big)\right\rVert_{\mathcal{L}\left(\mathrm{L}^{2}\right)}$ $=\left\lVert\mathrm{Op}_{1}\left(a\left(s,\sqrt{\varepsilon}y,\varepsilon\sigma,\sqrt{\varepsilon}\eta,y,\sqrt{\varepsilon}\sigma,\eta\right)\chi\left(\frac{\sqrt{\varepsilon}|(y,\sqrt{\varepsilon}\sigma,\eta)|}{\delta}\right)\right)\right\rVert_{\mathcal{L}\left(\mathrm{L}^{2}\right)}$
+$\leqslant C(a)$
+where $C$ does not depend of $\varepsilon$. Then
+$\mathcal{I}^{\varepsilon}_{f}\left(a_{\varepsilon,\delta}^{R}\right)=\Big\langle\mathrm{Op}_{1}\Big(\alpha(s,y,\sigma,\eta)\Big)f^{\varepsilon},\mathrm{Op}_{1}\Big(\beta(y,\sigma,\eta)\Big)f^{\varepsilon}\Big\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}+\mathcal{O}\left(\sqrt{\varepsilon}\right)+\mathcal{O}\left(\frac{1}{R}\right),$
+so,
+$\left|\mathcal{I}^{\varepsilon}_{f}\left(a_{\varepsilon,\delta}^{R}\right)\right|\leqslant$ $\ \left\lVert\mathrm{Op}_{1}\Big(\alpha(s,y,\sigma,\eta)\Big)f^{\varepsilon}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\left\lVert\mathrm{Op}_{1}\Big(\beta(y,\sigma,\eta)\Big)f^{\varepsilon}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}+\mathcal{O}\left(\sqrt{\varepsilon}\right)+\mathcal{O}\left(\frac{1}{R}\right),$
+$\leqslant$ $\ C(a)\left\lVert f^{\varepsilon}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\left\lVert\mathrm{Op}_{1}\left(1-\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{R\sqrt{\varepsilon}}\right)\right)f^{\varepsilon}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}+\mathcal{O}\left(\sqrt{\varepsilon}\right)+\mathcal{O}\left(\frac{1}{R}\right).$
+Then,
+$\limsup_{\delta\to 0}\limsup_{R\to+\infty}\limsup_{\varepsilon\to 0}\mathcal{I}^{\varepsilon}_{f}\left(a_{\varepsilon,\delta}^{R}\right)=0.$
+We conclude by using Remark 2.5. ∎
+Time dependency.
+As in Section 1.2, one can extend the definition of two-scale semiclassical measures to time-dependent families.
+More precisely, we consider families ${\left(f^{\varepsilon}_{t}\right)_{\varepsilon>0}}$ that are uniformly bounded in $\mathrm{L}^{\infty}\left(\mathbb{R},\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)$.
+We consider
+$\mathcal{I}^{\varepsilon}_{f}(\Xi,a)\coloneqq\int_{\mathbb{R}}\Xi(t)\left\langle\mathbf{op}^{\mathcal{C}}_{\varepsilon}(a)f_{t}^{\varepsilon},f_{t}^{\varepsilon}\right\rangle_{\mathrm{L}^{2}(\mathbb{R}^{2},\mathbb{C}^{2})}\mathrm{d}t,\quad a\in\mathcal{A},\quad\Xi\in\mathscr{C}_{c}^{\infty}(\mathbb{R},\mathbb{C}),$
+As before, we are interested in the limit of $\mathcal{I}^{\varepsilon}_{f}(\Xi,a)$ as $\varepsilon$ goes to 0.
+Adapting the context of Kammerer (2000) to time-dependent families, one defines two-scale semiclassical measures over the curve $\mathcal{C}$.
+Indeed, by the same argument of Section 1.2, we can extend Theorem 2.4 to time-dependent families.
+Therefore, as before, the limit measures will be absolutely continuous in time variable with respect to Lebesgue measure.
+2.4 Rescaling
+According to the expression of the rescaling operator $\Lambda_{\sqrt{\varepsilon}}$ defined in (2.4), we consider the family $\left(u^{\varepsilon}_{t}\right)_{\varepsilon>0}$ defined such that
+$\Lambda_{\sqrt{\varepsilon}}^{\dagger}\left[u^{\varepsilon}_{t}\right](s,y)\coloneqq\varphi_{t}^{\varepsilon}(s,y),\quad\quad\forall(t,s,y)\in\mathbb{R}\times\mathbb{R}\times\mathrm{I}_{1/2}.$
+Although the family $\left(u^{\varepsilon}_{t}\right)_{\varepsilon>0}$ is defined on $\mathbb{R}\times\mathrm{I}_{\frac{1}{2\sqrt{\varepsilon}}}$, we will study it on $\mathbb{R}\times\mathrm{I}_{\sqrt{\varepsilon}}^{\delta}$ with
+$\mathrm{I}_{\sqrt{\varepsilon}}^{\delta}\coloneqq\left(-\frac{\delta}{\sqrt{\varepsilon}},\frac{\delta}{\sqrt{\varepsilon}}\right)$
+[…]
+Lemma 2.8 (Double scaling operator).
+The family $\left(u^{\varepsilon}_{t}\right)_{\varepsilon>0}$ satisfies on $\mathbb{R}\times\mathrm{I}_{\sqrt{\varepsilon}}^{\delta}$,
+$\left(\sqrt{\varepsilon}\mathrm{D}_{t}+y\left|\nabla\mathrm{m}(\mathbf{t}(s))\right|\sigma_{1}+\mathrm{D}_{y}\sigma_{2}-\frac{\sqrt{\varepsilon}}{1+\sqrt{\varepsilon}y\kappa(s)}\mathrm{D}_{s}\sigma_{3}+\frac{\mathrm{i}\varepsilon y\kappa^{\prime}(s)}{2(1+\sqrt{\varepsilon}y\kappa(s))^{2}}\sigma_{3}\right)u_{t}^{\varepsilon}=-\sqrt{\varepsilon}\mathrm{R}^{\varepsilon}_{2}u_{t}^{\varepsilon},$ (2.12)
+where
+$\big(\mathrm{R}_{2}^{\varepsilon}u_{t}^{\varepsilon}\big)(s,y)\coloneqq y^{2}\int_{0}^{1}\nabla^{2}\mathrm{m}\Big(\Phi(s,\tau\sqrt{\varepsilon}y)\Big)\cdot\Big(\mathbf{n}(s),\mathbf{n}(s)\Big)\mathrm{d}\tau\sigma_{1}u_{t}^{\varepsilon}(s,y),\quad(s,y)\in\mathbb{R}\times\mathrm{I}_{\sqrt{\varepsilon}}^{\delta}.$
+Proof.
+On the one hand, we rewrite the equation (2.2) with the definition of $\left(u^{\varepsilon}_{t}\right)_{\varepsilon>0}$ on $\mathbb{R}\times\mathrm{I}_{\sqrt{\varepsilon}}^{\delta}$,
+$\left[\varepsilon\mathrm{D}_{t}+\mathrm{m}\Big(\Phi(s,\sqrt{\varepsilon}y)\Big)\sigma_{1}+\sqrt{\varepsilon}\mathrm{D}_{y}\sigma_{2}-\frac{\varepsilon}{1+\sqrt{\varepsilon}y\kappa(s)}\mathrm{D}_{s}\sigma_{3}+\frac{\mathrm{i}\varepsilon^{3/2}y\kappa^{\prime}(s)}{2(1+\sqrt{\varepsilon}y\kappa(s))^{2}}\sigma_{3}\right]u^{\varepsilon}_{t}=0.$
+On the other hand, for all $(s,y)\in\mathbb{R}\times\mathrm{I}_{\sqrt{\varepsilon}}^{\delta}$,
+$\mathrm{m}\left(\Phi(s,\sqrt{\varepsilon}y)\right)=$ $\ \underbrace{\mathrm{m}(\mathbf{t}(s))}_{=0}+\sqrt{\varepsilon}y\underbrace{\nabla\mathrm{m}(\mathbf{t}(s))\cdot\mathbf{n}(s)}_{=|\nabla\mathrm{m}(\mathbf{t}(s))|}$
+$+\ \varepsilon y^{2}\int_{0}^{1}(1-\tau)\nabla^{2}\mathrm{m}\Big(\Phi(s,\tau\sqrt{\varepsilon}y)\Big)\cdot\Big(\mathbf{n}(s),\mathbf{n}(s)\Big)\mathrm{d}\tau.$
+The operator $\sqrt{\varepsilon}\mathrm{R}^{\varepsilon}_{2}$ will play a role of remainder in the determination of the two-scale semiclassical measure at finite distance $M\mathrm{d}\nu$.
+For now, we make a few remarks about the two-scale operator.
+The quantity $\sqrt{\varepsilon}$ is in front of $\mathrm{D}_{s}$ while no power of $\varepsilon$ multiplies $\mathrm{D}_{y}$.
+It fits with the two-scale operators that we have introduced in Section 1.5.
+It leads us with a semiclassical operator in the variable $s$ at scale $\sqrt{\varepsilon}$ and a classical operator in the variable $y$.
+Because of the second-scale of the problem, we will consider the symbol of the principal term of equation (2.12) only with respect to the variable $s$, keeping the operator character over the variable $y$ as mentioned earlier in Section 1.5.
+In fact, we had introduced the operator $\mathrm{T}_{\mathrm{E}}$, the symbol of an operator acting on $\mathrm{L}^{2}\left(\mathbb{R}_{s},\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)\right)$ in (2.10), its eigenvalues in (1.16) with the associated normalized eigenvectors in (1.14) and the associated eigenprojectors in (1.15).
+Let us define the symbol of $\mathrm{R}^{\varepsilon}_{2}$ for the two-scale quantization (2.10) as the operator acting on $\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)$ such that
+$\mathfrak{r}^{\varepsilon}_{2}(s,\sigma)\coloneqq y^{2}\int_{0}^{1}\nabla^{2}\mathrm{m}\Big(\Phi(s,\tau\sqrt{\varepsilon}y)\Big)\cdot\Big(\mathbf{n}(s),\mathbf{n}(s)\Big)\mathrm{d}\tau\sigma_{1},\quad(s,\sigma)\in\mathbb{R}^{2}.$
+Moreover, we define the operator acting on $\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)$ such that
+$\mathfrak{r}^{\varepsilon}_{3}(s,\sigma)\coloneqq\frac{y\kappa(s)\sigma}{1+\sqrt{\varepsilon}y\kappa(s)}\sigma_{3},\quad(s,\sigma)\in\mathbb{R}^{2}.$
+We define $\mathfrak{r}^{\varepsilon}\coloneqq\mathfrak{r}^{\varepsilon}_{2}+\mathfrak{r}^{\varepsilon}_{3}$. With the notation (2.10), the equation (2.12) writes
+$\sqrt{\varepsilon}\mathrm{D}_{t}u^{\varepsilon}_{t}+\mathrm{Op}_{\sqrt{\varepsilon},1}\Big(\mathrm{T}_{\mathrm{E}}+\sqrt{\varepsilon}\mathfrak{r}^{\varepsilon}\Big)u^{\varepsilon}_{t}=0.$
+3 Proof of the main result
+In this section we describe the evolution of the two-scale semiclassical measures defined in Theorem 1.4 for $\left(\varphi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ solutions to equation (2.2).
+In Section 3.1.1, we establish the evolution of the measure at finite distance $M^{t}\mathrm{d}\nu^{t}$.
+In Section 3.1.2, we establish the evolution of the measure at infinity $M_{\infty}^{t}\mathrm{d}\nu_{\infty}^{t}$.
+In Section 3.2, we establish Theorem 1.5.
+In Section 3.3, we finally establish Corollaries 1.6, 1.7, 1.8 and 1.9.
+3.1 Two-scale semiclassical measures above the interface
+3.1.1 Two-scale Wigner measure at finite distance
+We prove here the first point of Theorem 1.5, which is the evolution of the measure at finite distance $M^{t}\mathrm{d}\nu^{t}$. We recall the following lemma of Gérard et al. (2000).
+Lemma 3.1 (Poisson bracket and projector properties).
+Let $\left(\Pi_{n}\right)_{n\in\mathbb{N}}$ be a family of orthogonal projectors such that for all $(n,k)\in\mathbb{N}^{2},\Pi_{n}\Pi_{k}=\delta_{k,n}$. Then
+$\Pi_{k}\Big(\left\{\Pi_{j},\Pi_{\ell}\right\}-\left\{\Pi_{\ell},\Pi_{j}\right\}\Big)\Pi_{k}=0,\quad(j,k,\ell)\in\mathbb{N}^{3}.$
+Proof of the first point of Theorem 1.5.
+Assume Assumption 1.1 and Assumption 1.3. For convenience, we assume that the sequence $\left(\varepsilon\right)_{\varepsilon>0}$ realizes the measure $M^{t}\mathrm{d}\nu^{t}$ of Theorem 1.4.
+Let $a\in\mathcal{A}$ and $R,\delta$ as defined in (2.5).
+We consider the part (2.8) of $a$ denoted $a_{\varepsilon,R}$ and $b(s,\sigma)\coloneqq a^{W}(s,0,0,0,y,\sigma,\mathrm{D}_{y})\chi\left(\frac{|(y,\sigma,\mathrm{D}_{y})|}{R}\right)$, the operator acting on $\mathcal{H}\coloneqq\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)$, obtained by Weyl quantization.
+The link between the second-scale quantization of $a_{\varepsilon,R}$ (1.11) and the two-scale quantization of $b$ (2.10) was established in (2.11).
+By considering $\left(u^{\varepsilon}_{t}\right)_{\varepsilon>0}$ solution to equation (2.12) as a family of $\mathrm{L}^{2}\left(\mathbb{R}_{s},\mathcal{H}\right)$, we have
+$\mathrm{i}\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b)u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle=\frac{1}{\sqrt{\varepsilon}}\left\langle\Big[\mathrm{Op}_{\sqrt{\varepsilon},1}(b),\mathrm{Op}_{\sqrt{\varepsilon},1}\Big(\mathrm{T}_{\mathrm{E}}\Big)\Big]u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle+\left\langle\Big[\mathrm{Op}_{\sqrt{\varepsilon},1}(b),\mathrm{Op}_{\sqrt{\varepsilon},1}\big(\mathfrak{r}^{\varepsilon}\big)\Big]u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle.$
+Because of (1.7), we have
+$\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b)u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle=$ $\ \frac{1}{\mathrm{i}\sqrt{\varepsilon}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big(\left[b,\mathrm{T}_{\mathrm{E}}\right]\Big)u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle$
+$+\frac{1}{\mathrm{i}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\left(\frac{1}{2\mathrm{i}}\left\{b,\mathrm{T}_{\mathrm{E}}\right\}-\frac{1}{2\mathrm{i}}\left\{\mathrm{T}_{\mathrm{E}},b\right\}+\left[b,\mathfrak{r}^{\varepsilon}\right]\right)u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle$
+$+\mathcal{O}_{R}\left(\sqrt{\varepsilon}\right).$
+  • $\bullet$
+    Step 1 : Structure of $M^{t}$
+Let $\Xi\in\mathscr{C}^{\infty}_{c}(\mathbb{R})$, then
+$\int_{\mathbb{R}}\Xi(t)\mathrm{i}\sqrt{\varepsilon}\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b)u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\mathrm{d}t=$ $\int_{\mathbb{R}}\Xi(t)\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big(\left[b,\mathrm{T}_{\mathrm{E}}\right]\Big)u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\mathrm{d}t$
+$+\frac{\sqrt{\varepsilon}}{2\mathrm{i}}\int_{\mathbb{R}}\Xi(t)\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big(\left\{b,\mathrm{T}_{\mathrm{E}}\right\}-\left\{\mathrm{T}_{\mathrm{E}},b\right\}\Big)u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\mathrm{d}t$
+$+\sqrt{\varepsilon}\int_{\mathbb{R}}\Xi(t)\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big(\left[b,\mathfrak{r}^{\varepsilon}\right]\Big)u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\mathrm{d}t+\mathcal{O}_{R}\left(\varepsilon\right)$
+$=$ $-\mathrm{i}\sqrt{\varepsilon}\int_{\mathbb{R}}\Xi^{\prime}(t)\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b)u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\mathrm{d}t+\mathcal{O}_{R}\left(\varepsilon\right).$
+But $b$ is compactly supported in $\left\{\left|(y,\sigma)\right|\leqslant R\right\}$ and for all $(s,\sigma,y)\in\mathbb{R}\times\left(-R,R\right)^{2}$,
+$y\underbrace{\frac{\kappa(s)}{1+\sqrt{\varepsilon}y\kappa(s)}}_{\in\mathrm{L}^{\infty}\big(\mathbb{R}\times\left(-R,R\right)\big)}\sigma=\mathcal{O}\left(R^{2}\right),\quad\quad y^{2}\underbrace{\int_{0}^{1}\nabla^{2}\mathrm{m}\Big(\Phi(s,\tau\sqrt{\varepsilon}y)\Big)\cdot\Big(\mathbf{n}(s),\mathbf{n}(s)\Big)\mathrm{d}\tau}_{\in\mathrm{L}^{\infty}\big(\mathbb{R}\times\left(-R,R\right)\big)}=\mathcal{O}\left(R^{2}\right).$
+Then
+$\sqrt{\varepsilon}\int_{\mathbb{R}}\Xi(t)\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big([b,\mathfrak{r}^{\varepsilon}]\Big)u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\mathrm{d}t=\mathcal{O}_{R}\left(\sqrt{\varepsilon}\right).$
+By passing to the limit, we have, for all $\Xi\in\mathscr{C}^{\infty}_{c}(\mathbb{R})$ and all $a\in\mathcal{A}$,
+$0$ $=\int_{\mathbb{R}^{3}}\Xi(t)\mathrm{Tr}_{\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)}\Bigg(\Big[a(s,0,0,y,\sigma,\mathrm{D}_{y}),\mathrm{T}_{\mathrm{E}}(s,\sigma)\Big]M^{t}(s,\sigma)\Bigg)\mathrm{d}\nu^{t}(s,\sigma)\mathrm{d}t$
+$=\int_{\mathbb{R}^{3}}\Xi(t)\mathrm{Tr}\Bigg(a(s,0,0,y,\sigma,\mathrm{D}_{y})\Big[\mathrm{T}_{\mathrm{E}}(s,\sigma),M^{t}(s,\sigma)\Big]\Bigg)\mathrm{d}\nu^{t}(s,\sigma)\mathrm{d}t.$
+So, $t$ a.e. $\nu^{t}$ a.e. $\Big[\mathrm{T}_{\mathrm{E}}(s,\sigma),M^{t}(s,\sigma)\Big]=0$, which implies (1.19), according to the spectral decomposition of the operator $\mathrm{T}_{\mathrm{E}}$.
+  • $\bullet$
+    Step 2 : Equation of $\left(\nu_{n}^{t}\right)_{n\in\mathbb{Z}}$
+By Step $1$, for all $(k,\ell)\in\mathbb{Z}^{2}$, with $k\neq\ell$ and for all $\beta$ function from $\mathbb{R}_{s}\times\mathbb{R}_{\sigma}$ to $\mathcal{L}\left(\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)\right)$ such that $\Pi_{k}\beta\Pi_{\ell}\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathcal{L}\left(\mathcal{H}\right)\right)$, we have
+$\limsup_{\delta\to 0}\limsup_{R\to+\infty}\lim_{\varepsilon\to 0}\int_{\mathbb{R}}\Xi(t)\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big(\Pi_{k}\beta\Pi_{\ell}\Big)u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\mathrm{d}t=0.$
+Then we can consider symbol of the form $\Pi_{\ell}\beta\Pi_{\ell}$. Since the projector $\Pi_{\ell}$ has rank one, we have
+$\Pi_{\ell}\beta\Pi_{\ell}=\mathrm{Tr}\Big(\Pi_{\ell}\beta\Pi_{\ell}\Big)\Pi_{\ell},$
+where $\mathrm{Tr}_{\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)}\big(\Pi_{\ell}\beta\Pi_{\ell}\big)\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathbb{C}\right)$.
+Therefore we will consider symbols of the form $\beta\Pi_{\ell}$ with $\beta\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathbb{C}\right)$.
+Let $n\in\mathbb{Z}$ and $\beta\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathbb{C}\right)$.
+Let us define
+$b_{n}(s,\sigma)\coloneqq\beta(s,\sigma)\Pi_{n}(s,\sigma),\quad(s,\sigma)\in\mathbb{R}^{2}.$
+The expression of $b_{n}$ implies $b_{n},\mathrm{T}_{\mathrm{E}}]=0$. Then
+$\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b_{n})u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle=$ $\ \frac{1}{2}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big(\{\mathrm{T}_{\mathrm{E}},b_{n}\}-\{b_{n},\mathrm{T}_{\mathrm{E}}\}\Big)u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle$
+$+\frac{1}{\mathrm{i}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big([b_{n},\mathfrak{r}^{\varepsilon}]\Big)u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle+\mathcal{O}_{R}\left(\sqrt{\varepsilon}\right).$
+We first show that the term with $\mathfrak{r}^{\varepsilon}$ will vanish as $\varepsilon$ goes to 0. By definition of $M^{t}\mathrm{d}\nu^{t}$, we have
+$\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\big([b_{n},\mathfrak{r}^{\varepsilon}]\big)u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle$ $\underset{\varepsilon\to 0}{\longrightarrow}\int_{\mathbb{R}^{2}}\mathrm{Tr}\Bigg(\Big[a(s,0,0,y,\sigma,D_{y}),\mathfrak{r}^{0}(s,\sigma)\Big]M^{t}(s,\sigma)\Bigg)\mathrm{d}\nu^{t}(s,\sigma)$
+$=\sum_{m\in\mathbb{Z}}\int_{\mathbb{R}^{2}}\mathrm{Tr}\Bigg(\Big[\beta(s,\sigma)\Pi_{n}(s,\sigma),\mathfrak{r}^{0}(s,\sigma)\Big]\Pi_{m}(s,\sigma)\Bigg)\mathrm{d}\nu^{t}_{m}(s,\sigma),$
+$=\sum_{m\in\mathbb{Z}}\int_{\mathbb{R}^{2}}\beta(s,\sigma)\mathrm{Tr}\Bigg(\Big[\Pi_{n}(s,\sigma),\mathfrak{r}^{0}(s,\sigma)\Big]\Pi_{m}(s,\sigma)\Bigg)\mathrm{d}\nu^{t}_{m}(s,\sigma),$
+where the second equality holds because of Step $1$ and the chosen shape of $a$ and the last equality holds because $\beta$ is scalar and $\mathfrak{r}^{0}$ is the value at zero of $\mathfrak{r}^{\varepsilon}$ given by for all $(s,\sigma)\in\mathbb{R}^{2}$,
+$\mathfrak{r}^{0}(s,\sigma)=\mathfrak{r}^{0}_{2}(s,\sigma)+\mathfrak{r}^{0}_{3}(s,\sigma)=y^{2}\nabla^{2}\mathrm{m}\Big(\Phi(s,0)\Big)\cdot\Big(\mathbf{n}(s),\mathbf{n}(s)\Big)\sigma_{1}+y\kappa(s)\sigma\sigma_{3}.$
+Then, we recall, for all $m\in\mathbb{Z}$,
+$\mathrm{Tr}\Big(\big[\Pi_{n},\mathfrak{r}^{0}\big]\Pi_{m}\Big)$ $=\mathrm{Tr}\Big(\Pi_{n}\mathfrak{r}^{0}\Pi_{m}\Big)-\mathrm{Tr}\Big(\mathfrak{r}^{0}\Pi_{n}\Pi_{m}\Big)=\mathrm{Tr}\Big(\mathfrak{r}^{0}\Pi_{m}\Pi_{n}\Big)-\mathrm{Tr}\Big(\mathfrak{r}^{0}\Pi_{n}\Pi_{m}\Big)$
+$=\delta_{n=m}\left(\mathrm{Tr}\Big(\mathfrak{r}^{0}\Pi_{n}\Big)-\mathrm{Tr}\Big(\mathfrak{r}^{0}\Pi_{n}\Big)\right)=0.$
+Let $k\in\mathbb{Z}$, let us compute $\Pi_{k}\{\mathrm{T}_{\mathrm{E}},b_{n}\}\Pi_{k}$,
+$\Pi_{k}\{\mathrm{T}_{\mathrm{E}},b_{n}\}\Pi_{k}=\Pi_{k}\left\{\sum_{m\in\mathbb{Z}}\lambda_{m}\Pi_{m}\,,\,\beta\Pi_{n}\right\}\Pi_{k}$ (3.1)
+then,
+$\Pi_{k}\{\mathrm{T}_{\mathrm{E}},b_{n}\}\Pi_{k}=\Pi_{k}\sum_{m\in\mathbb{Z}}\Bigg(\Pi_{m}\left\{\lambda_{m},\beta\right\}\Pi_{n}+\lambda_{m}\left\{\Pi_{m},\beta\right\}\Pi_{n}+\lambda_{m}\beta\left\{\Pi_{m},\Pi_{n}\right\}+\beta\Pi_{m}\left\{\lambda_{m},\Pi_{n}\right\}\Bigg)\Pi_{k}.$
+Let $m\in\mathbb{Z}$,
+$\Pi_{k}\Pi_{m}\left\{\lambda_{m},\beta\right\}\Pi_{n}\Pi_{k}$ $=\mathds{1}_{k=m=n}\left\{\lambda_{n},\beta\right\}\Pi_{n},$
+$\Pi_{k}\lambda_{m}\left\{\Pi_{m},\beta\right\}\Pi_{n}\Pi_{k}$ $=\mathds{1}_{n=k}\lambda_{m}\Pi_{n}\left\{\Pi_{m},\beta\right\}\Pi_{n}\ =\ 0,$
+$\Pi_{k}\lambda_{m}\beta\left\{\Pi_{m},\Pi_{n}\right\}\Pi_{k}$ $=\lambda_{m}\beta\Pi_{k}\left\{\Pi_{m},\Pi_{n}\right\}\Pi_{k},$
+$\Pi_{k}\beta\Pi_{m}\left\{\lambda_{m},\Pi_{n}\right\}\Pi_{k}$ $=\mathds{1}_{k=m}\beta\Pi_{m}\left\{\lambda_{m},\Pi_{n}\right\}\Pi_{m}\ =\ 0,$
+where the two vanishing equations hold because $\lambda_{n}$ and $\beta$ are scalar and $\Pi_{k}\partial\Pi_{m}\Pi_{k}$ is always equal to zero. Moreover, according to Lemma 3.1,
+$\lambda_{m}\beta\Pi_{k}\Big(\left\{\Pi_{m},\Pi_{n}\right\}-\left\{\Pi_{n},\Pi_{m}\right\}\Big)\Pi_{k}=0\,,\,\text{for all}\ m\in\mathbb{Z}.$
+Therefore,
+$\{\mathrm{T}_{\mathrm{E}},b_{n}\}-\{b_{n},\mathrm{T}_{\mathrm{E}}\}=\Big(\left\{\lambda_{n},\beta\right\}-\left\{\beta,\lambda_{n}\right\}\Big)\Pi_{n}.$
+Then,
+$\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b_{n})u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle=\frac{1}{2}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big(\Big(\left\{\lambda_{n},\beta\right\}-\left\{\beta,\lambda_{n}\right\}\Big)\Pi_{n}\Big)u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\ +\ \mathcal{O}_{R}\left(\sqrt{\varepsilon}\right).$
+By passing to the limit,
+$\int_{\mathbb{R}}\int_{\mathbb{R}^{2}}\Xi^{\prime}(t)$ $\mathrm{Tr}\Big(a^{W}(s,0,0,0,y,\sigma,\mathrm{D}_{y})M^{t}(s,\sigma)\Big)\mathrm{d}\nu^{t}(s,\sigma)\mathrm{d}t$
+$=\int_{\mathbb{R}}\int_{\mathbb{R}^{2}}\Xi(t)\mathrm{Tr}\Big(\left\{\lambda_{n},\beta\right\}\Pi_{n}(s,0,0,0,y,\sigma,\mathrm{D}_{y})M^{t}(s,\sigma)\Big)\mathrm{d}\nu^{t}(s,\sigma)\mathrm{d}t,$
+$=\int_{\mathbb{R}}\int_{\mathbb{R}^{2}}\Xi(t)\mathrm{Tr}\Big(\left\{\lambda_{n},\beta\right\}\Pi_{n}(s,0,0,0,y,\sigma,\mathrm{D}_{y})\Big)\mathrm{d}\nu^{t}_{n}(s,\sigma)\mathrm{d}t$
+$=\int_{\mathbb{R}}\int_{\mathbb{R}^{2}}\Xi(t)\left\{\lambda_{n},\beta\right\}(s,\sigma)\mathrm{d}\nu^{t}_{n}(s,\sigma)\mathrm{d}t,$
+$=-\int_{\mathbb{R}}\int_{\mathbb{R}^{2}}\Xi(t)\beta(s,\sigma)\left\{\lambda_{n},\mathrm{d}\nu^{t}_{n}\right\}(s,\sigma)\mathrm{d}t.$
+Then $\nu_{n}^{t}$ satisfies the equation (1.20) and is notably continuous on $\mathbb{R}$.
+It remains to prove that the value at zero of $\nu_{n}^{t}$ coincides with the measure of the initial condition $\nu_{n,0}$.
+Let $T\in\mathbb{R}$, on the one hand,
+$\int_{0}^{T}\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b_{n})u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\mathrm{d}t=\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b_{n})u^{\varepsilon}_{T},u^{\varepsilon}_{T}\right\rangle-\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b_{n})u^{\varepsilon}_{0},u^{\varepsilon}_{0}\right\rangle.$
+On the other hand,
+$\left|\int_{0}^{T}\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b_{n})u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\mathrm{d}t\right|$ $\leqslant\frac{1}{2}\int_{0}^{T}\left|\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big(\left\{\lambda_{n},\beta\right\}\Pi_{n}\Big)u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\right|\mathrm{d}t\ +\ \mathcal{O}_{T,R}\left(\sqrt{\varepsilon}\right),$
+$\leqslant\frac{C\left(\left\{\lambda_{n},\beta\right\}\Pi_{n}\right)T}{2}\left\lVert u^{\varepsilon}_{0}\right\rVert_{\mathrm{L}^{2}}^{2}\ +\ \mathcal{O}_{T,R}\left(\sqrt{\varepsilon}\right).$
+where the last inequality holds because of the Calderón-Vaillancourt Theorem (1.5) and the preservation of the $\mathrm{L}^{2}$-norm of the solution. Then, by passing to the limit, we have
+$\begin{split}\limsup_{T\to 0}\limsup_{\delta\to 0}\limsup_{R\to+\infty}\lim_{\varepsilon\to 0}\mathcal{I}^{\varepsilon}_{\varphi_{t}^{\varepsilon}}\left(a_{\varepsilon,R}\right)&=\int_{\mathbb{R}^{2}}\mathrm{Tr}\Big(a^{W}(s,0,0,0,y,\sigma,\mathrm{D}_{y})\Pi_{n}(s,\sigma)\Big)\mathrm{d}\nu_{n}^{0}(s,\sigma),\\ &=\limsup_{T\to 0}\limsup_{\delta\to 0}\limsup_{R\to+\infty}\lim_{\varepsilon\to 0}\mathcal{I}^{\varepsilon}_{\varphi_{0}^{\varepsilon}}\left(a_{\varepsilon,R}\right).\end{split}$
+Therefore, for all $n\in\mathbb{Z},\nu_{n}^{t}$ is continuous and the value at zero coincides with the initial condition.
+Moreover, the equality $M^{t}\mathrm{d}\nu^{t}=\sum_{n\in\mathbb{Z}}\Pi_{n}\mathrm{d}\nu_{n}^{t}$ holds as trace class operator so, thanks to Lebesgue dominated convergence Theorem, $M^{t}\mathrm{d}\nu^{t}$ is continuous on $\mathbb{R}$.
+∎
+Remark 3.2 (Zero value of the two-scale semiclassical measure at finite distance).
+We can construct the initial data inspired by the proof of the Lemma A.4 where we construct semiclassical measures for families valued in a separable Hilbert space.
+Let $\left(f^{\varepsilon}\right)_{\varepsilon>0}$ be a bounded family of $\mathrm{L}^{2}\left(\mathbb{R}_{s},\mathbb{C}\right)$ with the knowledge of $\mathrm{d}\rho$, the semiclassical measure of $\left(f^{\varepsilon}\right)_{\varepsilon>0}$ at the scale $\sqrt{\varepsilon}$.
+The function $(s,\sigma)\mapsto g_{1}^{s,\sigma}$ is smooth, bounded and has all derivatives bounded uniformly in $y$, therefore we can consider its Weyl $\sqrt{\varepsilon}$-quantization according to the variable $s$.
+Let us consider initial data $\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}$ such that for all $(s,y)\in\mathbb{R}\times\mathrm{I}_{1/2}$,
+$u_{0}^{\varepsilon}(s,y)=g_{0}^{s}(y)f(s)+\Big(\mathrm{Op}_{\sqrt{\varepsilon}}^{(s)}\big(g_{1}^{s,\sigma}(y)\big)f\Big)(s),$
+where the Weyl $\sqrt{\varepsilon}$-quantization of $g_{1}^{s,\sigma}$ is taken only in the $s$ variable.
+Then the two-scale semiclassical measure at finite distance of this initial data is for all $(s,\sigma)\in\mathbb{R}^{2}$,
+$\frac{1}{2\pi}\Big(\Pi_{0}(s)+\Pi_{1}(s,\sigma)+\Pi_{1,0}(s)+\Pi_{0,1}(s,\sigma)\Big)\mathrm{d}\rho(s,\sigma),$
+where for all $(k,\ell)\in\left\{0,1\right\}^{2}$ and for all $(s,\sigma)\in\mathbb{R}^{2},\Pi_{k,\ell}(s,\sigma)=g^{s,\sigma}_{k}\otimes g^{s,\sigma}_{\ell}$.
+Notably, it does not commute with $\mathrm{T}_{\mathrm{E}}(s,\sigma)$.
+3.1.2 Two-scale Wigner measure at infinity
+We prove here the second point of Theorem 1.5.
+We consider the symbol of $\mathrm{H}^{\mathrm{E}}_{\varepsilon}$ according to the second-scale quantization (1.11), given by for all $(s,y)\in\mathbb{R}^{2}$ and $z\coloneqq(z_{y},z_{\sigma},z_{\eta})\in\mathbb{R}^{3}$,
+$\mathcal{T}_{\mathrm{E}}^{\infty}(s,y,z)\coloneqq z_{y}\int_{0}^{1}\nabla\mathrm{m}\Big(\Phi(s,\tau y)\Big)\cdot\mathbf{n}(s)\mathrm{d}\tau\sigma_{1}+z_{\eta}\sigma_{2}-\frac{z_{\sigma}}{1+y\kappa(s)}\sigma_{3},$
+which has two eigenvalues :
+$\lambda_{\pm}^{\mathrm{E}}(s,y,z)\coloneqq\pm\sqrt{\frac{1}{(1+y\kappa(s))^{2}}z_{\sigma}^{2}+z_{y}^{2}\left|\int_{0}^{1}\nabla\mathrm{m}\Big(\Phi(s,\tau y)\Big)\cdot\mathbf{n}(s)\mathrm{d}\tau\right|^{2}+z_{\eta}^{2}},$
+which are homogeneous functions of order 1 in $z$. The eigenprojectors $\Pi_{\pm}^{\mathrm{E}}(s,y,z)$ are defined by
+$\Pi_{\pm}^{\mathrm{E}}(s,y,z)\coloneqq\frac{1}{2}\mathrm{Id}+\frac{1}{2\lambda_{\pm}^{\mathrm{E}}(s,y,z)}\mathcal{T}_{\mathrm{E}}^{\infty}(s,y,z)$
+and are homogeneous functions of order 0 in $z$.
+Moreover, we denote the principal symbol of the operator $\mathrm{H}_{\varepsilon}^{\mathrm{E}}$ associated with the second-scale quantization (1.11) by for all $(s,y)\in\mathbb{R}^{2}$ and for all $z\coloneqq(z_{y},z_{\sigma},z_{\eta})\in\mathbb{R}^{3}$,
+$\mathrm{T}^{\infty}_{\mathrm{E}}(s,z)=\mathcal{T}_{\mathrm{E}}^{\infty}(s,0,z)=\begin{pmatrix}z_{\sigma}&z_{y}r(s)^{2}-iz_{\eta}\\ z_{y}r(s)^{2}+iz_{\eta}&-z_{\sigma}\end{pmatrix}$,
+where $\mathrm{T}^{\infty}_{\mathrm{E}}$ was defined in (1.17). Its eigenvalues are
+$\lambda_{\pm}^{\infty}(s,z)\coloneqq\pm\sqrt{z_{\sigma}^{2}+z_{y}^{2}r(s)^{4}+z_{\eta}^{2}},$
+and the associated eigenprojectors $\Pi_{\pm}^{\infty}(s,z)$ were defined in (1.18).
+Then, the equation (2.2) writes on $\mathbb{R}_{t}\times\mathbb{R}_{s}\times\mathrm{I}_{1/2}$
+$\sqrt{\varepsilon}\mathrm{D}_{t}\varphi^{\varepsilon}_{t}+\mathbf{op}^{\mathcal{C}}_{\varepsilon}\Big(\mathcal{T}_{\mathrm{E}}^{\infty}\Big)\varphi^{\varepsilon}_{t}=0.$
+Proof of the second point of Theorem 1.5.
+Assume Assumption 1.1 and Assumption 1.3.
+For convenience, we assume that the sequence $\left(\varepsilon\right)_{\varepsilon>0}$ realizes the measure $M^{t}_{\infty}\mathrm{d}\nu^{t}_{\infty}$.
+Let $a\in\mathcal{A}$ and $R,\delta$ as defined in (2.5).
+We consider the part (2.7) of $a$ divided by $\lambda^{\infty}_{+}$.
+The symbol $\mathcal{T}_{\mathrm{E}}^{\infty}$ is homogeneous in $z$ of order $1$ and the symbol $a_{\varepsilon,\delta}^{R}/\lambda^{\infty}_{+}$ is homogeneous in $z$ of order 1-1 at infinity.
+Because of (2.7), we consider symbols on a subset of $\left\{z\in\mathbb{R}^{3}|\frac{R}{2}\leqslant\left|z\right|\leqslant\frac{2\delta}{\sqrt{\varepsilon}}\right\}$.
+Moreover, we have the following properties.
+  1. 1.
+    $\left\{\mathcal{T}_{\mathrm{E}}^{\infty},\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right\}_{z_{y},z_{\eta}}$ is homogeneous in $z$ of order 2-2.
+  2. 2.
+    $\left\{\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathcal{T}_{\mathrm{E}}^{\infty}\right\}_{s,z_{\sigma}}$ is homogeneous in $z$ of order 1-1.
+  3. 3.
+    $\left\{\mathcal{T}_{\mathrm{E}}^{\infty},\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right\}_{(s,y),(\sigma,\eta)}$ is homogeneous in $z$ of order 0.
+Then, according to standard semiclassical calculus,
+$\mathbf{op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\mathcal{T}_{\mathrm{E}}^{\infty},\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right\}_{z_{y},z_{\eta}}-\left\{\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathcal{T}_{\mathrm{E}}^{\infty}\right\}_{z_{y},z_{\eta}}\right)$ $=\mathcal{O}\left(\frac{1}{R^{2}}\right),$
+$\sqrt{\varepsilon}\mathbf{op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\mathcal{T}_{\mathrm{E}}^{\infty},\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right\}_{s,z_{\sigma}}-\left\{\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathcal{T}_{\mathrm{E}}^{\infty}\right\}_{s,z_{\sigma}}\right)$ $=\mathcal{O}_{R}\left(\sqrt{\varepsilon}\right),$
+$\varepsilon\mathbf{op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\mathcal{T}_{\mathrm{E}}^{\infty},\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right\}_{(s,y),(\sigma,\eta)}-\left\{\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathcal{T}_{\mathrm{E}}^{\infty}\right\}_{(s,y),(\sigma,\eta)}\right)$ $=\mathcal{O}_{R}\left(\varepsilon\right).$
+So
+$\left[\mathbf{op}^{\mathcal{C}}_{\varepsilon}\left(\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right),\mathbf{op}^{\mathcal{C}}_{\varepsilon}\left(\mathcal{T}_{\mathrm{E}}^{\infty}\right)\right]=$ $\ \mathbf{op}^{\mathcal{C}}_{\varepsilon}\left(\left[\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathcal{T}_{\mathrm{E}}^{\infty}\right]\right)$
+$+\frac{1}{2\mathrm{i}}\mathbf{op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathcal{T}_{\mathrm{E}}^{\infty}\right\}_{z_{y},z_{\eta}}-\left\{\mathcal{T}_{\mathrm{E}}^{\infty},\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right\}_{z_{y},z_{\eta}}\right)+\mathcal{O}\left(\frac{1}{R^{4}}\right)$
+$+\frac{\sqrt{\varepsilon}}{2\mathrm{i}}\mathbf{op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathcal{T}_{\mathrm{E}}^{\infty}\right\}_{s,z_{\sigma}}-\left\{\mathcal{T}_{\mathrm{E}}^{\infty},\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right\}_{s,z_{\sigma}}\right)+\mathcal{O}_{R}\left(\varepsilon\right).$
+  • $\bullet$
+    Step 1 : Structure of $M_{\infty}^{t}$
+Let $\Xi\in\mathscr{C}^{\infty}_{c}(\mathbb{R})$, then
+$\int_{\mathbb{R}}\Xi(t)\mathrm{i}\sqrt{\varepsilon}\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathbf{op}^{\mathcal{C}}_{\varepsilon}\left(\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right)\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle\mathrm{d}t=\int_{\mathbb{R}}\Xi(t)\left\langle\mathbf{op}^{\mathcal{C}}_{\varepsilon}\left(\left[\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathcal{T}_{\mathrm{E}}^{\infty}\right]\right)\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle\mathrm{d}t\\ +\frac{\sqrt{\varepsilon}}{2\mathrm{i}}\int_{\mathbb{R}}\Xi(t)\left\langle\mathbf{op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathcal{T}_{\mathrm{E}}^{\infty}\right\}_{s,z_{\sigma}}-\left\{\mathcal{T}_{\mathrm{E}}^{\infty},\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right\}_{s,z_{\sigma}}\right)\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle\mathrm{d}t+\ \mathcal{O}_{R}\left(\varepsilon\right)+\mathcal{O}\left(\frac{1}{R^{2}}\right)\\ =-\mathrm{i}\sqrt{\varepsilon}\int_{\mathbb{R}}\Xi^{\prime}(t)\left\langle\mathbf{op}^{\mathcal{C}}_{\varepsilon}\left(\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right)\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle\mathrm{d}t.$
+But $\left[\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathcal{T}_{\mathrm{E}}^{\infty}\right]=\left[a,\frac{\mathcal{T}_{\mathrm{E}}^{\infty}}{\lambda^{\infty}_{+}}\right]_{\varepsilon,\delta}^{R}$ and $\left[a,\frac{\mathcal{T}_{\mathrm{E}}^{\infty}}{\lambda^{\infty}_{+}}\right]\in\mathcal{A}$, then by passing to the limit, we deduce that for all $\Xi\in\mathscr{C}^{\infty}_{c}(\mathbb{R})$ and for all $a\in\mathcal{A}$,
+$0$ $=\int_{\mathbb{R}^{2}\times\mathbb{S}^{2}}\Xi(t)\mathrm{tr}\Bigg(\left[a,\frac{\mathcal{T}_{\mathrm{E}}^{\infty}}{\lambda^{\infty}_{+}}\right]_{\infty}(s,0,0,0,\omega)M_{\infty}^{t}(s,\omega)\Bigg)\mathrm{d}\nu_{\infty}^{t}(s,\omega)\mathrm{d}t$
+$=\int_{\mathbb{R}^{2}\times\mathbb{S}^{2}}\Xi(t)\mathrm{tr}\Bigg(\left[a_{\infty}(s,0,0,0,\omega),\frac{1}{\lambda^{\infty}_{+}(s,\omega)}\mathcal{T}_{\mathrm{E}}^{\infty}(s,0,\omega)\right]M_{\infty}^{t}(s,\omega)\Bigg)\mathrm{d}\nu_{\infty}^{t}(s,\omega)\mathrm{d}t$
+$=\int_{\mathbb{R}^{2}\times\mathbb{S}^{2}}\Xi(t)\frac{1}{\lambda^{\infty}_{+}(s,\omega)}\mathrm{tr}\Bigg(a_{\infty}(s,0,0,0,\omega)\Big[\mathrm{T}^{\infty}_{\mathrm{E}}(s,\omega),M_{\infty}^{t}(s,\omega)\Big]\Bigg)\mathrm{d}\nu_{\infty}^{t}(s,\omega)\mathrm{d}t.$
+So, for almost all $t\in\mathbb{R}$, we have $\nu_{\infty}^{t}$ a.e. $\Big[\mathrm{T}^{\infty}_{\mathrm{E}}(s,\omega),M_{\infty}^{t}(s,\omega)\Big]=0$ which implies (1.21).
+  • $\bullet$
+    Step 2 : Equation of $\mathrm{d}\nu_{\infty}^{t}$.
+Let $\beta\in\mathcal{A}$ and $a$ defined by for all $(s,y,\sigma,\eta,z)\in\mathbb{R}^{7}$,
+$a(s,y,\sigma,\eta,z)\coloneqq\lambda^{\infty}_{\pm}(s,z)\beta(s,y,\sigma,\eta,z)\Pi_{\pm}^{\mathrm{E}}(s,y,z).$
+The form of $a$ implies $\left[a,\mathcal{T}_{\mathrm{E}}^{\infty}\right]=0$. Then
+$\sqrt{\varepsilon}\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathbf{op}^{\mathcal{C}}_{\varepsilon}\left(a_{\varepsilon,\delta}^{R}\right)\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle=$ $\ \frac{1}{2}\left\langle\mathbf{op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\mathcal{T}_{\mathrm{E}}^{\infty},a_{\varepsilon,\delta}^{R}\right\}_{z_{y},z_{\eta}}-\left\{a_{\varepsilon,\delta}^{R},\mathcal{T}_{\mathrm{E}}^{\infty}\right\}_{z_{y},z_{\eta}}\right)\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle+\mathcal{O}\left(\frac{1}{R^{2}}\right)$
+$+\frac{\sqrt{\varepsilon}}{2}\left\langle\mathbf{op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\mathcal{T}_{\mathrm{E}}^{\infty},a_{\varepsilon,\delta}^{R}\right\}_{s,z_{\sigma}}-\left\{a_{\varepsilon,\delta}^{R},\mathcal{T}_{\mathrm{E}}^{\infty}\right\}_{s,z_{\sigma}}\right)\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle+\mathcal{O}_{R}\left(\varepsilon\right).$
+Because of the homogeneity in $z$, we have $\left\{\mathcal{T}_{\mathrm{E}}^{\infty},a\right\}_{z_{y},z_{\eta}}\in\mathcal{A}$. Moreover,
+$\mathbf{op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\mathcal{T}_{\mathrm{E}}^{\infty},a_{\varepsilon,\delta}^{R}\right\}_{z_{y},z_{\eta}}\right)=$ $\ \mathbf{op}^{\mathcal{C}}_{\varepsilon}\left(\left(\left\{\mathcal{T}_{\mathrm{E}}^{\infty},a\right\}_{z_{y},z_{\eta}}\right)_{\varepsilon,\delta}^{R}\right)+\frac{1}{R}\mathbf{op}^{\mathcal{C}}_{\varepsilon}\left(a\chi_{\frac{\delta}{\sqrt{\varepsilon}}}\left\{\mathcal{T}_{\mathrm{E}}^{\infty},\chi^{R}\right\}_{z_{y},z_{\eta}}\right)$
+$+\frac{\sqrt{\varepsilon}}{\delta}\mathbf{op}^{\mathcal{C}}_{\varepsilon}\left(a\chi^{R}\left\{\mathcal{T}_{\mathrm{E}}^{\infty},\chi_{\frac{\delta}{\sqrt{\varepsilon}}}\right\}_{z_{y},z_{\eta}}\right).$
+The function $\partial\chi^{R}$ is compactly supported in the ring $\left\{z\in\mathbb{R}^{3}|\frac{R}{2}\leqslant\left|z\right|\leqslant R\right\}$ and the function $a\chi_{\frac{\delta}{\sqrt{\varepsilon}}}\left\{\mathcal{T}_{\mathrm{E}}^{\infty},\chi^{R}\right\}_{z_{y},z_{\eta}}$ belongs to $\mathcal{A}$ so
+$\left(a\chi_{\frac{\delta}{\sqrt{\varepsilon}}}\left\{\mathcal{T}_{\mathrm{E}}^{\infty},\chi^{R}\right\}_{z_{y},z_{\eta}}\right)_{\infty}=0.$
+The operator $\mathbf{op}^{\mathcal{C}}_{\varepsilon}\left(a\chi^{R}\left\{\mathcal{T}_{\mathrm{E}}^{\infty},\chi_{\frac{\delta}{\sqrt{\varepsilon}}}\right\}_{z_{y},z_{\eta}}\right)$ is uniformly bounded in $\varepsilon$ then
+$\lim_{\varepsilon\to 0}\frac{\sqrt{\varepsilon}}{\delta}\left\langle\mathbf{op}^{\mathcal{C}}_{\varepsilon}\left(a\chi^{R}\left\{\mathcal{T}_{\mathrm{E}}^{\infty},\chi_{\frac{\delta}{\sqrt{\varepsilon}}}\right\}_{z_{y},z_{\eta}}\right)\varphi_{t}^{\varepsilon},\varphi_{t}^{\varepsilon}\right\rangle=0.$
+As in the computation of (3.1), we consider the quantity $\Pi_{+}^{\mathrm{E}}\left\{a,\mathcal{T}_{\mathrm{E}}^{\infty}\right\}_{z_{y},z_{\eta}}\Pi_{+}^{\mathrm{E}}$ then
+$\Pi_{+}^{\mathrm{E}}\left\{a,\mathcal{T}_{\mathrm{E}}^{\infty}\right\}_{z_{y},z_{\eta}}\Pi_{+}^{\mathrm{E}}=\lambda_{+}^{\infty}\left\{\beta,\lambda_{+}^{\infty}\right\}_{z_{y},z_{\eta}}\Pi_{+}^{\mathrm{E}}.$
+Then, by passing to the limit, we have
+$0$ $=\int_{\mathbb{R}^{2}\times\mathbb{S}^{2}}\Xi(t)\mathrm{tr}\Bigg(\left(\lambda_{+}^{\infty}\left\{\beta,\lambda_{+}^{\infty}\right\}_{z_{y},z_{\eta}}\Pi_{+}^{\mathrm{E}}\right)_{\infty}(s,0,0,0,\omega)M_{\infty}^{t}(s,\omega)\Bigg)\mathrm{d}\nu_{\infty}^{t}(s,\omega)\mathrm{d}t,$
+$=\int_{\mathbb{R}^{2}\times\mathbb{S}^{2}}\Xi(t)\mathrm{tr}\Bigg(\lambda_{+}^{\infty}\left\{\beta_{\infty},\lambda_{+}^{\infty}\right\}_{\omega_{y},\omega_{\eta}}(s,0,0,0,\omega)\Pi_{+}^{\infty}(s,\omega)\Bigg)\mathrm{d}\nu_{+}^{t}(s,\omega)\mathrm{d}t,$
+$=\int_{\mathbb{R}^{2}\times\mathbb{S}^{2}}\Xi(t)\lambda_{+}^{\infty}\left\{\beta_{\infty},\lambda_{+}^{\infty}\right\}_{\omega_{y},\omega_{\eta}}(s,0,0,0,\omega)\mathrm{d}\nu_{+}^{t}(s,\omega)\mathrm{d}t,$
+but
+$\lambda_{+}^{\infty}(s,\omega)\frac{\partial\lambda_{+}^{\infty}}{\partial z_{y}}(s,\omega)=\omega_{y}r(s)^{4},\quad\lambda_{+}^{\infty}(s,\omega)\frac{\partial\lambda_{+}^{\infty}}{\partial z_{\eta}}(s,\omega)=\omega_{\eta},\quad(s,\omega)\in\mathbb{R}\times\mathbb{S}^{2}.$
+Then
+$\int_{\mathbb{R}^{2}\times\mathbb{S}^{2}}\Xi(t)\Big(\vec{V}_{+}(s,\omega)\cdot\nabla_{\omega}\beta_{\infty}(s,\omega)\Big)\mathrm{d}\nu_{+}^{t}(s,\omega)\mathrm{d}t=0,$
+with
+$\vec{V}_{+}(s,\omega)=\begin{pmatrix}-\omega_{\eta}\\ 0\\ \omega_{y}r(s)^{4}\end{pmatrix}$
+and $\beta_{\infty}$ a function on the sphere with $\nabla_{\omega}\beta_{\infty}$ is a vector field over the sphere.
+Therefore we can rewrite $\vec{V}_{\pm}\cdot\nabla_{\omega}\beta=\vec{V}_{\pm}^{\infty}\cdot\nabla_{\omega}\beta$ where $V_{\pm}^{\infty}$ is the component of $\vec{V}_{\pm}$ which is tangential to the sphere.
+The vector $\vec{V}_{\pm}^{\infty}$ defines a vector field over the sphere and is given by for all $(s,\omega)\in\mathbb{R}\times\mathbb{S}^{2}$,
+$\vec{V}^{\infty}(s,\omega)\coloneqq\vec{V}_{+}(s,\omega)-\Big(\vec{V}_{+}(s,\omega)\cdot\omega\Big)\omega=\begin{pmatrix}\left(\left(1-r(s)^{4}\right)\omega_{y}^{2}-1\right)\omega_{\eta}\\ \left(1-r(s)^{4}\right)\omega_{y}\omega_{\eta}\omega_{\sigma}\\ \left(\left(1-r(s)^{4}\right)\omega_{\eta}^{2}+r(s)^{4}\right)\omega_{y}\end{pmatrix}$.
+Because of the equality $\lambda^{\infty}_{+}=-\lambda^{\infty}_{-}$, we have the same invariance for $\mathrm{d}\nu_{-}^{t}$. So, $t$ almost everywhere,
+$\int_{\mathbb{R}\times\mathbb{S}^{2}}\Big(\vec{V}^{\infty}(s,\omega)\cdot\nabla_{\omega}\beta_{\infty}(s,\omega)\Big)\mathrm{d}\nu_{\infty}^{t}(s,\omega)=0.$
+So, for almost all $(t,s,\omega)\in\mathbb{R}^{2}\times\mathbb{S}^{2}$, we have $\mathrm{div}_{\omega}\Big(\vec{V}^{\infty}(s,\omega)\mathrm{d}\nu^{t}(s,\omega)\Big)=0$. ∎
+3.2 The Wigner measure
+Proof of Theorem 1.5.
+Assume Assumption 1.1 and Assumption 1.3.
+Let $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ a uniformly bounded family in $\mathrm{L}^{\infty}\left(\mathbb{R},\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)$ solution to equation (1.1) with normalized initial condition $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$.
+Let us consider the family $\left(\varphi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ defined in (2.1) which is solution to equation (2.2), according to Lemma 2.2 and uniformly bounded in $\mathrm{L}^{\infty}\left(\mathbb{R},\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)$.
+Lemma 2.3 gives the link between any semiclassical measure of $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ over the curve $\mathcal{C}$ and the associated measure of $\left(\varphi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ on $\mathbb{R}_{s}$.
+According to Theorem 2.4, we can consider ${\left(M^{t}_{\infty}\mathrm{d}\nu^{t}_{\infty},M^{t}\mathrm{d}\nu^{t}\right)}$ a pair of two-scale semiclassical measure of $\left(\varphi^{\varepsilon}_{t}\right)_{\varepsilon>0}$.
+Finally, according to proofs of Section 3.1.1 and Section 3.1.2, we conclude the evolution of the semiclassical measure over the curve $\mathcal{C}$.
+∎
+3.3 Proof of applications
+Proof of Corollary 1.6.
+Assume Assumption 1.1 and Assumption 1.3. With the notation of Theorem 1.5, the evolution of $(\mathrm{d}\mu_{t},M^{t}\mathrm{d}\nu^{t})$ implies that for all $t\in\mathbb{R}$,
+$\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\mathrm{tr}\Big(\mathrm{d}\mu_{t}(x,\xi)\Big)$ $=\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\mathrm{tr}\Big(\mathrm{d}\mu_{0}(x,\xi)\Big),$
+$\int_{\mathbb{R}^{2}}\mathrm{Tr}\Big(M^{t}(s,\sigma)\Big)\mathrm{d}\nu^{t}(s,\sigma)$ $=\int_{\mathbb{R}^{2}}\mathrm{Tr}\Big(M^{0}(s,\sigma)\Big)\mathrm{d}\nu^{0}(s,\sigma).$
+By taking $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}\in\mathrm{L}^{2}(\mathbb{R}^{2},\mathbb{C}^{2})$ normalized such that
+$\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\mathrm{tr}\Big(\mathrm{d}\mu_{0}(x,\xi)\Big)+\int_{\mathbb{R}^{2}}\mathrm{Tr}\Big(M^{0}(s,\sigma)\Big)\mathrm{d}\nu^{0}(s,\sigma)=1=\left\lVert\psi^{\varepsilon}_{0}\right\rVert_{2}^{2},$
+we have the following equality because of the preservation of the $\mathrm{L}^{2}$-norm
+$\left\lVert\psi^{\varepsilon}_{t}\right\rVert_{2}^{2}=$ $\ \int_{\mathbb{R}^{4}\backslash\mathcal{C}}\mathrm{tr}\Big(\mathrm{d}\mu_{t}(x,\xi)\Big)+\int_{\mathbb{R}^{2}}\mathrm{Tr}\Big(M^{t}(s,\sigma)\Big)\mathrm{d}\nu^{t}(s,\sigma)+\int_{\mathbb{R}\times\mathbb{S}^{2}}\mathrm{tr}\Big(M^{t}_{\infty}(s,\omega)\Big)\mathrm{d}\nu^{t}_{\infty}(s,\omega),$
+$=$ $\ 1.$
+Therefore, for almost all $t$,
+$M^{t}_{\infty}(s,\omega)\mathrm{d}\nu^{t}_{\infty}(s,\omega)=0.$
+Proof of Corollary 1.7.
+Assume Assumption 1.1 and Assumption 1.3.
+According to Lemma 1.2, we know the evolution of any semiclassical measure of $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ outside the curve $\mathcal{C}$.
+According to Theorem 1.5, by applying Remark 2.6, we can describe the semiclassical measure above the interface so we conclude.
+∎
+Proof of Corollary 1.8.
+Assume Assumption 1.1 and Assumption 1.3.
 Let $\vec{f}\in\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$.
-If ${\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}}$ solves (1) with for all $\varepsilon\in(0,1]$ and for all $x\in\mathbb{R}^{2}$,
+Let ${\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}}$ be the solution to equation (1.1) with, for all $\varepsilon\in(0,1]$ and for all $x\in\mathbb{R}^{2}$,
 $\psi_{0}^{\varepsilon}(x)=\frac{1}{\sqrt{\varepsilon}}\vec{f}\left(\frac{x-x_{0}}{\sqrt{\varepsilon}}\right),$
-then, there exist a vanishing sequence ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ of positive numbers and a family of measurable maps ${t\mapsto\left(\mathrm{d}\gamma^{t}_{n}\right)_{n\in\mathbb{Z}}}$ valued in the set of scalar non negative Radon measures on $\mathbb{R}^{2}$, such that for all $(a,\Xi)\in\mathscr{C}_{c}^{\infty}\left(\Omega,\mathbb{C}^{2,2}\right)\times\mathscr{C}_{c}^{\infty}(\mathbb{R},\mathbb{C})$,
-$\int_{\mathbb{R}}\Xi(t)\left\langle a\psi^{\varepsilon_{k}}_{t},\psi^{\varepsilon_{k}}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\mathrm{d}t\underset{k\to+\infty}{\longrightarrow}\int_{\mathbb{R}}\int_{\mathbb{R}}\Xi(t)\mathrm{tr}\Big{(}a\big{(}\gamma(s)\big{)}\mathrm{d}\rho_{t}(s)\Big{)}\mathrm{d}t$
-with for all $s\in\mathbb{R}$,
-$\mathrm{d}\rho_{t}(s)=\frac{1}{2}\sum_{n\in\mathbb{Z}}\int_{\mathbb{R}}\begin{pmatrix}1&\frac{\mathrm{i}\sigma}{\lambda_{n}(s,\sigma)}e^{-\mathrm{i}\theta(s)}\\ \frac{-\mathrm{i}\sigma}{\lambda_{n}(s,\sigma)}e^{\mathrm{i}\theta(s)}&1\end{pmatrix}\mathrm{d}\gamma^{t}_{n}(s,\mathrm{d}\sigma),$
-where for all $n\in\mathbb{Z}$, $\mathrm{d}\gamma_{n}^{t}$ satisfies the following equation
-$\left\{\begin{array}[]{rll}\partial_{t}\mathrm{d}\gamma_{n}^{t}&=&\partial_{\sigma}\mathrm{d}\gamma_{n}^{t}\ \partial_{s}\lambda_{n}-\partial_{s}\mathrm{d}\gamma_{n}^{t}\ \partial_{\sigma}\lambda_{n}\\ \mathrm{d}\gamma_{n}^{t}|_{t=0}&=&\mathrm{d}\gamma_{n,0}\end{array}\right.,$ (6)
-and $\mathrm{d}\gamma_{n,0}$ is determined by the initial condition.
-We will see in Section 3.3 a more general result (Theorem 3.10), with initial data in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$.
-Remark (Link between Theorem 1.1 and Theorem 1.4).
-With the notation of Theorem 1.4, the measure $\mathrm{d}\gamma_{0}^{t}$ is the measure associated to the part $\mathrm{WP}^{\varepsilon}_{x_{t},0}$ of the solution of Theorem 1.1.
-Likewise, the measures $\left(\mathrm{d}\gamma_{n}^{t}\right)_{n\in\mathbb{Z}_{\neq 0}}$ are associated to the $\mathrm{L}^{\infty}$-remainder part of the solution of Theorem 1.1.
-Notably, if for all $n\in\mathbb{Z}_{\neq 0},\mathrm{d}\gamma_{n,0}=0$ and $\mathrm{d}\gamma_{0,0}=1$, then for all $(t,s,\sigma)\in\mathbb{R}^{3}$ and for all $n\in\mathbb{Z}_{\neq 0}$,
-$\mathrm{d}\gamma_{0}^{t}(s,\sigma)=\delta(s-t)\delta(\sigma)\quad,\quad\mathrm{d}\gamma_{n}^{t}(s,\sigma)=0,$
-where $\delta$ is the Dirac mass over 00.
-Theorem 1.4 gives a description of the evolution of the position density when initial data are wave packets.
-But Theorem 3.10 allow us to consider any initial data in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ concentrating near the curve $\mathcal{C}$ to describe their asymptotic position density behavior.
-A consequence of this theorem is the following one.
-Theorem 1.5.
-Assume Assumption 1.2 and 1.3.
-Let ${\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}}$ a uniformly bounded family of $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$, $\nu$ a non negative scalar Radon measure on $\mathbb{R}_{t}\times\mathbb{R}_{x}^{2}$ and ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ a vanishing sequence such that for all $(a,\Xi)\in\mathscr{C}_{c}^{\infty}\left(\Omega,\mathbb{C}\right)\times\mathscr{C}_{c}^{\infty}(\mathbb{R},\mathbb{C})$,
+then, applying Corollary 1.7 and Corollary 2.7, there exist a vanishing sequence of positive numbers ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ and a two-scale semiclassical measure at finite distance ${M^{t}\mathrm{d}\nu^{t}}$, such that for all observables $(a,\Xi)\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)\times\mathscr{C}_{c}^{\infty}(\mathbb{R},\mathbb{C})$,
+$\mathcal{I}^{\varepsilon_{k}}_{\psi}(\Xi,a)\underset{k\to+\infty}{\longrightarrow}\frac{1}{2}\sum_{n\in\mathbb{Z}}\int_{\mathbb{R}^{3}}\Xi(t)\mathrm{tr}\left(a\big(\mathbf{t}(s),0\big)\begin{pmatrix}1&\frac{-\sigma}{\lambda_{n}(s,\sigma)}e^{-\mathrm{i}\theta(s)}\\ \frac{-\sigma}{\lambda_{n}(s,\sigma)}e^{\mathrm{i}\theta(s)}&1\end{pmatrix}\mathrm{d}\nu^{t}_{n}(s,\sigma)\right)\mathrm{d}t,$
+with, for all $n\in\mathbb{Z}$, $\nu_{n}^{t}$ satisfies the following equation
+$\left\{\begin{array}[]{rll}\partial_{t}\ \nu_{n}^{t}&=&\left\{\nu_{n}^{t},\lambda_{n}\right\}\\ \nu_{n}^{t}|_{t=0}&=&\nu_{n,0}\end{array}\right$.,
+with $\nu_{n,0}=\mathrm{tr}\Big(\Pi_{n}M^{0}\Big)\nu^{0}$ and $M^{0}\mathrm{d}\nu^{0}$ is a two-scale semiclassical measure associated with the initial data.
+We conclude thanks to the conservation of $\varepsilon$-oscillation (Lemma A.2). ∎
+Proof of Corollary 1.9.
+Assume Assumption 1.1 and Assumption 1.3.
+Let ${\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}}$ a uniformly bounded family of $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$, $\nu$ a non-negative scalar Radon measure on $\mathbb{R}_{t}\times\mathbb{R}_{x}^{2}$ and a vanishing sequence of positive numbers ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ such that for all $(a,\Xi)\in\mathscr{C}_{c}^{\infty}\left(\Omega,\mathbb{C}\right)\times\mathscr{C}_{c}^{\infty}(\mathbb{R},\mathbb{C})$,
 $\int_{\mathbb{R}}\int_{\Omega}\Xi(t)a(x)\left|\psi_{t}^{\varepsilon_{k}}(x)\right|^{2}\mathrm{d}x\mathrm{d}t\underset{k\to+\infty}{\longrightarrow}\int_{\mathbb{R}}\int_{\Omega}\Xi(t)a(x)\mathrm{d}\nu(t,x),$
-where ${\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}}$ solves (1) with ${\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}}$ as initial condition, then,
-$\mathrm{d}\nu\mathds{1}_{\mathrm{E}}\geqslant\sum_{n\in\mathbb{Z}}\Phi^{*}\left(\mathrm{d}\gamma^{t}_{n}\right)\mathds{1}_{\mathrm{E}}\mathrm{d}t$
-where $\left(\mathrm{d}\gamma_{n}^{t}\right)_{n\in\mathbb{Z}}$ is defined by (6).
-This theorem gives a lower bound estimation of the position density at time $t$ of the evolution of any initial data concentrating over the curve $\mathrm{E}$.
-1.4 Main ideas of the proofs and organisation of the article
-There are two main ideas in the proof of Theorem 3.10.
-The first step is to transform the operator $\mathrm{H}_{\varepsilon}$ into a model operator as if the curve $\mathrm{E}$ were exactly a line. Let us discuss the case of the line.
-If the interface is a line defined by a mass $\mathrm{m}$ such that for all $x\coloneqq(x_{1},x_{2})\in\mathbb{R}^{2},\mathrm{m}(x)=\widetilde{m}(x_{1})x_{2}$, then the operator $\mathrm{H}_{\varepsilon}$ is unitary equivalent to the following one,
-$\sqrt{\varepsilon}\begin{pmatrix}\sqrt{\varepsilon}\mathrm{D}_{1}&x_{2}\widetilde{m}(x_{1})+i\mathrm{D}_{2}\\ x_{2}\widetilde{m}(x_{1})-i\mathrm{D}_{2}&-\sqrt{\varepsilon}\mathrm{D}_{1}\end{pmatrix}.$
-This operator is $\sqrt{\varepsilon}$-semiclassical with respect to $x_{1}$ only.
-Thus, it will be considered as the quantization of an operator-valued symbol.
-It leads us to the second idea of the proof.
-The second step is to build a two-scale quantization that fits the scale of our new operator.
-In fact, our analysis relies on a microlocal approach, replacing densities $\left(\left|\psi_{t}^{\varepsilon}(x)\right|^{2}\mathrm{d}x\right)_{\varepsilon>0}$ by the Wigner transform [26, 15].
-The Wigner transform provides us with a description of the dynamics for a very general class of equations and initial data, allowing us to deal with quantum observables.
-Therefore, we will study the quadratic quantity (3), with an observable $a$ depending of position and momentum, i.e.
-$a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2}_{x}\times\mathbb{R}^{2}_{\xi},\mathbb{C}^{2,2}\right)\longmapsto\left\langle a\left(\cdot{},\varepsilon\mathrm{D}_{x}\right)\psi^{\varepsilon}_{t},\psi^{\varepsilon}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}.$
-We recall basic facts about these notions in Section 2.
-However, to precisely analyze what happens near the curve $\mathcal{C}$, we have to take into account how $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ concentrates on the curve $\mathcal{C}$ with respect to the scale $\sqrt{\varepsilon}$.
-We introduce in Section 3 the two-scale Wigner measure that be of crucial use for proving our main results Theorems 3.8 and 3.10.
-Finally, we prove Theorems 1.41.53.8 and 3.10 in Section 4.
-1.5 Notation
-We denote by $\mathbb{S}^{2}$ the unit sphere of $\mathbb{R}^{3}$.
-We denote by $v\otimes w$ the $2$ by $2$ matrix $v\,^{t}w$ for $(v,w)\in\left(\mathbb{C}^{2}\right)^{2}$.
-We denote by ${\rm tr}$ the trace in $\mathcal{L}\left(\mathbb{C}^{2}\right)$ and by $\mathrm{Tr}$ the trace in $\mathcal{L}\left(\mathrm{L}^{2}\left(\mathbb{R},\mathbb{C}^{2}\right)\right)$.
-Let $\mathcal{H}$ a separable Hilbert space and $n$ an integer, for $f,g\in\mathrm{L}^{2}\left(\mathbb{R}^{n},\mathcal{H}\right)$, we denote the usual inner product by
-$\left\langle f,g\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{n},\mathcal{H}\right)}\coloneqq\int_{\mathbb{R}^{n}}\left\langle f(x),g(x)\right\rangle_{\mathcal{H}}\,\mathrm{d}x.$
-For $A$ an operator acting on $\mathrm{L}^{2}\left(\mathbb{R}^{n},\mathcal{H}\right)$, we denote by $A^{\dagger}$ the formal adjoint operator of $A$ associated to the usual scalar product of $\mathrm{L}^{2}\left(\mathbb{R}^{n},\mathcal{H}\right)$.
-We define the semiclassical Fourier transform for all $\varepsilon>0$ and for $f\in\mathrm{L}^{1}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ by
-$\mathcal{F}_{\varepsilon}f(\xi)\coloneqq\frac{1}{2\pi\varepsilon}\int_{\mathbb{R}^{2}}e^{-\mathrm{i}\frac{x\cdot\xi}{\varepsilon}}f(x)\mathrm{d}x,\quad\xi\in\mathbb{R}^{2},$
-and we denote by $\widehat{f}$ the Fourier transform of $f$ defined by $\widehat{f}\coloneqq\mathcal{F}_{1}f$.
-2 Semiclassical analysis
-Our problem requires an analysis in the phase space.
-We consider the Wigner transform of the solution $\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}$ of equation (1) and its weak limits in the space of distributions that are called Wigner measures.
-We recall the definition and the main properties of the Wigner transform in Section 2.1 below.
-Then, we apply these results to the solution of our problem in Section 2.2.
-2.1 Wigner transform and Wigner measure
-Let ${(f^{\varepsilon})_{\varepsilon>0}}$ uniformly bounded in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$.
-The Wigner transform of ${(f^{\varepsilon})_{\varepsilon>0}}$ is defined as follow, for all $(x,\xi)\in\mathbb{R}^{2}\times\mathbb{R}^{2}$,
-$\mathrm{W}^{\varepsilon}\left[f^{\varepsilon}\right](x,\xi)\coloneqq\int_{\mathbb{R}^{2}}e^{i\xi\cdot y}f^{\varepsilon}\left(x+\frac{\varepsilon y}{2}\right)\otimes\overline{f^{\varepsilon}\left(x-\frac{\varepsilon y}{2}\right)}\frac{\mathrm{d}y}{(2\pi)^{2}}.$
-It was introduced by Wigner in [25] and in [26].
-One can check that $\mathrm{W}^{\varepsilon}\left[f^{\varepsilon}\right]$ is valued in the set of Hermitian matrices and $\mathrm{W}^{\varepsilon}\left[f^{\varepsilon}\right]\in\mathrm{L}^{2}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)\cap\mathscr{C}^{0}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)$.
-Moreover, one recovers formally the density in position through the relation
-$\int_{\mathbb{R}^{2}}{\rm tr}\mathrm{W}^{\varepsilon}\left[f^{\varepsilon}\right](x,\xi)\mathrm{d}\xi=\left|f^{\varepsilon}(x)\right|_{\mathbb{C}^{2}}^{2}.$
-The function $\mathrm{W}^{\varepsilon}\left[f^{\varepsilon}\right]$ is usually considered as a distribution on the phase space $\mathbb{R}^{4}=\mathbb{R}^{2}_{x}\times\mathbb{R}^{2}_{\xi}$.
-Moreover, we have for all $a\in\mathscr{C}^{\infty}_{c}(\mathbb{R}^{4},\mathbb{C}^{2,2})$,
-$\mathrm{I}^{\varepsilon}_{f}(a)\coloneqq\int_{\mathbb{R}^{4}}{\rm tr}\Big{(}a(x,\xi)\mathrm{W}^{\varepsilon}\left[f^{\varepsilon}\right](x,\xi)\Big{)}\mathrm{d}x\mathrm{d}\xi=\left\langle\mathrm{Op}_{\varepsilon}(a)f^{\varepsilon},f^{\varepsilon}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)},$
-where $\mathrm{Op}_{\varepsilon}(a)$ is the semiclassical pseudo-differential operator of symbol $a$ obtained by the Weyl quantization
-$\mathrm{Op}_{\varepsilon}(a)f(x)\coloneqq\int_{\mathbb{R}^{4}}a\left(\frac{x+y}{2},\xi\right)f(y)e^{\mathrm{i}\frac{\xi\cdot(x-y)}{\varepsilon}}\frac{\mathrm{d}y\mathrm{d}\xi}{(2\pi\varepsilon)^{2}},\quad f\in\mathscr{S}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right),\ x\in\mathbb{R}^{2}.$
-The Calderón-Vaillancourt Theorem [4] asserts that the family of operators $\left(\mathrm{Op}_{\varepsilon}(a)\right)_{\varepsilon>0}$ is uniformly bounded with respect to $\varepsilon$ in $\mathcal{L}\left(L^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)$: there exists $(C,N)\in\mathbb{R}_{>0}\times\mathbb{N}_{\geqslant 0}$ such that for all $a=(a_{i,j})_{1\leqslant i,j\leqslant 2}\in\mathscr{S}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)$, for all $f\in\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right),$
-$\left\lVert\mathrm{Op}_{\varepsilon}(a)f\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\leqslant C\,\max_{1\leqslant i,j\leqslant 2}\,\max_{\left|\alpha\right|+\left|\beta\right|\leqslant N}\left\lVert\partial^{\alpha}_{x}\partial^{\beta}_{\xi}a_{i,j}\right\rVert_{\mathrm{L}^{\infty}\left(\mathbb{R}^{4}\right)}\left\lVert f\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}.$ (7)
-The Wigner transform is not a positive distribution.
-However, its weak limits are positive measures.
-This is due to the weak Gårding inequality.
-Adapting [11] Proposition 2.13 to the matrix case, for all $a\in\mathscr{C}^{\infty}_{c}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)$ non negative, for all $\delta>0$, there exists $C_{\delta}>0$, for all $f\in\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$,
-$\left\langle\mathrm{Op}_{\varepsilon}(a)f,f\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\geqslant-\left(\delta+C_{\delta}\varepsilon\right)\left\lVert f\right\rVert^{2}_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}.$
-There exists a subsequence $\left(f_{\varepsilon_{n}}\right)_{n\in\mathbb{N}}$ and a positive matrix of Radon measures $\mathrm{d}\mu=(\mathrm{d}\mu_{i,j})_{1\leqslant i,j\leqslant 2}$ such that for all $a\in\mathscr{S}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)$,
-$\mathrm{I}^{\varepsilon_{n}}_{f}(a)\underset{n\to+\infty}{\longrightarrow}\int_{\mathbb{R}^{4}}{\rm tr}\Big{(}a(x,\xi)\mathrm{d}\mu(x,\xi)\Big{)}.$ (8)
-The measure $\mathrm{d}\mu$ is called the Wigner measure or the semiclassical measure of $\left(f^{\varepsilon_{n}}\right)_{n\in\mathbb{N}}$.
-A proof of existence is recalled in Section B.1 for families valued in a separable Hilbert space.
-Example 2.1.
+where ${\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}}$ solves (1.1) with ${\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}}$ as initial condition.
+A consequence of Theorem 1.5 and Remark 2.6 is
+$\mathrm{d}\nu\mathds{1}_{\mathcal{C}}\geqslant\sum_{n\in\mathbb{Z}}\Phi^{*}\left(\mathrm{Tr}\Big(\Pi_{n}(s,\sigma)\Big)\mathrm{d}\nu^{t}_{n}\right)\mathds{1}_{\mathcal{C}}\mathrm{d}t$
+where $\left(\nu_{n}^{t}\right)_{n\in\mathbb{Z}}$ is defined by (1.20). We conclude with $\mathrm{Tr}\left(\Pi_{n}(s,\sigma)\right)=1$. ∎
+Appendix A Semiclassical analysis
+Our problem requires an analysis in phase space.
+We consider the Wigner transform of the solution $\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}$ of equation (1.1) and its weak limits in the space of distributions that are called Wigner measures.
+We apply the main properties of the Wigner transform (introduced in Section 1.2) to the solution of our problem in Section A.1.
+First, let us briefly introduce an important example and some definitions.
+Example A.1.
 Let us consider $\alpha\in(0,1]$, $f\in\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ and for all $x\in\mathbb{R}^{2}$,
@@
 $\widetilde{\nu}(\xi)=\mathrm{tr}\int_{\mathbb{R}^{2}}\mathrm{d}\mu(\mathrm{d}x,\xi).$
-Now, we apply these first properties on solutions to equation (1).
-2.2 Application to the Dirac equation
-In this section, we consider $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ the solution to equation (1) with the initial normalized condition $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$.
+Now, we apply these first properties on solutions to equation (1.1).
+A.1 Application to Dirac equation
+In this section, we consider $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ the solution to equation (1.1) with the normalized initial condition $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$.
 We revisit in this paragraph the notions introduced in the preceding one for this special family.
-2.2.1 $\varepsilon$-oscillation of solution to Dirac equation
-Lemma 2.2 (Conservation of $\varepsilon$-oscillating).
+Lemma A.2 (Conservation of $\varepsilon$-oscillating).
 Let $T>0$.
@@
 $\limsup_{\varepsilon\to 0}\int_{\left|\xi\right|>R/\varepsilon}\left|\widehat{\psi^{\varepsilon}_{t}}(\xi)\right|^{2}\mathrm{d}\xi\underset{R\to+\infty}{\longrightarrow}0.$
-Proof of Lemma 2.2.
+Proof of Lemma A.2.
 Let $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ be $\varepsilon$-oscillating. The proof relies on the analysis of
@@
 because $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ is uniformly bounded in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ and $\varepsilon$-oscillating. ∎
-The $\varepsilon$-oscillation conservation property justifies the semiclassical approach of solutions to equation (1) for analyzing the limits of the quantities (3).
-Note that similar arguments give that there is also no loss of mass at infinity in configuration space in the sense that if $\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}$ is compact at infinity, then $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ is compact at infinity uniformly in time $t\in[-T,T]$, for some $T>0$.
-The proof is in Appendix A.
-As a consequence, if $\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}$ is normalized, $\varepsilon$-oscillating and compact at infinity, then for all $\chi\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},[0,1]\right)$, with $\chi\equiv 1$ for $\left|x\right|<1$, $\chi\equiv 0$ for $\left|x\right|>2$,
-$\lim_{\varepsilon\to 0}\left\langle\mathrm{Op}_{\varepsilon}\left(\chi\left(\frac{x}{R}\right)\chi\left(\frac{\varepsilon\mathrm{D}_{x}}{R}\right)\right)\psi^{\varepsilon}_{0},\psi^{\varepsilon}_{0}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\;\underset{R\to+\infty}{\longrightarrow}\;\lim_{\varepsilon\to 0}\left\lVert\psi_{0}^{\varepsilon}\right\rVert^{2}_{\mathrm{L}^{2}\left(R^{2},\mathbb{C}^{2}\right)}=1.$
-So for all semiclassical measure $\mathrm{d}\mu$ of $\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}$ normalized, $\varepsilon$-oscillating and compact at infinity,
-$\mathrm{tr}\int_{\mathbb{R}^{4}}\mathrm{d}\mu(x,\xi)=1.$
-2.2.2 Time-averaged Wigner measures
-For taking into account the time dependence of the family $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$, we will consider time-averaged Wigner measure.
-We test the Wigner transform $\mathrm{W}^{\varepsilon}[\psi^{\varepsilon}_{t}]$ against time-dependent observables, so for all $\Xi\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R},\mathbb{C}\right)$,
-$\int_{\mathbb{R}}\int_{\mathbb{R}^{4}}\Xi(t){\rm tr}_{\mathbb{C}^{2}}\Big{(}a(x,\xi)\mathrm{W}^{\varepsilon}[\psi^{\varepsilon}_{t}](x,\xi)\Big{)}\mathrm{d}x\mathrm{d}\xi\mathrm{d}t\;=\;\int_{\mathbb{R}}\Xi(t)\left\langle\mathrm{Op}_{\varepsilon}(a)\psi^{\varepsilon}_{t},\psi^{\varepsilon}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\mathrm{d}t.$
-For the same reasons as before, we obtain the existence of a vanishing sequence $\left(\varepsilon_{k}\right)_{k\in\mathbb{N}}$ of positive numbers and of a measure $\mathrm{d}\mu$ on $\mathbb{R}^{5}=\mathbb{R}_{t}\times\mathbb{R}^{2}_{x}\times\mathbb{R}^{2}_{\xi}$ such that
-$\int_{\mathbb{R}}\Xi(t)\left\langle\mathrm{Op}_{\varepsilon_{k}}(a)\psi^{\varepsilon_{k}}_{t},\psi^{\varepsilon_{k}}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\mathrm{d}t\;\underset{k\to+\infty}{\longrightarrow}\;\int_{\mathbb{R}^{5}}\Xi(t)a(x,\xi)\mathrm{d}\mu(t,x,\xi).$
-Moreover, thanks to the control
-$\left|\int_{\mathbb{R}^{5}}\Xi(t)a(x,\xi)\mathrm{d}\mu(t,x,\xi)\right|\leqslant C\ \left\lVert\Xi\right\rVert_{\mathrm{L}^{1}(\mathbb{R})}\left\lVert\mathrm{Op}_{\varepsilon}(a)\right\rVert_{\mathcal{L}\left(\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)},$
-we can write $\mathrm{d}\mu(t,x,\xi)=\mathrm{d}\mu_{t}(x,\xi)\mathrm{d}t$. In other words, the measure $\mathrm{d}\mu$ is absolutely continuous in time with respect to Lebesgue measure.
-2.2.3 First properties of the Wigner measure of the solution of the Dirac equation
-It is known from [15, 11] that any time averaged Wigner measure, outside the curve $\mathcal{C}$, of $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ is completely characterized by those of the initial data.
-Let us first introduce some notations.
-Let us consider the symbol of the operator $\mathrm{H}_{\varepsilon}$ according to the Weyl quantization; for all $(x,\xi)\in\mathbb{R}^{4}$,
-$\mathrm{H}(x,\xi)\coloneqq\begin{pmatrix}\mathrm{m}(x)&\xi_{1}-\mathrm{i}\xi_{2}\\ \xi_{1}+\mathrm{i}\xi_{2}&-\mathrm{m}(x)\end{pmatrix}.$
-We remark that ${\det\mathrm{H}(x,\xi)=-\mathrm{m}(x)^{2}-|\xi|^{2}}$.
-Let $(x,\xi)\in\mathbb{R}^{4}$ such that $\det\mathrm{H}(x,\xi)\neq 0$, then $\mathrm{H}(x,\xi)$ has two different eigenvalues $\lambda_{+}(x,\xi)$ and $\lambda_{-}(x,\xi)$ defined by
-[…]
-Let us denote the projectors on the corresponding eigenspaces by
-$\Pi_{+}(x,\xi)\coloneqq\mathrm{Id}-\frac{1}{\lambda_{+}(x,\xi)}\mathrm{H}(x,\xi)\quad,\quad\Pi_{-}(x,\xi)\coloneqq\mathrm{Id}-\frac{1}{\lambda_{-}(x,\xi)}\mathrm{H}(x,\xi).$
-According to the definition of $\mathcal{C}$, we remark that
-$(x,\xi)\in\mathcal{C}\iff\det\mathrm{H}(x,\xi)\neq 0.$
-We recall the definition of the Poisson bracket
-$\left\{f,g\right\}\coloneqq\nabla_{\xi}f\nabla_{x}g-\nabla_{x}f\nabla_{\xi}g\;,\;(f,g)\in\mathscr{C}^{1}\left(\mathbb{R}^{2}_{x}\times\mathbb{R}^{2}_{\xi}\right)^{2}.$ (9)
-Lemma 2.3 (Evolution of semiclassical measure outside $\mathcal{C}$).
-Let $\mathrm{d}\mu_{t}$ be a semiclassical measure of $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ associated to $(\varepsilon_{n})_{n\in\mathbb{N}}$, then the measure $\mathrm{d}\mu_{t}$ can be decomposed outside $\mathcal{C}$ as follows : for all $(x,\xi)\in\mathbb{R}^{4}\setminus\mathcal{C}$
-$\mathrm{d}\mu_{t}(x,\xi)=\Pi_{+}(x,\xi)\mathrm{d}\mu_{t}^{+}(x,\xi)+\Pi_{-}(x,\xi)\mathrm{d}\mu_{t}^{-}(x,\xi),$
-where the scalar measures $\mathrm{d}\mu_{t}^{\pm}$ satisfy
-$\partial_{t}\ \mathrm{d}\mu_{t}^{+}=\left\{\mathrm{d}\mu_{t}^{+},\lambda_{+}\right\}$ $\;\;,\;\;\mathrm{d}\mu_{0}^{+}=\mathrm{tr}\Big{(}\Pi_{+}\mathrm{d}\mu_{0}\Big{)},$
-$\partial_{t}\ \mathrm{d}\mu_{t}^{-}=\left\{\mathrm{d}\mu_{t}^{-},\lambda_{-}\right\}$ $\;\;,\;\;\mathrm{d}\mu_{0}^{-}=\mathrm{tr}\Big{(}\Pi_{-}\mathrm{d}\mu_{0}\Big{)},$
-and $\mathrm{d}\mu_{0}$ is the semiclassical measure of $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ associated to $(\varepsilon_{n})_{n\in\mathbb{N}}$.
-This theorem is proved in [15, section 6].
-The proof uses two major arguments.
-First, we have the decomposition of the measure $\mathrm{d}\mu_{t}$ as a sum $\Pi_{+}\mathrm{d}\mu_{t}^{+}+\Pi_{-}\mathrm{d}\mu_{t}^{-}$.
-Secondly, the equation for $\mathrm{d}\mu_{t}^{\pm}$ comes from the analysis for all $a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{4},\mathbb{C}\right)$ of
-$\frac{1}{\mathrm{i}\varepsilon}\Big{[}\mathrm{Op}_{\varepsilon}\left(a\Pi_{\pm}\right),\mathrm{Op}_{\varepsilon}\left(\mathrm{H}_{\varepsilon}\right)\Big{]}.$
-This theorem implies that for all initial data $\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}$ localized around $(x_{0},\xi_{0})\notin\mathcal{C}$, the semiclassical measure evolved along the Hamiltonian trajectories associated to the functions $\lambda_{+}$ and $\lambda_{-}$.
-These trajectories are included in the hypersurface $\det\mathrm{H}(x,\xi)=\det\mathrm{H}(x_{0},\xi_{0})$.
-It means that if we localize the initial data far from the curve $\mathcal{C}$, then the semiclassical measure never approaches the curve $\mathcal{C}$ in the phase space.
-Our aim in the next sections is to describe the evolution of the measure $\mathrm{d}\mu_{t}$ above the curve $\mathcal{C}$.
-3 Two-scale analysis of concentrations
-To study the evolution of the measure $\mathrm{d}\mu_{t}$ near the curve $\mathcal{C}$, we use a two-scale analysis near $\mathcal{C}$.
-We use the geometric assumptions on $\mathcal{C}$ stated in the introduction to define our two-scale quantization.
-In geometric terms, the measure $\mathrm{d}\mu_{t}$ is a measure on the cotangent space to $\mathbb{R}^{4}$ and, as such, have the associated geometric invariance.
-In particular, we can operate a change of coordinates.
-We work locally in a tubular neighborhood of $\mathcal{C}$ where we can use normal geodesic coordinates that we introduce in Section 3.1.
-In these coordinates $(s,y)$, the curve $\mathrm{E}$ writes $\{y=0\}$.
-We use these coordinates to transform the operator and obtain a normal form of the operator $\mathrm{H}_{\varepsilon}$.
-In Section 3.2 we introduce two-scale Wigner measures.
-Finally, we state our main results, describing these two-scale semiclassical measures for families of solutions to equation (1) in Section 3.3.
-3.1 Normal geodesic coordinates
-We assume assumptions 1.2 and 1.3.
-We will work in the normal geodesic coordinates denoted $(s,y)$ defined by (4).
-The curve $\mathcal{C}$ in these new coordinates is reduced to $\left\{y=\sigma=\eta=0\right\}$ where $(\sigma,\eta)$ play the role of $\xi$.
-Let us define ${\mathrm{I}_{\delta_{0}}\coloneqq\left(-\frac{\delta_{0}}{2||\kappa||_{\infty}},\frac{\delta_{0}}{2||\kappa||_{\infty}}\right)}$ for $\delta_{0}\in\left(0,1\right]$ and let us consider the map
-$\begin{matrix}\Phi_{\delta_{0}}:&\mathbb{R}\times\mathrm{I}_{\delta_{0}}&\longrightarrow&\mathbb{R}^{2}\\ &(s,y)&\longmapsto&\gamma(s)+y\nu(s)\end{matrix}$
-A consequence of Assumption 1.3 is that for all $\delta_{0}\in\left(0,1\right]$, the function $\Phi_{\delta_{0}}$ is a diffeomorphism from ×$\mathbb{R}\times\mathrm{I}_{\delta_{0}}$ to
-$\Omega_{\delta_{0}}$$\coloneqq\left\{\gamma(s)+y\nu(s)|\ (s,y)\in\mathbb{R}\times\mathrm{I}_{\delta_{0}}\right\},$
-=$=\left\{x\in\mathbb{R}^{2}\Big{|}\ \mathrm{d}\left(x,\mathrm{E}\right)<\frac{\delta_{0}}{2||\kappa||_{\infty}}\right\}.$
-We now use $\Phi$ to obtain a normal form of the operator $\mathrm{H}_{\varepsilon}$ as in [3].
-Proposition 3.1 (Hamiltonian Normal form near the interface).
-There exists a unitary operator $\mathrm{U}$ from $\mathrm{L}^{2}(\Omega,\mathbb{C}^{2})$ to $\mathrm{L}^{2}(\mathbb{R}\times\mathrm{I},\mathbb{C}^{2})$ such that $\mathrm{U}\mathrm{H}_{\varepsilon}\mathrm{U}^{-1}=\mathrm{H}_{\varepsilon}^{\mathrm{E}}$ where the operator $\mathrm{H}_{\varepsilon}^{\mathrm{E}}$ is acting on $\mathrm{L}^{2}(\mathbb{R}\times\mathrm{I},\mathbb{C}^{2})$ and defined as
-$\mathrm{H}_{\varepsilon}^{\mathrm{E}}\coloneqq\mathrm{m}\Big{(}\Phi(s,y)\Big{)}\sigma_{1}-\varepsilon\mathrm{D}_{y}\sigma_{2}+\frac{\varepsilon}{1-y\kappa(s)}\mathrm{D}_{s}\sigma_{3}-\frac{\mathrm{i}\varepsilon y\kappa^{\prime}(s)}{2(1-y\kappa(s))^{2}}\sigma_{3}.$
-Remark.
-We have the following properties.
-  1. 1.
-    For all $(s,y)\in\mathbb{R}\times\mathrm{I},\left|1-y\kappa(s)\right|>1/2$ therefore the previous quantities are well-defined.
-  2. 2.
-    The operator $\mathrm{U}$ is a matrix-valued Fourier Integral Operator.
-It contains the pull-back by the change of variables $\Phi$ and the multiplication by an adequate $s$-dependent matrix (a rotation and a multiplication by a $s$-phase function).
-  3. 3.
-    In the following, we will more generally apply the operator $\mathrm{U}$ to function of $\mathrm{L}^{2}(\mathbb{R}^{2},\mathbb{C}^{2})$ but compactly supported in $\Omega$.
-Proof of Proposition 3.1.
-Let us define for all $(s,y)\in\mathbb{R}\times\mathrm{I}$ and for all $u\in\mathrm{L}^{2}(\Omega,\mathbb{C}^{2})$,
-$\mathrm{U}u(s,y)\coloneqq\sqrt{\frac{1-y\kappa(s)}{2}}\begin{pmatrix}e^{\mathrm{i}\frac{\pi}{4}+\mathrm{i}\frac{\theta(s)}{2}}&-e^{-\mathrm{i}\frac{\pi}{4}-\mathrm{i}\frac{\theta(s)}{2}}\\ e^{\mathrm{i}\frac{\pi}{4}+\mathrm{i}\frac{\theta(s)}{2}}&e^{-\mathrm{i}\frac{\pi}{4}-\mathrm{i}\frac{\theta(s)}{2}}\end{pmatrix}u\Big{(}\Phi(s,y)\Big{)}.$
-In order to facilitate the computation, we split the operator $\mathrm{U}$ into the composition of two operators.
-Let us define for all $(s,y)\in\mathbb{R}\times\mathrm{I}$, for all $\theta\in\mathbb{R}$, for all $u\in\mathrm{L}^{2}(\Omega,\mathbb{C}^{2})$ and for all $v\in\mathrm{L}^{2}\left(\mathbb{R}\times\mathrm{I},\mathbb{C}^{2}\right)$,
-$\mathrm{V}_{\Phi}u(s,y)$ $\coloneqq\sqrt{1-y\kappa(s)}u\Big{(}\Phi(s,y)\Big{)},$
-$\mathrm{U}_{\theta}v(s,y)$ $\coloneqq\frac{1}{\sqrt{2}}\begin{pmatrix}e^{\mathrm{i}\frac{\pi}{4}+\mathrm{i}\frac{\theta}{2}}&-e^{-\mathrm{i}\frac{\pi}{4}-\mathrm{i}\frac{\theta}{2}}\\ e^{\mathrm{i}\frac{\pi}{4}+\mathrm{i}\frac{\theta}{2}}&e^{-\mathrm{i}\frac{\pi}{4}-\mathrm{i}\frac{\theta}{2}}\end{pmatrix}v(s,y).$ (10)
-The operator $\mathrm{V}_{\Phi}$ is a unitary operator from $\mathrm{L}^{2}\left(\Omega,\mathbb{C}^{2}\right)$ to $\mathrm{L}^{2}\left(\mathbb{R}\times\mathrm{I},\mathbb{C}^{2}\right)$ and the operator $\mathrm{U}_{\theta}$ is a unitary operator from $\mathrm{L}^{2}\left(\mathbb{R}\times\mathrm{I},\mathbb{C}^{2}\right)$ to $\mathrm{L}^{2}\left(\mathbb{R}\times\mathrm{I},\mathbb{C}^{2}\right)$.
-Notably, $\mathrm{U}=\mathrm{U}_{\theta(s)}\mathrm{V}_{\Phi}$.
-Moreover,
-$\mathrm{V}_{\Phi}\mathrm{H}_{\varepsilon}\mathrm{V}_{\Phi}^{-1}=$ $\ \mathrm{m}\Big{(}\Phi(s,y)\Big{)}\sigma_{3}+\frac{\varepsilon}{1-y\kappa(s)}(\gamma^{\prime}\odot\sigma)\mathrm{D}_{s}+\varepsilon(\nu\odot\sigma)\mathrm{D}_{y}$
-$\ -\ \frac{\mathrm{i}\varepsilon y\kappa^{\prime}(s)}{2(1-y\kappa(s))^{2}}(\gamma^{\prime}\odot\sigma)-\frac{\mathrm{i}\varepsilon\kappa(s)}{2(1-y\kappa(s))}(\nu\odot\sigma),$
-with the notation $x\odot\sigma\coloneqq x_{1}\sigma_{1}+x_{2}\sigma_{2}$ for $x\in\mathbb{C}^{2}$ and $\sigma_{1},\sigma_{2}$ Pauli matrices. We conclude with the identities
-$\begin{split}\mathrm{U}_{\theta(s)}\sigma_{3}\mathrm{U}_{\theta(s)}^{-1}=\sigma_{1},\quad&\mathrm{U}_{\theta(s)}(\gamma^{\prime}\odot\sigma)\mathrm{U}_{\theta(s)}^{-1}=\sigma_{3},\\ \mathrm{U}_{\theta(s)}(\nu\odot\sigma)\mathrm{U}_{\theta(s)}^{-1}=-\sigma_{2},\quad&\mathrm{U}_{\theta(s)}(\gamma^{\prime}\odot\sigma)\left(\mathrm{D}_{s}\mathrm{U}_{\theta(s)}^{-1}\right)=-\frac{\kappa(s)}{2}\mathrm{i}\sigma_{2}.\end{split}$
-Let $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ solves (1) with the initial normalized condition $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$.
-Let $\chi\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},[0,1]\right)$, with $\chi\equiv 1$ for $\mathrm{d}(x,\mathrm{E})<\frac{1}{4||\kappa||_{\infty}}$ and $\chi\equiv 0$ for $\mathrm{d}(x,\mathrm{E})>\frac{1}{2||\kappa||_{\infty}}$, then we define for all $\varepsilon>0$,
-$\varphi^{\varepsilon}_{t}\coloneqq\mathrm{U}\chi\psi^{\varepsilon}_{t}.$ (11)
-Lemma 3.2 (Dirac equation in normal geodesic coordinates).
-The family $\left(\varphi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ satisfies on $\mathbb{R}_{t}\times\mathbb{R}_{s}\times\mathrm{I}_{1/2}$ the following equation
-$\varepsilon\mathrm{D}_{t}\varphi^{\varepsilon}_{t}+\mathrm{H}^{\mathrm{E}}_{\varepsilon}\varphi^{\varepsilon}_{t}=0.$ (12)
-Proof of Lemma 3.2.
-We apply $\mathrm{U}\chi$ to equation (1),
-$\varepsilon\mathrm{D}_{t}\varphi_{t}^{\varepsilon}+\mathrm{H}^{\mathrm{E}}_{\varepsilon}\varphi_{t}^{\varepsilon}=\mathrm{U}\left[\mathrm{H}_{\varepsilon},\chi\right]\psi_{t}^{\varepsilon}.$
-We compute
-$\left[\mathrm{H}_{\varepsilon},\chi\right]=-\mathrm{i}\varepsilon\begin{pmatrix}0&\partial_{1}\chi-\mathrm{i}\partial_{2}\chi\\ \partial_{1}\chi+\mathrm{i}\partial_{2}\chi&0\end{pmatrix},$
-where for all $j\in\left\{1,2\right\},\partial_{j}\chi\in\mathscr{C}_{c}^{\infty}(\mathbb{R}^{2},\mathbb{C})$ is supported in $\left\{x\in\mathbb{R}^{2}\ \Big{|}\ \frac{1}{4||\kappa||_{\infty}}\leqslant\mathrm{d}(x,\mathrm{E})\leqslant\frac{1}{2||\kappa||_{\infty}}\right\}$, then for all $(t,s,y)\in\mathbb{R}_{t}\times\mathbb{R}_{s}\times\mathrm{I}_{1/2}$,
-$\Big{(}\mathrm{U}\left[\mathrm{H}_{\varepsilon},\chi\right]\psi_{t}^{\varepsilon}\Big{)}(s,y)=0.$
-3.2 Two-scale Wigner measures
-In this section, we introduce observables that allow to concentrate on the set $\left\{y=0,\sigma=0,\eta=0\right\}$ at a precise scale, that we choose as $\sqrt{\varepsilon}$.
-As we will see later and as expected in view of [1], this scale is natural.
-3.2.1 Two-scale symbols and their quantization
-Following the course [11], we introduce
-Definition (Two-scale observable).
-We define the set $\mathfrak{A}$ as follows
-$\mathfrak{A}\coloneqq\left\{a\in\mathscr{C}^{\infty}\Big{(}\mathbb{R}^{4}\times\mathbb{R}^{3},\mathbb{C}^{2,2}\Big{)}\ \Bigg{|}\ \begin{array}[]{c}\exists K\subset\mathbb{R}^{4}\ \text{compact},\forall z\in\mathbb{R}^{3},a(\cdot,\cdot,z)\ \text{is supported in K}\\ \exists R_{0}>0,\forall|z|>R_{0},a(\cdot,\cdot,z)=a_{\infty}\left(\cdot,\cdot,\frac{z}{|z|}\right)\end{array}\right\}.$
-We can identify $\mathscr{C}^{\infty}_{c}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)$ as a subset of $\mathfrak{A}$.
-We define a second scale quantization near the interface $\mathcal{C}$ for these observable by
-$\begin{array}[]{ccccc}\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}&:&\mathfrak{A}&\to&\mathcal{L}\big{(}\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\big{)}\\ &&a&\mapsto&\mathrm{Op}_{1}\left(a\left(s,y,\varepsilon\sigma,\varepsilon\eta,\frac{y}{\sqrt{\varepsilon}},\sqrt{\varepsilon}\sigma,\sqrt{\varepsilon}\eta\right)\right)\end{array}$ (13)
-In particular, we have, for all $a\in\mathfrak{A}$,
-$\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}(a)$ $=\mathrm{Op}_{\varepsilon}\left(a\left(s,y,\sigma,\eta,\frac{(y,\sigma,\eta)}{\sqrt{\varepsilon}}\right)\right)$
-$=\mathfrak{T}_{\sqrt{\varepsilon}}^{\dagger}\mathrm{Op}_{1}\Big{(}a\left(s,\sqrt{\varepsilon}y,\varepsilon\sigma,\sqrt{\varepsilon}\eta,y,\sqrt{\varepsilon}\sigma,\eta\right)\Big{)}\mathfrak{T}_{\sqrt{\varepsilon}},\ \left(\text{semiclassical in }\sqrt{\varepsilon}\right)$ (14)
-with for all $f\in\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ and for all $(s,y)\in\mathbb{R}^{2},$
-$\mathfrak{T}_{\sqrt{\varepsilon}}^{\dagger}[f](s,y)=\frac{1}{\varepsilon^{1/4}}f\left(s,\frac{y}{\sqrt{\varepsilon}}\right).$ (15)
-Notably, $\mathfrak{T}_{\sqrt{\varepsilon}}$ is a unitary operator from $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ to $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$.
-Moreover, the quantization map $\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}$ associates to $a$ a bounded operator on $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ according to the Calderón-Vaillancourt Theorem (7) applied to expression (14).
-These observables allow us to differentiate three regions of the phase space $\mathbb{R}_{s}\times\mathrm{I}_{1/2}\times\mathbb{R}^{2}_{\sigma,\eta}$.
-Let us explain this fact.
-Let $a\in\mathfrak{A}$ and $\chi\in\mathscr{C}^{\infty}_{c}(\mathbb{R},[0,1])$ supported in $[-1,1]$ with $\chi\equiv 1$ on $[-1/2,1/2]$.
-Let $\delta>0$, $R>R_{0}$ with $R_{0}$ associated to $a$ and let us consider $\varepsilon$ small enough such that
-$2R\sqrt{\varepsilon}<\delta.$ (16)
-Let us denote $x\coloneqq(s,y),\xi\coloneqq(\sigma,\eta)$ and $z\coloneqq\left(y,\sigma,\eta\right)$ for simplicity. We now split $a$ into three observables,
-$a\left(x,\xi,z\right)=$ $\ a\left(x,\xi,z\right)\left(1-\chi\left(\frac{\sqrt{\varepsilon}\left|z\right|}{\delta}\right)\right)$ $\left(\eqqcolon a^{\delta}_{\varepsilon}(x,\xi,z)\right)$ (17)
-$+\ a\left(x,\xi,z\right)\chi\left(\frac{\sqrt{\varepsilon}\left|z\right|}{\delta}\right)\left(1-\chi\left(\frac{\left|z\right|}{R}\right)\right)$ $\left(\eqqcolon a_{\varepsilon,\delta}^{R}(x,\xi,z)\right)$ (18)
-$+\ a\left(x,\xi,z\right)\chi\left(\frac{\left|z\right|}{R}\right)$ $\left(\eqqcolon a_{\varepsilon,R}(x,\xi,z)\right).$ (19)
-Let us define $\chi_{R}\coloneqq\chi\left(\frac{\left|\cdot\right|}{R}\right)$ and $\chi^{R}\coloneqq 1-\chi_{R}$.
-The symbol $a_{\varepsilon}^{\delta}$ through the second scale quantization is supported outside $\mathcal{C}$, at a distance $\delta$ of $\mathcal{C}$.
-Because of (16), $\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a_{\varepsilon}^{\delta}\right)$ is a semiclassical operator which satisfies
-$\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a_{\varepsilon}^{\delta}\right)=\mathrm{Op}_{1}\left(a_{\infty}\left(s,y,\varepsilon\sigma,\varepsilon\eta,\frac{(y,\varepsilon\sigma,\varepsilon\eta)}{|(y,\varepsilon\sigma,\varepsilon\eta)|}\right)\left(1-\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{\delta}\right)\right)\right).$
-So, we can apply the semiclassical theory recalled in Section 2 to the observable $a_{\varepsilon}^{\delta}$.
-In the second scale quantization, the symbol $a_{\varepsilon,R}$ is supported at distance of order $\sqrt{\varepsilon}$ of $\mathcal{C}$ and $a_{\varepsilon,\delta}^{R}$ is supported at distance larger than $\sqrt{\varepsilon}$ of $\mathcal{C}$ but smaller than $\delta$.
-We are particularly interested by the two regions covered by $a_{\varepsilon,\delta}^{R}$ and $a_{\varepsilon,R}$.
-They will describe the measure near the curve $\mathcal{C}$.
-For the observable $a_{\varepsilon,\delta}^{R}$, we have
-$\begin{split}\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a_{\varepsilon,\delta}^{R}\right)=&\ \mathrm{Op}_{1}\left(a\left(s,0,0,0,\frac{y}{\sqrt{\varepsilon}},\sqrt{\varepsilon}\sigma,\sqrt{\varepsilon}\eta\right)\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{\delta}\right)\left(1-\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{R\sqrt{\varepsilon}}\right)\right)\right)\\ &+\mathcal{O}\left(\delta\right).\end{split}$ (20)
-But we choose $R>R_{0}$ then
-$\begin{split}\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a_{\varepsilon,\delta}^{R}\right)=&\ \mathrm{Op}_{1}\left(a_{\infty}\left(s,0,0,0,\frac{(y,\varepsilon\sigma,\varepsilon\eta)}{\left|(y,\varepsilon\sigma,\varepsilon\eta)\right|}\right)\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{\delta}\right)\left(1-\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{R\sqrt{\varepsilon}}\right)\right)\right)\\ &+\mathcal{O}\left(\delta\right).\end{split}$
-By using (14) on $a_{\varepsilon,R}$, we have
-$\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a_{\varepsilon,R}\right)$ $=\mathrm{Op}_{1}\left(a(s,0,0,0,\frac{y}{\sqrt{\varepsilon}},\sqrt{\varepsilon}\sigma,\sqrt{\varepsilon}\eta)\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{R\sqrt{\varepsilon}}\right)\right)\ +\ \mathcal{O}\left(\sqrt{\varepsilon}\right),$
-$=\mathfrak{T}_{\sqrt{\varepsilon}}^{\dagger}\mathrm{Op}_{1}\left(a(s,0,0,0,y,\sqrt{\varepsilon}\sigma,\eta)\chi\left(\frac{|(y,\sqrt{\varepsilon}\sigma,\eta)|}{R}\right)\right)\mathfrak{T}_{\sqrt{\varepsilon}}\ +\ \mathcal{O}\left(\sqrt{\varepsilon}\right).$ (21)
-So we will consider the separable Hilbert space $\mathcal{H}\coloneqq\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)$ and the setting of Appendix B: the functional space $\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathcal{L}\left(\mathcal{H}\right)\right)$, and the two-scale quantization defined for $b\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathcal{L}\left(\mathcal{H}\right)\right)$,
-$\mathrm{Op}_{\sqrt{\varepsilon},1}(b)\coloneqq\mathrm{Op}_{1}\Big{(}b\left(s,\sqrt{\varepsilon}\sigma\right)\Big{)}.$ (22)
-Let us denote, for all $(s,\sigma)\in\mathbb{R}^{2}$, the bounded operator on $\mathcal{H}$ obtained by the Weyl quantization of $(y,\eta)\mapsto a(s,0,0,0,y,\sigma,\eta)$ by $a^{W}(s,0,0,0,y,\sigma,\mathrm{D}_{y})$.
-Then, for the operator $b(s,\sigma)$ defined by for all $(s,\sigma)\in\mathbb{R}^{2}$, $b(s,\sigma)\coloneqq a^{W}(s,0,0,0,y,\sigma,\mathrm{D}_{y})\chi\left(\frac{|(y,\sigma,\mathrm{D}_{y})|}{R}\right)$, we have
-$\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a_{\varepsilon,R}\right)=\mathfrak{T}_{\sqrt{\varepsilon}}^{\dagger}\mathrm{Op}_{\sqrt{\varepsilon},1}(b)\mathfrak{T}_{\sqrt{\varepsilon}}\ +\ \mathcal{O}\left(\sqrt{\varepsilon}\right).$ (23)
-In the following sections, we will take advantage of these properties of $a_{\varepsilon,R}$ and $a_{\varepsilon,\delta}^{R}$.
-3.2.2 Two-scale semiclassical measures
-Let us first consider the stationary case.
-We use the observable of $\mathfrak{A}$ to analyze the oscillations at the scale $\sqrt{\varepsilon}$ and $\delta$ of uniformly bounded families $\left(f^{\varepsilon}\right)_{\varepsilon>0}$ of $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ through $a_{\varepsilon,R}$ and $a_{\varepsilon,\delta}^{R}$, respectively.
-Let us define for $a\in\mathfrak{A}$ the quantity
-$\mathcal{I}^{\varepsilon}_{f}(a)\coloneqq\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}(a)f^{\varepsilon},f^{\varepsilon}\right\rangle_{\mathrm{L}^{2}(\mathbb{R}^{2},\mathbb{C}^{2})}.$
-Theorem 3.3 (Existence of the two-scale semiclassical measure).
-Let ${(f^{\varepsilon})_{\varepsilon>0}}$ uniformly bounded in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$, then, there exist ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ a vanishing sequence of positive numbers and
-  • $\bullet$
-    Two measurable map $\mathrm{d}\nu_{\infty}$ and $\mathrm{d}\gamma$, valued in the set of scalar non negative measures on $\mathcal{C}\times\mathbb{S}^{2}$ and $\mathbb{R}^{2}$, respectively,
-  • $\bullet$
-    A $\mathrm{d}\nu_{\infty}$-measurable map $(s,\omega)\mapsto M_{\infty}(s,\omega)\in\mathbb{C}^{2,2}$, valued in the set of Hermitian trace $1$ positive matrices,
-  • $\bullet$
-    A $\mathrm{d}\gamma(s,\sigma)$ measurable map $(s,\sigma)\mapsto\Gamma(s,\sigma)$, valued in the set of trace $1$ positive operator on $\mathrm{L}^{2}\left(\mathbb{R},\mathbb{C}^{2}\right)$,
-such that for all $a\in\mathfrak{A}$ we have
-$\mathcal{I}^{\varepsilon_{k}}_{f}(a)\underset{k\to+\infty}{\longrightarrow}$ $\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\mathrm{tr}\left(a_{\infty}\left(s,y,\sigma,\eta,\frac{(y,\sigma,\eta)}{|(y,\sigma,\eta)|}\right)\mathrm{d}\mu(s,y,\sigma,\eta)\right)$ (24)
-+×$+\ \int_{\mathbb{R}\times\mathbb{S}^{2}}\mathrm{tr}\Big{(}a_{\infty}(s,0,0,0,\omega)M_{\infty}(s,\omega)\Big{)}\mathrm{d}\nu_{\infty}(s,\omega)$ (25)
-$+\ \int_{\mathbb{R}^{2}}\mathrm{Tr}\Big{(}a^{W}(s,0,0,0,y,\sigma,\mathrm{D}_{y})\Gamma(s,\sigma)\Big{)}\mathrm{d}\gamma(s,\sigma),$ (26)
-where $\mathrm{d}\mu$ the semiclassical measure of ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$ associated to ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ and for all $(s,\sigma)\in\mathbb{R}^{2}$, $a^{W}(s,0,0,0,y,\sigma,\mathrm{D}_{y})$ denotes the bounded operator on $\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)$ obtained by the Weyl quantization of $(y,\eta)\mapsto a(s,0,0,0,y,\sigma,\eta)$.
-The proof of this theorem is in Appendix B.
-We denote by $\Gamma\mathrm{d}\gamma$ the pair $\left(\Gamma,\mathrm{d}\gamma\right)$ with $\Gamma$ a trace one positive operator $\mathrm{d}\gamma$-measurable and $\mathrm{d}\gamma$ a non negative measure.
-We denote by $M_{\infty}\mathrm{d}\nu_{\infty}$ the pair $\left(M_{\infty},\mathrm{d}\nu_{\infty}\right)$ with $M_{\infty}$ a Hermitian trace one positive matrix $\mathrm{d}\nu_{\infty}$-measurable and $\mathrm{d}\nu_{\infty}$ a non negative measure.
-Definition (Two-scale semiclassical measure).
-We call two-scale semiclassical measures, associated with the concentration of ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$ near the curve $\mathcal{C}$ of order between $\sqrt{\varepsilon}$ and $\delta$ and of order $\sqrt{\varepsilon}$, respectively, the pair ${\left(M_{\infty}\mathrm{d}\nu_{\infty},\Gamma\mathrm{d}\gamma\right)}$.
-We will say that $\Gamma\mathrm{d}\gamma$ is a two-scale semiclassical measure at finite distance and $M_{\infty}\mathrm{d}\nu_{\infty}$ is a two-scale semiclassical measure at infinity.
-Remark 3.4 (Link with the limit of the previous observable).
-Let ${(f^{\varepsilon})_{\varepsilon>0}}$ uniformly bounded in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ and ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ a vanishing sequence given by Theorem 3.3, then, with the same notations, for all $a\in\mathfrak{A}$,
-$\limsup_{\delta\to 0}\limsup_{R\to+\infty}\lim_{k\to+\infty}\mathcal{I}^{\varepsilon_{k}}_{f}\left(a_{\varepsilon_{k}}^{\delta}\right)=$ $\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\mathrm{tr}\left(a_{\infty}\left(s,y,\sigma,\eta,\frac{(y,\sigma,\eta)}{|(y,\sigma,\eta)|}\right)\mathrm{d}\mu(s,y,\sigma,\eta)\right),$
-$\limsup_{\delta\to 0}\limsup_{R\to+\infty}\lim_{k\to+\infty}\mathcal{I}^{\varepsilon_{k}}_{f}\left(a_{\varepsilon_{k},\delta}^{R}\right)=$$\int_{\mathbb{R}\times\mathbb{S}^{2}}\mathrm{tr}\Big{(}a_{\infty}(s,0,0,0,\omega)M_{\infty}(s,\omega)\Big{)}\mathrm{d}\nu_{\infty}(s,\omega),$
-$\limsup_{\delta\to 0}\limsup_{R\to+\infty}\lim_{k\to+\infty}\mathcal{I}^{\varepsilon_{k}}_{f}\left(a_{\varepsilon_{k},R}\right)=$ $\int_{\mathbb{R}^{2}}\mathrm{Tr}\Big{(}a^{W}(s,0,0,0,y,\sigma,\mathrm{D}_{y})\Gamma(s,\sigma)\Big{)}\mathrm{d}\gamma(s,\sigma).$
-Example.
-Let us consider, for all $(s,y)\in\mathbb{R}^{2}$,
-$\phi_{\varepsilon}(s,y)=\frac{1}{\varepsilon^{1/4}}u_{\varepsilon}\left(s\right)f\left(\frac{y}{\sqrt{\varepsilon}}\right)\vec{V}$
-with $\vec{V}\in\mathbb{C}^{2}$, $(f,u_{\varepsilon})\in\mathscr{S}(\mathbb{R},\mathbb{C})^{2}$ and the knowledge of $\mathrm{d}\rho$ the semiclassical measure of $(u_{\varepsilon})_{\varepsilon>0}$ at the scale $\sqrt{\varepsilon}$, then for all $a\in\mathscr{C}^{\infty}_{c}(\mathbb{R}^{7},\mathbb{C}^{2,2})$,
-$\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}(a)\phi_{\varepsilon},\phi_{\varepsilon}\right\rangle\underset{\varepsilon\to 0}{\longrightarrow}\int_{\mathbb{R}^{2}}\mathrm{Tr}\Big{(}a^{W}(s,0,0,0,y,\sigma,\mathrm{D}_{y})\Gamma(s,\sigma)\Big{)}\mathrm{d}\gamma(s,\sigma),$
-where
-$\Gamma(s,\sigma)\mathrm{d}\gamma(s,\sigma)=\frac{1}{2\pi}\Pi_{f\vec{V}}\mathrm{d}\rho(s,\sigma),$
-with $\Pi_{f\vec{V}}$ the projector on $\mathrm{L}^{2}\left(\mathbb{R},\mathbb{C}^{2}\right)$ over the function $f\vec{V}$.
-Notably, if $u_{\varepsilon}$ is a wave packet, which means for all $s\in\mathbb{R}$,
-$u_{\varepsilon}\left(s\right)\coloneqq\frac{1}{\varepsilon^{1/4}}u\left(\frac{s}{\sqrt{\varepsilon}}\right)$
-with $u\in\mathscr{S}(\mathbb{R},\mathbb{C})$, then the semiclassical measure of $(u_{\varepsilon})_{\varepsilon>0}$ at the scale $\sqrt{\varepsilon}$ is $\mathrm{d}\rho(\sigma)=\left|\mathcal{F}u(\sigma)\right|^{2}\mathrm{d}\sigma$ (according to the case $\alpha=1$ of Example 2.1 but with a $\sqrt{\varepsilon}$-Weyl quantization) so
-$\Gamma(s,\sigma)\mathrm{d}\gamma(s,\sigma)=\left|\mathcal{F}u(\sigma)\right|^{2}\Pi_{f\vec{V}}\frac{\mathrm{d}\sigma}{2\pi}.$
-Remark 3.5.
-Let ${(f^{\varepsilon})_{\varepsilon>0}}$ uniformly bounded in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ and ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ a vanishing sequence given by Theorem 3.3.
-According to the notations of Theorem 3.3, the two-scale semiclassical measures ${\left(M_{\infty}\mathrm{d}\nu_{\infty},\Gamma\mathrm{d}\gamma\right)}$ satisfy for all $a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)$,
-$\mathcal{I}^{\varepsilon_{k}}_{f}(a)\underset{k\to+\infty}{\longrightarrow}$ $\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\mathrm{tr}\Big{(}a\left(s,y,\sigma,\eta\right)\mathrm{d}\mu(s,y,\sigma,\eta)\Big{)}$
-+×$+\ \int_{\mathbb{R}\times\mathbb{S}^{2}}\mathrm{tr}\Big{(}a(s,0,0,0)M_{\infty}(s,\omega)\Big{)}\mathrm{d}\nu_{\infty}(s,\omega)$
-$+\ \int_{\mathbb{R}^{2}}\mathrm{Tr}\Big{(}a(s,0,0,0)\Gamma(s,\sigma)\Big{)}\mathrm{d}\gamma(s,\sigma).$
-Corollary 3.6 (A condition to have $M_{\infty}\mathrm{d}\nu_{\infty}=0$).
-Let ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$ uniformly bounded in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ such that there exists $\chi$ a cut off function with $\chi\in\mathscr{C}^{\infty}_{c}(\mathbb{R},\mathbb{R})$ supported in $[-1,1]$ and $\chi\equiv 1$ on $[-1/2,1/2]$ such that
-$\limsup_{\varepsilon\to 0}\left\lVert\mathrm{Op}_{1}\left(1-\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{R\sqrt{\varepsilon}}\right)\right)f^{\varepsilon}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\underset{R\to+\infty}{\longrightarrow}0.$
-Then, for all two-scale semiclassical measure $M_{\infty}\mathrm{d}\nu_{\infty}$ at infinity of ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$,
-$M_{\infty}\mathrm{d}\nu_{\infty}=0.$
-Proof of Corollary 3.6.
-Let ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$ uniformly bounded in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ and $\chi$ a cut off function, $\chi\in\mathscr{C}^{\infty}_{c}(\mathbb{R},[0,1])$ supported in $[-1,1]$ and $\chi\equiv 1$ on $[-1/2,1/2]$ such that
-$\limsup_{\varepsilon\to 0}\left\lVert\mathrm{Op}_{1}\left(1-\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{R\sqrt{\varepsilon}}\right)\right)f^{\varepsilon}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\underset{R\to+\infty}{\longrightarrow}0.$
-For all $a\in\mathfrak{A}$, we consider,
-$\alpha(s,y,\sigma,\eta)$ $\coloneqq a\left(s,y,\varepsilon\sigma,\varepsilon\eta,\frac{y}{\sqrt{\varepsilon}},\sqrt{\varepsilon}\sigma,\sqrt{\varepsilon}\eta\right)\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{\delta}\right),$
-$\beta(y,\sigma,\eta)$ $\coloneqq 1-\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{R\sqrt{\varepsilon}}\right).$
-According to pseudo-differential calculus and of the inequality (16), we fix $\delta$ and let $\varepsilon$ goes to 00 then $R$ goes to +$\infty$+ ∞, so
-$\mathrm{Op}_{1}\Big{(}\alpha(s,y,\sigma,\eta)\beta(y,\sigma,\eta)\Big{)}=$ $\ \mathrm{Op}_{1}\Big{(}\alpha(s,y,\sigma,\eta)\Big{)}\mathrm{Op}_{1}\Big{(}\beta(y,\sigma,\eta)\Big{)}$
-$+\ \mathcal{O}_{\mathcal{L}\left(\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)}\left(\sqrt{\varepsilon}\right)+\mathcal{O}_{\mathcal{L}\left(\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)}\left(\frac{1}{R}\right).$
-Moreover,
-$\mathrm{Op}_{1}\Big{(}\alpha(s,y,\sigma,\eta)\Big{)}=\mathfrak{T}_{\sqrt{\varepsilon}}^{\dagger}\mathrm{Op}_{1}\left(a\left(s,\sqrt{\varepsilon}y,\varepsilon\sigma,\sqrt{\varepsilon}\eta,y,\sqrt{\varepsilon}\sigma,\eta\right)\chi\left(\frac{\sqrt{\varepsilon}|(y,\sqrt{\varepsilon}\sigma,\eta)|}{\delta}\right)\right)\mathfrak{T}_{\sqrt{\varepsilon}}.$
-By application of the Calderón-Vaillancourt Theorem (7),
-$\left\lVert\mathrm{Op}_{1}\Big{(}\alpha(s,y,\sigma,\eta)\Big{)}\right\rVert_{\mathcal{L}\left(\mathrm{L}^{2}\right)}$ $=\left\lVert\mathrm{Op}_{1}\left(a\left(s,\sqrt{\varepsilon}y,\varepsilon\sigma,\sqrt{\varepsilon}\eta,y,\sqrt{\varepsilon}\sigma,\eta\right)\chi\left(\frac{\sqrt{\varepsilon}|(y,\sqrt{\varepsilon}\sigma,\eta)|}{\delta}\right)\right)\right\rVert_{\mathcal{L}\left(\mathrm{L}^{2}\right)}$
-$\leqslant C(a)$
-where $C$ does not depend of $\varepsilon$. Then
-$\mathcal{I}^{\varepsilon}_{f}\left(a_{\varepsilon,\delta}^{R}\right)=$ $\ \Big{\langle}\mathrm{Op}_{1}\Big{(}\alpha(s,y,\sigma,\eta)\Big{)}f^{\varepsilon},\mathrm{Op}_{1}\Big{(}\beta(y,\sigma,\eta)\Big{)}f^{\varepsilon}\Big{\rangle}_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}$
-$+\ \mathcal{O}\left(\sqrt{\varepsilon}\right)+\mathcal{O}\left(\frac{1}{R}\right),$
-$\left|\mathcal{I}^{\varepsilon}_{f}\left(a_{\varepsilon,\delta}^{R}\right)\right|\leqslant$ $\ \left\lVert\mathrm{Op}_{1}\Big{(}\alpha(s,y,\sigma,\eta)\Big{)}f^{\varepsilon}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\left\lVert\mathrm{Op}_{1}\Big{(}\beta(y,\sigma,\eta)\Big{)}f^{\varepsilon}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}$
-$+\ \mathcal{O}\left(\sqrt{\varepsilon}\right)+\mathcal{O}\left(\frac{1}{R}\right),$
-$\leqslant$ $\ C(a)\left\lVert f^{\varepsilon}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\left\lVert\mathrm{Op}_{1}\left(1-\chi\left(\frac{|(y,\varepsilon\sigma,\varepsilon\eta)|}{R\sqrt{\varepsilon}}\right)\right)f^{\varepsilon}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}$
-$+\ \mathcal{O}\left(\sqrt{\varepsilon}\right)+\mathcal{O}\left(\frac{1}{R}\right).$
-Then,
-$\limsup_{\delta\to 0}\limsup_{R\to+\infty}\limsup_{\varepsilon\to 0}\mathcal{I}^{\varepsilon}_{f}\left(a_{\varepsilon,\delta}^{R}\right)=0.$
-We conclude by using Remark 3.4. ∎
-3.2.3 Time dependent two-scale semiclassical measures
-As in Section 2.2.2, one can extend the definition of two-scale semiclassical measure to time dependent families.
-More precisely, we consider families ${\left(f^{\varepsilon}_{t}\right)_{\varepsilon>0}}$ that are uniformly bounded in $\mathrm{L}^{\infty}\left(\mathbb{R},\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)$.
-Let $a\in\mathfrak{A},\Xi\in\mathscr{C}_{c}^{\infty}(\mathbb{R},\mathbb{C})$, we define the quantity
-$\mathcal{I}^{\varepsilon}_{f}(\Xi,a)\coloneqq\int_{\mathbb{R}}\Xi(t)\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}(a)f_{t}^{\varepsilon},f_{t}^{\varepsilon}\right\rangle_{\mathrm{L}^{2}(\mathbb{R}^{2},\mathbb{C}^{2})}\mathrm{d}t.$
-As before, we are interested by the limit of $\mathcal{I}^{\varepsilon}_{f}(\Xi,a)$ when $\varepsilon$ goes to 00.
-Moreover, if $a$ does not depend of $z$, then the limit will be the semiclassical measure as defined in (8).
-Adapting the context of  [10] to time dependent families, one defines two-scale semiclassical measures over the curve $\mathcal{C}$.
-Indeed, by the same argument of Section 2.2.2, we can extend Theorem 3.3 to time families.
-Therefore, such as before, the limit measures will be absolutely continuous in time variable with respect to Lebesgue measure.
-Theorem 3.7 (Existence of the time-averaged two-scale semiclassical measure).
-Let ${(f^{\varepsilon}_{t})_{\varepsilon>0}}$ uniformly bounded in $\mathrm{L}^{\infty}\left(\mathbb{R},\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)$, then, there exist ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ a vanishing sequence and
-  • $\bullet$
-    Two measurables map $t\mapsto\mathrm{d}\nu^{t}_{\infty}$ and $t\mapsto\mathrm{d}\gamma^{t}$, valued in the set of scalar non negative measures on $\mathcal{C}\times\mathbb{S}^{2}$ and $\mathbb{R}^{2}$, respectively,
-  • $\bullet$
-    A $\mathrm{d}\nu^{t}_{\infty}\otimes\mathrm{d}t$-measurable map $(t,s,\omega)\mapsto M^{t}_{\infty}(s,\omega)\in\mathbb{C}^{2,2}$, valued in the set of Hermitian trace $1$ positive matrices,
-  • $\bullet$
-    A $\mathrm{d}\gamma^{t}(s,\sigma)\otimes\mathrm{d}t$ measurable map $(t,s,\sigma)\mapsto\Gamma^{t}(s,\sigma)$, valued in the set of trace $1$ positive operator on $\mathrm{L}^{2}\left(\mathbb{R},\mathbb{C}^{2}\right)$,
-such that for all $(a,\Xi)\in\mathfrak{A}\times\mathscr{C}_{c}^{\infty}(\mathbb{R},\mathbb{C})$,
-$\mathcal{I}^{\varepsilon_{k}}_{f}(\Xi,a)\underset{k\to+\infty}{\longrightarrow}$ $\int_{\mathbb{R}}\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\Xi(t)\mathrm{tr}\left(a_{\infty}\left(s,y,\sigma,\eta,\frac{(y,\sigma,\eta)}{|(y,\sigma,\eta)|}\right)\mathrm{d}\mu_{t}(s,y,\sigma,\eta)\right)\mathrm{d}t$
-+×$+\ \int_{\mathbb{R}}\int_{\mathbb{R}\times\mathbb{S}^{2}}\Xi(t)\mathrm{tr}\Big{(}a_{\infty}(s,0,0,0,\omega)M_{\infty}^{t}(s,\omega)\Big{)}\mathrm{d}\nu_{\infty}^{t}(s,\omega)\mathrm{d}t$
-$+\ \int_{\mathbb{R}}\int_{\mathbb{R}^{2}}\Xi(t)\mathrm{Tr}\Big{(}a^{W}(s,0,0,0,y,\sigma,\mathrm{D}_{y})\Gamma^{t}(s,\sigma)\Big{)}\mathrm{d}\gamma^{t}(s,\sigma)\mathrm{d}t,$
-where $\mathrm{d}\mu_{t}$ is the semiclassical measure of ${(f^{\varepsilon}_{t})_{\varepsilon>0}}$ associated to ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ and for all $(s,\sigma)\in\mathbb{R}^{2}$, $a^{W}(s,0,0,0,y,\sigma,\mathrm{D}_{y})$ denotes the bounded operator on $\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)$ obtained by the Weyl quantization of $(y,\eta)\mapsto a(s,0,0,0,y,\sigma,\eta)$.
-Remark (Geometrical point of view).
-In the parametrization $(s,y)$ of ×$\mathbb{R}\times\mathrm{I}$, the variables $(s,y,\sigma,\eta)$ parametrized $\mathrm{T}^{*}\mathbb{R}^{2}$.
-Moreover the variable $s$ parametrizes the curve $\mathrm{E}\subset\mathbb{R}^{2}$.
-The variable $y$ parametrizes the normal fiber to the curve $\mathrm{E}$ at the point $\gamma(s)$.
-In the computation of $\mathrm{Tr}$, the space $\mathrm{L}^{2}(\mathbb{R}_{y},\mathbb{C}^{2})$ can be understand as the space $\mathrm{L}^{2}(N_{\gamma(s)}\mathrm{E},\mathbb{C}^{2})$.
-The variable $\sigma$ denotes the variable of the fiber of the cotangent bundle $\mathrm{T}^{*}\mathrm{E}$ over $\mathrm{E}$.
-For $\gamma(s)\in\mathrm{E}$ and $\mathrm{T}_{s}\mathrm{E}$, the tangent to $\mathrm{E}$ above $s$, the fiber of the cotangent bundle $\mathrm{T}^{*}\mathrm{E}$ over $\gamma(s)$ is $(\mathrm{T}_{s}\mathrm{E})^{*}\sim\mathbb{R}$.
-The measure $\mathrm{d}\gamma^{t}$ defines a measure on $\mathrm{T}^{*}\mathrm{E}$.
-Since $\mathcal{C}=\{y=\eta=\sigma=0\}\subset\mathrm{T}^{*}\mathbb{R}^{2}$, $\mathcal{C}\times\mathbb{S}^{2}$ coincides to the normal bundle in sphere to $\mathcal{C}$: $\mathrm{S}\mathcal{C}$.
-The fibers of $\mathrm{S}\mathcal{C}$ over $\rho\in\mathcal{C}$ is obtained as follows : let $\mathrm{T}_{\rho}\mathcal{C}$ be the tangent space to $\mathcal{C}$ over $\rho$, define $\mathrm{N}_{\rho}\mathcal{C}$ the subspace of $\left(\mathrm{T}_{\rho}\left(\mathrm{T}^{*}\mathbb{R}^{2}\right)\right)^{*}$ consisting of the linear forms on $\mathrm{T}_{\rho}\left(\mathrm{T}^{*}\mathbb{R}^{2}\right)$ which vanish on $\mathrm{T}_{\rho}\mathcal{C}$, $S_{\rho}\mathcal{C}$ is obtained by taking the quotient of $\mathrm{N}_{\rho}\mathcal{C}$ by the action of $\mathbb{R}_{\geqslant 0}$ by homotheties.
-3.3 Main result
-Now we will describe the semiclassical measure of Theorem 3.7 for the solution to equation (1).
-Let us define some quantities involved in the main theorem.
-Let us define the Hermite functions $\left(\mathfrak{h}_{m}\right)_{m\in\mathbb{N}_{\geqslant 0}}$ as the Hilbertian basis of $\mathrm{L}^{2}\left(\mathbb{R},\mathbb{C}\right)$ satisfying for all $y\in\mathbb{R}$ and for all $m\in\mathbb{N}_{\geqslant 0}$,
-$-\frac{\mathrm{d}^{2}\mathfrak{h}_{m}}{\mathrm{d}y^{2}}(y)+y^{2}\mathfrak{h}_{m}(y)=(2m+1)\mathfrak{h}_{m}(y).$ (27)
-Let us consider for all $(s,y,\sigma)\in\mathbb{R}^{3}$,
-$g_{0}^{s}(y)\coloneqq\frac{1}{\sqrt{r(s)}}\begin{pmatrix}0\\ \mathfrak{h}_{0}\big{(}r(s)y\big{)}\end{pmatrix}\;\;$ $,\;\;g_{n}^{s,\sigma}(y)\coloneqq\frac{1}{\sqrt{2r(s)}}\sqrt{1+\frac{\sigma}{\lambda_{n}(s,\sigma)}}\begin{pmatrix}\mathfrak{h}_{|n|-1}\big{(}r(s)y\big{)}\\ \frac{\sqrt{2|n|}r(s)}{\lambda_{n}(s,\sigma)+\sigma}\mathfrak{h}_{|n|}\big{(}r(s)y\big{)}\end{pmatrix},$ (28)
-$\Pi_{0}(s)\coloneqq g_{0}^{s}\otimes g_{0}^{s}\;\;$ $,\;\;\Pi_{n}(s,\sigma)\coloneqq g_{n}^{s,\sigma}\otimes g_{n}^{s,\sigma},$ (29)
-where $\left(\lambda_{n}\right)_{n\in\mathbb{Z}}$ were defined in (5), $r(s)\coloneqq|\nabla\mathrm{m}(\gamma(s))|^{1/2}$.
-Although the functions $g_{0}^{s}$ and $\Pi_{0}$ do not depend on $\sigma$, we also will denote them by $g_{0}^{s,\sigma}$ and $\Pi_{0}(s,\sigma)$.
-Let us consider for all $s\in\mathbb{R}$ and for all $z\in\mathbb{R}^{3}$,
-$\Pi_{\pm}^{\infty}(s,z)\coloneqq\frac{1}{2}\mathrm{Id}\pm\frac{1}{2\sqrt{z_{\sigma}^{2}+z_{y}^{2}r(s)^{4}+z_{\eta}^{2}}}\begin{pmatrix}z_{\sigma}&z_{y}r(s)^{2}+iz_{\eta}\\ z_{y}r(s)^{2}-iz_{\eta}&-z_{\sigma}\end{pmatrix}$ (30)
-We are going to use the two-scale semiclassical measures of the family $\left(\varphi_{t}^{\varepsilon}\right)_{\varepsilon>0}$ to describe the semiclassical measure of $\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}$ near the curve $\mathcal{C}$.
-We already know its semiclassical measure and its evolution outside $\mathcal{C}$.
-Therefore, we will focus on terms (25) and (26).
-Theorem 3.8 (Evolution of the two-scale semiclassical measure in normal geodesic coordinates).
-Assume Assumption 1.2 and 1.3.
-Let $\left(\varphi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ uniformly bounded in $\mathrm{L}^{\infty}\left(\mathbb{R},\mathrm{L}^{2}\left(\mathbb{R}\times\mathrm{I},\mathbb{C}^{2}\right)\right)$ defined in (11).
-Let ${\left(M^{t}_{\infty}\mathrm{d}\nu^{t}_{\infty},\Gamma^{t}\mathrm{d}\gamma^{t}\right)}$ be a pair of two-scale semiclassical measure of $\left(\varphi^{\varepsilon}_{t}\right)_{\varepsilon>0}$, then we have the following properties.
-  1. 1.
-    There exists a family of non negative scalar Radon measure $\left(\mathrm{d}\gamma_{n}^{t}\right)_{n\in\mathbb{Z}}$ such that
-    $\Gamma^{t}\mathrm{d}\gamma^{t}=\sum_{n\in\mathbb{Z}}\Pi_{n}\mathrm{d}\gamma_{n}^{t}.$ (31)
-    Moreover, for all $n\in\mathbb{Z}$, $\mathrm{d}\gamma_{n}^{t}$ satisfies the following equation
-    $\left\{\begin{array}[]{rll}\partial_{t}\ \mathrm{d}\gamma_{n}^{t}&=&\left\{\mathrm{d}\gamma_{n}^{t},\lambda_{n}\right\}\\ \mathrm{d}\gamma_{n}^{t}|_{t=0}&=&\mathrm{d}\gamma_{n,0}\end{array}\right.,$
-    with $\mathrm{d}\gamma_{n,0}=\mathrm{tr}\Big{(}\Pi_{n}\Gamma^{0}\Big{)}\mathrm{d}\gamma^{0}$ and $\Gamma^{0}\mathrm{d}\gamma^{0}$ is a two-scale semiclassical measure associated to the initial data.
-  2. 2.
-    There exists a family of non negative scalar Radon measure $\mathrm{d}\nu^{t}_{\pm}$ such that
-    $M^{t}_{\infty}\mathrm{d}\nu^{t}_{\infty}=\Pi_{+}^{\infty}\mathrm{d}\nu^{t}_{+}+\Pi_{-}^{\infty}\mathrm{d}\nu^{t}_{-}.$ (32)
-    Moreover, $\mathrm{d}\nu_{\pm}^{t}$ satisfies the invariance $\mathrm{div}_{\omega}\left(\vec{V}^{\infty}(s,\omega)\mathrm{d}\nu_{\pm}^{t}(s,\omega)\right)=0,$ with for all $s\in\mathbb{R}$ and for all $\omega\coloneqq\left(\omega_{y},\omega_{\sigma},\omega_{\eta}\right)\in\mathbb{S}^{2}$,
-    $\vec{V}^{\infty}(s,\omega)\coloneqq\begin{pmatrix}\left(\omega_{y}\omega_{\eta}\left(1-r(s)^{4}\right)-1\right)\omega_{\eta}\\ \left(1-r(s)^{4}\right)\omega_{y}\omega_{\eta}\omega_{\sigma}\\ \left(\omega_{y}\omega_{\eta}\left(1-r(s)^{4}\right)+r(s)^{4}\right)\omega_{y}\end{pmatrix}.$
-The proof of this theorem is in Section 4.3.
-Remark (Continuity of the two-scale semiclassical measure at finite distance).
-From the evolution (6), we can deduce that the measure $t\mapsto\Gamma^{t}\mathrm{d}\gamma^{t}$ is continuous and entirely determined by the evolution equation and the initial data.
-Remark (Value at zero of the two-scale semiclassical measure at finite distance).
-The value at zero of $\Gamma^{t}\mathrm{d}\gamma^{t}$ is not necessarily the two-scale semiclassical measure at finite distance of the initial condition.
-As we will see in the proof, $\Gamma^{t}(s,\sigma)$ commutes with the operator $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$ because of the PDE satisfied by $\left(\varphi^{\varepsilon}_{t}\right)_{\varepsilon>0}$.
-Nevertheless, there exists initial condition such that its two-scale semiclassical measure at finite distance does not commute with the operator $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$.
-We will see an example in Remark 4.4.
-Corollary 3.9 (Evolution consequence).
-With the notation of Theorem 3.8, if
-$\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\mathrm{tr}\Big{(}\mathrm{d}\mu_{0}(x,\xi)\Big{)}+\int_{\mathbb{R}^{2}}\mathrm{Tr}\Big{(}\Gamma^{0}(s,\sigma)\Big{)}\mathrm{d}\gamma^{0}(s,\sigma)=1,$
-with $\mathrm{d}\mu_{0}$ a semiclassical measure of $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ associated to the same vanishing sequence $\left(\varepsilon_{n}\right)_{n\in\mathbb{N}}$ of the two-scale semiclassical measure $\left(M^{0}_{\infty}\mathrm{d}\nu^{0}_{\infty},\Gamma^{0}\mathrm{d}\gamma^{0}\right)$, then, for almost all $t$,
-$M^{t}_{\infty}(s,\omega)\mathrm{d}\nu^{t}_{\infty}(s,\omega)=0.$
-Proof of Corollary 3.9.
-With the notation of Theorem 3.8, the evolution of $(\mathrm{d}\mu_{t},\Gamma^{t}\mathrm{d}\gamma^{t})$ implies that for all $t\in\mathbb{R}$,
-$\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\mathrm{tr}\Big{(}\mathrm{d}\mu_{t}(x,\xi)\Big{)}$ $=\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\mathrm{tr}\Big{(}\mathrm{d}\mu_{0}(x,\xi)\Big{)},$
-$\int_{\mathbb{R}^{2}}\mathrm{Tr}\Big{(}\Gamma^{t}(s,\sigma)\Big{)}\mathrm{d}\gamma^{t}(s,\sigma)$ $=\int_{\mathbb{R}^{2}}\mathrm{Tr}\Big{(}\Gamma^{0}(s,\sigma)\Big{)}\mathrm{d}\gamma^{0}(s,\sigma).$
-By taking $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}\in\mathrm{L}^{2}(\mathbb{R}^{2},\mathbb{C}^{2})$ normalized such that
-$\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\mathrm{tr}\Big{(}\mathrm{d}\mu_{0}(x,\xi)\Big{)}+\int_{\mathbb{R}^{2}}\mathrm{Tr}\Big{(}\Gamma^{0}(s,\sigma)\Big{)}\mathrm{d}\gamma^{0}(s,\sigma)=1=\left\lVert\psi^{\varepsilon}_{0}\right\rVert_{2}^{2},$
-we have the following equality because of the preservation of the $\mathrm{L}^{2}$-norm
-$\left\lVert\psi^{\varepsilon}_{t}\right\rVert_{2}^{2}=$ $\ \int_{\mathbb{R}^{4}\backslash\mathcal{C}}\mathrm{tr}\Big{(}\mathrm{d}\mu_{t}(x,\xi)\Big{)}+\int_{\mathbb{R}\times\mathbb{S}^{2}}\mathrm{tr}\Big{(}M^{t}_{\infty}(s,\omega)\Big{)}\mathrm{d}\nu^{t}_{\infty}(s,\omega)$
-$+\int_{\mathbb{R}^{2}}\mathrm{Tr}\Big{(}\Gamma^{t}(s,\sigma)\Big{)}\mathrm{d}\gamma^{t}(s,\sigma),$
-$=$ $\ 1$
-Therefore, for almost all $t$,
-$M^{t}_{\infty}(s,\omega)\mathrm{d}\nu^{t}_{\infty}(s,\omega)=0.$
-The evolution of the two-scale semiclassical measure in normal geodesic coordinates allows us to describe the general evolution of any initial normalized condition for equation (1).
-Theorem 3.10 (Evolution of the semiclassical measure of solution to equation (1)).
-Assume Assumption 1.2 and 1.3.
-Let $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ uniformly bounded in $\mathrm{L}^{\infty}\left(\mathbb{R},\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)$ solves (1) with initial normalized condition $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$.
-Then, there exist ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ a vanishing sequence of positive numbers and
-  • $\bullet$
-    a semiclassical measure $\mathrm{d}\mu_{t}$, evolving as described in Lemma 2.3,
-  • a family of non negative scalar Radon measure $\left(\mathrm{d}\gamma_{n}^{t}\right)_{n\in\mathbb{Z}}$ satisfying (6),
-  • $\bullet$
-    a two-scale semiclassical measure ${M^{t}_{\infty}\mathrm{d}\nu^{t}_{\infty}}$ at infinity described in Theorem 3.8,
-such that for all $(a,\Xi)\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)\times\mathscr{C}_{c}^{\infty}(\mathbb{R},\mathbb{C})$,
-$\mathcal{I}^{\varepsilon_{k}}_{\psi}(\Xi,a)\underset{k\to+\infty}{\longrightarrow}\ \int_{\mathbb{R}}\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\Xi(t)\mathrm{tr}\Big{(}a(x,\xi)\mathrm{d}\mu_{t}(x,\xi)\Big{)}\mathrm{d}t+\int_{\mathbb{R}}\int_{\mathbb{R}}\Xi(t)\mathrm{tr}\Big{(}a\big{(}\gamma(s),0\big{)}\mathrm{d}\rho_{t}(s)\Big{)}\mathrm{d}t,$
-with for all $s\in\mathbb{R}$,
-$\mathrm{d}\rho_{t}(s)=\int_{\mathbb{S}^{2}}\mathrm{U}_{\theta(s)}^{\dagger}M^{t}_{\infty}(s,\omega)\mathrm{U}_{\theta(s)}\mathrm{d}\nu^{t}_{\infty}(s,\omega)+\frac{1}{2}\sum_{n\in\mathbb{Z}}\int_{\mathbb{R}}\begin{pmatrix}1&\frac{\mathrm{i}\sigma}{\lambda_{n}(s,\sigma)}e^{-\mathrm{i}\theta(s)}\\ \frac{-\mathrm{i}\sigma}{\lambda_{n}(s,\sigma)}e^{\mathrm{i}\theta(s)}&1\end{pmatrix}\mathrm{d}\gamma^{t}_{n}(s,\mathrm{d}\sigma),$
-where $\mathrm{U}_{\theta(s)}$ is defined in (10).
-The proof of this theorem is in Section 4.3.
-4 Proofs of the main result
-4.1 Change of variable on semiclassical measure
-First of all, we will see the consequences of the change of variable from Proposition 3.1 on the semiclassical measure of $\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}$ solution to equation (1) and the link with the two-scale semiclassical measure of the family $\left(\varphi_{t}^{\varepsilon}\right)_{\varepsilon>0}$ defined in (11).
-Lemma 4.1 (Semiclassical measure and normal geodesic coordinates).
-All semiclassical measure $\mathrm{d}\mu_{t}\mathrm{d}t$ of $(\psi^{\varepsilon}_{t})_{\varepsilon>0}$ solution to equation (1) associated to $(\varepsilon_{k})_{k\in\mathbb{N}}$ satisfies
-$\mathrm{d}\mu_{t}\mathrm{d}t\mathds{1}_{\mathcal{C}}=\Phi^{*}\Big{(}\mathrm{U}_{\theta(s)}^{\dagger}\mathrm{d}\rho_{t}\mathrm{U}_{\theta(s)}\Big{)}\mathrm{d}t\mathds{1}_{\mathcal{C}}$
-where $\mathrm{U}_{\theta(s)}$ is defined in (10) and $\mathrm{d}\rho_{t}$ is a semiclassical measure on $\mathbb{R}_{s}\times\mathrm{I}_{1/2}\times\mathbb{R}_{\sigma,\eta}^{2}\times\mathbb{R}_{t}$ associated to $\left(\varphi_{t}^{\varepsilon}\right)_{\varepsilon>0}$ and $\left(\varepsilon_{k}\right)_{k\in\mathbb{N}}$.
-Proof of Lemma 4.1.
-Let $a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)$ such that $\mathrm{supp}(a)\subset\Omega_{1/2}\times\mathbb{R}^{2}$ and $\chi$ the cut-off function taken in (11), then for all $N\in\mathbb{N}$,
-$\left\langle\mathrm{Op}_{\varepsilon}\left(a\right)\psi^{\varepsilon}_{t},\psi^{\varepsilon}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}$ $=\left\langle\mathrm{Op}_{\varepsilon}\left(a\right)\chi\psi^{\varepsilon}_{t},\chi\psi^{\varepsilon}_{t}\right\rangle_{\mathrm{L}^{2}\left(\Omega_{1/2},\mathbb{C}^{2}\right)}+\mathcal{O}\left(\varepsilon^{N}\right)$
-$=\left\langle\mathrm{U}\mathrm{Op}_{\varepsilon}\left(a\right)\mathrm{U}^{\dagger}\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}\times\mathrm{I}_{1/2},\mathbb{C}^{2}\right)}+\mathcal{O}\left(\varepsilon^{N}\right).$
-Since $\rm U=U_{\theta(s)}V_{\Phi}$, $\rm U_{\theta(s)}$ being the multiplication by a function depending only of the variable $s$ and $\rm V_{\Phi}$ a change of variables, standard results on semiclassical pseudodifferential calculus ([27, Theorem 9.3]) imply
-$\left\langle\mathrm{Op}_{\varepsilon}\left(a\right)\psi^{\varepsilon}_{t},\psi^{\varepsilon}_{t}\right\rangle=$ $\ \left\langle\mathrm{U}_{\theta(s)}\mathrm{Op}_{\varepsilon}\Big{(}\mathrm{V}_{\Phi}a\mathrm{V}_{\Phi}^{\dagger}\Big{)}\mathrm{U}_{\theta(s)}^{\dagger}\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}\times\mathrm{I}_{1/2},\mathbb{C}^{2}\right)}+\mathcal{O}\left(\varepsilon\right)$
-$=$ $\ \left\langle\mathrm{Op}_{\varepsilon}\left(\mathrm{U}_{\theta(s)}a\mathrm{U}_{\theta(s)}^{\dagger}\left(\Phi(s,y),{}^{t}\nabla\left(\Phi^{-1}\right)\Big{(}\Phi(s,y)\Big{)}\left[\sigma,\eta\right]\right)\right)\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}\times\mathrm{I}_{1/2},\mathbb{C}^{2}\right)}$
-$+\mathcal{O}\left(\varepsilon\right).$
-For simplicity, we assume $\left(\varepsilon\right)_{\varepsilon>0}$ realizes the semiclassical measure $\mathrm{d}\mu_{t}\mathrm{d}t$ of $(\psi^{\varepsilon}_{t})_{\varepsilon>0}$.
-Passing to the limit, we obtain
-$\left\langle\mathrm{Op}_{\varepsilon}\left(a\right)\psi^{\varepsilon}_{t},\psi^{\varepsilon}_{t}\right\rangle$ $\underset{\varepsilon\to 0}{\longrightarrow}\int_{\mathbb{R}\times\mathrm{I}_{1/2}\times\mathbb{R}^{2}}\mathrm{Tr}\left(\mathrm{U}_{\theta(s)}a\mathrm{U}_{\theta(s)}^{\dagger}\left(\Phi(s,y),{}^{t}\nabla\left(\Phi^{-1}\right)\Big{(}\Phi(s,y)\Big{)}\left[\sigma,\eta\right]\right)\mathrm{d}\rho_{t}(s,y,\sigma,\eta)\right)$
-=×$=\int_{\mathbb{R}\times\mathrm{I}_{1/2}\times\mathbb{R}^{2}}\mathrm{Tr}\left(a\left(\Phi(s,y),{}^{t}\nabla\left(\Phi^{-1}\right)\Big{(}\Phi(s,y)\Big{)}\left[\sigma,\eta\right]\right)\Big{(}\mathrm{U}_{\theta(s)}^{\dagger}\mathrm{d}\rho_{t}(s,y,\sigma,\eta)\mathrm{U}_{\theta(s)}\Big{)}\right)$
-$=\int_{\Omega_{1/2}\times\mathbb{R}^{2}}\mathrm{Tr}\left(a(x,\xi)\Phi^{*}\Big{(}\mathrm{U}_{\theta(s)}^{\dagger}\mathrm{d}\rho_{t}\mathrm{U}_{\theta(s)}\Big{)}(x,\xi)\right).$
-We conclude by taking the restriction of the semiclassical measure $\mathrm{d}\rho_{t}$ to the curve $\mathcal{C}$. ∎
-According to the definition (15) of the rescaling operator $\mathfrak{T}_{\sqrt{\varepsilon}}$, we define the family $\left(u^{\varepsilon}_{t}\right)_{\varepsilon>0}$ by
-$\mathfrak{T}_{\sqrt{\varepsilon}}^{\dagger}\left[u^{\varepsilon}_{t}\right](s,y)=\varphi_{t}^{\varepsilon}(s,y)\quad,\quad\forall(t,s,y)\in\mathbb{R}\times\mathbb{R}\times\mathrm{I}_{1/2}.$
-The family $\left(u^{\varepsilon}_{t}\right)_{\varepsilon>0}$ is defined on ×$\mathbb{R}\times\mathrm{I}_{\frac{1}{2\sqrt{\varepsilon}}}$ but we will study it on ×$\mathbb{R}\times\mathrm{I}_{\sqrt{\varepsilon}}^{\delta}$ with
-$\mathrm{I}_{\sqrt{\varepsilon}}^{\delta}\coloneqq\left(-\frac{\delta}{\sqrt{\varepsilon}},\frac{\delta}{\sqrt{\varepsilon}}\right)$
-where we consider $\delta\leqslant\frac{1}{4\left|\left|\kappa\right|\right|_{\infty}}$.
-Lemma 4.2 (Double scaling operator).
-The family $\left(u^{\varepsilon}_{t}\right)_{\varepsilon>0}$ satisfies on ×$\mathbb{R}\times\mathrm{I}_{\sqrt{\varepsilon}}^{\delta}$,
-$\begin{split}\left(\sqrt{\varepsilon}\mathrm{D}_{t}+y\left|\nabla\mathrm{m}(\gamma(s))\right|\sigma_{1}-\mathrm{D}_{y}\sigma_{2}+\frac{\sqrt{\varepsilon}}{1-\sqrt{\varepsilon}y\kappa(s)}\mathrm{D}_{s}\sigma_{3}-\varepsilon y\frac{\mathrm{i}\kappa^{\prime}(s)}{2(1-\sqrt{\varepsilon}y\kappa(s))^{2}}\sigma_{3}\right)u_{t}^{\varepsilon}=-\sqrt{\varepsilon}\mathfrak{R}^{\varepsilon}_{2}u_{t}^{\varepsilon},\end{split}$ (33)
-where for all $(s,y)\in\mathbb{R}\times\mathrm{I}_{\sqrt{\varepsilon}}^{\delta},$
-$\begin{split}\big{(}\mathfrak{R}_{2}^{\varepsilon}u_{t}^{\varepsilon}\big{)}(s,y)\coloneqq&y^{2}\int_{0}^{1}\nabla^{2}\mathrm{m}\Big{(}\Phi(s,\tau\sqrt{\varepsilon}y)\Big{)}\cdot\Big{(}\nu(s),\nu(s)\Big{)}\mathrm{d}\tau\sigma_{1}u_{t}^{\varepsilon}(s,y).\end{split}$
-Proof of Lemma 4.2.
-On the one hand, we rewrite the equation (12) with the definition of $\left(u^{\varepsilon}_{t}\right)_{\varepsilon>0}$ on ×$\mathbb{R}\times\mathrm{I}_{\sqrt{\varepsilon}}^{\delta}$,
-$\left[\varepsilon\mathrm{D}_{t}+\mathrm{m}\left(\Phi(s,\sqrt{\varepsilon}y)\right)\sigma_{1}-\sqrt{\varepsilon}\mathrm{D}_{y}\sigma_{2}+\frac{\varepsilon}{1-\sqrt{\varepsilon}y\kappa(s)}\mathrm{D}_{s}\sigma_{3}-\frac{\mathrm{i}\varepsilon^{3/2}y\kappa^{\prime}(s)}{2(1-\sqrt{\varepsilon}y\kappa(s))^{2}}\sigma_{3}\right]u^{\varepsilon}_{t}=0.$
-On the other hand, for all $(s,y)\in\mathbb{R}\times\mathrm{I}_{\sqrt{\varepsilon}}^{\delta}$,
-$\mathrm{m}\left(\Phi(s,\sqrt{\varepsilon}y)\right)=$ $\ \underbrace{\mathrm{m}(\gamma(s))}_{=0}+\sqrt{\varepsilon}y\underbrace{\nabla\mathrm{m}(\gamma(s))\cdot\nu(s)}_{=|\nabla\mathrm{m}(\gamma(s))|}$
-$+\ \varepsilon y^{2}\int_{0}^{1}(1-\tau)\nabla^{2}\mathrm{m}\Big{(}\Phi(s,\tau\sqrt{\varepsilon}y)\Big{)}\cdot\Big{(}\nu(s),\nu(s)\Big{)}\mathrm{d}\tau.$
-The operator $\sqrt{\varepsilon}\mathfrak{R}^{\varepsilon}_{2}$ will play a role of rest in the determination of $\Gamma\mathrm{d}\gamma$, the two-scale semiclassical measure at finite distance.
-For now, we do some remarks about the two-scale operator.
-Indeed, we remark that $\sqrt{\varepsilon}$ is in front of $\mathrm{D}_{s}$ but there is no $\sqrt{\varepsilon}$ in front of $\mathrm{D}_{y}$.
-It fits with the two-scale operators that we have introduced in Section 3.2.
-It is a semiclassical operator for the scale $\sqrt{\varepsilon}$ in the variable $s$, without scale in $y$.
-Because of the second scale of the problem, we will consider the symbol of the principal term of equation (33) according only to the first variable $s$, keeping the operator character over the variable $y$.
-Let us define, for all $(s,\sigma)\in\mathbb{R}^{2}$,
-$\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)\coloneqq\begin{pmatrix}\sigma&y|\nabla\mathrm{m}(\gamma(s))|+\mathrm{i}\mathrm{D}_{y}\\ y|\nabla\mathrm{m}(\gamma(s))|-\mathrm{i}\mathrm{D}_{y}&-\sigma\end{pmatrix},$ (34)
-the symbol of an operator acting on $\mathrm{L}^{2}\left(\mathbb{R}_{s},\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)\right)$ as in the end of the Section 3.2.1 in (22).
-From a spectral analysis (Appendix C), the eigenvalues of the operator $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$ were defined in (5) with the associated normalized eigenvectors (28) and the associated eigenprojectors (29).
-We also define the symbol of $\mathfrak{R}^{\varepsilon}_{2}$ for the two-scale quantization (22) as the operator acting on $\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)$ such that for all $(s,\sigma)\in\mathbb{R}^{2}$,
-$\mathfrak{r}^{\varepsilon}_{2}(s,\sigma)\coloneqq y^{2}\int_{0}^{1}\nabla^{2}\mathrm{m}\Big{(}\Phi(s,\tau\sqrt{\varepsilon}y)\Big{)}\cdot\Big{(}\nu(s),\nu(s)\Big{)}\mathrm{d}\tau\sigma_{1}.$
-Moreover, we define the operator acting on $\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)$ such that for all $(s,\sigma)\in\mathbb{R}^{2}$,
-$\mathfrak{r}^{\varepsilon}_{3}(s,\sigma)\coloneqq\frac{y\kappa(s)\sigma}{1-\sqrt{\varepsilon}y\kappa(s)}\sigma_{3}.$
-We define $\mathfrak{r}^{\varepsilon}\coloneqq\mathfrak{r}^{\varepsilon}_{2}+\mathfrak{r}^{\varepsilon}_{3}$. With the notation (22), the equation (33) writes
-$\sqrt{\varepsilon}\mathrm{D}_{t}u^{\varepsilon}_{t}+\mathrm{Op}_{\sqrt{\varepsilon},1}\Big{(}\mathrm{T}_{\mathrm{E}}^{\rm ds}+\sqrt{\varepsilon}\mathfrak{r}^{\varepsilon}\Big{)}u^{\varepsilon}_{t}=0.$
-4.2 The two-scale semiclassical measures in normal geodesic coordinates
-In this section we will describe the evolution of the two-scale semiclassical measures defined in Theorem 3.7 for $\left(\varphi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ solution to equation (12).
-In fact, in Section 4.2.1, we describe the evolution of the measure at finite distance $\Gamma^{t}\mathrm{d}\gamma^{t}$ and in Section 4.2.2, we describe the evolution of the measure at infinity $M_{\infty}^{t}\mathrm{d}\nu_{\infty}^{t}$.
-In Section 4.3, we prove Theorems 3.8 and 3.10.
-In Section 4.4, we prove Theorem 1.4.
-4.2.1 Two-scale Wigner measure at finite distance
-We prove here the part one of Theorem 3.8. We recall the following lemma of [16].
-Lemma 4.3 (Poisson bracket and projector properties).
-Let $\left(\Pi_{n}\right)_{n\in\mathbb{N}}$ be a family of orthogonal projectors such that for all $(n,k)\in\mathbb{N}^{2},\Pi_{n}\Pi_{k}=\delta_{k,n}$. Then for all $(j,k,\ell)\in\mathbb{N}^{3}$
-$\Pi_{k}\Big{(}\left\{\Pi_{j},\Pi_{\ell}\right\}-\left\{\Pi_{\ell},\Pi_{j}\right\}\Big{)}\Pi_{k}=0$
-Proof of part one of Theorem 3.8.
-Assume Assumption 1.2 and 1.3.
-For convenience, we assume that the sequence $\left(\varepsilon\right)_{\varepsilon>0}$ realizes the measure $\Gamma^{t}\mathrm{d}\gamma^{t}$.
-Let $a\in\mathfrak{A}$ and $R,\delta$ as defined in Section 3.2.
-We consider the part (19) of $a$ denoted $a_{\varepsilon,R}$.
-Let us consider $b(s,\sigma)\coloneqq a(s,0,0,0,y,\sigma,\mathrm{D}_{y})\chi\left(\frac{|(y,\sigma,\mathrm{D}_{y})|}{R}\right)$, the operator acting on $\mathcal{H}\coloneqq\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)$ and obtained by Weyl quantization.
-The link between the second scale quantization of $a_{\varepsilon,R}$ (13) and the two-scale quantization of $b$ (22) was established in (23).
-By considering $\left(u^{\varepsilon}_{t}\right)_{\varepsilon>0}$ solution to equation (33) as a family of $\mathrm{L}^{2}\left(\mathbb{R}_{s},\mathcal{H}\right)$, we have
-$\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b)u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle=$ $\ \frac{1}{\mathrm{i}\sqrt{\varepsilon}}\left\langle\Big{[}\mathrm{Op}_{\sqrt{\varepsilon},1}(b),\mathrm{Op}_{\sqrt{\varepsilon},1}\Big{(}\mathrm{T}_{\mathrm{E}}^{\rm ds}\Big{)}\Big{]}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle$
-$+\frac{1}{\mathrm{i}}\left\langle\Big{[}\mathrm{Op}_{\sqrt{\varepsilon},1}(b),\mathrm{Op}_{\sqrt{\varepsilon},1}\big{(}\mathfrak{r}^{\varepsilon}\big{)}\Big{]}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle$
-Because of (9), we have
-$\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b)u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle=$ $\ \frac{1}{\mathrm{i}\sqrt{\varepsilon}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big{(}\left[b,\mathrm{T}_{\mathrm{E}}^{\rm ds}\right]\Big{)}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle$
-$+\frac{1}{\mathrm{i}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\left(\frac{1}{2\mathrm{i}}\left\{b,\mathrm{T}_{\mathrm{E}}^{\rm ds}\right\}-\frac{1}{2\mathrm{i}}\left\{\mathrm{T}_{\mathrm{E}}^{\rm ds},b\right\}+\left[b,\mathfrak{r}^{\varepsilon}\right]\right)u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle$
-$+\mathcal{O}_{R}\left(\sqrt{\varepsilon}\right).$
-  • $\bullet$
-    Step 1 : Structure of $\Gamma^{t}$
-Let $\Xi\in\mathscr{C}^{\infty}_{c}(\mathbb{R})$, then
-$\int_{\mathbb{R}}\Xi(t)\mathrm{i}\sqrt{\varepsilon}\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b)u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\mathrm{d}t=$ $\int_{\mathbb{R}}\Xi(t)\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big{(}\left[b,\mathrm{T}_{\mathrm{E}}^{\rm ds}\right]\Big{)}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\mathrm{d}t$
-$+\frac{\sqrt{\varepsilon}}{2\mathrm{i}}\int_{\mathbb{R}}\Xi(t)\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big{(}\left\{b,\mathrm{T}_{\mathrm{E}}^{\rm ds}\right\}-\left\{\mathrm{T}_{\mathrm{E}}^{\rm ds},b\right\}\Big{)}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\mathrm{d}t$
-$+\sqrt{\varepsilon}\int_{\mathbb{R}}\Xi(t)\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big{(}\left[b,\mathfrak{r}^{\varepsilon}\right]\Big{)}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\mathrm{d}t+\mathcal{O}_{R}\left(\varepsilon\right)$
-$=$ $-\mathrm{i}\sqrt{\varepsilon}\int_{\mathbb{R}}\Xi^{\prime}(t)\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b)u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\mathrm{d}t+\mathcal{O}_{R}\left(\varepsilon\right).$
-But $b$ is compactly supported in $\left\{\left|(y,\sigma)\right|\leqslant R\right\}$ and for all $(s,\sigma,y)\in\mathbb{R}\times\left(-R,R\right)^{2}$,
-$y\underbrace{\frac{\kappa(s)}{1-\sqrt{\varepsilon}y\kappa(s)}}_{\in\mathrm{L}^{\infty}\big{(}\mathbb{R}\times\left(-R,R\right)\big{)}}\sigma=\mathcal{O}\left(R^{2}\right),\quad\quad y^{2}\underbrace{\int_{0}^{1}\nabla^{2}\mathrm{m}\Big{(}\Phi(s,\tau\sqrt{\varepsilon}y)\Big{)}\cdot\Big{(}\nu(s),\nu(s)\Big{)}\mathrm{d}\tau}_{\in\mathrm{L}^{\infty}\big{(}\mathbb{R}\times\left(-R,R\right)\big{)}}=\mathcal{O}\left(R^{2}\right).$
-Then
-$\sqrt{\varepsilon}\int_{\mathbb{R}}\Xi(t)\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big{(}[b,\mathfrak{r}^{\varepsilon}]\Big{)}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\mathrm{d}t=\mathcal{O}_{R}\left(\sqrt{\varepsilon}\right).$
-By passing to the limit, we have, for all $\Xi\in\mathscr{C}^{\infty}_{c}(\mathbb{R})$ and for all $a\in\mathfrak{A}$,
-00 0 $=\int_{\mathbb{R}^{3}}\Xi(t)\mathrm{Tr}_{\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)}\Bigg{(}\Big{[}a(s,0,0,y,\sigma,\mathrm{D}_{y}),\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)\Big{]}\Gamma^{t}(s,\sigma)\Bigg{)}\mathrm{d}\gamma^{t}(s,\sigma)\mathrm{d}t$
-$=\int_{\mathbb{R}^{3}}\Xi(t)\mathrm{Tr}\Bigg{(}a(s,0,0,y,\sigma,\mathrm{D}_{y})\Big{[}\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma),\Gamma^{t}(s,\sigma)\Big{]}\Bigg{)}\mathrm{d}\gamma^{t}(s,\sigma)\mathrm{d}t.$
-So, $t$ a.e. $\mathrm{d}\gamma^{t}$ a.e. $\Big{[}\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma),\Gamma^{t}(s,\sigma)\Big{]}=0$, which implies (31).
-  • $\bullet$
-    Step 2 : Equation of $\left(\mathrm{d}\gamma_{n}^{t}\right)_{n\in\mathbb{Z}}$
-By Step $1$, for all $(k,\ell)\in\mathbb{Z}^{2}$, with $k\neq\ell$ and for all $\beta$ function from $\mathbb{R}_{s}\times\mathbb{R}_{\sigma}$ to $\mathcal{L}\left(\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)\right)$ such that $\Pi_{k}\beta\Pi_{\ell}\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathcal{L}\left(\mathcal{H}\right)\right)$, we have
-$\limsup_{\delta\to 0}\limsup_{R\to+\infty}\lim_{\varepsilon\to 0}\int_{\mathbb{R}}\Xi(t)\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big{(}\Pi_{k}\beta\Pi_{\ell}\Big{)}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\mathrm{d}t=0.$
-Then we can consider symbol of the form $\Pi_{\ell}\beta\Pi_{\ell}$. Since the projector $\Pi_{\ell}$ has rank one, we have
-$\Pi_{\ell}\beta\Pi_{\ell}=\mathrm{Tr}\Big{(}\Pi_{\ell}\beta\Pi_{\ell}\Big{)}\Pi_{\ell},$
-where $\mathrm{Tr}_{\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)}\big{(}\Pi_{\ell}\beta\Pi_{\ell}\big{)}\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathbb{C}\right)$.
-Therefore we will consider symbols of the form $\beta\Pi_{\ell}$ with $\beta\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathbb{C}\right)$.
-Let $n\in\mathbb{Z}$ and $\beta\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathbb{C}\right)$.
-Let us define for all $(s,\sigma)\in\mathbb{R}^{2}$,
-$b_{n}(s,\sigma)\coloneqq\beta(s,\sigma)\Pi_{n}(s,\sigma).$
-The form of $b_{n}$ implies $[b_{n},\mathrm{T}_{\mathrm{E}}^{\rm ds}]=0$. Then
-$\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b_{n})u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle=$ $\ \frac{1}{2}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big{(}\{\mathrm{T}_{\mathrm{E}}^{\rm ds},b_{n}\}-\{b_{n},\mathrm{T}_{\mathrm{E}}^{\rm ds}\}\Big{)}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle$
-$+\frac{1}{\mathrm{i}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big{(}[b_{n},\mathfrak{r}^{\varepsilon}]\Big{)}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle+\mathcal{O}_{R}\left(\sqrt{\varepsilon}\right).$
-We first show that the term with $\mathfrak{r}^{\varepsilon}$ will vanish as $\varepsilon$ goes to 00. By definition of $\Gamma^{t}\mathrm{d}\gamma^{t}$, we have
-$\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\big{(}[b_{n},\mathfrak{r}^{\varepsilon}]\big{)}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle$ $\underset{\varepsilon\to 0}{\longrightarrow}\int_{\mathbb{R}^{2}}\mathrm{Tr}\Bigg{(}\Big{[}a(s,0,0,y,\sigma,D_{y}),\mathfrak{r}^{0}(s,\sigma)\Big{]}\Gamma^{t}(s,\sigma)\Bigg{)}\mathrm{d}\gamma^{t}(s,\sigma)$
-$=\sum_{m\in\mathbb{Z}}\int_{\mathbb{R}^{2}}\mathrm{Tr}\Bigg{(}\Big{[}\beta(s,\sigma)\Pi_{n}(s,\sigma),\mathfrak{r}^{0}(s,\sigma)\Big{]}\Pi_{m}(s,\sigma)\Bigg{)}\mathrm{d}\gamma^{t}_{m}(s,\sigma),$
-$=\sum_{m\in\mathbb{Z}}\int_{\mathbb{R}^{2}}\beta(s,\sigma)\mathrm{Tr}\Bigg{(}\Big{[}\Pi_{n}(s,\sigma),\mathfrak{r}^{0}(s,\sigma)\Big{]}\Pi_{m}(s,\sigma)\Bigg{)}\mathrm{d}\gamma^{t}_{m}(s,\sigma),$
-where the second equality holds because of Step $1$ and the chosen shape of $a$ and the last equality holds because $\beta$ is scalar and $\mathfrak{r}^{0}$ is the value at zero of $\mathfrak{r}^{\varepsilon}$ given by for all $(s,\sigma)\in\mathbb{R}^{2}$,
-$\mathfrak{r}^{0}(s,\sigma)=\mathfrak{r}^{0}_{2}(s,\sigma)+\mathfrak{r}^{0}_{3}(s,\sigma)=y^{2}\nabla^{2}\mathrm{m}\Big{(}\Phi(s,0)\Big{)}\cdot\Big{(}\nu(s),\nu(s)\Big{)}\sigma_{1}+y\kappa(s)\sigma\sigma_{3}.$
-Then, we recall, for all $m\in\mathbb{Z}$,
-$\mathrm{Tr}\Big{(}\big{[}\Pi_{n},\mathfrak{r}^{0}\big{]}\Pi_{m}\Big{)}=\mathrm{Tr}\Big{(}\Pi_{n}\mathfrak{r}^{0}\Pi_{m}\Big{)}-\mathrm{Tr}\Big{(}\mathfrak{r}^{0}\Pi_{n}\Pi_{m}\Big{)}=\mathrm{Tr}\Big{(}\mathfrak{r}^{0}\Pi_{m}\Pi_{n}\Big{)}-\mathrm{Tr}\Big{(}\mathfrak{r}^{0}\Pi_{n}\Pi_{m}\Big{)}\\ =\delta_{n=m}\left(\mathrm{Tr}\Big{(}\mathfrak{r}^{0}\Pi_{n}\Big{)}-\mathrm{Tr}\Big{(}\mathfrak{r}^{0}\Pi_{n}\Big{)}\right)=0.\$
-Let $k\in\mathbb{Z}$, let us compute $\Pi_{k}\{\mathrm{T}_{\mathrm{E}}^{\rm ds},b_{n}\}\Pi_{k}$,
-$\Pi_{k}\{\mathrm{T}_{\mathrm{E}}^{\rm ds},b_{n}\}\Pi_{k}=$ $\ \Pi_{k}\left\{\sum_{m\in\mathbb{Z}}\lambda_{m}\Pi_{m}\,,\,\beta\Pi_{n}\right\}\Pi_{k}$ (35)
-$=$ $\ \Pi_{k}\sum_{m\in\mathbb{Z}}\Bigg{(}\Pi_{m}\left\{\lambda_{m},\beta\right\}\Pi_{n}+\lambda_{m}\left\{\Pi_{m},\beta\right\}\Pi_{n}+\lambda_{m}\beta\left\{\Pi_{m},\Pi_{n}\right\}\Bigg{)}\Pi_{k}$
-$\ +\Pi_{k}\sum_{m\in\mathbb{Z}}\Big{(}\beta\Pi_{m}\left\{\lambda_{m},\Pi_{n}\right\}\Big{)}\Pi_{k}.$
-Let $m\in\mathbb{Z}$,
-$\Pi_{k}\Pi_{m}\left\{\lambda_{m},\beta\right\}\Pi_{n}\Pi_{k}$ $=\mathds{1}_{k=m=n}\left\{\lambda_{n},\beta\right\}\Pi_{n},$
-[…]
-$\Pi_{k}\lambda_{m}\beta\left\{\Pi_{m},\Pi_{n}\right\}\Pi_{k}$ $=\lambda_{m}\beta\Pi_{k}\left\{\Pi_{m},\Pi_{n}\right\}\Pi_{k},$
-[…]
-where the two vanishing equations hold because $\lambda_{n}$ and $\beta$ are scalar and $\Pi_{k}\partial\Pi_{m}\Pi_{k}$ is always equal to zero. Moreover, according to Lemma 4.3,
-$\lambda_{m}\beta\Pi_{k}\Big{(}\left\{\Pi_{m},\Pi_{n}\right\}-\left\{\Pi_{n},\Pi_{m}\right\}\Big{)}\Pi_{k}=0\,,\,\text{for all}\ m\in\mathbb{Z}.$
-Therefore,
-$\{\mathrm{T}_{\mathrm{E}}^{\rm ds},b_{n}\}-\{b_{n},\mathrm{T}_{\mathrm{E}}^{\rm ds}\}=\Big{(}\left\{\lambda_{n},\beta\right\}-\left\{\beta,\lambda_{n}\right\}\Big{)}\Pi_{n}.$
-Then,
-$\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b_{n})u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle=\frac{1}{2}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big{(}\Big{(}\left\{\lambda_{n},\beta\right\}-\left\{\beta,\lambda_{n}\right\}\Big{)}\Pi_{n}\Big{)}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\ +\ \mathcal{O}_{R}\left(\sqrt{\varepsilon}\right).$
-By passing to the limit,
-$\int_{\mathbb{R}}\int_{\mathbb{R}^{2}}\Xi^{\prime}(t)\mathrm{Tr}\Big{(}a^{W}(s,0,0,0,y,\sigma,\mathrm{D}_{y})\Gamma^{t}(s,\sigma)\Big{)}\mathrm{d}\gamma^{t}(s,\sigma)\mathrm{d}t\\ =\int_{\mathbb{R}}\int_{\mathbb{R}^{2}}\Xi(t)\mathrm{Tr}\Big{(}\left\{\lambda_{n},\beta\right\}\Pi_{n}(s,0,0,0,y,\sigma,\mathrm{D}_{y})\Gamma^{t}(s,\sigma)\Big{)}\mathrm{d}\gamma^{t}(s,\sigma)\mathrm{d}t\\ =\int_{\mathbb{R}}\int_{\mathbb{R}^{2}}\Xi(t)\mathrm{Tr}\Big{(}\left\{\lambda_{n},\beta\right\}\Pi_{n}(s,0,0,0,y,\sigma,\mathrm{D}_{y})\Big{)}\mathrm{d}\gamma^{t}_{n}(s,\sigma)\mathrm{d}t\\ =\int_{\mathbb{R}}\int_{\mathbb{R}^{2}}\Xi(t)\left\{\lambda_{n},\beta\right\}(s,\sigma)\mathrm{d}\gamma^{t}_{n}(s,\sigma)\mathrm{d}t.\\ =-\int_{\mathbb{R}}\int_{\mathbb{R}^{2}}\Xi(t)\beta(s,\sigma)\left\{\lambda_{n},\mathrm{d}\gamma^{t}_{n}\right\}(s,\sigma)\mathrm{d}t.\$
-Then $\mathrm{d}\gamma_{n}^{t}$ satisfies the equation (6) and is notably continuous on $\mathbb{R}$.
-It remains to prove that the value at zero of $\mathrm{d}\gamma_{n}^{t}$ coincides with the measure of the initial condition $\mathrm{d}\gamma_{n,0}$.
-Let $T\in\mathbb{R}$, on the one hand,
-$\int_{0}^{T}\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b_{n})u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\mathrm{d}t=\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b_{n})u^{\varepsilon}_{T},u^{\varepsilon}_{T}\right\rangle-\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b_{n})u^{\varepsilon}_{0},u^{\varepsilon}_{0}\right\rangle.$
-On the other hand,
-$\left|\int_{0}^{T}\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}(b_{n})u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\mathrm{d}t\right|$ $\leqslant\frac{1}{2}\int_{0}^{T}\left|\left\langle\mathrm{Op}_{\sqrt{\varepsilon},1}\Big{(}\left\{\lambda_{n},\beta\right\}\Pi_{n}\Big{)}u^{\varepsilon}_{t},u^{\varepsilon}_{t}\right\rangle\right|\mathrm{d}t\ +\ \mathcal{O}_{T,R}\left(\sqrt{\varepsilon}\right),$
-$\leqslant\frac{C\left(\left\{\lambda_{n},\beta\right\}\Pi_{n}\right)T}{2}\left\lVert u^{\varepsilon}_{0}\right\rVert_{\mathrm{L}^{2}}^{2}\ +\ \mathcal{O}_{T,R}\left(\sqrt{\varepsilon}\right).$
-where the last inequality holds because of the Calderón-Vaillancourt Theorem (7) and the preservation of the $\mathrm{L}^{2}$-norm of the solution. Then, by passing to the limit, we have
-$\begin{split}\limsup_{T\to 0}\limsup_{\delta\to 0}\limsup_{R\to+\infty}\lim_{\varepsilon\to 0}\mathcal{I}^{\varepsilon}_{\varphi_{t}^{\varepsilon}}\left(a_{\varepsilon,R}\right)&=\int_{\mathbb{R}^{2}}\mathrm{Tr}\Big{(}a^{W}(s,0,0,0,y,\sigma,\mathrm{D}_{y})\Pi_{n}(s,\sigma)\Big{)}\mathrm{d}\gamma_{n}^{0}(s,\sigma),\\ &=\limsup_{T\to 0}\limsup_{\delta\to 0}\limsup_{R\to+\infty}\lim_{\varepsilon\to 0}\mathcal{I}^{\varepsilon}_{\varphi_{0}^{\varepsilon}}\left(a_{\varepsilon,R}\right).\end{split}$
-Therefore, for all $n\in\mathbb{Z},\mathrm{d}\gamma_{n}^{t}$ is continuous and the value at zero coincides with the initial condition.
-Moreover, the equality $\Gamma^{t}\mathrm{d}\gamma^{t}=\sum_{n\in\mathbb{Z}}\Pi_{n}\mathrm{d}\gamma_{n}^{t}$ holds as trace class operator so, thanks to Lebesgue dominated convergence Theorem, $\Gamma^{t}\mathrm{d}\gamma^{t}$ is continuous on $\mathbb{R}$.
-∎
-Remark 4.4 (Zero value of the two-scale semiclassical measure at finite distance).
-We can construct the initial data with inspiration of the proof of the Lemma B.1 where we construct semiclassical measures for families valued in a separable Hilbert space.
-Let $\left(f^{\varepsilon}\right)_{\varepsilon>0}$ be a bounded family of $\mathrm{L}^{2}\left(\mathbb{R}_{s},\mathbb{C}\right)$ with the knowledge of $\mathrm{d}\rho$, the semiclassical measure of $\left(f^{\varepsilon}\right)_{\varepsilon>0}$ at the scale $\sqrt{\varepsilon}$.
-The function $(s,\sigma)\mapsto g_{1}^{s,\sigma}$ is smooth, bounded and with all its derivatives bounded uniformly in $y$, therefore we can consider its Weyl $\sqrt{\varepsilon}$-quantization according to the variable $s$.
-Let us consider initial data $\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}$ such that for all $(s,y)\in\mathbb{R}\times\mathrm{I}_{1/2}$,
-$u_{0}^{\varepsilon}(s,y)=g_{0}^{s}(y)f(s)+\Big{(}\mathrm{Op}_{\sqrt{\varepsilon}}^{(s)}\big{(}g_{1}^{s,\sigma}(y)\big{)}f\Big{)}(s),$
-where the Weyl $\sqrt{\varepsilon}$-quantization of $g_{1}^{s,\sigma}$ is taken only in the $s$ variable.
-Then the two-scale semiclassical measure at finite distance of this initial data is for all $(s,\sigma)\in\mathbb{R}^{2}$,
-$\frac{1}{2\pi}\Big{(}\Pi_{0}(s)+\Pi_{1}(s,\sigma)+\Pi_{1,0}(s)+\Pi_{0,1}(s,\sigma)\Big{)}\mathrm{d}\rho(s,\sigma),$
-where for all $(k,\ell)\in\left\{0,1\right\}^{2}$ and for all $(s,\sigma)\in\mathbb{R}^{2},\Pi_{k,\ell}(s,\sigma)=g^{s,\sigma}_{k}\otimes g^{s,\sigma}_{\ell}$.
-Notably, it does not commute with $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$.
-4.2.2 Two-scale Wigner measure at infinity
-We prove here the part two of Theorem 3.8.
-We consider the symbol of $\mathrm{H}^{\mathrm{E}}_{\varepsilon}$ according to the second scale quantization (13), given by for all $(s,y)\in\mathbb{R}^{2}$ and $z\coloneqq(z_{y},z_{\sigma},z_{\eta})\in\mathbb{R}^{3}$,
-$\mathrm{T}^{\mathrm{E}}(s,y,z)\coloneqq z_{y}\int_{0}^{1}\nabla\mathrm{m}\Big{(}\Phi(s,\tau y)\Big{)}\cdot\nu(s)\mathrm{d}\tau\sigma_{1}-z_{\eta}\sigma_{2}+\frac{z_{\sigma}}{1-y\kappa(s)}\sigma_{3}$
-which has two eigenvalues :
-$\lambda_{\pm}^{\mathrm{E}}(s,y,z)\coloneqq\pm\sqrt{\frac{1}{(1-y\kappa(s))^{2}}z_{\sigma}^{2}+z_{y}^{2}\left|\int_{0}^{1}\nabla\mathrm{m}\Big{(}\Phi(s,\tau y)\Big{)}\cdot\nu(s)\mathrm{d}\tau\right|^{2}+z_{\eta}^{2}}$
-which are homogeneous functions of order 1 in $z$. The eigenprojector $\Pi_{\pm}^{\mathrm{E}}(s,y,z)$ defined by
-$\Pi_{\pm}^{\mathrm{E}}(s,y,z)\coloneqq\frac{1}{2}\mathrm{Id}+\frac{1}{2\lambda_{\pm}^{\mathrm{E}}(s,y,z)}\mathrm{T}^{\mathrm{E}}(s,y,z)$
-are homogeneous functions of order 0 in $z$.
-Moreover, we denote the principal symbol of the operator $\mathrm{H}_{\varepsilon}^{\mathrm{E}}$ associated to the second scale quantization (13) by for all $(s,y)\in\mathbb{R}^{2}$ and for all $z\coloneqq(z_{y},z_{\sigma},z_{\eta})\in\mathbb{R}^{3}$,
-$\mathcal{T}^{\infty}_{\mathrm{E}}(s,z)\coloneqq\mathrm{T}^{\mathrm{E}}(s,0,z)=\begin{pmatrix}z_{\sigma}&z_{y}r(s)^{2}+iz_{\eta}\\ z_{y}r(s)^{2}-iz_{\eta}&-z_{\sigma}\end{pmatrix}.$
-Its eigenvalues are
-$\lambda_{\pm}^{\infty}(s,z)\coloneqq\pm\sqrt{z_{\sigma}^{2}+z_{y}^{2}r(s)^{4}+z_{\eta}^{2}},$
-and the associated eigenprojectors $\Pi_{\pm}^{\infty}(s,z)$ were defined in (30).
-Then, the equation (12) writes on $\mathbb{R}_{t}\times\mathbb{R}_{s}\times\mathrm{I}_{1/2}$
-$\sqrt{\varepsilon}\mathrm{D}_{t}\varphi^{\varepsilon}_{t}+\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\Big{(}\mathrm{T}^{\mathrm{E}}\Big{)}\varphi^{\varepsilon}_{t}=0.$
-Proof of part two of Theorem 3.8.
-Assume Assumption 1.2 and 1.3.
-For convenience, we assume that the sequence $\left(\varepsilon\right)_{\varepsilon>0}$ realizes the measure $M^{t}_{\infty}\mathrm{d}\nu^{t}_{\infty}$.
-Let $a\in\mathfrak{A}$ and $R,\delta$ as defined in Section 3.2.
-We consider the part (18) of $a$ divided by $\lambda^{\infty}_{+}$.
-The symbol $\mathrm{T}^{\mathrm{E}}$ is homogeneous in $z$ of order $1$ and the symbol $a_{\varepsilon,\delta}^{R}/\lambda^{\infty}_{+}$ is homogeneous in $z$ of order $-1$.
-Because of (18), we consider symbols on a subset of $\left\{z\in\mathbb{R}^{3}|\frac{R}{2}\leqslant\left|z\right|\leqslant\frac{2\delta}{\sqrt{\varepsilon}}\right\}$.
-Moreover, we have the following properties.
-  1. 1.
-    $\left\{\mathrm{T}^{\mathrm{E}},\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right\}_{z_{y},z_{\eta}}$ is homogeneous in $z$ of order $-2$.
-  2. 2.
-    $\left\{\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathrm{T}^{\mathrm{E}}\right\}_{s,z_{\sigma}}$ is homogeneous in $z$ of order $-1$.
-  3. 3.
-    $\left\{\mathrm{T}^{\mathrm{E}},\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right\}_{(s,y),(\sigma,\eta)}$ is homogeneous in $z$ of order 00.
-Then, according to standard semiclassical calculus,
-$\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\mathrm{T}^{\mathrm{E}},\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right\}_{z_{y},z_{\eta}}-\left\{\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathrm{T}^{\mathrm{E}}\right\}_{z_{y},z_{\eta}}\right)$ $=\mathcal{O}\left(\frac{1}{R^{2}}\right),$
-$\sqrt{\varepsilon}\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\mathrm{T}^{\mathrm{E}},\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right\}_{s,z_{\sigma}}-\left\{\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathrm{T}^{\mathrm{E}}\right\}_{s,z_{\sigma}}\right)$ $=\mathcal{O}_{R}\left(\sqrt{\varepsilon}\right),$
-$\varepsilon\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\mathrm{T}^{\mathrm{E}},\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right\}_{(s,y),(\sigma,\eta)}-\left\{\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathrm{T}^{\mathrm{E}}\right\}_{(s,y),(\sigma,\eta)}\right)$ $=\mathcal{O}_{R}\left(\varepsilon\right).$
-So
-$\left[\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right),\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\mathrm{T}^{\mathrm{E}}\right)\right]=$ $\ \mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left[\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathrm{T}^{\mathrm{E}}\right]\right)$
-$+\frac{1}{2\mathrm{i}}\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathrm{T}^{\mathrm{E}}\right\}_{z_{y},z_{\eta}}-\left\{\mathrm{T}^{\mathrm{E}},\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right\}_{z_{y},z_{\eta}}\right)+\mathcal{O}\left(\frac{1}{R^{4}}\right)$
-$+\frac{\sqrt{\varepsilon}}{2\mathrm{i}}\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathrm{T}^{\mathrm{E}}\right\}_{s,z_{\sigma}}-\left\{\mathrm{T}^{\mathrm{E}},\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right\}_{s,z_{\sigma}}\right)+\mathcal{O}_{R}\left(\varepsilon\right).$
-Considering $\left(\varphi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ solution to (12), we have
-$\mathrm{i}\sqrt{\varepsilon}\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right)\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle=\left\langle\left[\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right),\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\mathrm{T}^{\mathrm{E}}\right)\right]\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle,\\ =\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left[\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathrm{T}^{\mathrm{E}}\right]\right)\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle+\frac{\sqrt{\varepsilon}}{2\mathrm{i}}\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathrm{T}^{\mathrm{E}}\right\}_{s,z_{\sigma}}-\left\{\mathrm{T}^{\mathrm{E}},\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right\}_{s,z_{\sigma}}\right)\\ +\mathcal{O}_{R}\left(\varepsilon\right)+\mathcal{O}\left(\frac{1}{R^{2}}\right).\$
-  • $\bullet$
-    Step 1 : Structure of $M_{\infty}^{t}$
-Let $\Xi\in\mathscr{C}^{\infty}_{c}(\mathbb{R})$, then
-$\int_{\mathbb{R}}\Xi(t)\mathrm{i}\sqrt{\varepsilon}\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right)\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle\mathrm{d}t=\int_{\mathbb{R}}\Xi(t)\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left[\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathrm{T}^{\mathrm{E}}\right]\right)\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle\mathrm{d}t\\ +\frac{\sqrt{\varepsilon}}{2\mathrm{i}}\int_{\mathbb{R}}\Xi(t)\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathrm{T}^{\mathrm{E}}\right\}_{s,z_{\sigma}}-\left\{\mathrm{T}^{\mathrm{E}},\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right\}_{s,z_{\sigma}}\right)\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle\mathrm{d}t+\ \mathcal{O}_{R}\left(\varepsilon\right)+\mathcal{O}\left(\frac{1}{R^{2}}\right)\\ =-\mathrm{i}\sqrt{\varepsilon}\int_{\mathbb{R}}\Xi^{\prime}(t)\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}}\right)\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle\mathrm{d}t.\$
-But $\left[\frac{a_{\varepsilon,\delta}^{R}}{\lambda^{\infty}_{+}},\mathrm{T}^{\mathrm{E}}\right]=\left[a,\frac{\mathrm{T}^{\mathrm{E}}}{\lambda^{\infty}_{+}}\right]_{\varepsilon,\delta}^{R}$ and $\left[a,\frac{\mathrm{T}^{\mathrm{E}}}{\lambda^{\infty}_{+}}\right]\in\mathfrak{A}$, then by passing to the limit, we deduce that for all $\Xi\in\mathscr{C}^{\infty}_{c}(\mathbb{R})$ and for all $a\in\mathfrak{A}$,
-00 0 $=\int_{\mathbb{R}^{2}\times\mathbb{S}^{2}}\Xi(t)\mathrm{tr}\Bigg{(}\left[a,\frac{\mathrm{T}^{\mathrm{E}}}{\lambda^{\infty}_{+}}\right]_{\infty}(s,0,0,0,\omega)M_{\infty}^{t}(s,\omega)\Bigg{)}\mathrm{d}\nu_{\infty}^{t}(s,\omega)\mathrm{d}t$
-$=\int_{\mathbb{R}^{2}\times\mathbb{S}^{2}}\Xi(t)\mathrm{tr}\Bigg{(}\left[a_{\infty}(s,0,0,0,\omega),\frac{1}{\lambda^{\infty}_{+}(s,\omega)}\mathrm{T}^{\mathrm{E}}(s,0,\omega)\right]M_{\infty}^{t}(s,\omega)\Bigg{)}\mathrm{d}\nu_{\infty}^{t}(s,\omega)\mathrm{d}t$
-$=\int_{\mathbb{R}^{2}\times\mathbb{S}^{2}}\Xi(t)\frac{1}{\lambda^{\infty}_{+}(s,\omega)}\mathrm{tr}\Bigg{(}a_{\infty}(s,0,0,0,\omega)\Big{[}\mathcal{T}^{\infty}_{\mathrm{E}}(s,\omega),M_{\infty}^{t}(s,\omega)\Big{]}\Bigg{)}\mathrm{d}\nu_{\infty}^{t}(s,\omega)\mathrm{d}t.$
-So, for almost all $t\in\mathbb{R}$, we have $\mathrm{d}\nu_{\infty}^{t}\ \mathrm{a.e.}$ $\Big{[}\mathcal{T}^{\infty}_{\mathrm{E}}(s,\omega),M_{\infty}^{t}(s,\omega)\Big{]}=0$ which implies (32).
-  • $\bullet$
-    Step 2 : Equation of $\mathrm{d}\nu_{\infty}^{t}$.
-Let $\beta\in\mathfrak{A}$ and $a$ defined by for all $(s,y,\sigma,\eta,z)\in\mathbb{R}^{7}$,
-$a(s,y,\sigma,\eta,z)\coloneqq\lambda^{\infty}_{\pm}(s,z)\beta(s,y,\sigma,\eta,z)\Pi_{\pm}^{\mathrm{E}}(s,y,z).$
-The form of $a$ implies $\left[a,\mathrm{T}^{\mathrm{E}}\right]=0$. Then
-$\sqrt{\varepsilon}\frac{\mathrm{d}}{\mathrm{dt}}\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a_{\varepsilon,\delta}^{R}\right)\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle=$ $\ \frac{1}{2}\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\mathrm{T}^{\mathrm{E}},a_{\varepsilon,\delta}^{R}\right\}_{z_{y},z_{\eta}}-\left\{a_{\varepsilon,\delta}^{R},\mathrm{T}^{\mathrm{E}}\right\}_{z_{y},z_{\eta}}\right)\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle+\mathcal{O}\left(\frac{1}{R^{2}}\right)$
-$+\frac{\sqrt{\varepsilon}}{2}\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\mathrm{T}^{\mathrm{E}},a_{\varepsilon,\delta}^{R}\right\}_{s,z_{\sigma}}-\left\{a_{\varepsilon,\delta}^{R},\mathrm{T}^{\mathrm{E}}\right\}_{s,z_{\sigma}}\right)\varphi^{\varepsilon}_{t},\varphi^{\varepsilon}_{t}\right\rangle+\mathcal{O}_{R}\left(\varepsilon\right).$
-Because of the homogeneity in $z$, we have $\left\{\mathrm{T}^{\mathrm{E}},a\right\}_{z_{y},z_{\eta}}\in\mathfrak{A}$. Moreover,
-$\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left\{\mathrm{T}^{\mathrm{E}},a_{\varepsilon,\delta}^{R}\right\}_{z_{y},z_{\eta}}\right)=$ $\ \mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(\left(\left\{\mathrm{T}^{\mathrm{E}},a\right\}_{z_{y},z_{\eta}}\right)_{\varepsilon,\delta}^{R}\right)+\frac{1}{R}\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a\chi_{\frac{\delta}{\sqrt{\varepsilon}}}\left\{\mathrm{T}^{\mathrm{E}},\chi^{R}\right\}_{z_{y},z_{\eta}}\right)$
-$+\frac{\sqrt{\varepsilon}}{\delta}\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a\chi^{R}\left\{\mathrm{T}^{\mathrm{E}},\chi_{\frac{\delta}{\sqrt{\varepsilon}}}\right\}_{z_{y},z_{\eta}}\right).$
-The function $\partial\chi^{R}$ is compactly supported in the ring $\left\{z\in\mathbb{R}^{3}|\frac{R}{2}\leqslant\left|z\right|\leqslant R\right\}$ and the function $a\chi_{\frac{\delta}{\sqrt{\varepsilon}}}\left\{\mathrm{T}^{\mathrm{E}},\chi^{R}\right\}_{z_{y},z_{\eta}}$ belongs to $\mathfrak{A}$ so
-$\left(a\chi_{\frac{\delta}{\sqrt{\varepsilon}}}\left\{\mathrm{T}^{\mathrm{E}},\chi^{R}\right\}_{z_{y},z_{\eta}}\right)_{\infty}=0.$
-The operator $\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a\chi^{R}\left\{\mathrm{T}^{\mathrm{E}},\chi_{\frac{\delta}{\sqrt{\varepsilon}}}\right\}_{z_{y},z_{\eta}}\right)$ is uniformly bounded in $\varepsilon$ then
-$\lim_{\varepsilon\to 0}\frac{\sqrt{\varepsilon}}{\delta}\left\langle\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a\chi^{R}\left\{\mathrm{T}^{\mathrm{E}},\chi_{\frac{\delta}{\sqrt{\varepsilon}}}\right\}_{z_{y},z_{\eta}}\right)\varphi_{t}^{\varepsilon},\varphi_{t}^{\varepsilon}\right\rangle=0.$
-As in the computation of (35), we consider the quantity $\Pi_{+}^{\mathrm{E}}\left\{a,\mathrm{T}^{\mathrm{E}}\right\}_{z_{y},z_{\eta}}\Pi_{+}^{\mathrm{E}}$ then
-$\Pi_{+}^{\mathrm{E}}\left\{a,\mathrm{T}^{\mathrm{E}}\right\}_{z_{y},z_{\eta}}\Pi_{+}^{\mathrm{E}}=\lambda_{+}^{\infty}\left\{\beta,\lambda_{+}^{\infty}\right\}_{z_{y},z_{\eta}}\Pi_{+}^{\mathrm{E}}.$
-Then, by passing to the limit, we have
-00 0 $=\int_{\mathbb{R}^{2}\times\mathbb{S}^{2}}\Xi(t)\mathrm{tr}\Bigg{(}\left(\lambda_{+}^{\infty}\left\{\beta,\lambda_{+}^{\infty}\right\}_{z_{y},z_{\eta}}\Pi_{+}^{\mathrm{E}}\right)_{\infty}(s,0,0,0,\omega)M_{\infty}^{t}(s,\omega)\Bigg{)}\mathrm{d}\nu_{\infty}^{t}(s,\omega)\mathrm{d}t,$
-$=\int_{\mathbb{R}^{2}\times\mathbb{S}^{2}}\Xi(t)\mathrm{tr}\Bigg{(}\lambda_{+}^{\infty}\left\{\beta_{\infty},\lambda_{+}^{\infty}\right\}_{\omega_{y},\omega_{\eta}}(s,0,0,0,\omega)\Pi_{+}^{\infty}(s,\omega)\Bigg{)}\mathrm{d}\nu_{+}^{t}(s,\omega)\mathrm{d}t,$
-$=\int_{\mathbb{R}^{2}\times\mathbb{S}^{2}}\Xi(t)\lambda_{+}^{\infty}\left\{\beta_{\infty},\lambda_{+}^{\infty}\right\}_{\omega_{y},\omega_{\eta}}(s,0,0,0,\omega)\mathrm{d}\nu_{+}^{t}(s,\omega)\mathrm{d}t,$
-but for all $(s,\omega)\in\mathbb{R}\times\mathbb{S}^{2}$,
-$\lambda_{+}^{\infty}(s,\omega)\frac{\partial\lambda_{+}^{\infty}}{\partial z_{y}}(s,\omega)=\omega_{y}r(s)^{4},\quad\quad\quad\lambda_{+}^{\infty}(s,\omega)\frac{\partial\lambda_{+}^{\infty}}{\partial z_{\eta}}(s,\omega)=\omega_{\eta}.$
-Then
-$\int_{\mathbb{R}^{2}\times\mathbb{S}^{2}}\Xi(t)\Big{(}\vec{V}_{+}(s,\omega)\cdot\nabla_{\omega}\beta_{\infty}(s,\omega)\Big{)}\mathrm{d}\nu_{+}^{t}(s,\omega)\mathrm{d}t=0,$
-with
-$\vec{V}_{+}(s,\omega)=\begin{pmatrix}-\omega_{\eta}\\ 0\\ \omega_{y}r(s)^{4}\end{pmatrix}$
-and $\beta_{\infty}$ a function on the sphere with $\nabla_{\omega}\beta_{\infty}$ is a vector field over the sphere.
-Therefore we can rewrite $\vec{V}_{\pm}\cdot\nabla_{\omega}\beta=\vec{V}_{\pm}^{\infty}\cdot\nabla_{\omega}\beta$ where $V_{\pm}^{\infty}$ is the component of $\vec{V}_{\pm}$ which is tangential to the sphere.
-The vector $\vec{V}_{\pm}^{\infty}$ defines a vector field over the sphere and is given by for all $(s,\omega)\in\mathbb{R}\times\mathbb{S}^{2}$,
-$\vec{V}^{\infty}(s,\omega)\coloneqq\vec{V}_{+}(s,\omega)-\Big{(}\vec{V}_{+}(s,\omega)\cdot\omega\Big{)}\omega=\begin{pmatrix}\left(\left(1-r(s)^{4}\right)\omega_{y}^{2}-1\right)\omega_{\eta}\\ \left(1-r(s)^{4}\right)\omega_{y}\omega_{\eta}\omega_{\sigma}\\ \left(\left(1-r(s)^{4}\right)\omega_{\eta}^{2}+r(s)^{4}\right)\omega_{y}\end{pmatrix}.$
-Because of the equality $\lambda^{\infty}_{+}=-\lambda^{\infty}_{-}$, we have the same invariance for $\mathrm{d}\nu_{-}^{t}$. So, $t$ almost everywhere,
-×$\int_{\mathbb{R}\times\mathbb{S}^{2}}\Big{(}\vec{V}^{\infty}(s,\omega)\cdot\nabla_{\omega}\beta_{\infty}(s,\omega)\Big{)}\mathrm{d}\nu_{\infty}^{t}(s,\omega)=0.$
-So, for almost all $(t,s,\omega)\in\mathbb{R}^{2}\times\mathbb{S}^{2}$, we have $\mathrm{div}_{\omega}\Big{(}\vec{V}^{\infty}(s,\omega)\mathrm{d}\nu^{t}(s,\omega)\Big{)}=0$. ∎
-4.3 The Wigner measure
-Proof of Theorem 3.8.
-Assume Assumption 1.2 and 1.3.
-Let $\left(\varphi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ a uniformly bounded family in $\mathrm{L}^{\infty}\left(\mathbb{R},\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)$ be a solution to equation (12) with initial normalized condition $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$.
-According to Theorem 3.7, we can consider ${\left(M^{t}_{\infty}\mathrm{d}\nu^{t}_{\infty},\Gamma^{t}\mathrm{d}\gamma^{t}\right)}$ a pair of two-scale semiclassical measure of $\left(\varphi^{\varepsilon}_{t}\right)_{\varepsilon>0}$.
-Then, according to proof of Section 4.2.1 and Section 4.2.2, we conclude.
-∎
-Proof of Theorem 3.10.
-Assume Assumption 1.2 and 1.3.
-Let $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ a uniformly bounded family in $\mathrm{L}^{\infty}\left(\mathbb{R},\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)$ a solution to equation (1) with initial normalized condition $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$.
-On the one hand, according to Lemma 2.3, we know the evolution of any semiclassical measure of $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ outside the curve $\mathcal{C}$.
-On the other hand, according to Lemma 3.2, we consider the family $\left(\varphi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ uniformly bounded in $\mathrm{L}^{\infty}\left(\mathbb{R},\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)$ solution to equation (12) with initial normalized condition $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ a family of $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$.
-Then, by applying the Lemma 4.1, we have the link between any semiclassical measure of $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ over the curve $\mathcal{C}$ and the associated measure of $\left(\varphi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ on $\mathbb{R}_{s}$.
-Then, by applying Theorem 3.8, we have the evolution of the semiclassical measure over the curve $\mathcal{C}$.
-Finally, by applying Remark 3.5 and by calculating the shape of the two-scale semiclassical measure at finite distance according to expressions (5) and (28), we conclude. ∎
-4.4 Proof of Theorem 1.4 and Theorem 1.5
-Proof of Theorem 1.4.
-Assume Assumption 1.2 and 1.3.
-Let $\vec{f}\in\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$.
-Let ${\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}}$ be the solution to equation (1) with for all $\varepsilon\in(0,1]$ and for all $x\in\mathbb{R}^{2}$,
-$\psi_{0}^{\varepsilon}(x)=\frac{1}{\sqrt{\varepsilon}}\vec{f}\left(\frac{x-x_{0}}{\sqrt{\varepsilon}}\right),$
-then, applying Theorem 3.10, according to Remark 3.5 and Corollary 3.6, there exist ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ a vanishing sequence and ${\Gamma^{t}\mathrm{d}\gamma^{t}}$ a two-scale semiclassical measure at finite distance, such that for all $(a,\Xi)\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)\times\mathscr{C}_{c}^{\infty}(\mathbb{R},\mathbb{C})$,
-$\mathcal{I}^{\varepsilon_{k}}_{\psi}(\Xi,a)\underset{k\to+\infty}{\longrightarrow}\frac{1}{2}\sum_{n\in\mathbb{Z}}\int_{\mathbb{R}^{3}}\Xi(t)\mathrm{tr}\left(a\big{(}\gamma(s)\big{)}\begin{pmatrix}1&\frac{\mathrm{i}\sigma}{\lambda_{n}(s,\sigma)}e^{-\mathrm{i}\theta(s)}\\ \frac{-\mathrm{i}\sigma}{\lambda_{n}(s,\sigma)}e^{\mathrm{i}\theta(s)}&1\end{pmatrix}\mathrm{d}\gamma^{t}_{n}(s,\sigma)\right)\mathrm{d}t,$
-with for all $n\in\mathbb{Z}$, $\mathrm{d}\gamma_{n}^{t}$ satisfies the following equation
-$\left\{\begin{array}[]{rll}\partial_{t}\ \mathrm{d}\gamma_{n}^{t}&=&\left\{\mathrm{d}\gamma_{n}^{t},\lambda_{n}\right\}\\ \mathrm{d}\gamma_{n}^{t}|_{t=0}&=&\mathrm{d}\gamma_{n,0}\end{array}\right.,$
-with $\mathrm{d}\gamma_{n,0}=\mathrm{tr}\Big{(}\Pi_{n}\Gamma^{0}\Big{)}\mathrm{d}\gamma^{0}$ and $\Gamma^{0}\mathrm{d}\gamma^{0}$ is a two-scale semiclassical measure associated to the initial data.
-We conclude thanks to the conservation of $\varepsilon$-oscillation (Lemma 2.2). ∎
-Proof of Theorem 1.5.
-Assume Assumption 1.2 and 1.3.
-Let ${\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}}$ a uniformly bounded family of $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$, $\nu$ a non negative scalar Radon measure on $\mathbb{R}_{t}\times\mathbb{R}_{x}^{2}$ and ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ a vanishing sequence such that for all $(a,\Xi)\in\mathscr{C}_{c}^{\infty}\left(\Omega,\mathbb{C}\right)\times\mathscr{C}_{c}^{\infty}(\mathbb{R},\mathbb{C})$,
-$\int_{\mathbb{R}}\int_{\Omega}\Xi(t)a(x)\left|\psi_{t}^{\varepsilon_{k}}(x)\right|^{2}\mathrm{d}x\mathrm{d}t\underset{k\to+\infty}{\longrightarrow}\int_{\mathbb{R}}\int_{\Omega}\Xi(t)a(x)\mathrm{d}\nu(t,x),$
-where ${\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}}$ solves (1) with ${\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}}$ as initial condition.
-A consequence of Theorem 3.8 and Remark 3.5 is
-$\mathrm{d}\nu\mathds{1}_{\mathcal{C}}\geqslant\sum_{n\in\mathbb{Z}}\Phi^{*}\left(\mathrm{Tr}\Big{(}\Pi_{n}(s,\sigma)\Big{)}\mathrm{d}\gamma^{t}_{n}\right)\mathds{1}_{\mathcal{C}}\mathrm{d}t$
-where $\left(\mathrm{d}\gamma_{n}^{t}\right)_{n\in\mathbb{Z}}$ is defined by (6). We conclude with $\mathrm{Tr}\left(\Pi_{n}(s,\sigma)\right)=1$. ∎
-Appendix A Proof of compactness at infinity
+The $\varepsilon$-oscillation conservation property justifies the semiclassical approach to solutions to equation (1.1) for analyzing the limits of the quantities (1.4).
+Note that similar arguments show that there is also no loss of mass at infinity in configuration space in the sense that if $\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}$ is compact at infinity, then $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ is compact at infinity uniformly in time $t\in[-T,T]$, for some $T>0$.
 Lemma A.3 (Conservation of compactness at infinity).
 Let $T>0$.
-If $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ is compact at infinity then $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ solution to (1) is compact at infinity uniformly for $t\in[-T,T]$.
+If $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ is compact at infinity then $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ solution to (1.1) is compact at infinity uniformly for $t\in[-T,T]$.
 Proof of Lemma A.3.
@@
 $\underset{R\to+\infty}{\longrightarrow}0.$
-Appendix B Construction of semiclassical measures
-B.1 Semiclassical measures for families valued in a separable Hilbert space
-We will use the more general framework of [14] and consider families valued in a separable Hilbert space $\mathcal{H}$.
+As a consequence, if $\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}$ is normalized, $\varepsilon$-oscillating and compact at infinity, then for all $\chi\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},[0,1]\right)$, with $\chi\equiv 1$ for $\left|x\right|<1$, $\chi\equiv 0$ for $\left|x\right|>2$,
+$\lim_{\varepsilon\to 0}\left\langle\mathrm{Op}_{\varepsilon}\left(\chi\left(\frac{x}{R}\right)\chi\left(\frac{\varepsilon\mathrm{D}_{x}}{R}\right)\right)\psi^{\varepsilon}_{0},\psi^{\varepsilon}_{0}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\;\underset{R\to+\infty}{\longrightarrow}\;\lim_{\varepsilon\to 0}\left\lVert\psi_{0}^{\varepsilon}\right\rVert^{2}_{\mathrm{L}^{2}\left(R^{2},\mathbb{C}^{2}\right)}=1.$
+So for all semiclassical measure $\mathrm{d}\mu$ of $\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}$ normalized, $\varepsilon$-oscillating and compact at infinity,
+$\mathrm{tr}\int_{\mathbb{R}^{4}}\mathrm{d}\mu(x,\xi)=1.$
+A.2 Semiclassical measures for families valued in a separable Hilbert space
+We will use the more general framework of Gérard (1991) and consider families valued in a separable Hilbert space $\mathcal{H}$.
 More precisely, in this section, $\mathcal{H}=\mathrm{L}^{2}\left(\mathbb{R},\mathbb{C}^{2}\right)$.
 Note that in Section 1.2, $\mathcal{H}=\mathbb{C}^{2}$.
-Lemma B.1 (Operator valued measure).
-Let ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$ a bounded family in $\mathrm{L}^{2}\left(\mathbb{R},\mathcal{H}\right)$, then there exists ${(\varepsilon_{k})_{k>0}}$ a vanishing sequence and a positive operator valued measure $\Gamma\mathrm{d}\gamma$ such that for all $a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathcal{K}\left(\mathcal{H}\right)\right)$,
-$\left\langle\mathrm{Op}_{\varepsilon_{k}}(a)f^{\varepsilon_{k}},f^{\varepsilon_{k}}\right\rangle\underset{k\to+\infty}{\longrightarrow}\int_{\mathbb{R}^{2}}\mathrm{Tr}_{\mathcal{L}(\mathcal{H})}\Big{(}a(x,\xi)\Gamma(x,\xi)\Big{)}\mathrm{d}\gamma(x,\xi),$
+Lemma A.4 (Operator valued measure).
+Let ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$ a bounded family in $\mathrm{L}^{2}\left(\mathbb{R},\mathcal{H}\right)$, then there exists a vanishing sequence of positive numbers ${(\varepsilon_{k})_{k>0}}$ and a positive operator valued measure $M\mathrm{d}\nu$ such that for all $a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathcal{K}\left(\mathcal{H}\right)\right)$,
+$\left\langle\mathrm{Op}_{\varepsilon_{k}}(a)f^{\varepsilon_{k}},f^{\varepsilon_{k}}\right\rangle\underset{k\to+\infty}{\longrightarrow}\int_{\mathbb{R}^{2}}\mathrm{Tr}_{\mathcal{L}(\mathcal{H})}\Big(a(x,\xi)M(x,\xi)\Big)\mathrm{d}\nu(x,\xi),$
 with $\mathcal{K}\left(\mathcal{H}\right)$ the set of compact operators on $\mathcal{H}$.
-Proof of Lemma B.1.
+Proof of Lemma A.4.
 Let $\mathcal{H}$ be a separable Hilbert space, ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$ uniformly bounded in $\mathrm{L}^{2}\left(\mathbb{R},\mathcal{H}\right)$, $(h_{n})_{n\in\mathbb{N}}$ a Hermitian basis of $\mathcal{H}$ and $\left(\left(f^{\varepsilon}_{n}\right)_{\varepsilon>0}\right)_{n\in\mathbb{N}}$ families of $\mathrm{L}^{2}\left(\mathbb{R}\right)$ such that $x$ a.e.
@@
 Then the families $\left(f^{\varepsilon}_{n}\right)_{\varepsilon>0}$ satisfy, for some constant $C>0$,
-$\sum_{n\in\mathbb{N}}\left\lVert f^{\varepsilon}_{n}\right\rVert_{\mathrm{L}^{2}(\mathbb{R})}^{2}=\left\lVert f^{\varepsilon}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R},\mathcal{H}\right)}^{2}\leqslant C.$ (36)
+$\sum_{n\in\mathbb{N}}\left\lVert f^{\varepsilon}_{n}\right\rVert_{\mathrm{L}^{2}(\mathbb{R})}^{2}=\left\lVert f^{\varepsilon}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R},\mathcal{H}\right)}^{2}\leqslant C.$ (A.1)
 For all $n\in\mathbb{N}$, $\left(f^{\varepsilon}_{n}\right)_{\varepsilon>0}$ are bounded in $\mathrm{L}^{2}\left(\mathbb{R}\right)$.
@@
 Moreover, for all $(n,m)\in\mathbb{N}^{2}$, $\mathrm{d}\mu_{n,n}$ is a positive Radon measure and $\mathrm{d}\mu_{m,n}=\overline{\mathrm{d}\mu_{n,m}}$.
-First of all, we construct the operator valued measure $\Gamma\mathrm{d}\gamma$. By (36),
+First of all, we construct the operator valued measure $M\mathrm{d}\nu$. By (A.1),
 $\limsup_{k\to+\infty}\sum_{n\in\mathbb{N}}\left\lVert f^{\varepsilon_{k}}_{n}\right\rVert_{\mathrm{L}^{2}}^{2}<+\infty,$
-then ${\mathrm{d}\gamma\coloneqq\sum_{n\in\mathbb{N}}\mathrm{d}\mu_{n,n}}$ defines a positive Radon measure of finite mass.
-By the Radon-Nykodym Theorem, there exists a sequence of $\mathrm{d}\gamma$-measurable functions $\left(\Gamma_{n,m}\right)_{(n,m)\in\mathbb{N}^{2}}$ such that $\forall(n,m)\in\mathbb{N}^{2}$,
-$\mathrm{d}\mu_{n,m}=\Gamma_{n,m}\mathrm{d}\gamma$
-with $0\leqslant\Gamma_{n,n}\leqslant 1$ and $\Gamma_{n,m}=\overline{\Gamma_{m,n}}$ $\mathrm{d}\gamma$-almost everywhere.
-Let us define the operator valued function $\Gamma:(x,\xi)\mapsto\Gamma(x,\xi)$ such that, for all $(n,m)\in\mathbb{N}^{2}$, for all $(x,\xi)\in\mathbb{R}^{2}$,
-$\left\langle\Gamma(x,\xi)h_{n},h_{m}\right\rangle_{\mathcal{H}}=\Gamma_{n,m}(x,\xi).$
-Because of the positivity of $\left(\mathrm{d}\mu_{n,n}\right)_{n\in\mathbb{N}}$ and the adjoint relation of $\left(\mathrm{d}\mu_{n,m}\right)_{(n,m)\in\mathbb{N}^{2}}$, $\Gamma(x,\xi)$ is a positive self-adjoint operator and its trace satisfies
-$\int_{\mathbb{R}^{2}}\mathrm{Tr}_{\mathcal{L}(\mathcal{H})}\Big{(}\Gamma(x,\xi)\Big{)}\mathrm{d}\gamma(x,\xi)=\sum_{n\in\mathbb{N}}\int_{\mathbb{R}^{2}}\Gamma_{n,n}(x,\xi)\mathrm{d}\gamma(x,\xi)=\sum_{n\in\mathbb{N}}\mathrm{d}\gamma_{n,n}\left(\mathbb{R}^{2}\right)<+\infty.$
-Therefore, the pair $\left(\Gamma,\mathrm{d}\gamma\right)$ generates an operator measure $\Gamma\mathrm{d}\gamma$. We now have to prove that $\Gamma\mathrm{d}\gamma$ plays the expected role.
+then ${\nu\coloneqq\sum_{n\in\mathbb{N}}\mu_{n,n}}$ defines a positive Radon measure of finite mass.
+By the Radon-Nykodym Theorem, there exists a sequence of $\mathrm{d}\nu$-measurable functions $\left(M_{n,m}\right)_{(n,m)\in\mathbb{N}^{2}}$ such that $\forall(n,m)\in\mathbb{N}^{2}$,
+$\mathrm{d}\mu_{n,m}=M_{n,m}\mathrm{d}\nu$
+with $0\leqslant M_{n,n}\leqslant 1$ and $M_{n,m}=\overline{M_{m,n}}$ $\mathrm{d}\nu$-almost everywhere.
+Let us define the operator valued function $M:(x,\xi)\mapsto M(x,\xi)$ such that, for all $(n,m)\in\mathbb{N}^{2}$, for all $(x,\xi)\in\mathbb{R}^{2}$,
+$\left\langle M(x,\xi)h_{n},h_{m}\right\rangle_{\mathcal{H}}=M_{n,m}(x,\xi).$
+Because of the positivity of $\left(\mathrm{d}\mu_{n,n}\right)_{n\in\mathbb{N}}$ and the adjoint relation of $\left(\mathrm{d}\mu_{n,m}\right)_{(n,m)\in\mathbb{N}^{2}}$, $M(x,\xi)$ is a positive self-adjoint operator and its trace satisfies
+$\int_{\mathbb{R}^{2}}\mathrm{Tr}_{\mathcal{L}(\mathcal{H})}\Big(M(x,\xi)\Big)\mathrm{d}\nu(x,\xi)=\sum_{n\in\mathbb{N}}\int_{\mathbb{R}^{2}}M_{n,n}(x,\xi)\mathrm{d}\nu(x,\xi)=\sum_{n\in\mathbb{N}}\mathrm{d}\nu_{n,n}\left(\mathbb{R}^{2}\right)<+\infty.$
+Therefore, the pair $\left(M,\mathrm{d}\nu\right)$ generates an operator measure $M\mathrm{d}\nu$. We now have to prove that $M\mathrm{d}\nu$ plays the expected role.
 Let $a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathcal{K}\left(\mathcal{H}\right)\right)$, $\Pi_{\leqslant n}$ the projector on the finite dimensional subspace of $\mathcal{H}$ generated by $(h_{m})_{m\leqslant n}$ and consider the finite rank operator-valued function $a_{n}\coloneqq\Pi_{\leqslant n}a\Pi_{\leqslant n}$ which belongs to $\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathcal{K}\left(\mathcal{H}\right)\right)$, then
-$\sup_{(x,\xi)\in\mathrm{supp}(a)}\left\lVert a(x,\xi)-a_{n}(x,\xi)\right\rVert_{\mathcal{L}\left(\mathcal{H}\right)}\underset{n\to+\infty}{\longrightarrow}0.$ (37)
+$\sup_{(x,\xi)\in\mathrm{supp}(a)}\left\lVert a(x,\xi)-a_{n}(x,\xi)\right\rVert_{\mathcal{L}\left(\mathcal{H}\right)}\underset{n\to+\infty}{\longrightarrow}0.$ (A.2)
 By passing to the limit, thanks to the construction of $\left(\varepsilon_{k}\right)_{k\in\mathbb{N}}$,
 $\left\langle\mathrm{Op}_{\varepsilon_{k}}\left(a_{n}\right)f^{\varepsilon_{k}},f^{\varepsilon_{k}}\right\rangle$ $=\sum_{m,\ell\leqslant n}\left\langle\mathrm{Op}_{\varepsilon_{k}}\Big(\left\langle a_{n}h_{m},h_{\ell}\right\rangle_{\mathcal{H}}\Big)f^{\varepsilon_{k}}_{m},f^{\varepsilon_{k}}_{\ell}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R},\mathbb{C}^{2}\right)}$
-$\underset{k\to+\infty}{\longrightarrow}\sum_{m,\ell\leqslant n}\int_{\mathbb{R}^{2}}\left\langle a_{n}(x,\xi)h_{m},h_{\ell}\right\rangle_{\mathcal{H}}\mathrm{d}\gamma_{m,\ell}(x,\xi)$
-$=\sum_{m,\ell\leqslant n}\int_{\mathbb{R}^{2}}\left\langle a_{n}h_{m},h_{\ell}\right\rangle_{\mathcal{H}}\Gamma_{m,\ell}\mathrm{d}\gamma(x,\xi)$
-$=\int_{\mathbb{R}^{2}}\mathrm{Tr}_{\mathcal{L}(\mathcal{H})}\Big{(}a_{n}(x,\xi)\Gamma(x,\xi)\Big{)}\mathrm{d}\gamma(x,\xi).$
-Then, with (37) and Lebesgue dominated convergence Theorem, we have
-$\int_{\mathbb{R}^{2}}\mathrm{Tr}_{\mathcal{L}(\mathcal{H})}\Big{(}\big{(}a(x,\xi)-a_{n}(x,\xi)\big{)}\Gamma(x,\xi)\Big{)}\mathrm{d}\gamma(x,\xi)\underset{n\to+\infty}{\longrightarrow}0.$
-B.2 Two-scale semiclassical measures
-In this section, we prove Theorem 3.3.
-Proof of Theorem 3.3.
-Let ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$ uniformly bounded in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$.
-For simplicity, we assume that the sequence $\left(\varepsilon\right)_{\varepsilon>0}$ realizes the semiclassical measure $\mathrm{d}\mu$ of ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$.
-We use the decomposition (17).
-First we analysis of the concentration of $a^{\delta}$.
-  • $\bullet$
-    Construction of $\mathrm{d}\mu$ outside $\mathcal{C}$
-For $a\in\mathfrak{A}$ such that $a=0$ in a neighborhood of $\mathcal{C}$, then for $\varepsilon$ small enough, we have
-$\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a\right)$ $=\mathrm{Op}_{\varepsilon}\left(a\left(s,y;\sigma,\eta;\frac{y}{\sqrt{\varepsilon}},\frac{\sigma}{\sqrt{\varepsilon}},\frac{\eta}{\sqrt{\varepsilon}}\right)\right)$
-$=\mathrm{Op}_{1}\left(a_{\infty}\left(s,y,\varepsilon\sigma,\varepsilon\eta,\frac{(y,\varepsilon\sigma,\varepsilon\eta)}{|(y,\varepsilon\sigma,\varepsilon\eta)|}\right)\right).$
-After extraction of a subsequence that we still denote by $(\varepsilon)_{\varepsilon>0}$, we have
-$\limsup_{\delta\to 0}\limsup_{R\to+\infty}\limsup_{\varepsilon\to 0}\mathcal{I}^{\varepsilon}_{f}\left(a^{\delta}\right)=\int_{\mathbb{R}^{4}\backslash\mathcal{C}}\mathrm{tr}\left(a_{\infty}\left(s,y,\sigma,\eta,\frac{(y,\sigma,\eta)}{|(y,\sigma,\eta)|}\right)\mathrm{d}\mu(s,y,\sigma,\eta)\right).$
-Then we construct the pair $\Gamma\mathrm{d}\gamma$.
-  • $\bullet$
-    Construction of $\Gamma\mathrm{d}\gamma$
-Let $a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{7},\mathbb{C}^{2,2}\right)$, since $a$ is compactly supported, we recall (21)
-$\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a\right)=\mathfrak{T}_{\sqrt{\varepsilon}}^{\dagger}\mathrm{Op}_{\sqrt{\varepsilon},1}(b)\mathfrak{T}_{\sqrt{\varepsilon}}+\mathcal{O}\left(\sqrt{\varepsilon}\right),$
-The function $b$ is valued in the set of compact operators on $\mathcal{H}$.
-We can see ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$ as a bounded family of $\mathrm{L}^{2}\left(\mathbb{R}_{s},\mathcal{H}\right)$.
-We are thus doing classical calculus on $\mathrm{L}^{2}\left(\mathbb{R}_{s},\mathcal{H}\right)$.
-We apply Lemma B.1, then there exists $\left(\varepsilon_{k}\right)_{k\in\mathbb{N}}$ a vanishing sequence such that for all $a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{7},\mathbb{C}^{2,2}\right)$,
-$\lim_{k\to+\infty}\left\langle\mathfrak{Op}_{\varepsilon_{k}}\left(a\right)f^{\varepsilon_{k}},f^{\varepsilon_{k}}\right\rangle_{\mathrm{L}^{2}(\mathbb{R}^{2},\mathbb{C}^{2})}=\int_{\mathbb{R}^{2}}\mathrm{Tr}\Big{(}a^{W}(s,0;0,0;y,\sigma,\mathrm{D}_{y})\Gamma(s,\sigma)\Big{)}\mathrm{d}\gamma(s,\sigma).$
-Then, we apply this property to $a_{\varepsilon,R}$. Finally we construct the pair $M_{\infty}\mathrm{d}\nu_{\infty}$.
-  • $\bullet$
-    Construction of $M_{\infty}\mathrm{d}\nu_{\infty}$
-Let $a\in\mathfrak{A}$, then we recall (20)
-$\begin{split}\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a_{\varepsilon,\delta}^{R}\right)=&\ \mathfrak{T}_{\sqrt{\varepsilon}}^{\dagger}\mathrm{Op}_{1}\Bigg{(}a(s,\sqrt{\varepsilon}y,\varepsilon\sigma,\sqrt{\varepsilon}\eta,y,\sqrt{\varepsilon}\sigma,\eta)\chi\left(\frac{|(\sqrt{\varepsilon}y,\varepsilon\sigma,\sqrt{\varepsilon}\eta)|}{\delta}\right)\\ &\times\left(1-\chi\left(\frac{|(y,\sqrt{\varepsilon}\sigma,\eta)|}{R}\right)\right)\Bigg{)}\mathfrak{T}_{\sqrt{\varepsilon}}\\ =&\ \mathfrak{T}_{\sqrt{\varepsilon}}^{\dagger}\mathrm{Op}_{1}\underbrace{\left(a(s,0,0,0,y,\sqrt{\varepsilon}\sigma,\eta)\chi\left(\frac{|(\sqrt{\varepsilon}y,\varepsilon\sigma,\sqrt{\varepsilon}\eta)|}{\delta}\right)\left(1-\chi\left(\frac{|(y,\sqrt{\varepsilon}\sigma,\eta)|}{R}\right)\right)\right)}_{\eqqcolon\widetilde{a}^{R}_{\delta}(s,0,0,0,y,\sqrt{\varepsilon}\sigma,\eta)}\mathfrak{T}_{\sqrt{\varepsilon}}\\ &+\mathcal{O}\left(\delta\right).\end{split}$
-therefore we consider the family $\widetilde{f}^{\varepsilon}\coloneqq\mathfrak{T}_{\sqrt{\varepsilon}}f^{\varepsilon}$ which is bounded in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$.
-Then the quantity
-$\mathscr{I}^{\varepsilon_{k}}_{\widetilde{f}}\left(\widetilde{a}^{R}_{\delta}\right)\coloneqq\left\langle\mathrm{Op}_{1}\left(\widetilde{a}^{R}_{\delta}(s,0,0,0,y,\sqrt{\varepsilon_{k}}\sigma,\eta)\right)\widetilde{f}^{\varepsilon_{k}},\widetilde{f}^{\varepsilon_{k}}\right\rangle_{\mathrm{L}^{2}(\mathbb{R}^{2},\mathbb{C}^{2})}.$
-is uniformly bounded in $\varepsilon>0$ and $R>1$. $a^{R}_{\delta}$ is supported at distance at least $R\sqrt{\varepsilon}$ of $\mathcal{C}$, so
-$\mathscr{I}^{\varepsilon_{k}}_{\widetilde{f}}\left(\widetilde{a}^{R}_{\delta}\right)=\left\langle\mathrm{Op}_{1}\left(a_{\infty}(s,0,0,0,y,\sqrt{\varepsilon_{k}}\sigma,\eta)\right)\widetilde{f}^{\varepsilon_{k}},\widetilde{f}^{\varepsilon_{k}}\right\rangle_{\mathrm{L}^{2}(\mathbb{R}^{2},\mathbb{C}^{2})},$
-as soon as $R$ is large enough, depending of $\varepsilon$.
-We then deduce by a diagonal extraction argument that there exists a subsequence $\left(\varepsilon_{n}\right)_{n\in\mathbb{N}}$ of $\left(\varepsilon_{k}\right)_{k\in\mathbb{N}}$, a sequence $\left(R_{n}\right)_{n\in\mathbb{N}}$ and a sequence $\left(\delta_{n}\right)_{n\in\mathbb{N}}$ such that for all $a\in\mathfrak{A}$,
-$\mathscr{I}^{\varepsilon_{n}}_{\widetilde{f}}\left(\widetilde{a}^{R_{n}}_{\delta_{n}}\right)\underset{n\to+\infty}{\longrightarrow}\mathscr{I}\left(a_{\infty}\right).$
-It remains to prove that $a_{\infty}\mapsto\mathscr{I}\left(a_{\infty}\right)$ is a measure, which will allow us to define $M_{\infty}\mathrm{d}\nu_{\infty}$.
-First, we prove that $a_{\infty}\mapsto\mathscr{I}\left(a_{\infty}\right)$ is a non negative matrix-valued distribution.
-We observe that there exists $C>0$ such that for all $a\in\mathfrak{A}$, for all $n\in\mathbb{N}$,
-$\mathscr{I}^{\varepsilon_{n}}_{\widetilde{f}}\left(\widetilde{a}^{R_{n}}_{\delta_{n}}\right)\leqslant C\mathrm{N}_{d}\left(\widetilde{a}^{R_{n}}_{\delta_{n}}\right),$
-with $\mathrm{N}_{d}$ a Schwarz semi-norm.
-We have $\mathrm{N}_{d}\left(\widetilde{a}^{R_{n}}_{\delta_{n}}\right)\to\mathrm{N}_{d}\left(a_{\infty}\right)$ then $\mathscr{I}\left(a_{\infty}\right)\leqslant C\mathrm{N}_{d}\left(a_{\infty}\right)$.
-Therefore $\mathscr{I}$ is a matrix-valued distribution.
-Secondly, we observe that the operators $a\mapsto\mathrm{Op}_{1}\left(a^{R}_{\delta}(s,0,0,0,y,\sqrt{\varepsilon}\sigma,\eta)\right)$ satisfy a semiclassical calculus in the parameters $\varepsilon$, $1/R$ and $\delta$.
-We have for all $a,a_{1},a_{2}\in\mathfrak{A}$,
-  • $\bullet$
-    Adjoint : $\mathrm{Op}_{1}\left(\widetilde{a}^{R}_{\delta}(s,0,0,0,y,\sqrt{\varepsilon}\sigma,\eta)\right)^{\dagger}=\mathrm{Op}_{1}\left(\overline{\widetilde{a}^{R}_{\delta}}(s,0,0,0,y,\sqrt{\varepsilon}\sigma,\eta)\right)$
-  • $\bullet$
-    Symbolic Calculus : as operator of $\mathcal{L}\left(\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)$,
-    $\mathrm{Op}_{1}\Big{(}\widetilde{a}^{R}_{1,\delta}\Big{)}\mathrm{Op}_{1}\Big{(}\widetilde{a}^{R}_{2,\delta}\Big{)}=\mathrm{Op}_{1}\Big{(}\left(\widetilde{a}_{1}\widetilde{a}_{2}\right)^{R}_{1,\delta}\Big{)}+\mathcal{O}\left(\sqrt{\varepsilon}\right)+\mathcal{O}\left(\frac{1}{R}\right),$
-  • $\bullet$
-    Weak Gårding inequality : if $a\geqslant 0$, then for all $\widetilde{\delta}$, there exists $C_{\widetilde{\delta}}>0$, such that for all $f\in\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$,
-    $\left\langle\mathrm{Op}_{1}\Big{(}\widetilde{a}^{R}_{\delta}\Big{)}\widetilde{f},\widetilde{f}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\geqslant-\left(\widetilde{\delta}+C_{\widetilde{\delta}}\left(\varepsilon+\frac{1}{R}\right)\right)\left\lVert f\right\rVert^{2}_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)},$
-Therefore $a_{\infty}\mapsto\mathscr{I}\left(a_{\infty}\right)$ is a non negative matrix-valued measure defined on ×$\mathbb{R}\times\mathbb{S}^{2}$ that we denote $M\mathrm{d}\nu$. ∎
-Appendix C Spectral Analysis of the operator $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$
-In this appendix, we study the spectrum of the operator $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$ introduced in (34).
-Theorem C.1 (Spectrum of $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$).
-For all $(s,\sigma)\in\mathbb{R}^{2}$, the spectrum of $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$ consists in multiplicity one eigenvalues
-$\mathrm{Sp}\Big{(}\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)\Big{)}=\left\{\lambda_{0}(\sigma)\right\}\bigcup\left\{\lambda_{n}(s,\sigma)\ |\ n\in\mathbb{Z}_{\neq 0}\right\},$
-with for all $n\in\mathbb{Z}_{\neq 0}$,
-$\lambda_{0}(\sigma)\coloneqq-\sigma\;,\;\lambda_{n}(s,\sigma)\coloneqq\mathrm{sgn}(n)\sqrt{\sigma^{2}+2|n|r(s)^{2}}.$
-Moreover, an Hilbertian basis of $\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)$, composed of eigenfunctions of $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$, is given by ${\left(g_{n}^{s,\sigma}\right)_{n\in\mathbb{Z}}}$
-$g_{0}^{s}(y)$ $\coloneqq\frac{1}{\sqrt{r(s)}}\begin{pmatrix}0\\ \mathfrak{h}_{0}\big{(}r(s)y\big{)}\end{pmatrix},$
-$g_{n}^{s,\sigma}(y)$ $\coloneqq\alpha_{n}(s,\sigma)\begin{pmatrix}\mathfrak{h}_{|n|-1}\big{(}r(s)y\big{)}\\ \frac{\sqrt{2|n|}r(s)}{\lambda_{n}(\sigma)+\sigma}\mathfrak{h}_{|n|}\big{(}r(s)y\big{)}\end{pmatrix},\quad\quad n\in\mathbb{Z}_{\neq 0}$
-$\mbox{with}\quad\quad\alpha_{n}(s,\sigma)$ $\coloneqq\frac{1}{\sqrt{2r(s)}}\sqrt{1+\frac{\sigma}{\lambda_{n}(s,\sigma)}},$
-and $\left(\mathfrak{h}_{n}\right)_{n\in\mathbb{N}}$ are Hilbert functions defined in (27). Moreover, the projectors over the eigenspaces of $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$ defined by
+$\underset{k\to+\infty}{\longrightarrow}\sum_{m,\ell\leqslant n}\int_{\mathbb{R}^{2}}\left\langle a_{n}(x,\xi)h_{m},h_{\ell}\right\rangle_{\mathcal{H}}\mathrm{d}\nu_{m,\ell}(x,\xi)$
+$=\sum_{m,\ell\leqslant n}\int_{\mathbb{R}^{2}}\left\langle a_{n}h_{m},h_{\ell}\right\rangle_{\mathcal{H}}M_{m,\ell}\mathrm{d}\nu(x,\xi)$
+$=\int_{\mathbb{R}^{2}}\mathrm{Tr}_{\mathcal{L}(\mathcal{H})}\Big(a_{n}(x,\xi)M(x,\xi)\Big)\mathrm{d}\nu(x,\xi).$
+Then, with (A.2) and Lebesgue dominated convergence Theorem, we have
+$\int_{\mathbb{R}^{2}}\mathrm{Tr}_{\mathcal{L}(\mathcal{H})}\Big(\big(a(x,\xi)-a_{n}(x,\xi)\big)M(x,\xi)\Big)\mathrm{d}\nu(x,\xi)\underset{n\to+\infty}{\longrightarrow}0.$
+Appendix B Spectral Analysis of the principal normal form operator
+In this appendix, we study the operator $\mathrm{T}_{\mathrm{E}}(s,\sigma)$ introduced in (1.12).
+First, we rescale the operator by $r(s)$.
+This is possible because $r$ is smooth ($\mathrm{m}$ is in $\mathscr{C}^{\infty}\left(\mathbb{R}^{2},\mathbb{R}\right)$) and strictly non-negative (Assumption 1.1).
+Let us define
+$\mathscr{T}\coloneqq\begin{pmatrix}\zeta&\mathfrak{a}^{\dagger}\\ \mathfrak{a}&-\zeta\end{pmatrix}$,
+where $\mathfrak{a}\coloneqq y+\mathrm{i}\mathrm{D}_{y}$ denotes the usual annihilation operator on the variable $y$.
+Remark B.1.
+By denoting by $\zeta$ the ratio $\sigma/r(s)$, we have the following, for all $f\in\mathrm{L}^{2}\left(\mathbb{R},\mathbb{C}^{2}\right)$,
+$\mathrm{T}_{\mathrm{E}}(s,\sigma)\Bigg[f\Big(r(s)y\Big)\Bigg]$ $=\left[\begin{pmatrix}\sigma&r(s)(y-\mathrm{i}\mathrm{D}_{y})\\ r(s)(y+\mathrm{i}\mathrm{D}_{y})&-\sigma\end{pmatrix}f\right]\Big(r(s)y\Big)$
+$=r(s)\left[\begin{pmatrix}\zeta&y-\mathrm{i}\mathrm{D}_{y}\\ y+\mathrm{i}\mathrm{D}_{y}&-\zeta\end{pmatrix}f\right]\Big(r(s)y\Big).$
+Then the link between spectra is
+$\mathrm{Sp}\Big(\mathrm{T}_{\mathrm{E}}(s,\sigma)\Big)=r(s)\mathrm{Sp}(\mathscr{T}).$
+So we can study $\mathscr{T}$ instead of $\mathrm{T}_{\mathrm{E}}(s,\sigma)$.
+Let us recall usual properties of creation and annihilation operators.
+Proposition B.2.
+We have the following.
+  • $\mathfrak{a}\mathfrak{a}^{\dagger}=-\partial_{x}^{2}+x^{2}+1$.
+  • $\operatorname{sp}\left(\mathfrak{a}\mathfrak{a}^{\dagger}\right)=\{2n\mid n\in\mathbb{N}_{>0}\}$.
+  • The Hermite functions $(\mathfrak{h}_{n})_{n\in\mathbb{N}}$, defined in (1.13), form an Hilbertian basis of $\mathrm{L}^{2}(\mathbb{R},\mathbb{C})$, of eigenfunctions of $\mathfrak{a}\mathfrak{a}^{\dagger}$, and we have
+    $\mathfrak{a}\mathfrak{a}^{\dagger}\mathfrak{h}_{n}=2(n+1)\mathfrak{h}_{n},\quad n\in\mathbb{N}.$
+  • Moreover $\mathfrak{a}$ and $\mathfrak{a}^{\dagger}$ satisfies the following identities
+    $\mathfrak{a}\mathfrak{h}_{n}=\sqrt{2n}\mathfrak{h}_{n-1},\quad\mathfrak{a}^{\dagger}\mathfrak{h}_{n}=\sqrt{2(n+1)}\mathfrak{h}_{n+1}.$
+The next Proposition consists to diagonalize, at fixed $(s,\sigma)$, the operator $\mathrm{T}_{\mathrm{E}}(s,\sigma)$ on $\mathrm{L}^{2}(\mathbb{R}_{y},\mathbb{C}^{2})$ since $(\mathrm{T}_{\mathrm{E}}(s,\sigma),\mathcal{B}^{1}(\mathbb{R}_{y},\mathbb{C}^{2}))$ is self-adjoint with compact resolvent.
+Theorem B.3 (Spectrum of $\mathrm{T}_{\mathrm{E}}(s,\sigma)$).
+The spectrum of $\mathrm{T}_{\mathrm{E}}(s,\sigma)$ consists in multiplicity one eigenvalues
+$\mathrm{Sp}\Big(\mathrm{T}_{\mathrm{E}}(s,\sigma)\Big)=\left\{\lambda_{0}(\sigma)\right\}\bigcup\left\{\lambda_{n}(s,\sigma)\ |\ n\in\mathbb{Z}_{\neq 0}\right\},\quad(s,\sigma)\in\mathbb{R}^{2},$
+with, for all $n\in\mathbb{Z}_{\neq 0}$,
+$\lambda_{0}(\sigma)\coloneqq\sigma,\quad\lambda_{n}(s,\sigma)\coloneqq\mathrm{sgn}(n)\sqrt{\sigma^{2}+2|n|r(s)^{2}}.$
+Moreover, an Hilbertian basis of $\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)$, composed of eigenfunctions of $\mathrm{T}_{\mathrm{E}}(s,\sigma)$, is given by ${\left(g_{n}^{s,\sigma}\right)_{n\in\mathbb{Z}}}$
+$g_{0}^{s}(y)$ $\coloneqq\frac{1}{\sqrt{r(s)}}\begin{pmatrix}\mathfrak{h}_{0}\big(r(s)y\big)\\ 0\end{pmatrix}$,
+$g_{n}^{s,\sigma}(y)$ $\coloneqq\alpha_{n}(s,\sigma)\begin{pmatrix}\frac{\sqrt{2|n|}r(s)}{\lambda_{n}(\sigma)-\sigma}\mathfrak{h}_{|n|}\big(r(s)y\big)\\ \mathfrak{h}_{|n|-1}\big(r(s)y\big)\end{pmatrix},\quad\quad n\in\mathbb{Z}_{\neq 0}$
+$\mbox{with}\quad\quad\alpha_{n}(s,\sigma)$ $\coloneqq\frac{1}{\sqrt{2r(s)}}\sqrt{1-\frac{\sigma}{\lambda_{n}(s,\sigma)}},$
+and $\left(\mathfrak{h}_{n}\right)_{n\in\mathbb{N}}$ are Hermite functions defined in (1.13). Moreover, the projectors over the eigenspaces of $\mathrm{T}_{\mathrm{E}}(s,\sigma)$ defined by
 $\Pi_{0}(s,\sigma)\coloneqq g_{0}^{s}\otimes g_{0}^{s},\quad\quad\Pi_{n}(s,\sigma)\coloneqq g_{n}^{s,\sigma}\otimes g_{n}^{s,\sigma},$
@@
 Notably,
-$\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)=\sum_{n\in\mathbb{Z}}\lambda_{n}(s,\sigma)\Pi_{n}(s,\sigma).$
-Proof of Theorem C.1.
-First, we will rescale the function in order to have the form $f\left(s,r(s)y\right)$.
-We can do it because $r$ is smooth ($\mathrm{m}$ is in $\mathscr{C}^{\infty}\left(\mathbb{R}^{2},\mathbb{R}\right)$) and strictly non negative (Assumption 1.2).
-Then we dilate with $r(s)^{-1}$ and we denote by $\zeta$ the ratio $\sigma/r(s)$.
-For all $f\in\mathrm{L}^{2}\left(\mathbb{R},\mathbb{C}^{2}\right)$,
-$\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)\Bigg{[}f\Big{(}r(s)y\Big{)}\Bigg{]}$ $=\left[\begin{pmatrix}\sigma&r(s)(y+\mathrm{i}\mathrm{D}_{y})\\ r(s)(y-\mathrm{i}\mathrm{D}_{y})&-\sigma\end{pmatrix}f\right]\Big{(}r(s)y\Big{)}$
-$=r(s)\left[\begin{pmatrix}\zeta&y+\mathrm{i}\mathrm{D}_{y}\\ y-\mathrm{i}\mathrm{D}_{y}&-\zeta\end{pmatrix}f\right]\Big{(}r(s)y\Big{)}.$
-Let us define
-$\mathscr{T}\coloneqq\begin{pmatrix}\zeta&\mathfrak{a}\\ \mathfrak{a}^{\dagger}&-\zeta\end{pmatrix},$
-where $\mathfrak{a}\coloneqq y+\mathrm{i}\mathrm{D}_{y}$ denotes the usual annihilation operator on the variable $y$. Then the link between the spectrum is
-$\mathrm{Sp}\Big{(}\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)\Big{)}=r(s)\mathrm{Sp}(\mathscr{T})$
-So we can study $\mathscr{T}$ instead of $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$.
-Let $\lambda_{\zeta}\in\mathrm{Sp}(\mathscr{T}),\begin{pmatrix}f\\ g\end{pmatrix}\in E_{\lambda_{\zeta}}(\mathscr{T})$ if and only if
-$\left\{\begin{matrix}2\zeta f&+&\mathfrak{a}g&=&\left(\lambda_{\zeta}+\zeta\right)f\\ &&\mathfrak{a}^{\dagger}f&=&\left(\lambda_{\zeta}+\zeta\right)g\end{matrix}\right.$
-If $\lambda_{\zeta}=-\zeta$, the injectivity of $\mathfrak{a}^{\dagger}$ on $\mathrm{L}^{2}\left(\mathbb{R}\right)$ implies $f=0$ whence $\mathfrak{a}g=0$ so $g=A\mathfrak{h}_{0}$ for some $A\in\mathbb{C}$.
-So $-\zeta\in\mathrm{Sp}(\mathscr{T})$ and is a simple eigenvalue with ${\begin{pmatrix}0\\ \mathfrak{h}_{0}\end{pmatrix}}$ a normalized eigenfunction.
-If $\lambda_{\zeta}\neq-\zeta$, then
-$\left\{\begin{matrix}\mathfrak{a}\mathfrak{a}^{\dagger}f&=&\left(\lambda_{\zeta}^{2}-\zeta^{2}\right)f\\ g&=&\frac{1}{\lambda_{\zeta}+\zeta}\mathfrak{a}^{\dagger}f\end{matrix}\right.$
-and eigenvalues of $\mathfrak{a}\mathfrak{a}^{\dagger}$ are known so we deduce that there exists $n\in\mathbb{N}_{>0},\lambda_{\zeta}^{2}-\zeta^{2}=2n$ so
-$\lambda_{\zeta}\in\left\{\mathrm{sgn}(n)\sqrt{\zeta^{2}+2|n|}\ |\ n\in\mathbb{Z}_{\neq 0}\right\}=\mathrm{Sp}\left(\mathfrak{a}\mathfrak{a}^{\dagger}\right),$
+$\mathrm{T}_{\mathrm{E}}(s,\sigma)=\sum_{n\in\mathbb{Z}}\lambda_{n}(s,\sigma)\Pi_{n}(s,\sigma).$
+Proof.
+According to Remark B.1, we study $\mathscr{T}$. Let $\lambda_{\zeta}\in\mathrm{Sp}(\mathscr{T}),\begin{pmatrix}f\\ g\end{pmatrix}\in E_{\lambda_{\zeta}}(\mathscr{T})$ if and only if
+$\left\{\begin{matrix}&&\mathfrak{a}^{\dagger}g&=&\left(\lambda_{\zeta}-\zeta\right)f\\ \mathfrak{a}f&-&2\zeta g&=&\left(\lambda_{\zeta}-\zeta\right)g\end{matrix}\right$.
+If $\lambda_{\zeta}=\zeta$, then the injectivity of $\mathfrak{a}^{\dagger}$ on $\mathrm{L}^{2}\left(\mathbb{R}\right)$ implies $g=0$ whence $\mathfrak{a}f=0$ so $f=A\mathfrak{h}_{0}$ for some $A\in\mathbb{C}$.
+So $\zeta\in\mathrm{Sp}(\mathscr{T})$ and is a simple eigenvalue with ${\begin{pmatrix}\mathfrak{h}_{0}\\ 0\end{pmatrix}}$ a normalized eigenfunction.
+If $\lambda_{\zeta}\neq\zeta$, then
+$\left\{\begin{matrix}f&=&\frac{1}{\lambda_{\zeta}-\zeta}\mathfrak{a}^{\dagger}g\\ \mathfrak{a}\mathfrak{a}^{\dagger}g&=&\left(\lambda_{\zeta}^{2}-\zeta^{2}\right)g\end{matrix}\right$.
+so, according to Proposition B.2, we deduce that there exists $n\in\mathbb{N}_{>0},\lambda_{\zeta}^{2}-\zeta^{2}=2n$ so
+$\lambda_{\zeta}\in\left\{\mathrm{sgn}(n)\sqrt{\zeta^{2}+2|n|}\ |\ n\in\mathbb{Z}_{\neq 0}\right\},$
 and for some $B\in\mathbb{C}$,
-$\forall(\zeta,y)\in\mathbb{R}^{2},\ \begin{pmatrix}f_{n}\\ g_{n}\end{pmatrix}(\zeta,y)\coloneqq B\begin{pmatrix}\mathfrak{h}_{|n|-1}(y)\\ \frac{\sqrt{2|n|}}{\mathrm{sgn}(n)\sqrt{\zeta^{2}+2|n|}+\zeta}\mathfrak{h}_{|n|}(y)\end{pmatrix}.$
+$\forall(\zeta,y)\in\mathbb{R}^{2},\ \begin{pmatrix}f_{n}\\ g_{n}\end{pmatrix}(\zeta,y)\coloneqq B\begin{pmatrix}\frac{\sqrt{2|n|}}{\mathrm{sgn}(n)\sqrt{\zeta^{2}+2|n|}-\zeta}\mathfrak{h}_{|n|}(y)\\ \mathfrak{h}_{|n|-1}(y)\end{pmatrix}$.
 Then
-$\mathrm{Sp}\Big{(}\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)\Big{)}=\left\{\lambda_{0}(\sigma)\right\}\bigcup\left\{\lambda_{n}(s,\sigma)\ |\ n\in\mathbb{Z}_{\neq 0}\right\}.$
-The associated projectors $\left(\Pi_{n}\right)_{n\in\mathbb{Z}}$ are smooth because $r$ is smooth ($\mathrm{m}$ is in $\mathscr{C}^{\infty}\left(\mathbb{R}^{2},\mathbb{R}\right)$) and strictly non negative (Assumption 1.2); for all $n\in\mathbb{Z}_{\neq 0}$, $\lambda_{n}$ and $\lambda_{n}+\sigma$ is smooth and never vanishes.
+$\mathrm{Sp}\Big(\mathrm{T}_{\mathrm{E}}(s,\sigma)\Big)=\left\{\lambda_{0}(\sigma)\right\}\bigcup\left\{\lambda_{n}(s,\sigma)\ |\ n\in\mathbb{Z}_{\neq 0}\right\}.$
+The associated projectors $\left(\Pi_{n}\right)_{n\in\mathbb{Z}}$ are smooth because $r$ is smooth ($\mathrm{m}$ is in $\mathscr{C}^{\infty}\left(\mathbb{R}^{2},\mathbb{R}\right)$) and strictly non-negative (Assumption 1.1); for all $n\in\mathbb{Z}_{\neq 0}$, $\lambda_{n}$ and $\lambda_{n}+\sigma$ is smooth and never vanishes.
 ∎
+Acknowledgments
+The author is very grateful to his advisors Clotilde Fermanian-Kammerer and Nicolas Raymond for many insightful discussions and valuable suggestions.
+The author acknowledges the support of the Région Pays de la Loire via the Connect Talent Project HiFrAn 2022 07750, and from the France 2030 program, Centre Henri Lebesgue ANR-11-LABX-0020-01.